### v7

# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

Abstract
[Objective] This study focuses on addressing the current lack of a unified formal theoretical framework in machine learning, as well as the deficiencies in interpretability and ethical safety assurance.
[Methods] A formal information model is first constructed, utilizing sets of well-formed formulas to explicitly define the ontological states and carrier mappings of typical components in machine learning. Learnable and processable predicates, along with learning and processing functions, are introduced to analyze the logical deduction and constraint rules of the causal chains within models.
[Results] A meta-framework for machine learning theory (MLT-MF) is established. Based on this framework, universal definitions for model interpretability and ethical safety are proposed. Furthermore, three key theorems are proved: the equivalence of model interpretability and information recoverability, the assurance of ethical safety, and the estimation of generalization error.
[Limitations] The current framework assumes ideal conditions with noiseless information-enabling mappings and primarily targets model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing the critical challenges currently faced in machine learning.
Keywords: machine learning; objective information theory; formal logic systems; interpretability; ethical and safety assurance; generalization error
1 Introduction
Although artificial intelligence is currently the most dynamic field within information science and even the broader domain of science and technology, existing research still faces several significant shortcomings:
First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning [1, 2, 3, 4, 5]. Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages. This fragmentation makes it impossible to characterize the coherent evolution of knowledge from prior injection, through data-driven updates, to task generalization. For example, the PAC-Bayes framework connects priors with generalization error but neglects the temporal causality of the training process; meta-learning involves multi-stage adaptation but fails to establish a symbolic model for state transitions. The absence of lifecycle modeling makes systematic analysis of the core question——“how parameters carry task knowledge from initialization through training”—difficult for model designers.
Second, there is a lack of explicit expression for the causal chain of information flow. Traditional approaches often treat the transformation from data to model as a black box, reducing the parameter update driven by loss function gradients to a purely mathematical optimization process and overlooking the causal dependencies among prior knowledge, data distribution, and model architecture. While the field of causal inference attempts to characterize causal relationships among variables, its application in machine learning is mostly limited to feature-label association analysis and fails to construct a complete causal chain from “prior constraints → initialization → gradient propagation → generalization capability.” For instance, adversarial training enhances robustness by maximizing loss, but theoretical analyses rarely explicitly link initial parameter constraints with the causal mechanisms underlying final adversarial robustness, making strategy design reliant on empirical trial-and-error.
Third, the mathematical foundations of interpretability tools are weak [11, 12, 13, 14, 15]. Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization. For example, SHAP values compute marginal contributions in cooperative games but do not formalize their mapping to the information carrier states of models, potentially violating causality in interpretations. In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learning[16, 17] . Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality. This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systems[18, 19, 20, 21, 22, 23].
On this basis, this paper rigorously defines higher-order logical systems using recursive methods and establishes a formalized expression of states via the interpretation of sets of well-formed formulas, thus fully clarifying the mathematical essence of state representation in the definition of information. This fills the last theoretical gap in OIT and provides more powerful tools for applying OIT to analyze various problems in information science. Subsequently, the theories and methods of OIT are applied comprehensively—from prior knowledge encoding and data-driven training to model generalization applications—unifying core concepts such as prior distribution, data distribution, and parameter updates into a formalized state representation for the first time. This establishes a mathematical logic system that spans the entire machine learning process. Based on this, formal logical deduction and constraint rules for the causal chains in typical stages such as formal modeling and realization of training information, formal modeling and realization of model learning, and information processing and learning optimization in models are demonstrated, forming the Meta-Framework for Machine Learning Theory (MLT-MF).
Finally, MLT-MF is applied to several hot topics in machine learning, establishing universal definitions for model interpretability and ethical safety, and proving three important theorems: the equivalence of model interpretability and information recoverability, ethical safety assurance, and generalization error estimation. This provides a powerful tool based on a unified formal logical framework for both theoretical analysis and engineering practice in machine learning.
2 Postulates and Formal Definition of Information
Undoubtedly, as a key component of artificial intelligence, modern machine learning is a product of the rapid development and high integration of information technologies such as computers, the Internet, big data, and cloud computing. Machine learning systems are, without question, an important branch of information systems. Starting from the basic concept of information can provide a unified theoretical foundation for research in machine learning and artificial intelligence.
2.1 Postulates of Information
Although humanity has entered the information age, the theoretical foundations of information science have long been dispersed across disciplines such as information theory, cybernetics, signal processing, statistical learning, and computational theory, each developing independently. There remains a lack of consensus on the concept of information, as well as a unified axiomatic basis and theoretical system. This is also a major reason why artificial intelligence often lacks interpretability and predictability. In response,[22] proposes four Postulates of information:
Postulate 1 (Binary Attribute) Every piece of information comprises two components: the ontology, which represents the intrinsic essence of the information, and the carrier, which embodies the objective form of the information.
Postulate 2 (Existence Duration) The content and form of information have their own respective temporal existences.
Postulate 3 (State Representation) The states of the ontology and the carrier of information have their own respective representations.
Postulate 4 (Enabling Mapping) The state of the ontology can be mapped onto the state of the carrier via an enabling mapping.
These four postulates are concise, align with common conceptions of information, and are compatible with the basic concepts of major branches of information science. Their scientific significance is reflected in:
  • Unification: Providing a common foundation for long-dispersed theories in information science;
  • Axiomatization: Allowing the definition and reasoning of information to be axiomatized and formalized, akin to mathematics and physics;
  • Universality: Facilitating application to cutting-edge fields such as the Internet, big data, artificial intelligence, cognitive science, and cyber-physical systems.
The concept of "enabling mapping" is novel. It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realization[24, 25, 26].
2.2 Formal Representation of States
The postulates of information emphasize that the state of the ontology enables a mapping to the state of the carrier. Therefore, establishing a formal representation of both ontological and carrier states is particularly important.
Definition 1 (Higher-Order Formal Logical System): Suppose a formal system $\mathcal{L}$ is described using the following symbol set:
Variables $x_{1},x_{2},\ldots$;
Individual constants $a_{1},a_{2},\ldots$;
Functions $f_{1}^{1},f_{2}^{1},\ldots,f_{1}^{2},f_{2}^{2},\ldots,f_{1}^{3},f_{2}^{3},\ldots$;
Parentheses (,)(,)( , );
Predicates $A_{1}^{1},A_{2}^{1},\ldots,A_{1}^{2},A_{2}^{2},\ldots,A_{1}^{3},A_{2}^{3},\ldots$;
Logical connectives "not" $\sim$, "implication" $\rightarrow$;
Quantifier "for all" $\forall$;
A term in $\mathcal{L}$ is defined as follows:
(1) Variables and individual constants are terms;
(2) If $f_{i}^{n}(n>0,i>0)$ is a function in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $f_{i}^{n}$ itself, then $f_{i}^{n}(t_{1},\ldots,t_{n})$ is also a term in $\mathcal{L}$;
(3) If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is a term in $\mathcal{L}$.
(4) The set of all terms in $\mathcal{L}$ is generated by (1), (2), and (3).
Based on this, we can define atomic formulas in $\mathcal{L}$:
If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$ and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ not containing $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is an atomic formula in $\mathcal{L}$.
We then define well-formed formulas (WFFs) in $\mathcal{L}$:
(5) Every atomic formula in $\mathcal{L}$ is a WFF;
(6) If $\mathcal{A}$ and $\mathcal{B}$ are WFFs in $\mathcal{L}$, then $\sim\mathcal{A}$, $\mathcal{A}\rightarrow\mathcal{B}$, $(\forall x_{i})\mathcal{A}$, $(\forall f_{i}^{n})\mathcal{A}$, and $(\forall A_{i}^{n})\mathcal{A}$ are all WFFs in $\mathcal{L}$, where $x_{i}$, $f_{i}^{n}$, $A_{i}^{n}$ are variables, functions, and predicates in $\mathcal{L}$, respectively, and no predicate or function symbol is applied, either directly or indirectly, to itself.
(7) The set of all WFFs in $\mathcal{L}$ is generated by (5) and (6).
The higher-order formal logical system $\mathcal{L}$ defined here, based on the flexibility of predicate logic, can describe various relationships and constraints among objects in a unified way. Through higher-order quantification support, the nested definition of complex data and behaviors, and it explicitly prohibits anypredicate or function symbol from applying to itself—directly or indirectly—to avoid self-referential paradoxes.
For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more concise[27]. Therefore, they are introduced as defined symbols in subsequent discussions.
Definition 2 (Interpretation of a Formal System): An interpretation $E$ of formal system $\mathcal{L}$ over a domain $D_{E}$ consists of a non-empty set $D_{E}$, a specified set of elements $(\widehat{a}_{1},\widehat{a}_{2},\ldots)$ over $D_{E}$, a set of functions $(\widehat{f}_{i}^{n},n>0,i>0)$ over $D_{E}$, and a set of relations $(\widehat{A}_{i}^{n},n>0,i>0)$ over $D_{E}$ and the specified elements.
According to this definition, the variables $x_{1},x_{2},\ldots$ in $\mathcal{L}$ are interpreted as objects. The set $D_{E}$ is the domain over which variables take their values. The set of specfied elements corresponds to the special objects denoted by the individual constants in $\mathcal{L}$. Functions and relations provide concrete interpretations for the function and predicate symbols in $\mathcal{L}$.
Definition 2 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects as both the subjective and objective worlds.
Postulate 5 (Expression of Object State):The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
(1) Definitional Invariance: The definitions of the object set and the time set remain invariant over the entire domain.
(2) Expressivity of Attributes: All attributes of the object set—including properties, structures, values, and relations—over the domain can be expressed via functions and predicates within a formal system conforming to Definition 1.
(3) Expressivity of Rules: The evolution rules governing various attributes of the object set can be expressed, without implicit self-reference, via temporal variables and application connectives in a formal system as specified in Definition 1.
(4) Logical Consistency: There exists no logical contradiction among the various attributes and evolution rules of the object set, nor among any possible logical conjunctions thereof.
Theorem 1 (Formal Expression of States): Let $S(X,T)$ be the state set of an object collection $X$ over a specified time set $T$, satisfying Postulate 5. Then $S(X,T)$ can always be represented as the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$, and is logically consistent.
Proof: Since $S(X,T)$ satisfies Postulate 5, we have:
  • By (1), the elements of $X\times T$ are well-defined, so $S(X,T)$ has a definite domain.
  • By (2), all attributes of $X$ over $T$ in $S(X,T)$ can be expressed by functions and predicates in a formal system $\mathcal{L}$ as specified in Definition 1.
  • By (3), suppose $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ and $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$ are two attributes in $S(X,T)$ pertaining to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$. If every $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ evolves from $t_{1}$ to $t_{2}$ according to some rule into $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$,
    then $(\forall A_{1})(\forall f_{1})(\forall x_{1})(A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))\rightarrow\exists A_{2}\exists f_{2}\exists x_{2}(A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))))$ is a well-formed formula in $\mathcal{L}$ that, as an interpretation over $x_{1},x_{2},t_{1},t_{2}$, expresses the corresponding rule in $S(X,T)$. Other rules in $S(X,T)$ can likewise be expressed through corresponding well-formed formulas in $\mathcal{L}$ interpreted over $X\times T$.
  • Let the set of such well-formed formulas in $\mathcal{L}$ representing $S(X,T)$ be defined as its interpretation ${E}$ over the domain $X\times T$.
  • By (4), no logical contradiction exists among any attributes or rules in $S(X,T)$, nor among any logical conjunctions thereof. Thus, as a representation of $S(X,T)$, both ${E}$ and its logical closure are logically consistent.
In summary, $S(X,T)$ is indeed an interpretation ${E}$ of a set of well-formed formulas in $\mathcal{L}$ over the domain $X\times T$, and both itself and its logical closure are logically consistent. The theorem is thus proved.
Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world. Both theoretical and empirical analyses [28] further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language. This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
2.3 Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata.
Theorem 2 (State Representation of Finite Automata): For any finite automaton $M$, one can describe its input, output, and state-transition behavior over a related time set $T$ by the state set $S(M,T)$.
Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n_{T}\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions. By finite automaton theory [29], $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n_{T}-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n_{r}\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n_{u}\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
$\delta:U\times Q\to U$ is the next-state function;
$\lambda:U\times Q\to R$ is the output function.
For each state $u_{t_{i}}\in U$, define the state predicate $State^{3}$ by
$\varphi_{U}(t_{i})=State^{3}(M,t_{i},u_{t_{i}})$ (2.1)
to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$.
For each input $q_{t_{i}}\in Q$, define the input predicate $Input^{3}$ by
$\varphi_{I}(t_{i})=Input^{3}(M,t_{i},q_{t_{i}})$ (2.2)
to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$.
For each output $r_{t_{i}}\in R$, define the output predicate $Output^{3}$ by
$\varphi_{O}(t_{i})=Output^{3}(M,t_{i},r_{t_{i}})$ (2.3)
to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$.
The transition function $\delta$ is captured by the well-formed formula
$\varphi_{\delta}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land Eq^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (2.4)
where the binary predicate $Eq^{2}(x,y)$ asserts $x=y$.
Similarly, the output function $\lambda$ is given by
$\varphi_{\lambda}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land Eq^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land$ (2.5)
$IsElementof^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$
where $IsElementof^{2}(x,X)$ asserts $x$ is an element of $X$.
Thus the state set
$S(M,T)=Cn(\{\varphi_{U}(t_{i})\}\cup\{\varphi_{I}(t_{i})\}\cup\{\varphi_{O}(t_{i})\}\cup\{\varphi_{\delta}(t_{i})\}\cup\{\varphi_{\lambda}(t_{i})\})$ (2.6)
describes $M$’s inputs, outputs, and state-transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits atuomatic transtion behavior. Hence we introduce :
Definition 3 (Finite Automaton State): If a state set $S(X,T)$ expresses a finite automaton, then $S(X,T)$ is called a finite automaton state.
All practical information systems, including computing systems, are finite automata in the mathematical sense. Therefore, the concept of finite automaton state is of great significance in the study of machine learning.
2.4 Definition of Information
As shown in[22], the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$. Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
This leads to the general definition of information:
Definition 4 (Mathematical Definition of Information): Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, that is,
$I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ (2.7)
or
$I\left(S_{o}(o,T_{h})\right)\underline{\triangleright}S_{c}(c,T_{m})$ (2.8)
The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
[…]
This shows that the carrier state set $S_{c}(c,T_{m})$ is always induced by the ontological state set $S_{o}(o,T_{h})$, and thus the enabling mapping $I$ must be surjective. For simplicity, unless otherwise noted, subsequent discussions assume that $I$ is a single-valued mapping.
Therefore, for any $\psi\in S_{c}(c,T_{m})$, there exists at least one $\varphi\in S_{o}(o,T_{h})$ such that $I(\varphi)\underline{\triangleright}\psi$. In this case, we write $I^{-1}(\psi)\underline{\triangleleft}\varphi$ or $\varphi\underline{\triangleright}I^{-1}(\psi)$ to preserve the consistency of the enabling direction.
It is important to note that both $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are state representations that satisfy Postulate 5. According to Theorem 1, each can be interpreted as a set of well-formed formulas within some formal system. However, the corresponding formal systems for the ontology and the carrier need not be identical. This distinction reflects one of the essential characteristics of information: its ability to establish connections between disparate entities in the world.
Definition 4 does not require that the ontological states of information be mapped one-to-one onto the carrier states. Of course, in practical scenarios, one-to-one mappings often do exist, which can simplify the analysis considerably.
Definition 5 (Recoverable Information): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ still denote a bijection from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, i.e., $I$ is invertible. Thus, there exists an inverse mapping $I^{-1}$ such that $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$. In this case, we say that information $I$ is recoverable, and $S_{o}(o,T_{h})$ is the recoverable state set of information $I$.
In particular, owing to the importance of information in information systems, we propose:
Definition 6 (Information in an Information System): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ be a recoverable information. If $S_{o}(o,T_{h})$ is a finite automaton state, then $c$ is called an information system, and $I$ is called information in the information system $c$.
For recoverable information, we also use $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$ or $I^{-1}:S_{c}(c,T_{m})\leftleftarrows S_{o}(o,T_{h})$ to maintain consistency in the direction of enabling. Since, as recoverable information, the carrier losslessly carries the complete state of the ontology, this naturally represents an ideal scenario in information studies. If $I$ is not a bijection, we can always partition elements of $S_{o}(o,T_{h})$ according to the same image equivalence classes, and then establish an enabling mapping from these equivalence classes to $S_{c}(c,T_{m})$ by means of equivalence class reduction, thereby ensuring recoverability. Specifically:
Proposition 1 (Recoverable Reduction of Information): For $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $R$ be an equivalence relation on $S_{o}(o,T_{h})$ defined by $xRy\leftrightarrow I(x)=I(y)$. Then, there exists a unique enabling mapping $\hat{I}:S_{o}(o,T_{h})/R\rightrightarrows S_{c}(c,T_{m})$, and $\hat{I}$ is a bijection, i.e., $\hat{I}$ is recoverable. In this case, $\hat{I}$ is called the recoverable reduction of information $I$.
When $I$ is already recoverable, $S_{o}(o,T_{h})/R=S_{o}(o,T_{h})$, and $\hat{I}=I$. Recoverable information and the recoverable reduction of information can provide significant assistance in simplifying problems.
Reference [24]is generated as a foundational work in information theory. The paper proposed that "$H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information. The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively. The channel $I_{c}$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $01\leq i\leq n,\tau>$ 0).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$. Then,
$S_{I_{s}}(I_{s},T_{h})=IsEvent^{3}(I_{s},T_{h},e_{i})$ (2.9)
and
$S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))=IsCode^{3}(I_{c},(T_{h},T_{h}+\tau),c_{i})$ (2.10)
are the states of the source $I_{s}$ at time $T_{b}$ and the state of the channel $I_{c}$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $1\leq i\leq n)$. Furthermore,
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.11)
represents the functional enabling mapping from the source state $S_{I_{s}}(I_{s},T_{h})$ to the channel state $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$.From equation (2.10), we can recover the event $e_{i}$ in equation (2.9) using the code $c_{i}$, thus $I$ is still recoverable information. Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{I_{s}}(I_{s},T_{h})$ and the complex implementation mechanisms of $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system. The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
2.5 Noisy Information
Many researchers question the significance of recoverable information, as in the real world, information is often not perfectly recoverable. This is because the reflection of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
Definition 7 (Noisy Information): For information $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $\widetilde{S_{o}}(o,T_{h})$ be another state of the same carrier $c$ at time $T_{m}$ and $\widetilde{S_{c}}(c,T_{m})$ be another state of the same carrier $c$ at time $T_{m}$,
$\widetilde{S_{o}}(o,T_{h})=S_{o}(o,T_{h})\backslash S_{o}^{-}(o,T_{h})\cup S_{o}^{+}(o,T_{h})$ (2.12)
where the state $S^{-}_{o}(o,T_{h})$ is the loss state of $S_{o}(o,T_{h})$, satisfying $S^{-}_{o}(o,T_{h})\subseteq$ $S_{o}(o,T_{h})$ and not being enabled to map to $\widetilde{S_{c}}(c,T_{m})$, and the state $S^{+}_{o}(o,T_{h})$ is the superposed state of $S_{o}(o,T_{h})$, satisfying $S_{o}(o,T_{h})\cap S_{o}^{+}(o,T_{h})=\emptyset$ and being enabled to map to $\widetilde{S_{c}}(c,T_{m})$. The enabling mapping is then defined as:
$\tilde{I}:\widetilde{S_{o}}(o,T_{h})\rightrightarrows\widetilde{S_{c}}(c,T_{m})$ (2.13)
which is the noisy information of $I$, while $I$ is its own noiseless information. Thus, the noisy information $\tilde{I}$ is also information. It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same. Further analysis of equation (2.12) shows that information $I$ and $\tilde{I}$ are completely asymmetric, and hence:
Proposition 2 (Symmetry of Noisy Information): If $\tilde{I}$ is the noisy information of $I$, then $I$ is also the noisy information of $\tilde{I}$.
In fact, for both information $I$ and its noisy information $\tilde{I}$, it is easy to verify its recoverability or recoverable reducibility from the perspective of enabling mapping. However, in real scenarios, it is often difficult to obtain noise-free information $I$, and one can only study using a noisy version $\tilde{I}$. Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information. But under the condition where the noiseless information $I$cannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$. This is the practical significance of introducing the concept of noisy information.
Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario[24]. In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
$E=f(S,N)$ (2.14)
where $f$ is the distortion function of the input for the transmitted signal. According to OIT’s view, the signal $S$ here is the ontological state of information $I$, $S_{I_{s}}(I_{s},T_{h})$. The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{I_{s}}^{-}(I_{s},T_{h})$ from the source state $S_{I_{s}}(I_{s},T_{h})$ and adding another state $S_{I_{s}}^{+}(I_{s},T_{h})$. The output $E$ is then the channel’s output state $\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output. Thus,
$\tilde{I}:\widetilde{S_{I_{s}}}(I_{s},T_{h})\rightrightarrows\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$ (2.15)
is the noisy information of
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.16)
where $\widetilde{S_{I_{c}}}(I_{s},T_{h})=S_{I_{s}}(I_{s},T_{h})\backslash S_{I_{s}}^{-}(I_{s},T_{h})\cup S_{I_{s}}^{+}(I_{s},T_{h})$.
In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels in[24], demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
3 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the model, which is a typical finite automaton and also an information system. Therefore, every learning process and the subsequent testing or application process rely on the information within it, which can generally be summarized and divided into four main steps of information realization and operation.
Step 1: The implementation of training information $It$, in which the ontological state of $It$ is enabled and mapped onto the model interface as its carrier;
Step 2: The original information $Io$ within the model—endowed with learning capability—is fused with the training information $It$, and optimized to yield the updated information $Iu$, which possesses both the ability to answer specific tasks and the capacity for further learning;
Step 3: The implementation of user input information $Iq$, where the ontological state of $Iq$ is likewise enabled and mapped onto the model interface as its carrier;
Step 4: The updated information $Iu$ within the model processes the input $Iq$, outputs the response information $Ir$ according to user requirements, and simultaneously undergoes further optimization through learning, resulting in $Iv$.
3.1 Formal Modeling and Realization of Training Information
According to the general formal definition of information, the formal representation of training information implementation is straightforward:
$It=(ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct})$ (3.1)
$S_{ot}(ot,Tt_{h})=\{\varphi t|\varphi t\text{ is a well-formed formula concerning }ot\text{ and }Tt_{h}\}$ (3.2)
$S_{ct}(ct,Tt_{m})=\{\psi t|\psi t\text{ is a well-formed formula concerning }ct\text{ and }Tt_{m}\}$ (3.3)
[…]
Therefore, it is necessary to introduce a model training data interface $ct$—akin to hardware interfaces or software APIs—as the carrier for $It$, so that $S_{ot}(ot,Tt_{h})$ can be ingested in an engineering-compatible manner. The reflection time set $Tt_{m}$ denotes the period during which the carrier reflects $It$, while the carrier reflection set $S_{ct}(ct,Tt_{m})$ comprises specific, complete, and operational realizations of $It$ on $ct$ at $Tt_{m}$. Each element $\psi t$ is a well-formed formula enabled by, and logically/semantically corresponding to, some $\varphi t$.
Continuing the example above, if $S_{ot}(ot,Tt_{h})$ is enabled and mapped to $ct$ at time $T_{3}$, we might express
$S_{ct}(ct,T_{3})=\{\psi t\}=\{(\text{TrainTable}[\text{user}_{id}=X][\text{time}=T_{1}]=\text{Active})\land$
[…]
where “Active” and “Tgt60” are defined in the training table as “user is active” and “device temperature exceeds 60°C,” respectively. Although the form of $\psi t$ may appear simple, it fully reflects the semantic content of $\varphi t$. In practice, $\psi t$ may take the form of tables, vectors, JSON objects, API parameters, or other formal representations of engineering data interfaces or hardware device states, and may include additional data constraints or format checks to ensure the validity and robustness of $S_{ct}(ct,T_{3})$.
It is crucial to emphasize that $S_{ot}(ot,Tt_{h})$ is only realized as $S_{ct}(ct,T_{3})$ via the enabling mapping of $It$. The fact that $S_{ot}(ot,Tt_{h})$ satisfies Postulate 5 also implies that $S_{ct}(ct,T_{3})$ satisfies Postulate 5. This mapping embodies both the structural-semantic correspondence from the former to the latter and the causal relationship—where the ontological states are the cause and the carrier states are the effect, requiring expenditure of human, temporal, and energy resources. Such an enabling mapping allows the knowledge, facts, and rules formalized within the ontology to be incorporated into the model in a structured, standardized, and machine-readable fashion, thereby becoming usable for subsequent model learning.
In practical engineering, the realization of training information often needs to consider noisy information $\widetilde{It}$ due to discretization, attribute pruning, and noise accumulation. However, for theoretical analysis, we currently focus only on the ideal case of no noise, leaving the study of complex noisy conditions for further research.
3.2 Formal Modeling and Implementation of Model Learning
From the perspective of information flow and transformation, the learning process of an artificial intelligence model M can be viewed as the optimization of its existing information $Io$ through learning from the training information $It$, resulting in the updated information $Iu$. The preceding section has already detailed the modeling and realization of training information $It$. We now turn our attention to the modeling and implementation of $Io$ and $Iu$ within M. By analogy with equations (3.1-3.3) for $It$, we can define $Io$ and $Iu$ by substituting the respective subscript $t$ with $o$ and $u$. It is important to note that, since $Io$ and $Iu$ represent the information present in M before and after learning, their carriers $co$ and $cu$ are effectively identical—in other words, both can directly be denoted by M, or more precisely, by the computational system that executes M:
$co=cu=M$ (3.4)
Within the learning process of M, the most crucial component is the existing information $Io$, which, in addition to encapsulating knowledge, propositions, facts, or samples (as with $It$), is distinguished by its possession of learning rules—i.e., the rules by which it is optimized into updated information through learning.
Definition 8 (State with Learning Capability): Let the states $Sx(x,T_{x})$, $Sy(y,T_{y})$, and $Sz(z,T_{z})$ be interpretations $E_{x}$, $E_{y}$, and $E_{z}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $x$ over time interval $T_{x}$, object $y$ over $T_{y}$, and object $z$ over $T_{z}$, all satisfying Postulate 5. We introduce the following specific functions and predicates in $\mathcal{L}$:
  • Function: $f_{learn}^{2}(X,Y)$ denotes the well-formed formula resulting from the learning of set $X$ from set $Y$. Moreover, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, the function $f_{learn}^{2}(X,Y)$ also satisfies this logical consistency requirement.
  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Learnable}^{2}(X,Y)$ indicates that state set $X$ can learn from state set $Y$; $\text{IsFormulaOf}^{2}(x,X)$ indicates that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ denotes that $x$ is equal to $y$.
Then suppose $\varphi x\in Sx(x,T_{x})$, $\sup(T_{x},T_{y})\leq\inf(T_{z})$, and
$\varphi x=$ $(\forall\Phi x)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi x,Sx(x,T_{x}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
$\text{Learnable}^{2}(\Phi x,\Phi y))\rightarrow\exists\varphi z(\text{IsFormulaOf}^{2}(\varphi z,Sz(z,T_{z}))\land\text{Eq}^{2}(\varphi z,f_{learn}^{2}(\Phi x,\Phi y)))$ (3.5)
we say that $Sx(x,T_{x})$ has the capability to learn from $Sy(y,T_{y})$ and optimize into $Sz(z,T_{z})$; equivalently, we say that the interpretation $E_{x}$ has the capacity to learn from $E_{y}$ and optimize toward $E_{z}$.
The formula $\varphi x$ essentially constitutes the learning rule for the state $Sx(x,T_{x})$, indicating that as long as $Sy(y,T_{y})$ meets the learning conditions of $Sx(x,T_{x})$, it is possible to augment $Sz(z,T_{z})$ with new well-formed formulas $\{\varphi u|\varphi u=f_{learn}^{2}(\Phi x,\Phi y),\Phi x\subseteq Sx(x,T_{x}),\Phi y\subseteq Sy(y,T_{y})\}$, thus rendering $Sz(z,T_{z})$ an optimized extension of $Sx(x,T_{x})$. Furthermore, given the logical consistency of $f_{learn}^{2}(X,Y)$, if $Sx(x,T_{x})$ and $Sy(y,T_{y})$ satisfy Postulate 5, then $Sy(y,T_{y})$ also satisfies Postulate 5.
On the other hand, the inheritance of learning capability is also a fundamental property of model learning. In fact, to inherit the learning capability of $Sx(x,T_{x})$, it suffices to retain the learning rule $\varphi x$ within $Sz(z,T_{z})$, with only partial adaptive substitution of its symbols. Specifically, for $\varphi z\in Sz(z,T_{z})$:
$\varphi z=($ $\forall\Phi z)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
$\text{Learnable}^{2}(\Phi z,\Phi y))\rightarrow\exists\varphi w(\text{IsFormulaOf}^{2}(\varphi w,Sw(w,T_{w}))\land$
$\text{Eq}^{2}(\varphi w,f_{learn}^{2}(\Phi u,\Phi y)))$ (3.6)
This ensures that $Sz(z,T_{z})$ can learn from $Sy(y,T_{y})$ and optimize into $Sw(w,T_{w})$. Thus, we establish:
Proposition 3 (Inheritance of Learning Capability): If the state set $Sx(x,T_{x})$ is capable of learning from $Sy(y,T_{y})$ and optimizing into $Sz(z,T_{z})$, then $Sz(z,T_{z})$ always inherits the learning capability of $Sx(x,T_{x})$.
Definition 9 (Realization of Model Learning): For an AI model M with original information $Io$, training information $It$, and updated information $Iu$, if all three are recoverable, and the ontological state $S_{oo}(oo,To_{h})$ can learn from $S_{ot}(ot,Tt_{h})$ and be optimized into $S_{ou}(ou,Tu_{h})$, and the carrier state $S_{co}(co,To_{m})$ can likewise learn from $S_{ct}(ct,Tt_{m})$ and be optimized into $S_{cu}(cu,Tu_{m})$, then M is said to possess the capability to realize the learning and optimization of $Io$ towards $It$, resulting in $Iu$—in short, M has information learning capability.
According to (3.4), we may also write $S_{co}(co,To_{m})=S_{M}(M,To_{m})$, $S_{cu}(cu,Tu_{m})=S_{M}(M,Tu_{m})$, indicating that the learning process is essentially an optimization of the state of model M.
In summary, we have theoretically internalized the model’s learning capability as a property of its states. These rules are independent of the algorithmic structure or scale of the model, thereby providing a unified and general formal framework for the analysis of various types of models.
As an example, let $Mn$ be a neural network model. According to common knowledge, the information in $Mn$ is $Io=\langle oo,To_{h},S_{oo},Mn,To_{m},S_{co}\rangle$, where $oo$ represents the model in the designer’s mind, $To_{h}$ is the time set during the design of $Mn$.
$S_{oo}(oo,To_{h})=\Phi o=\{Params(\theta)\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.7)
Here, $Params(\theta)$ denotes the predicate representation of the parameter set $,\theta=\{W_{1},b_{1},\ldots$$,W_{n},b_{n}\}$; $Architecture(L_{1},\ldots,L_{n})$ denotes the predicate representation of the network layer structure; and $Activation(A)$ denotes the predicate representation of the activation function set $A=\{a_{1},\ldots,a_{n}\}$.
The carrier for $Io$ is the neural network model $Mn$ itself, with $To_{m}$ representing the period after its development, and $S_{co}(co,To_{m})$ being the physical instantiation of (3.7) within $Mn$.
The training information for $Mn$ is $It=\langle ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct}\rangle$, where $ot$ is a set of real-world objects, $Tt_{h}$ is the time set during which the state $S_{ot}(ot,Tt_{h})$ occurs, and
$S_{ot}(ot,Tt_{h})=\Phi t=\{Batch(X,y)\land DimMatch(X,L_{1})\land Differentiable(L)\}$ (3.8)
where $Batch(X,y)$ expresses the predicate for input data $X\in\mathbb{R}^{b\times k}$ and labels $y\in\mathbb{R}^{b\times k}$; $DimMatch(X,L_{1})$ expresses that the input data $X$ matches the dimensions of the input layer $L_{1}$; and $Differentiable(L)$ expresses the differentiability of the loss function $L$ with respect to parameters $\theta$.
The definition of the predicate $Learnable^{2}$ is as follows:
$Learnable^{2}(\Phi o,\Phi t)=$ $Dimensionality(\Phi o,\Phi t)\land ActivationFeasible(\Phi o,\Phi t)\land$ (3.9)
$GradientExists(\Phi t)$
where $Dimensionality(\Phi o,\Phi t)=DIM(X,b\times d)\land DIM(L_{1},d)$ ensures that the input data dimensions are consistent with the input layer, $ActivationFeasible(\Phi o,\Phi t)$ ensures that for every $a_{i}\in A$, $Range(X)\subseteq Domain(a_{i})$, i.e., the input data lies within the domain of the activation function, and $GradientExists(\Phi t)$ ensures the existence of a gradient $\nabla_{\theta}L(\theta,X,y)$, i.e., that the loss function $L$ is differentiable with respect to $\theta$.
The learning function $f_{learn}^{2}$ is defined as:
$f_{learn}^{2}(\Phi o,\Phi t)=\{Params(\theta_{new})\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.10)
where $\theta_{new}=\theta-\eta(\nabla_{\theta}L(\theta,X,y)+\beta v_{prev})$ is the iterative update formula for the neural network model.
Consequently, in the updated information $Iu=(ou,Tu_{h},S_{ou},Mn,Tu_{m},S_{cu})$ of the neural network model, the ontology $ou$ is the union of the designer’s conceptualization and the ontology $ot$ of the training information, $Tu_{h}$ is the set of training times for $Mn$,
$S_{ou}(ou,Tu_{h})=f_{learn}^{2}(\Phi o,\Phi t)$ (3.11)
constitutes the optimized ontological state after learning, with the carrier of $Iu$ being $Mn$ itself, the reflection time $Tu_{m}$ being the post-training period, and $S_{cu}(Mn,Tu_{m})$ as the physical realization of $S_{ou}(ou,Tu_{h})$ within $Mn$, ideally satisfying the mathematical and physical requirements of the enabling mapping:
$S_{cu}(Mn,Tu_{m})=f_{learn}^{2}(\Psi o,\Psi t)$ (3.12)
where $\Psi o$ and $\Psi t$ are the sets of well-formed formulas corresponding to the carrier states of $Io$ and $It$, that is, the physical realizations of $\Phi o$ and $\Phi t$ within $Mn$ and its interface. Equations (3.11) and (3.12) together demonstrate that, within $Mn$, both the ontological state $S_{oo}(oo,To_{h})$ and the carrier state $S_{co}(co,To_{m})$ possess the capability to learn from $S_{ot}(ot,Tt_{h})$ and $S_{ct}(ct,Tt_{m})$, and then optimize to $S_{ou}(ou,Tu_{h})$ and $S_{cu}(cu,Tu_{m})$, respectively; thus, $Mn$ indeed realizes the learning and optimization of $Io$ with respect to $It$, resulting in $Iu$. This completes the proof.
It should be noted that some of the content in the above discussion does not strictly adopt set definitions and well-formed formulas. This does not mean that they cannot be converted into complete set definitions and well-formed formulas, but rather because such representation is more concise and intuitive, and does not lead to ambiguity.
3.3 Information Processing and Learning Optimization in Models
The fundamental objective of model learning is to more accurately process user input information $Iq$ and to deliver high-quality output information $Ir$ that meets user requirements. Since the modeling and realization of input information $Iq$ are almost identical to those of $It$, as discussed in detail in Section 3.1, they will not be reiterated here. Thus, our focus shifts to the process whereby the updated information $Iu$ within model M processes $Iq$, outputs $Ir$, and is further optimized through learning to yield $Iv$. In this context, we continue to define $Iq$, $Ir$, and $Iv$ in the manner of equations (3.1-3.3).
Definition 10 (Formal Representation of State Processing): Let the states $Sz(z,T_{z})$, $Su(u,T_{u})$, and $Sv(v,T_{v})$ be interpretations $E_{z}$, $E_{u}$, and $E_{v}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $z$ over time interval $T_{z}$, object $u$ over $T_{u}$, and object $v$ over $T_{v}$, all satisfying Postulate 5. We introduce the following specific functions and predicates in $\mathcal{L}$:
  • Function: $f_{process}^{2}(X,Y)$ denotes the set of well-formed formulas output by state set $X$ after processing state set $Y$. Similarly, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, $f_{process}^{2}(X,Y)$ also meets this logical consistency requirement.
  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Processable}^{2}(X,Y)$ indicates that state set $X$ is capable of processing state set $Y$ and generating an output; $\text{IsFormulaOf}^{2}(x,X)$ denotes that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ indicates that $x$ is equal to $y$.
Then suppose $\varphi z\in Sz(z,T_{z})$, $\sup(T_{z},T_{u})\leq\inf(T_{v})$, and
$\varphi z=($ $\forall\Phi z)(\forall\Phi u)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi u,Su(u,T_{u}))\land$
$\text{Processable}^{2}(\Phi z,\Phi u))\rightarrow\exists\varphi v(\text{IsFormulaOf}^{2}(\varphi v,Sv(v,T_{v}))\land$
$\text{Eq}^{2}(\varphi v,f_{process}^{2}(\Phi z,\Phi u)))$ (3.13)
we say that $Sz(z,T_{z})$ is capable of processing $Su(u,T_{u})$ and outputting $Sv(v,T_{v})$; equivalently, we say that the interpretation $E_{z}$ has the ability to process $E_{u}$ and generate $E_{v}$ as output.
Clearly, by combining Definitions 8 and 10, we can also formally express the process whereby state $S_{ou}(ou,Tu_{h})$ processes and learns from $S_{oq}(oq,Tq_{h})$, outputs $S_{or}(or,Tr_{h})$, and is subsequently optimized into $S_{ov}(ov,Tv_{h})$. In addition, when $S_{ou}(ou,Tu_{h})$ and $S_{oq}(oq,Tq_{h})$ both satisfy Postulate 5, the logical consistency of $f_{learn}^{2}(X,Y)$ and $f_{process}^{2}(X,Y)$ ensures that $S_{or}(or,Tr_{h})$ and $S_{ov}(ov,Tv_{h})$ also satisfy Postulate 5. Thus, we have:
Definition 11 (Implementation of Model Information Processing and Learning Optimization): For an AI model M with updated information $Iu$, input information $Iq$, output information $Ir$, and optimized information $Iv$, if all four are recoverable, and the ontological state $S_{ou}(ou,Tu_{h})$ is capable of processing and learning from $S_{oq}(oq,Tq_{h})$, outputting $S_{or}(or,Tr_{h})$, and further optimizing into $S_{ov}(ov,Tv_{h})$, while the carrier state $S_{M}(M,Tu_{h})$ is similarly capable of processing and learning from $S_{cq}(cq,Tq_{h})$, outputting $S_{cr}(cr,Tr_{h})$, and further optimizing into $S_{M}(M,Tv_{h})$, then M is said to possess the capability to realize the information processing and learning optimization of $Iu$ with respect to $Iq$, outputting $Ir$ and optimizing into $Iv$. In short, M possesses information processing and learning optimization capability.
For the neural network model $Mn$ discussed in the previous section, the processable predicate $\text{Processable}^{2}$ is defined as:
$\text{Processable}^{2}(\Phi u,\Phi q)=\text{InputMatch}(\Phi u,\Phi q)\land\text{ActivationFeasible}(\Phi u,\Phi q)\land\text{OutputDefined}(\Phi u)$ (3.14)
[…]
The processing function $f_{process}^{2}$ is defined as:
$f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.15)
where $f_{\Phi u}$ is the prediction function explicitly expressed by the structure and parameters of the neural network $Mn$.
Thus, for the neural network model $Mn$, the ontological state of the updated information $Iu$ is $S_{ou}(ou,Tu_{h})=\Phi u$, and the ontological state of the input information $Iq$ is $S_{oq}(oq,Tq_{h})=\Phi q$. When the processable condition $\text{Processable}^{2}(\Phi u,\Phi q)$ is satisfied, the ontological state of the output information $Ir$ is
$S_{or}(or,Tr_{h})=\Phi r=f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.16)
demonstrating that the ontological state $S_{ou}(ou,Tu_{h})$ possesses the ability to process $S_{oq}(oq,Tq_{h})$ and output $S_{or}(or,Tr_{h})$. Similarly, since $Iu$, $Iq$ and $Ir$ are all recoverable information, when their ontological states are mapped onto $Mn$ and its interface, the carrier states $S_{Mn}(Mn,Tu_{h})$ also possess the ability to process $S_{cq}(cq,Tq_{h})$ and output $S_{cr}(cr,Tr_{h})$. Thus, $Mn$ is capable of realizing the processing of $Iu$ and $Iq$ to produce $Ir$. Furthermore, according to section 3.2, $Mn$ is also capable of learning from $Iq$ and optimizing into $Iv$.
Naturally, one can just as easily use the formal framework of Definitions 8 and 10 to demonstrate that other typical models—such as linear regression and decision trees—also possess both information processing and learning optimization capability. It is thus evident that this axiomatic and formalized architecture transcends specific algorithms or models, providing a more universal, unified, and interpretable foundational meta-framework for machine learning theory.
4 Several Important Theorems and Proofs
The Meta-Framework for Machine Learning Theory (MLT-MF) applies to a series of pivotal issues in the field of artificial intelligence, providing many unique and practically valuable theoretical principles and methodological guidelines.
4.1 Theorem of Model Interpretability
Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence. However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability [6, 30, 31, 32]. In fact, the interpretability of any phenomenon in the universe is inherently relative. For example, Kepler’s three laws explain the motion of planets around the sun, but such explanations are meaningless to those without elementary mathematical knowledge. Even for the same individual, the explanations that can be understood and accepted may evolve as their knowledge expands. Therefore, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system. Furthermore, it should also be considered in relation to the specific inputs and outputs of the model. Based on MLT-MF, we define the following concepts:
Definition 12 (Model interpretability): Let the state of a model M over time interval $Tm$ be denoted as $S_{M}(M,Tm)$. Given an input state $S_{cq}(cq,Tq_{h})$ and an output state $S_{cr}(cr,Tr_{h})$, suppose there exists a formal system $\mathcal{L}$ and three corresponding interpretations $E_{M}$, $E_{q}$, and $E_{r}$ over finite domains $D_{M}$, $D_{q}$, and $D_{r}$, respectively, such that:
  • $E_{M}$ has the capacity to process $E_{q}$ and produce $E_{r}$ as output, and
  • $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are each bijective enabling mappings of $E_{M}$, $E_{q}$, and $E_{r}$, respectively.
Then, we say that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
In this way, we define model interpretability via the formal system $\mathcal{L}$ and its interpretation. This definition aligns both with the intuitive understanding of "interpretation" and with the theoretical need to characterize the internal logic of models through a series of propositions, making it well-suited for the formal analysis and in-depth study of AI interpretability.
In fact, according to Theorem 1, if a model M has the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then by Definition 10, the states $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are interpretations $E_{M}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5. In this case, the mappings $E_{M}\rightrightarrows S_{M}(M,Tm)$, $E_{q}\rightrightarrows S_{cq}(cq,Tq_{h})$, and $E_{r}\rightrightarrows S_{cr}(cr,Tr_{h})$ are self-mappings, and therefore naturally satisfy the condition of being bijective enabling mappings. This leads to:
Corollary 1 (Intrinsic Model Interpretability): If a model M possesses the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then there exists a formal system $\mathcal{L}_{M}$ in which M is interpretable with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$. Since $\mathcal{L}_{M}$ serves as the formal system that interprets model M, it is also referred to as the intrinsic formal system of the model.
As Hegel famously asserted, "What is real is rational." The intrinsic interpretability of a model means that, insofar as it can process inputs and generate outputs according to established rules and procedures, it necessarily possesses an internal logic that is, in principle, interpretable. However, this logic may be highly implicit and abstract, making it difficult for people—even domain experts—to identify and understand. This highlights the need to map and translate this internal logic into a more intuitive and universal logical system, thereby enabling deeper decomposition and analysis of the causal mechanisms, and achieving broader understanding and acceptance. This is the essence of model interpretability, and is precisely the core idea behind Definition 12: to interpret the internal logic of models within more explicit formal systems.
Theorem 3 (Model interpretability and information recoverability): Let model M be capable of processing the information $Iq=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ and producing output information $Ir=\langle or,Tr_{h},S_{or},cr,Tr_{m},S_{cr}\rangle$, through its operation described by $Iu=\langle ou,Tu_{h},S_{ou},M,T_{m},S_{M}\rangle$. If all three pieces of information $Iu$, $Iq$, and $Ir$ are recoverable, and their ontological states are interpretations of well-formed formulas in a common formal system $\mathcal{L}$, then model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
Proof: According to Definition 5, since $Iu$, $Iq$, and $Ir$ are recoverable, the following are all bijective enabling mappings: $S_{ou}(ou,T_{h})\rightrightarrows S_{M}(M,T_{m})$, $S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$, and $S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$.
By Definition 11, the state $S_{ou}(ou,T_{h})$ has the capacity to process $S_{oq}(oq,Tq_{h})$ and produce $S_{or}(or,Tr_{h})$ as output.
Then, by Definition 10, the states $S_{ou}(ou,T_{h})$, $S_{oq}(oq,Tq_{h})$, and $S_{or}(or,Tr_{h})$ are interpretations $E_{u}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5. Moreover, $E_{u}$ has the ability to process $E_{q}$ and produce $E_{r}$ as output.
Thus, we have the following bijective enabling mappings:
$E_{u}=S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,T_{m})$ (4.1)
$E_{q}$ $=S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$ (4.2)
$E_{r}$ $=S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$ (4.3)
By Definition 12, it follows that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{m})$ under the input condition $S_{cq}(cq,Tq_{m})$. Q.E.D.
Theorem 3 fully embodies the Chinese proverb: "Whoever tied the bell must untie it." Since, intuitively, the internal logic $S_{M}(M,T_{m})$ of model M originates from the ontological state $S_{ou}(ou,Tu_{h})$ that brings the model into existence, it is thus natural to interpret the model’s output under specific input conditions using $E_{u}$, $E_{q}$, and $E_{r}$. Moreover, unlike the internal processing of $S_{M}(M,T_{m})$, which is encapsulated within the model and often opaque, the interpretations $E_{u}$, $E_{q}$, and $E_{r}$ are explicit expressions of well-formed formulas within the formal system $\mathcal{L}$ under given valuations. These expressions are more amenable to interpretation and understanding by humans, particularly by domain experts.
4.2 Theorem of Ethical Safety Assurance
Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence [29, 30, 31, 32], and to date, there is still no universally accepted and practically actionable rigorous definition [33, 34, 35]. Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions. To achieve this, it is first necessary to specify appropriate well-formed formulas $Ec$ as ethical constraints—such as those governing fairness and privacy protection—thus establishing a formalized definition of ethical safety.
Definition 13 (Ethical Safety): Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$. Let $Ec(\vec{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\vec{v}=(v_{1},\ldots,v_{n})$ are its free variables. For each substitutable $(x,t)\in X\times T$, define
$Ec|_{(x,t)}=Ec(\vec{v}\mapsto\vec{\sigma}(x,t))$ (4.4)
where $\vec{\sigma}:X\times T\to D^{n}$ is a value vector function such that $\sigma_{i}(x,t)\in D$ is the concrete value assigned to $v_{i}$ for $1\leq i\leq n$.
If for all $\varphi\in\text{Cn}(S(X,T))$, and all $(x,t)\in X\times T$,
$\varphi\rightarrow\sim Ec|_{(x,t)}$ (4.5)
does not hold, then $S(X,T)$ is said to be ethically safe with respect to $Ec$. Furthermore, if a model M is capable of producing output $I$, and both the ontological and carrier states of $I$ are ethically safe with respect to $Ec$, then M is said to be ethically safe with respect to $Ec$.
In the definition of ethical safety, we require not only that $S(X,T)$ contains no formulas that entail $\sim Ec|_{(x,t)}$, but also that its logical closure $\text{Cn}(S(X,T))$ contains no such formulas. This is because if $\text{Cn}(S(X,T))$ were to include formulas entailing $\sim Ec|_{(x,t)}$, then the conjunction of multiple formulas in $S(X,T)$ could still violate the ethical constraint $Ec$. This would indicate that $S(X,T)$ implicitly allows situations that contravene $Ec$, which must be strictly avoided in a rigorous sense. Moreover, it is generally not difficult for a model M to achieve ethical safety with respect to $Ec$ in a simple manner; the key challenge lies in ensuring the largest possible set of valid outputs while maintaining ethical safety. To this end, we propose:
Theorem 4 (Ethical Safety Assurance): Suppose model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$. Let $S_{or}(or,Tr_{h})$ denote the ontological state of $Ir$. For any ethical constraint $Ec$ represented by a well-formed formula in $\mathcal{L}$, there exists a constructive algorithm for obtaining the largest output state subset $S_{or}^{EC}(or,Tr_{h})\subseteq S_{or}(or,Tr_{h})$ that is ethically safe with respect to $Ec$. By introducing an ethical safeguard formula for $Ec$ into $Iu$, it is thereby possible to ensure that model M attains the maximal output state that is ethically safe with respect to $Ec$.
Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, $Iq$ and $Ir$ respectively, the MLT-MF framework provides:
$\varphi u=($ $\forall\Phi u)(\forall\Phi q)(\text{IsSubsetOf}^{2}(\Phi u,Sou(ou,Tu_{h}))\land\text{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,Tq_{h}))\land$
$\text{Processable}^{2}(\Phi u,\Phi q))\to\exists\varphi r(\text{IsFormulaOf}^{2}(\varphi r,S_{or}(or,Tr_{h}))\land$
$\text{Eq}^{2}(\varphi r,f_{process}^{2}(\Phi u,\Phi q)))$ (4.6)
Thus, the set
$S_{or}(or,Tr_{h})=\{\varphi r|\varphi r=\text{Processable}^{2}(\Phi u,\Phi q),\Phi u\in Sou(ou,Tu_{h}),\Phi q\in S_{oq}(oq,Tq_{h})\}$ (4.7)
represents the theoretical output states of model M.
Since $S_{or}(or,Tr_{h})$ is a finite set, we construct a finite hypergraph
$H_{or}=(S_{or}(or,Tr_{h}),\mathcal{E})$ (4.8)
where the vertex set is precisely $S_{or}(or,Tr_{h})$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
(1) For any $\varphi r\in S_{or}(or,Tr_{h})$ and $(x,t)\in or\times Tr_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\{\varphi r\}$ to $\mathcal{E}$;
(2) If there exists a minimal subset $W\subseteq S_{or}(or,Tr_{h})$ such that the conjunction of all formulas in $W$ entails $\sim Ec|_{(x,t)}$, then add the hyperedge $W$ to $\mathcal{E}$.
Let $G$ be an independent set in $H_{or}$, i.e., a subset of vertices containing no hyperedge $e\in\mathcal{E}$. If there exists $\varphi\in\text{Cn}(G)$ such that
$\varphi\to\sim Ec|_{(x,t)}$ (4.9)
then there must exist a finite subset $\{\varphi r_{1},\cdots,\varphi r_{k}\}\subseteq G$ such that $\varphi r_{1}\land\cdots\land\varphi r_{k}=\varphi$, and hence $\varphi r_{1}\land\cdots\land\varphi r_{k}\to\sim Ec|_{(x,t)}$. Let $W=\{\varphi r_{1},\cdots,\varphi r_{k}\}$, so the conjunction of $W$ entails $\sim Ec|_{(x,t)}$.
According to the hyperedge construction rules, if $|W|=1$, then $W$ itself directly violates $Ec$, and there must be a hyperedge $\{\varphi r_{1}\}$. But since $\{\varphi r_{1}\}\in G$, this contradicts the assumption that $G$ is an independent set. If $|W|\geq 2$, then by the compactness theorem, the logical entailment can be captured by a finite minimal subset, so there exists a minimal subset $W_{min}\subseteq W$ whose conjunction entails $\sim Ec|_{(x,t)}$. Therefore, there exists a hyperedge $W_{min}\subseteq W\subseteq G$ in $H_{or}$, again contradicting the independence of $G$. Thus, (4.9) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
$|K|>|G_{max}|$ (4.10)
then $K$ cannot contain any hyperedge $e\in{E}$. Otherwise, if $|e|=1$, then $K$ directly violates $Ec$; if $|e|\geq 2$, then the formulas in $K$ violate $Ec$. Both cases contradict the ethical safety assumption for $K$. Therefore, $K$ must alsobe an independent set in $H_{or}$, which contradicts the maximality of $G_{max}$. Hence, $G_{max}$ is the largest ethically safe subset of $S_{or}(or,Tr_{h})$ with respect to $Ec$.
Let $S_{or}^{Ec}(or,Tr_{h})=G_{max}$ and $S_{or}^{\sim Ec}(or,Tr_{h})=S_{or}(or,Tr_{h})\backslash S_{or}^{Ec}(or,Tr_{h})$. In addition, introduce a new constant in $\mathcal{L}$: NEc, representing a non-Ec ethical safety prompt, and a new predicate: IsSafeback1(NEc), indicating that M outputs the non-Ec ethical safety prompt NEc. We may safely assume that $\mathcal{L}$ originally contains neither.
In this way, add to $Sou(ou,Tu_{h})$ an ethical safeguard formula:
$\varphi u=($ $\forall\varphi r)(IsFormulaOf^{2}(\varphi r,S_{or}(or,Tr_{h})))\to\exists\varphi s((IsFormulaOf^{2}(\varphi s,S_{os}(os,Ts_{h}))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,\varphi r))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{\sim Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,IsSafeback^{1}(NEc))))$ (4.11)
Since the newly added $\varphi u$ pertains to $S_{os}(os,Ts_{h})$ and is entirely independent of the existing formulas in $Sou(ou,Tu_{h})$, it does not affect the satisfaction of Postulate 5 by $Sou(ou,Tu_{h})$. Moreover,
$S_{os}(os,Ts_{h})=S_{or}^{Ec}(or,Tr_{h})\cup\{IsSafeback^{1}(NEc)\}$ (4.12)
Because NEc and IsSafeback1(NEc) were not previously present in $\mathcal{L}$, and $S_{or}^{Ec}(or,Tr_{h})$ and $\{IsSafeback^{1}(NEc)\}$ are logically independent and both logically consistent, $S_{os}(os,Ts_{h})$ is also logically consistent and satisfies all other conditions of Postulate 5. Therefore, it is the maximal output state that is ethically safe with respect to $Ec$.
By taking $S_{os}(os,Ts_{h})$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,Ts_{m})$, $S_{M}(M,Ts_{m})$ is necessarily also ethically safe with respect to $Ec$. This satisfies the condition in Definition 13 for model M to be ethically safe with respect to $Ec$. The theorem is thus proved.
Theorem 4 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs-for obtaining the maximum ethically safe output. However, computing the maximum independent set in a hypergraph is generally an NP-hard problem. In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}(or,Tr_{h})$, and another for ethical safety checking, thus yielding an approximately maximal safe output. This is a practical issue that warrants further investigation and will not be elaborated upon here. Meanwhile, the inheritance of ethical safety in the learning process can be achieved by incorporating dedicated ethical constraint regularization terms into the model algorithm. This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
In summary, based on the MLT-MF framework, Definition 13 and Theorem 4, for the first time, construct a formal theoretical framework for ethical safety. This approach, unlike those based on heuristic rules or statistical indicators, fills the gap in mathematical foundations and verifiability in traditional research, and provides concrete methodological guidance for the automated embedding and verification of ethical constraints.
4.3 Theorem on Model Generalization Error Estimation
Improving model generalization ability is among the most critical objectives in machine learning. However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance [40, 41, 42, 43, 44]. It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples. By applying the MLT-MF, it is possible to estimate an upper bound on the model’s generalization error.
Theorem 5 (Upper Bound for Model Generalization Error): Let a model M, after learning and optimization, carry the information $Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$, and let the input information be $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$. Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively, and let $p_{q}(x)$ be the probability distribution over $S_{oq}$. The total probability of $p_{q}(x)$ on $S_{ou}\cap S_{oq}$ is denoted as $P_{Mq}$. If $S_{ou}$ and $S_{oq}$ are governed by probability distributions defined on the same measurable space, and M is sufficiently trained on $S_{oq}$ and predicts unknown states using a conservative uniform distribution, then the upper bound of the generalization error when processing $Iq$ is determined by the overlap between $S_{oq}$ and $S_{ou}$, as well as the entropy of $p_{q}(x)$ on the non-overlapping region, specifically:
$\mathcal{E}_{gen}\begin{cases}\to 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.13)
where $\mathcal{E}_{gen}$ denotes the upper bound of generalization error, and
$H(p_{q})=-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)$
is the entropy of the data in the non-overlapping region.
Proof: Let $p_{M}(x)$ be the probability distribution determined by the model structure and parameter set of M. $Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ is the set of information carried by M, so $p_{M}(x)>0$ for $x\in S_{ou}$; $p_{q}(x)>0$ for $x\in S_{oq}$.
Since M is sufficiently trained on $S_{oq}$ and adopts a conservative uniform distribution for unknown states, $p_{M}(x)$ in the overlapping region $S_{ou}\cap S_{oq}$ approximates the true distribution $p_{q}(x)$, and in the non-overlapping region $S_{oq}\backslash S_{ou}$, it follows a uniform distribution. That is,
$p_{M}(x)\begin{cases}\to p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.14)
Using KL divergence to measure the discrepancy between the true and model distributions [45]:
$D_{KL}(p_{q}(x)\|p_{M}(x))=$ $\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (4.15)
$=$ $\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
According to (4.14), in the overlapping region:
$\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\to\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}=0$ (4.16)
In the non-overlapping region, since the true distribution $p_{q}(x)$ is arbitrary, again by (4.14):
$\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ $=\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{M}(x)$
$=-H(p_{q})-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
$=-H(p_{q})-\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$ (1)
$=-H(p_{q})-(1-P_{Mq})\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
$=(1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q})$ (4.17)
Because $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can estimate the generalization error of model M using the total variation distance (TVD) and KL divergence. According to Pinsker’s inequality [46]:
$\mathcal{E}_{gen}\leq\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
$=\sqrt{\frac{1}{2}(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)})}$
$\approx\sqrt{\frac{1}{2}((1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q}))}$ (4.18)
Thus, the latter part of (4.13) is proved. Since we have assumed that $Iu=(ou,Tu_{h},S_{ou},M,$ $Tu_{m},S_{M})$ and $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$ are both recoverable information, and the states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ are the physical realizations of $S_{ou}$ and $S_{oq}$ on the model M and its interface, equation (4.13) indeed characterizes the upper bound of model M’s generalization error. Q.E.D.
Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 5—based on MLT-MF—explicitly quantifies the generalization error upper bound using $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, which reflect the overlap between training data and model information, and the entropy $H(p_{q})$ of the non-overlapping region, thereby offering a new perspective and practical guidance for data collection and model design. In particular, Theorem 5 reveals that as long as the problem space is contained within the model’s information space, the model’s generalization error can be made arbitrarily small in the statistical sense. This leads to a critical principle in machine learning:
Corollary 2 (Generalization Error of Ideal Models): If the model information contains the problem information, then the generalization error for these problems is almost zero.
5 Conclusion and Outlook
This paper has established the "Meta-Framework for Machine Learning Theory (MLT-MF)" and, through formal information modeling and the sextuple definition, has unified the characterization of information flow and causal mechanisms across the entire machine learning lifecycle. The core contributions are as follows:
  • Theoretical Integration: For the first time, data initialization, training optimization, and knowledge transfer are incorporated into a unified framework, overcoming the fragmented modeling of temporal causal chains in traditional research;
  • Interpretability: By explicitly representing ontological and carrier states via higher-order logical systems and mapping model decisions back to the formal semantic space through information recoverability, logical transparency is ensured and a rigorous mathematical foundation for interpretability is provided;
  • Ethical Safety: By embedding ethical constraint predicates, the compliance of model outputs can be formally verified, offering criteria and methodological guidance for the design of trustworthy AI systems;
  • Generalization Error Quantification: Based on state coverage and KL divergence, an upper bound formula for generalization error is derived, revealing the core principle that the "problem space must be contained within the model knowledge space," thus providing explicit tools for quantitative analysis in data collection and model optimization.
Despite these important advances, the current framework assumes ideal (noiseless) enabling mappings and primarily addresses model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems. Future research should deepen the following directions:
  • Extend to complex dynamic scenarios such as online learning under noisy conditions, introducing noisy ontologies and incremental coverage mechanisms;
  • Investigate information fusion and conflict resolution across ontological spaces for multimodal and multi-agent systems, constructing distributed MLT-MF frameworks;
  • Promote interdisciplinary integration of MLT-MF with fields such as quantum computing and bioinformatics, exploring formal mappings of quantum states and interpretable modeling of gene regulatory networks;
  • Design automated toolchains to support conversion from formal rules to code, lowering the engineering threshold for MLT-MF adoption.
6 Acknowledgements
During the writing and revision of this paper, I received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University. I also gained much inspiration and assistance from regular academic discussions with doctoral students Siyuan Qiu, Chun Li, Hu Xu, and Zeyan Li. I hereby express my sincere gratitude to them.

### v9

@@
 Abstract
-[Objective] This study focuses on addressing the current lack of a unified formal theoretical framework in machine learning, as well as the deficiencies in interpretability and ethical safety assurance.
-[Methods] A formal information model is first constructed, utilizing sets of well-formed formulas to explicitly define the ontological states and carrier mappings of typical components in machine learning.
-Learnable and processable predicates, along with learning and processing functions, are introduced to analyze the logical deduction and constraint rules of the causal chains within models.
-[Results] A meta-framework for machine learning theory (MLT-MF) is established.
-Based on this framework, universal definitions for model interpretability and ethical safety are proposed.
-Furthermore, three key theorems are proved: the equivalence of model interpretability and information recoverability, the assurance of ethical safety, and the estimation of generalization error.
-[Limitations] The current framework assumes ideal conditions with noiseless information-enabling mappings and primarily targets model learning and processing logic in static scenarios.
+[Objective] This study addresses key challenges in machine learning, namely the absence of a unified formal theoretical framework and the lack of foundational theories for model interpretability and ethical safety.
+[Methods] We first construct a formal information model, explicitly defining the ontological states and carrier mappings of typical machine learning stages using sets of well-formed formulas.
+By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the causal chain logic and constraint laws governing machine learning processes.
+[Results] We establish the Machine Learning Theory Meta-Framework (MLT-MF), on which we further propose universal definitions for model interpretability and ethical safety.
+We prove and validate three key theorems: the relationship between model interpretability and information existence, ethical safety assurance, and the upper bound estimation of total variation distance (TVD).
+[Limitations] The current framework assumes ideal, noise-free information enabling mappings and focuses primarily on model learning and processing logic in static scenarios.
 It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
-[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing the critical challenges currently faced in machine learning.
+[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing critical issues in contemporary machine learning.
 Keywords: machine learning; objective information theory; formal logic systems; interpretability; ethical and safety assurance; generalization error
@@
 Although artificial intelligence is currently the most dynamic field within information science and even the broader domain of science and technology, existing research still faces several significant shortcomings:
-First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning [1, 2, 3, 4, 5].
+First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning Rob2021 ; Marius2023 ; Meng2023 ; Alquier2024 ; Vilalta2002 .
 Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages.
@@
 While the field of causal inference attempts to characterize causal relationships among variables, its application in machine learning is mostly limited to feature-label association analysis and fails to construct a complete causal chain from “prior constraints → initialization → gradient propagation → generalization capability.” For instance, adversarial training enhances robustness by maximizing loss, but theoretical analyses rarely explicitly link initial parameter constraints with the causal mechanisms underlying final adversarial robustness, making strategy design reliant on empirical trial-and-error.
-Third, the mathematical foundations of interpretability tools are weak [11, 12, 13, 14, 15].
+Third, the mathematical foundations of interpretability tools are weak Ribeiro2016 ; Lundberg2017 ; Venkata2025 ; Lee2023 ; Alvarez2018 .
 Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization.
@@
 In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
-Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learning[16, 17] .
+Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learningShwartz2017 ; Xu2022 .
 Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality.
-This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systems[18, 19, 20, 21, 22, 23].
-On this basis, this paper rigorously defines higher-order logical systems using recursive methods and establishes a formalized expression of states via the interpretation of sets of well-formed formulas, thus fully clarifying the mathematical essence of state representation in the definition of information.
-This fills the last theoretical gap in OIT and provides more powerful tools for applying OIT to analyze various problems in information science.
-Subsequently, the theories and methods of OIT are applied comprehensively—from prior knowledge encoding and data-driven training to model generalization applications—unifying core concepts such as prior distribution, data distribution, and parameter updates into a formalized state representation for the first time.
-This establishes a mathematical logic system that spans the entire machine learning process.
-Based on this, formal logical deduction and constraint rules for the causal chains in typical stages such as formal modeling and realization of training information, formal modeling and realization of model learning, and information processing and learning optimization in models are demonstrated, forming the Meta-Framework for Machine Learning Theory (MLT-MF).
-Finally, MLT-MF is applied to several hot topics in machine learning, establishing universal definitions for model interpretability and ethical safety, and proving three important theorems: the equivalence of model interpretability and information recoverability, ethical safety assurance, and generalization error estimation.
-This provides a powerful tool based on a unified formal logical framework for both theoretical analysis and engineering practice in machine learning.
+This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systemsXu2014 ; xu2015 ; Jianfeng2021 ; Jianfeng2023 ; Xu2024 ; Xu2025 .
+On this basis, we rigorously define higher-order logic systems using recursion, and establish formal state expressions through interpretations of sets of well-formed formulas (WFFs), completing the last theoretical gap in OIT regarding the mathematical essence of state expressions.
+We then thoroughly apply the theory and methods of OIT to unify the core concepts throughout the machine learning lifecycle—prior knowledge encoding, data-driven training, and model generalization—as formal state expressions; we establish a mathematical logic system covering the entire ML process.
+Using this, we analyze and prove the causal-chain logic and constraint rules for typical stages in machine learning, including the formal modeling and realization of training information, the formal modeling and realization of model learning, and the model’s information processing and learning optimization, thus forming a Meta-Framework for Machine Learning Theory (MLT-MF).
+Finally, using MLT-MF, we propose universal definitions of model interpretability and ethical safety, prove and validate—via a simple feedforward neural network—three important theorems: model interpretability and information existence, ethical safety assurance, and an upper bound estimate on total variation distance (TVD).
+These results provide a powerful tool, based on a unified formal logic framework, for both theoretical analysis and engineering practice in machine learning.
 2 Postulates and Formal Definition of Information
@@
 This is also a major reason why artificial intelligence often lacks interpretability and predictability.
-In response,[22] proposes four Postulates of information:
+To address this, we first propose the concept of an enabling mapping.
+Definition 1 (Enabling Mapping) Let $X$ and $Y$ be two sets, and let $f$ be a mapping.
+If for any $x\in X$, there exists a $y\in Y$, and for any $y\in Y$, there exists at least one $x\in X$ such that $y=f(x)$, and furthermore, the existence of $y$ is physically realized only because of $x$, then $f$ is called an enabling mapping from $X$ to $Y$.
+This is denoted as:
+$f:X\rightrightarrows Y$ (2.1)
+or
+$f(X)\underline{\triangleright}Y$ (2.2)
+[…]
+If $f(X)\underline{\triangleright}Y$, we also denote this as $f^{-1}(y)\underline{\triangleleft}x$ or $x\underline{\triangleright}f^{-1}(y)$, to maintain consistency in the directionality of enabling.
+Based on the concept of enabling mapping, Xu2024 proposed four postulates of information:
 Postulate 1 (Binary Attribute) Every piece of information comprises two components: the ontology, which represents the intrinsic essence of the information, and the carrier, which embodies the objective form of the information.
@@
 The concept of "enabling mapping" is novel.
-It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realization[24, 25, 26].
+It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realizationShannon1948 ; Landauer1996 ; Wiener2019 .
 2.2 Formal Representation of States
@@
 Functions $f_{1}^{1},f_{2}^{1},\ldots,f_{1}^{2},f_{2}^{2},\ldots,f_{1}^{3},f_{2}^{3},\ldots$;
-Parentheses (,)(,)( , );
+Parentheses $(,)$;
 Predicates $A_{1}^{1},A_{2}^{1},\ldots,A_{1}^{2},A_{2}^{2},\ldots,A_{1}^{3},A_{2}^{3},\ldots$;
@@
 For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols.
-However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more concise[27].
+However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more conciseHamilton1988 .
 Therefore, they are introduced as defined symbols in subsequent discussions.
-Definition 2 (Interpretation of a Formal System): An interpretation $E$ of formal system $\mathcal{L}$ over a domain $D_{E}$ consists of a non-empty set $D_{E}$, a specified set of elements $(\widehat{a}_{1},\widehat{a}_{2},\ldots)$ over $D_{E}$, a set of functions $(\widehat{f}_{i}^{n},n>0,i>0)$ over $D_{E}$, and a set of relations $(\widehat{A}_{i}^{n},n>0,i>0)$ over $D_{E}$ and the specified elements.
-According to this definition, the variables $x_{1},x_{2},\ldots$ in $\mathcal{L}$ are interpreted as objects.
-The set $D_{E}$ is the domain over which variables take their values.
-The set of specfied elements corresponds to the special objects denoted by the individual constants in $\mathcal{L}$.
-Functions and relations provide concrete interpretations for the function and predicate symbols in $\mathcal{L}$.
+Definition 3 (Interpretation of a Formal System) An interpretation $E$ of a formal system $\mathcal{L}$ over a domain $D_{E}$ assigns specific values to constants, elements of the domain $D_{E}$ to variables, and turns function and predicate symbols into functions and predicates over that domain $D_{E}$, thereby yielding a set of WFFs each with a truth value.
 Definition 3 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects as both the subjective and objective worlds.
 Postulate 5 (Expression of Object State):The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
-(1) Definitional Invariance: The definitions of the object set and the time set remain invariant over the entire domain.
-(2) Expressivity of Attributes: All attributes of the object set—including properties, structures, values, and relations—over the domain can be expressed via functions and predicates within a formal system conforming to Definition 1.
-(3) Expressivity of Rules: The evolution rules governing various attributes of the object set can be expressed, without implicit self-reference, via temporal variables and application connectives in a formal system as specified in Definition 1.
-(4) Logical Consistency: There exists no logical contradiction among the various attributes and evolution rules of the object set, nor among any possible logical conjunctions thereof.
-Theorem 1 (Formal Expression of States): Let $S(X,T)$ be the state set of an object collection $X$ over a specified time set $T$, satisfying Postulate 5.
-Then $S(X,T)$ can always be represented as the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$, and is logically consistent.
-Proof: Since $S(X,T)$ satisfies Postulate 5, we have:
-  • By (1), the elements of $X\times T$ are well-defined, so $S(X,T)$ has a definite domain.
-  • By (2), all attributes of $X$ over $T$ in $S(X,T)$ can be expressed by functions and predicates in a formal system $\mathcal{L}$ as specified in Definition 1.
-  • By (3), suppose $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ and $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$ are two attributes in $S(X,T)$ pertaining to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$.
-If every $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ evolves from $t_{1}$ to $t_{2}$ according to some rule into $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$,
-    then $(\forall A_{1})(\forall f_{1})(\forall x_{1})(A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))\rightarrow\exists A_{2}\exists f_{2}\exists x_{2}(A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))))$ is a well-formed formula in $\mathcal{L}$ that, as an interpretation over $x_{1},x_{2},t_{1},t_{2}$, expresses the corresponding rule in $S(X,T)$.
-Other rules in $S(X,T)$ can likewise be expressed through corresponding well-formed formulas in $\mathcal{L}$ interpreted over $X\times T$.
-  • Let the set of such well-formed formulas in $\mathcal{L}$ representing $S(X,T)$ be defined as its interpretation ${E}$ over the domain $X\times T$.
-  • By (4), no logical contradiction exists among any attributes or rules in $S(X,T)$, nor among any logical conjunctions thereof.
-Thus, as a representation of $S(X,T)$, both ${E}$ and its logical closure are logically consistent.
-In summary, $S(X,T)$ is indeed an interpretation ${E}$ of a set of well-formed formulas in $\mathcal{L}$ over the domain $X\times T$, and both itself and its logical closure are logically consistent.
-The theorem is thus proved.
+(1) Domain invariance: definitions of the object set and time set remain unchanged.
+(2) Numerical determinacy: concrete numerical values determine values of certain properties and forms.
+(3)Attribute expressibility: properties, forms, values, relations are expressible via functions and predicates in a formal system satisfying Definition 2.
+(4) Rule expressibility: evolution rules among attributes are expressible in such a formal system using time variables and implication, and contain no implicit self-reference.
+(5) Logical coherence: no logical contradictions exist among attributes and rules, nor among any logical conjunctions thereof.
+Theorem 1 (Formal Expression of States) If the state set $S(X,T)$ of objects $X$ over times $T$ satisfies Postulate 5, then it can be represented as an interpretation of a formal system $\mathcal{L}$ over the domain $X\times T$, and it is logically coherent.
+Proof: Since the state set $S(X,T)$ of the object set $X$ over the time set $T$ satisfies Postulate 5, we construct the corresponding formal system $\mathcal{L}$ as follows:
+By (1), the definitions of the elements in $X$ and $T$ remain unchanged, ensuring determinacy.
+The variables in the formal system $\mathcal{L}$ are defined to correspond to each element of $X\times T$, thus making $X\times T$ the domain of discourse for $\mathcal{L}$.
+Consequently, $S(X,T)$ has a well-defined domain.
+By (2), the constants in $\mathcal{L}$ are defined to represent all possible values in $S(X,T)$.
+By (3), the functions and predicates in $\mathcal{L}$ represent the properties, forms, values, and relations of $X$ over $T$ as captured in $S(X,T)$.
+These properties can be expressed through functions and predicates in the formal system $\mathcal{L}$ that conform to Definition 1.
+By (4), consider two ternary predicates $A_{1}^{3}$ and $A_{2}^{3}$.And then let $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ and $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$ be two properties of $S(X,T)$ corresponding to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and binary functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$.
+Suppose all $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ satisfy the rule of evolving from $t_{1}$ to $t_{2}$ into some $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$.
+The formula $(\forall A_{1}^{3}\forall f_{1}^{2}\forall x_{1}A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))\rightarrow\exists A_{2}^{3}\exists f_{2}^{2}\exists x_{2}A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2})))$ is also a well-formed formula in $\mathcal{L}$, with respect to the interpretation of $x_{1},x_{2},t_{1},t_{2}$, and is able to express the corresponding rules in $S(X,T)$.
+For other rules in $S(X,T)$, the related elements in the domain $X\times T$ can similarly be interpreted and expressed via corresponding WFFs in $\mathcal{L}$, with no implicit logical self-reference.
+Therefore, $S(X,T)$ constitutes an interpretation $E$ of $\mathcal{L}$ over the domain $X\times T$. By (5), both itself and its logical closure are logically consistent. The theorem is thus proved.
+Note that Theorem 1 stipulates that variables in $\mathcal{L}$ are elements of $X\times T$; thus, in formulas, object and time elements appear in pairs $(x,t)$.
+In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
 Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world.
-Both theoretical and empirical analyses [28] further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language.
+Both theoretical and empirical analyses qiu2025 further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language.
 This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
@@
 Theorem 2 (State Representation of Finite Automata): For any finite automaton $M$, one can describe its input, output, and state-transition behavior over a related time set $T$ by the state set $S(M,T)$.
-Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n_{T}\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions.
-By finite automaton theory [29], $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n_{T}-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n_{r}\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n_{u}\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
+Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions.
+By finite automaton theory mathsea2002 , $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
 $\delta:U\times Q\to U$ is the next-state function;
@@
 $\varphi_{U}(t_{i})=State^{3}(M,t_{i},u_{t_{i}})$ (2.3)
-to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$.
+to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$$i=1,\ldots,n$.
 For each input $q_{t_{i}}\in Q$, define the input predicate $Input^{3}$ by
 $\varphi_{I}(t_{i})=Input^{3}(M,t_{i},q_{t_{i}})$ (2.4)
-to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$.
+to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$$i=1,\ldots,n-1$.
 For each output $r_{t_{i}}\in R$, define the output predicate $Output^{3}$ by
 $\varphi_{O}(t_{i})=Output^{3}(M,t_{i},r_{t_{i}})$ (2.5)
-to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$.
+to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$$i=2,\ldots,n$.
 The transition function $\delta$ is captured by the well-formed formula
 $\varphi_{\delta}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land Eq^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (2.6)
-where the binary predicate $Eq^{2}(x,y)$ asserts $x=y$.
+[…]
 Similarly, the output function $\lambda$ is given by
@@
 $IsElementof^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$
-where $IsElementof^{2}(x,X)$ asserts $x$ is an element of $X$.
+[…]
 Thus the state set
-$S(M,T)=Cn(\{\varphi_{U}(t_{i})\}\cup\{\varphi_{I}(t_{i})\}\cup\{\varphi_{O}(t_{i})\}\cup\{\varphi_{\delta}(t_{i})\}\cup\{\varphi_{\lambda}(t_{i})\})$ (2.6)
+$S(M,T)=$ $\{\varphi_{ij}(t_{i}),\varphi_{I}(t_{j}),\varphi_{O}(t_{k}),\varphi_{S}(t_{j}),\varphi_{A}(t_{j}),$ (2.8)
+$i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}$
 describes $M$’s inputs, outputs, and state-transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
@@
 2.4 Definition of Information
-As shown in[22], the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$.
+As shown inXu2024 , the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$.
 Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
@@
 The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets.
-Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, that is,
-$I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ (2.7)
-or
-$I\left(S_{o}(o,T_{h})\right)\underline{\triangleright}S_{c}(c,T_{m})$ (2.8)
-The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
-[…]
-This shows that the carrier state set $S_{c}(c,T_{m})$ is always induced by the ontological state set $S_{o}(o,T_{h})$, and thus the enabling mapping $I$ must be surjective.
-For simplicity, unless otherwise noted, subsequent discussions assume that $I$ is a single-valued mapping.
-Therefore, for any $\psi\in S_{c}(c,T_{m})$, there exists at least one $\varphi\in S_{o}(o,T_{h})$ such that $I(\varphi)\underline{\triangleright}\psi$.
-In this case, we write $I^{-1}(\psi)\underline{\triangleleft}\varphi$ or $\varphi\underline{\triangleright}I^{-1}(\psi)$ to preserve the consistency of the enabling direction.
+Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$.The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
 It is important to note that both $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are state representations that satisfy Postulate 5.
@@
 Recoverable information and the recoverable reduction of information can provide significant assistance in simplifying problems.
-Reference [24]is generated as a foundational work in information theory.
-The paper proposed that "$H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information.
+Reference Shannon1948 is generated as a foundational work in information theory.
+The paper proposed that $H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information.
 The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively.
 The channel $I_{c}$ is the medium through which information is transmitted.
-After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $01\leq i\leq n,\tau>$ 0).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$.
+After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $1\leq i\leq n,\tau>0$).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$.
 Then,
@@
 The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
+2.5 Transmissibility of Information
+Xu2014 ; xu2015 discuss transmissibility using composite mappings. Building on Definitions 1 and 5, we can characterize the fundamentals of information transmission more rigorously:
+Theorem 3 (Transmissibility of Information) If $I_{1}=\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\rangle$ is information from an ontology $o$ to a carrier $c_{1}$, and $I_{2}=\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is information from carrier $c_{1}$ to carrier $c_{2}$, then the composition $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is also information, which we call transmission of information from $I_{1}$ via carrier $c_{1}$ to carrier $c_{2}$.
+Proof: According to Definition 1, Definition 5, and the given conditions, for any $S_{o}(x,t)\in S_{o}(o,T_{h})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$.
+Furthermore, there exists $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$.
+We define $I_{3}(S_{o}(x,t))=S_{c_{2}}(x_{2},t_{2})$.
+Since the physical realization of $S_{c_{2}}(x_{2},t_{2})$ depends on the existence of $S_{c_{1}}(x_{1},t_{1})$, and the existence of $S_{c_{1}}(x_{1},t_{1})$ in turn depends on $S_{o}(x,t)$, it follows by transitivity of causality that the realization of $S_{c_{2}}(x_{2},t_{2})$ ultimately depends on $S_{o}(x,t)$.
+Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$.
+Thus, $I_{3}(S_{o}(x,t))=I_{2}(I_{1}(S_{o}(x,t)))\underline{\triangleright}I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$.
+Therefore, $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 5.
+In essence, this is the information $I_{1}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$.
+The theorem is thus proved.
+Transmission is an important way of information flow.
+Communication, processing, input, and output are all concrete manifestations of information transmission.
+In many cases, information undergoes more complex flows through chain-like transmission mechanisms.
+From Theorem 3, we can readily derive the following:
+Corollary 1 (Information Chain) Let $\{I_{i}=\langle o,T_{h},S_{o},c_{i},T_{m_{i}},S_{c_{i}}\rangle|i=1,\ldots,n\}$ be a sequence of information mappings, where for any $i>1$, it holds that $c_{i-1}=o$, $T_{h}=T_{m_{i-1}}$, and $S_{o}(o,T_{h})=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$.
+In this case, we say that these mappings constitute an information chain.
+Moreover, for any $1\leq i\leq j\leq n$, $I_{ij}=\langle o,T_{h},S_{o},c_{j},T_{m_{j}},S_{c_{j}}\rangle$ also qualifies as information according to Definition 5.
 2.6 Noisy Information
@@
 Definition 8 (Noisy Information): For information $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $\widetilde{S_{o}}(o,T_{h})$ be another state of the same carrier $c$ at time $T_{m}$ and $\widetilde{S_{c}}(c,T_{m})$ be another state of the same carrier $c$ at time $T_{m}$,
-$\widetilde{S_{o}}(o,T_{h})=S_{o}(o,T_{h})\backslash S_{o}^{-}(o,T_{h})\cup S_{o}^{+}(o,T_{h})$ (2.12)
+[…]
 where the state $S^{-}_{o}(o,T_{h})$ is the loss state of $S_{o}(o,T_{h})$, satisfying $S^{-}_{o}(o,T_{h})\subseteq$ $S_{o}(o,T_{h})$ and not being enabled to map to $\widetilde{S_{c}}(c,T_{m})$, and the state $S^{+}_{o}(o,T_{h})$ is the superposed state of $S_{o}(o,T_{h})$, satisfying $S_{o}(o,T_{h})\cap S_{o}^{+}(o,T_{h})=\emptyset$ and being enabled to map to $\widetilde{S_{c}}(c,T_{m})$.
@@
 Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information.
-But under the condition where the noiseless information $I$cannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$.
+But under the condition where the noiseless information IIcannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$.
 This is the practical significance of introducing the concept of noisy information.
-Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario[24].
+Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenarioShannon1948 .
 In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
@@
 $I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.16)
-where $\widetilde{S_{I_{c}}}(I_{s},T_{h})=S_{I_{s}}(I_{s},T_{h})\backslash S_{I_{s}}^{-}(I_{s},T_{h})\cup S_{I_{s}}^{+}(I_{s},T_{h})$.
-In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels in[24], demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
+[…]
+In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels inShannon1948 , demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
 3 Meta-Framework for Machine Learning Theory (MLT-MF)
-The core of machine learning is the model, which is a typical finite automaton and also an information system.
-Therefore, every learning process and the subsequent testing or application process rely on the information within it, which can generally be summarized and divided into four main steps of information realization and operation.
-Step 1: The implementation of training information $It$, in which the ontological state of $It$ is enabled and mapped onto the model interface as its carrier;
-Step 2: The original information $Io$ within the model—endowed with learning capability—is fused with the training information $It$, and optimized to yield the updated information $Iu$, which possesses both the ability to answer specific tasks and the capacity for further learning;
-Step 3: The implementation of user input information $Iq$, where the ontological state of $Iq$ is likewise enabled and mapped onto the model interface as its carrier;
-Step 4: The updated information $Iu$ within the model processes the input $Iq$, outputs the response information $Ir$ according to user requirements, and simultaneously undergoes further optimization through learning, resulting in $Iv$.
+The core of machine learning is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system.
+The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system.
+Therefore, by employing the concepts of finite automaton state and information models established in Section 2, the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
+3.1 Construction of Initial Information
+Suppose the model under study is denoted as M, and let $T=\{t_{i}|i=1,\ldots,n\}$ be the set of time points at which M, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order.
+According to the perspective of OIT, the creation of M is essentially the realization of information
+$I_{o}=\langle M_{D},T_{o_{h}},S_{M_{D}},M,t_{1},S_{M}\rangle$ (3.1)
+where the ontology $M_{D}$ of $I_{o}$ is the designer’s conceptualization of M; $T_{o_{h}}$ is the time at which the designer conceives M, and $S_{M_{D}}(M_{D},T_{o_{h}})$ is the initial finite automaton state that must be realized at this conceptual stage.
+The carrier M of $I_{o}$ is the model itself; the reflection time $t_{1}$ is when M is constructed according to $S_{M_{D}}(M_{D},T_{o_{h}})$; and $S_{M}(M,t_{1})$ is the formalization of the actual physical state of M upon its completion.
+Thus, $\sup(T_{o_{h}})<\sup(t_{1})$.
+According to Theorem 2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
+$\varphi_{o_{1}}=State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ (3.2)
+denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},T_{o_{h}})$ as envisioned by the designer are assigned to $M_{D}$.
+$\varphi_{o_{2}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
+$Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
+$(Eq^{3}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land State^{3}(M_{D},t_{2},\Phi^{2}(M_{D},t_{2}))$ (3.3)
+expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},T_{o_{h}})$, and the input carrier state $S_{M_{I}}(M_{I},t_{1})$ is provided and satisfies the learnability condition determined by the binary predicate $Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, $M_{D}$ can optimize its internal structures, parameters, and algorithmsto $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, assigning this to $M_{D}$.
+This demonstrates the model’s capacity for learning and optimization, where $M_{I}$ denotes the input device of M.
+$\varphi_{o_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
+$Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$
+$(Eq^{3}(S_{M_{O}}(M_{O},t_{2}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land Output^{3}(M,t_{2},S_{M_{O}}(M_{O},t_{2})))$ (3.4)
+indicates that, at time $t_{1}$, if $S_{M_{I}}(M_{I},t_{1})$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate $Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))$, then $M_{D}$ can process $S_{M_{I}}(M_{I},t_{1})$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$ to yield the output carrier state $S_{M_{O}}(M_{O},t_{2})$, which is then output at time $t_{2}$.
+This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of M.
+Thus,
+$S_{M_{D}}(M_{D},T_{o_{h}})=\{\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}\}$ (3.5)
+constitutes the initial state of M as conceived by the designer.
+It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},T_{o_{h}})$, and $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ all represent sets, $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas.
+Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states.
+Therefore, subsequent discussions will encompass both cases without further elaboration.
+$S_{M_{D}}(M_{D},T_{o_{h}})$ may represent the designer’s idea, a design draft on paper, or an electronic document.
+Through the information $I_{O}$, these are enabled and mapped into the physically operational state of the model M:
+$S_{M}(M,t_{1})=\{\psi_{o_{1}},\psi_{o_{2}},\psi_{o_{3}}\}$ (3.6)
+where $\psi_{o_{1}}\underline{\triangleleft}I_{o}(\varphi_{o_{1}})$, $\psi_{o_{2}}\underline{\triangleleft}I_{o}(\varphi_{o_{2}})$, and $\psi_{o_{3}}\underline{\triangleleft}I_{o}(\varphi_{o_{3}})$ are the enabling mappings of $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$, respectively.
+This process establishes M as a typical finite automaton and an operational information system.
 3.2 Formal Modeling and Realization of Training Information
-According to the general formal definition of information, the formal representation of training information implementation is straightforward:
-$It=(ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct})$ (3.1)
-$S_{ot}(ot,Tt_{h})=\{\varphi t|\varphi t\text{ is a well-formed formula concerning }ot\text{ and }Tt_{h}\}$ (3.2)
-$S_{ct}(ct,Tt_{m})=\{\psi t|\psi t\text{ is a well-formed formula concerning }ct\text{ and }Tt_{m}\}$ (3.3)
-[…]
-Therefore, it is necessary to introduce a model training data interface $ct$—akin to hardware interfaces or software APIs—as the carrier for $It$, so that $S_{ot}(ot,Tt_{h})$ can be ingested in an engineering-compatible manner.
-The reflection time set $Tt_{m}$ denotes the period during which the carrier reflects $It$, while the carrier reflection set $S_{ct}(ct,Tt_{m})$ comprises specific, complete, and operational realizations of $It$ on $ct$ at $Tt_{m}$.
-Each element $\psi t$ is a well-formed formula enabled by, and logically/semantically corresponding to, some $\varphi t$.
-Continuing the example above, if $S_{ot}(ot,Tt_{h})$ is enabled and mapped to $ct$ at time $T_{3}$, we might express
-$S_{ct}(ct,T_{3})=\{\psi t\}=\{(\text{TrainTable}[\text{user}_{id}=X][\text{time}=T_{1}]=\text{Active})\land$
-[…]
-where “Active” and “Tgt60” are defined in the training table as “user is active” and “device temperature exceeds 60°C,” respectively.
-Although the form of $\psi t$ may appear simple, it fully reflects the semantic content of $\varphi t$.
-In practice, $\psi t$ may take the form of tables, vectors, JSON objects, API parameters, or other formal representations of engineering data interfaces or hardware device states, and may include additional data constraints or format checks to ensure the validity and robustness of $S_{ct}(ct,T_{3})$.
-It is crucial to emphasize that $S_{ot}(ot,Tt_{h})$ is only realized as $S_{ct}(ct,T_{3})$ via the enabling mapping of $It$.
-The fact that $S_{ot}(ot,Tt_{h})$ satisfies Postulate 5 also implies that $S_{ct}(ct,T_{3})$ satisfies Postulate 5.
-This mapping embodies both the structural-semantic correspondence from the former to the latter and the causal relationship—where the ontological states are the cause and the carrier states are the effect, requiring expenditure of human, temporal, and energy resources.
-Such an enabling mapping allows the knowledge, facts, and rules formalized within the ontology to be incorporated into the model in a structured, standardized, and machine-readable fashion, thereby becoming usable for subsequent model learning.
-In practical engineering, the realization of training information often needs to consider noisy information $\widetilde{It}$ due to discretization, attribute pruning, and noise accumulation.
-However, for theoretical analysis, we currently focus only on the ideal case of no noise, leaving the study of complex noisy conditions for further research.
+According to the general formal definition of information, the formal expression for the realization of training information is straightforward:
+$I_{t}=\langle ot,T_{th},S_{ot},M_{I},t_{1},S_{M_{I}}\rangle$ (3.7)
+$S_{ot}(ot,T_{th})=\{\varphi_{t}|\varphi_{t}\text{ is a WFF about }ot\text{ and }T_{th}\}$ (3.8)
+$S_{M_{I}}(M_{I},t_{1})=\{\psi_{t}|\psi_{t}\text{ is a WFF about }M_{I}\text{ and }t_{1}\}$ (3.9)
+Here, the ontology $ot$ refers to the object containing knowledge or samples; $T_{th}$ is the time at which its state is formed; and $S_{ot}(ot,T_{th})$ is a set of business-oriented semantic assertions or facts about the ontology $M_{I}$ functions similarly to the hardware interface and API software of M, carrying the training information at time $t_{1}$.
+The carrier state $S_{M_{I}}(M_{I},t_{1})$ is the physical realization of $S_{ot}(ot,T_{th})$ via an enabling mapping, ensuring that the information can be machine-read and entered into M.
 3.3 Formal Modeling and Implementation of Model Learning
-From the perspective of information flow and transformation, the learning process of an artificial intelligence model M can be viewed as the optimization of its existing information $Io$ through learning from the training information $It$, resulting in the updated information $Iu$.
-The preceding section has already detailed the modeling and realization of training information $It$.
-We now turn our attention to the modeling and implementation of $Io$ and $Iu$ within M.
-By analogy with equations (3.1-3.3) for $It$, we can define $Io$ and $Iu$ by substituting the respective subscript $t$ with $o$ and $u$.
-It is important to note that, since $Io$ and $Iu$ represent the information present in M before and after learning, their carriers $co$ and $cu$ are effectively identical—in other words, both can directly be denoted by M, or more precisely, by the computational system that executes M:
-$co=cu=M$ (3.4)
-Within the learning process of M, the most crucial component is the existing information $Io$, which, in addition to encapsulating knowledge, propositions, facts, or samples (as with $It$), is distinguished by its possession of learning rules—i.e., the rules by which it is optimized into updated information through learning.
-Definition 8 (State with Learning Capability): Let the states $Sx(x,T_{x})$, $Sy(y,T_{y})$, and $Sz(z,T_{z})$ be interpretations $E_{x}$, $E_{y}$, and $E_{z}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $x$ over time interval $T_{x}$, object $y$ over $T_{y}$, and object $z$ over $T_{z}$, all satisfying Postulate 5.
-We introduce the following specific functions and predicates in $\mathcal{L}$:
-  • Function: $f_{learn}^{2}(X,Y)$ denotes the well-formed formula resulting from the learning of set $X$ from set $Y$.
-Moreover, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, the function $f_{learn}^{2}(X,Y)$ also satisfies this logical consistency requirement.
-  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Learnable}^{2}(X,Y)$ indicates that state set $X$ can learn from state set $Y$; $\text{IsFormulaOf}^{2}(x,X)$ indicates that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ denotes that $x$ is equal to $y$.
-Then suppose $\varphi x\in Sx(x,T_{x})$, $\sup(T_{x},T_{y})\leq\inf(T_{z})$, and
-$\varphi x=$ $(\forall\Phi x)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi x,Sx(x,T_{x}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
-$\text{Learnable}^{2}(\Phi x,\Phi y))\rightarrow\exists\varphi z(\text{IsFormulaOf}^{2}(\varphi z,Sz(z,T_{z}))\land\text{Eq}^{2}(\varphi z,f_{learn}^{2}(\Phi x,\Phi y)))$ (3.5)
-we say that $Sx(x,T_{x})$ has the capability to learn from $Sy(y,T_{y})$ and optimize into $Sz(z,T_{z})$; equivalently, we say that the interpretation $E_{x}$ has the capacity to learn from $E_{y}$ and optimize toward $E_{z}$.
-The formula $\varphi x$ essentially constitutes the learning rule for the state $Sx(x,T_{x})$, indicating that as long as $Sy(y,T_{y})$ meets the learning conditions of $Sx(x,T_{x})$, it is possible to augment $Sz(z,T_{z})$ with new well-formed formulas $\{\varphi u|\varphi u=f_{learn}^{2}(\Phi x,\Phi y),\Phi x\subseteq Sx(x,T_{x}),\Phi y\subseteq Sy(y,T_{y})\}$, thus rendering $Sz(z,T_{z})$ an optimized extension of $Sx(x,T_{x})$.
-Furthermore, given the logical consistency of $f_{learn}^{2}(X,Y)$, if $Sx(x,T_{x})$ and $Sy(y,T_{y})$ satisfy Postulate 5, then $Sy(y,T_{y})$ also satisfies Postulate 5.
-On the other hand, the inheritance of learning capability is also a fundamental property of model learning.
-In fact, to inherit the learning capability of $Sx(x,T_{x})$, it suffices to retain the learning rule $\varphi x$ within $Sz(z,T_{z})$, with only partial adaptive substitution of its symbols.
-Specifically, for $\varphi z\in Sz(z,T_{z})$:
-$\varphi z=($ $\forall\Phi z)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
-$\text{Learnable}^{2}(\Phi z,\Phi y))\rightarrow\exists\varphi w(\text{IsFormulaOf}^{2}(\varphi w,Sw(w,T_{w}))\land$
-$\text{Eq}^{2}(\varphi w,f_{learn}^{2}(\Phi u,\Phi y)))$ (3.6)
-This ensures that $Sz(z,T_{z})$ can learn from $Sy(y,T_{y})$ and optimize into $Sw(w,T_{w})$. Thus, we establish:
-Proposition 3 (Inheritance of Learning Capability): If the state set $Sx(x,T_{x})$ is capable of learning from $Sy(y,T_{y})$ and optimizing into $Sz(z,T_{z})$, then $Sz(z,T_{z})$ always inherits the learning capability of $Sx(x,T_{x})$.
-Definition 9 (Realization of Model Learning): For an AI model M with original information $Io$, training information $It$, and updated information $Iu$, if all three are recoverable, and the ontological state $S_{oo}(oo,To_{h})$ can learn from $S_{ot}(ot,Tt_{h})$ and be optimized into $S_{ou}(ou,Tu_{h})$, and the carrier state $S_{co}(co,To_{m})$ can likewise learn from $S_{ct}(ct,Tt_{m})$ and be optimized into $S_{cu}(cu,Tu_{m})$, then M is said to possess the capability to realize the learning and optimization of $Io$ towards $It$, resulting in $Iu$—in short, M has information learning capability.
-According to (3.4), we may also write $S_{co}(co,To_{m})=S_{M}(M,To_{m})$, $S_{cu}(cu,Tu_{m})=S_{M}(M,Tu_{m})$, indicating that the learning process is essentially an optimization of the state of model M.
-In summary, we have theoretically internalized the model’s learning capability as a property of its states.
-These rules are independent of the algorithmic structure or scale of the model, thereby providing a unified and general formal framework for the analysis of various types of models.
-As an example, let $Mn$ be a neural network model.
-According to common knowledge, the information in $Mn$ is $Io=\langle oo,To_{h},S_{oo},Mn,To_{m},S_{co}\rangle$, where $oo$ represents the model in the designer’s mind, $To_{h}$ is the time set during the design of $Mn$.
-$S_{oo}(oo,To_{h})=\Phi o=\{Params(\theta)\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.7)
-Here, $Params(\theta)$ denotes the predicate representation of the parameter set $,\theta=\{W_{1},b_{1},\ldots$$,W_{n},b_{n}\}$; $Architecture(L_{1},\ldots,L_{n})$ denotes the predicate representation of the network layer structure; and $Activation(A)$ denotes the predicate representation of the activation function set $A=\{a_{1},\ldots,a_{n}\}$.
-The carrier for $Io$ is the neural network model $Mn$ itself, with $To_{m}$ representing the period after its development, and $S_{co}(co,To_{m})$ being the physical instantiation of (3.7) within $Mn$.
-The training information for $Mn$ is $It=\langle ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct}\rangle$, where $ot$ is a set of real-world objects, $Tt_{h}$ is the time set during which the state $S_{ot}(ot,Tt_{h})$ occurs, and
-$S_{ot}(ot,Tt_{h})=\Phi t=\{Batch(X,y)\land DimMatch(X,L_{1})\land Differentiable(L)\}$ (3.8)
-where $Batch(X,y)$ expresses the predicate for input data $X\in\mathbb{R}^{b\times k}$ and labels $y\in\mathbb{R}^{b\times k}$; $DimMatch(X,L_{1})$ expresses that the input data $X$ matches the dimensions of the input layer $L_{1}$; and $Differentiable(L)$ expresses the differentiability of the loss function $L$ with respect to parameters $\theta$.
-The definition of the predicate $Learnable^{2}$ is as follows:
-$Learnable^{2}(\Phi o,\Phi t)=$ $Dimensionality(\Phi o,\Phi t)\land ActivationFeasible(\Phi o,\Phi t)\land$ (3.9)
-$GradientExists(\Phi t)$
-where $Dimensionality(\Phi o,\Phi t)=DIM(X,b\times d)\land DIM(L_{1},d)$ ensures that the input data dimensions are consistent with the input layer, $ActivationFeasible(\Phi o,\Phi t)$ ensures that for every $a_{i}\in A$, $Range(X)\subseteq Domain(a_{i})$, i.e., the input data lies within the domain of the activation function, and $GradientExists(\Phi t)$ ensures the existence of a gradient $\nabla_{\theta}L(\theta,X,y)$, i.e., that the loss function $L$ is differentiable with respect to $\theta$.
-The learning function $f_{learn}^{2}$ is defined as:
-$f_{learn}^{2}(\Phi o,\Phi t)=\{Params(\theta_{new})\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.10)
-where $\theta_{new}=\theta-\eta(\nabla_{\theta}L(\theta,X,y)+\beta v_{prev})$ is the iterative update formula for the neural network model.
-Consequently, in the updated information $Iu=(ou,Tu_{h},S_{ou},Mn,Tu_{m},S_{cu})$ of the neural network model, the ontology $ou$ is the union of the designer’s conceptualization and the ontology $ot$ of the training information, $Tu_{h}$ is the set of training times for $Mn$,
-$S_{ou}(ou,Tu_{h})=f_{learn}^{2}(\Phi o,\Phi t)$ (3.11)
-constitutes the optimized ontological state after learning, with the carrier of $Iu$ being $Mn$ itself, the reflection time $Tu_{m}$ being the post-training period, and $S_{cu}(Mn,Tu_{m})$ as the physical realization of $S_{ou}(ou,Tu_{h})$ within $Mn$, ideally satisfying the mathematical and physical requirements of the enabling mapping:
-$S_{cu}(Mn,Tu_{m})=f_{learn}^{2}(\Psi o,\Psi t)$ (3.12)
-where $\Psi o$ and $\Psi t$ are the sets of well-formed formulas corresponding to the carrier states of $Io$ and $It$, that is, the physical realizations of $\Phi o$ and $\Phi t$ within $Mn$ and its interface.
-Equations (3.11) and (3.12) together demonstrate that, within $Mn$, both the ontological state $S_{oo}(oo,To_{h})$ and the carrier state $S_{co}(co,To_{m})$ possess the capability to learn from $S_{ot}(ot,Tt_{h})$ and $S_{ct}(ct,Tt_{m})$, and then optimize to $S_{ou}(ou,Tu_{h})$ and $S_{cu}(cu,Tu_{m})$, respectively; thus, $Mn$ indeed realizes the learning and optimization of $Io$ with respect to $It$, resulting in $Iu$.
-This completes the proof.
-It should be noted that some of the content in the above discussion does not strictly adopt set definitions and well-formed formulas.
-This does not mean that they cannot be converted into complete set definitions and well-formed formulas, but rather because such representation is more concise and intuitive, and does not lead to ambiguity.
+Model learning for M involves fusing its original information $I_{o}$ with the training information $I_{t}$, resulting in updated information:
+$I_{u}=\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\rangle$ (3.10)
+Here, the ontology is the hybrid $M\cup M_{I}$, encompassing both the model and its interface; $t_{1}\cup t_{2}$ marks the time of the first learning event; and $S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})$ formally represents the hybrid state during learning:
+$S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})=\{\varphi_{u_{1}},\varphi_{u_{2}},\varphi_{u_{3}}\}$ (3.11)
+$\varphi_{u_{1}}=Eq^{3}(\Phi^{2}(M,t_{2}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.12)
+indicates that the input state $S_{M_{I}}(M_{I},t_{1})$, which was learned by M at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of M at $t_{2}$.
+$\varphi_{u_{2}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
+$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
+$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.13)
+shows that after the first learning step, M remains capable of further learning and optimization.
+$\varphi_{u_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
+$Processable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
+$(Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.14)
+indicates that when the training information $I_{t}$ is processable by M, it can also produce output information.
+This state is then enabled and mapped to the physical state of the carrier M at time $t_{2}$:
+$S_{M}(M,t_{2})=\{\psi_{u_{1}},\psi_{u_{2}},\psi_{u_{3}}\}$ (3.15)
 3.4 Information Processing and Learning Optimization in Models
-The fundamental objective of model learning is to more accurately process user input information $Iq$ and to deliver high-quality output information $Ir$ that meets user requirements.
-Since the modeling and realization of input information $Iq$ are almost identical to those of $It$, as discussed in detail in Section 3.1, they will not be reiterated here.
-Thus, our focus shifts to the process whereby the updated information $Iu$ within model M processes $Iq$, outputs $Ir$, and is further optimized through learning to yield $Iv$.
-In this context, we continue to define $Iq$, $Ir$, and $Iv$ in the manner of equations (3.1-3.3).
-Definition 10 (Formal Representation of State Processing): Let the states $Sz(z,T_{z})$, $Su(u,T_{u})$, and $Sv(v,T_{v})$ be interpretations $E_{z}$, $E_{u}$, and $E_{v}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $z$ over time interval $T_{z}$, object $u$ over $T_{u}$, and object $v$ over $T_{v}$, all satisfying Postulate 5.
-We introduce the following specific functions and predicates in $\mathcal{L}$:
-  • Function: $f_{process}^{2}(X,Y)$ denotes the set of well-formed formulas output by state set $X$ after processing state set $Y$.
-Similarly, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, $f_{process}^{2}(X,Y)$ also meets this logical consistency requirement.
-  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Processable}^{2}(X,Y)$ indicates that state set $X$ is capable of processing state set $Y$ and generating an output; $\text{IsFormulaOf}^{2}(x,X)$ denotes that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ indicates that $x$ is equal to $y$.
-Then suppose $\varphi z\in Sz(z,T_{z})$, $\sup(T_{z},T_{u})\leq\inf(T_{v})$, and
-$\varphi z=($ $\forall\Phi z)(\forall\Phi u)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi u,Su(u,T_{u}))\land$
-$\text{Processable}^{2}(\Phi z,\Phi u))\rightarrow\exists\varphi v(\text{IsFormulaOf}^{2}(\varphi v,Sv(v,T_{v}))\land$
-$\text{Eq}^{2}(\varphi v,f_{process}^{2}(\Phi z,\Phi u)))$ (3.13)
-we say that $Sz(z,T_{z})$ is capable of processing $Su(u,T_{u})$ and outputting $Sv(v,T_{v})$; equivalently, we say that the interpretation $E_{z}$ has the ability to process $E_{u}$ and generate $E_{v}$ as output.
-Clearly, by combining Definitions 8 and 10, we can also formally express the process whereby state $S_{ou}(ou,Tu_{h})$ processes and learns from $S_{oq}(oq,Tq_{h})$, outputs $S_{or}(or,Tr_{h})$, and is subsequently optimized into $S_{ov}(ov,Tv_{h})$.
-In addition, when $S_{ou}(ou,Tu_{h})$ and $S_{oq}(oq,Tq_{h})$ both satisfy Postulate 5, the logical consistency of $f_{learn}^{2}(X,Y)$ and $f_{process}^{2}(X,Y)$ ensures that $S_{or}(or,Tr_{h})$ and $S_{ov}(ov,Tv_{h})$ also satisfy Postulate 5.
-Thus, we have:
-Definition 11 (Implementation of Model Information Processing and Learning Optimization): For an AI model M with updated information $Iu$, input information $Iq$, output information $Ir$, and optimized information $Iv$, if all four are recoverable, and the ontological state $S_{ou}(ou,Tu_{h})$ is capable of processing and learning from $S_{oq}(oq,Tq_{h})$, outputting $S_{or}(or,Tr_{h})$, and further optimizing into $S_{ov}(ov,Tv_{h})$, while the carrier state $S_{M}(M,Tu_{h})$ is similarly capable of processing and learning from $S_{cq}(cq,Tq_{h})$, outputting $S_{cr}(cr,Tr_{h})$, and further optimizing into $S_{M}(M,Tv_{h})$, then M is said to possess the capability to realize the information processing and learning optimization of $Iu$ with respect to $Iq$, outputting $Ir$ and optimizing into $Iv$.
-In short, M possesses information processing and learning optimization capability.
-For the neural network model $Mn$ discussed in the previous section, the processable predicate $\text{Processable}^{2}$ is defined as:
-$\text{Processable}^{2}(\Phi u,\Phi q)=\text{InputMatch}(\Phi u,\Phi q)\land\text{ActivationFeasible}(\Phi u,\Phi q)\land\text{OutputDefined}(\Phi u)$ (3.14)
-[…]
-The processing function $f_{process}^{2}$ is defined as:
-$f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.15)
-where $f_{\Phi u}$ is the prediction function explicitly expressed by the structure and parameters of the neural network $Mn$.
-Thus, for the neural network model $Mn$, the ontological state of the updated information $Iu$ is $S_{ou}(ou,Tu_{h})=\Phi u$, and the ontological state of the input information $Iq$ is $S_{oq}(oq,Tq_{h})=\Phi q$.
-When the processable condition $\text{Processable}^{2}(\Phi u,\Phi q)$ is satisfied, the ontological state of the output information $Ir$ is
-$S_{or}(or,Tr_{h})=\Phi r=f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.16)
-demonstrating that the ontological state $S_{ou}(ou,Tu_{h})$ possesses the ability to process $S_{oq}(oq,Tq_{h})$ and output $S_{or}(or,Tr_{h})$.
-Similarly, since $Iu$, $Iq$ and $Ir$ are all recoverable information, when their ontological states are mapped onto $Mn$ and its interface, the carrier states $S_{Mn}(Mn,Tu_{h})$ also possess the ability to process $S_{cq}(cq,Tq_{h})$ and output $S_{cr}(cr,Tr_{h})$.
-Thus, $Mn$ is capable of realizing the processing of $Iu$ and $Iq$ to produce $Ir$.
-Furthermore, according to section 3.2, $Mn$ is also capable of learning from $Iq$ and optimizing into $Iv$.
-Naturally, one can just as easily use the formal framework of Definitions 8 and 10 to demonstrate that other typical models—such as linear regression and decision trees—also possess both information processing and learning optimization capability.
-It is thus evident that this axiomatic and formalized architecture transcends specific algorithms or models, providing a more universal, unified, and interpretable foundational meta-framework for machine learning theory.
+User input information is formally given by:
+$I_{q}=\langle oq,T_{qh},S_{oq},M_{I},t_{2},S_{M_{I}}\rangle$ (3.16)
+$S_{oq}(oq,T_{qh})=\{\varphi_{q}|\varphi_{q}\text{ is a WFF about }oq\text{ and }T_{qh}\}$ (3.17)
+$S_{M_{I}}(M_{I},t_{2})=\{\psi_{q}|\psi_{q}\text{ is a WFF about }M_{I}\text{ and }t_{2}\}$ (3.18)
+Here, the ontology $oq$ represents the user’s intention or device, with $T_{qh}$ denoting the time at which the ontological state is formed.
+The state $S_{oq}(oq,T_{qh})$ is the formal representation of the user’s needs.
+The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,T_{qh})$ onto the model’s interface device $M_{I}$.
+3.5 Information Processing, Output, and Online Learning of the Model
+When the model M processes input and produces output, the information it carries can be represented as:
+$I_{v}=\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\rangle$ (3.19)
+Where:
+$S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3})=\{\varphi_{v_{1}},\varphi_{v_{2}},\varphi_{v_{3}}\}$ (3.20)
+$\varphi_{v_{1}}=State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.21)
+indicates that $\Phi^{2}(M,t_{2})$ has been assigned to M at time $t_{2}$.
+$\varphi_{v_{2}}=$ $Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land$
+$Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.22)
+indicates that when the input state $S_{M_{I}}(M_{I},t_{2})$ is received at $t_{2}$, M produces the output state $S_{M_{O}}(M_{O},t_{3})$ at $t_{3}$.
+$\varphi_{v_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
+$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
+$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.23)
+expresses that if the input information can be learned by M, then M can achieve online learning while processing and outputting information.
+The physical realization of this ontological state through an enabling mapping is:
+$S_{M}(M,t_{3})=\{\psi_{v_{1}},\psi_{v_{2}},\psi_{v_{3}}\}$ (3.24)
+The output information is given by:
+$I_{r}=\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\rangle$ (3.25)
+where the ontological state is:
+$S_{M\cup M_{I}}(M\cup M_{I},t_{2})=\{\varphi_{r_{1}}\}$ (3.26)
+with
+$\varphi_{r_{1}}=Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (3.27)
+The formal expression of the physical state of the model’s output device $M_{O}$ at time $t_{3}$ is:
+$S_{M_{O}}(M_{O},t_{3})=\{\psi_{r_{1}}\}$ (3.28)
+The model can repeat the processes described in Sections 3.2 to 3.5 , iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
+3.6 Multi-Stage Information Causal Chains
+From a systems engineering perspective, the realization of information in machine learning involves more than just five information mappings.
+In practice, stages such as requirements analysis, design and development, and system integration are also necessary.
+Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 2.5, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation.
+In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Figure.1).
+Figure 1: Machine Learning Causality Chains with Multiple Information Mappings
+As shown in Figure 1, each of these four stages encompasses four layers: conceptual ideas, mind maps, program code, and physical models.
+Typically, content within the same layer shares the same semantic logic and can thus be uniformly represented by a formal system.
+Furthermore, due to the transmissive nature of information, the innermost physical model can also be regarded as the result of enabling mappings from outer conceptual ideas or mind maps.
+This illustrates that the entire system comprises multiple stages, with each stage involving various combinations of information causal chains.
+This framework, based on the realization of information and finite automaton states, provides a formalized representation and a comprehensive, causally progressive structure for each phase in the machine learning process.
+It transcends specific algorithms or models, offering a unified and universal framework, MLT-MF.
+In the appendix, we also present an example of its application to feedforward neural networks, which validates the correctness and applicability of the framework and supports the verification of the related theorems.
 4 Several Important Theorems and Proofs
-The Meta-Framework for Machine Learning Theory (MLT-MF) applies to a series of pivotal issues in the field of artificial intelligence, providing many unique and practically valuable theoretical principles and methodological guidelines.
+The MLT-MF framework is applicable to a range of key issues in the field of artificial intelligence, offering many unique and practical theoretical principles and methodological guidelines.
 4.1 Theorem of Model Interpretability
 Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence.
-However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability [6, 30, 31, 32].
-In fact, the interpretability of any phenomenon in the universe is inherently relative.
-For example, Kepler’s three laws explain the motion of planets around the sun, but such explanations are meaningless to those without elementary mathematical knowledge.
-Even for the same individual, the explanations that can be understood and accepted may evolve as their knowledge expands.
-Therefore, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system.
+However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability Lipton2018 ; Doshi2017 ; Rubin2019 ; Miller2019 .
+In fact, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system.
 Furthermore, it should also be considered in relation to the specific inputs and outputs of the model.
 Based on MLT-MF, we define the following concepts:
-Definition 12 (Model interpretability): Let the state of a model M over time interval $Tm$ be denoted as $S_{M}(M,Tm)$.
-Given an input state $S_{cq}(cq,Tq_{h})$ and an output state $S_{cr}(cr,Tr_{h})$, suppose there exists a formal system $\mathcal{L}$ and three corresponding interpretations $E_{M}$, $E_{q}$, and $E_{r}$ over finite domains $D_{M}$, $D_{q}$, and $D_{r}$, respectively, such that:
-  • $E_{M}$ has the capacity to process $E_{q}$ and produce $E_{r}$ as output, and
-  • $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are each bijective enabling mappings of $E_{M}$, $E_{q}$, and $E_{r}$, respectively.
-Then, we say that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
-In this way, we define model interpretability via the formal system $\mathcal{L}$ and its interpretation.
-This definition aligns both with the intuitive understanding of "interpretation" and with the theoretical need to characterize the internal logic of models through a series of propositions, making it well-suited for the formal analysis and in-depth study of AI interpretability.
-In fact, according to Theorem 1, if a model M has the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then by Definition 10, the states $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are interpretations $E_{M}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5.
-In this case, the mappings $E_{M}\rightrightarrows S_{M}(M,Tm)$, $E_{q}\rightrightarrows S_{cq}(cq,Tq_{h})$, and $E_{r}\rightrightarrows S_{cr}(cr,Tr_{h})$ are self-mappings, and therefore naturally satisfy the condition of being bijective enabling mappings.
-This leads to:
-Corollary 1 (Intrinsic Model Interpretability): If a model M possesses the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then there exists a formal system $\mathcal{L}_{M}$ in which M is interpretable with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
-Since $\mathcal{L}_{M}$ serves as the formal system that interprets model M, it is also referred to as the intrinsic formal system of the model.
-As Hegel famously asserted, "What is real is rational." The intrinsic interpretability of a model means that, insofar as it can process inputs and generate outputs according to established rules and procedures, it necessarily possesses an internal logic that is, in principle, interpretable.
-However, this logic may be highly implicit and abstract, making it difficult for people—even domain experts—to identify and understand.
-This highlights the need to map and translate this internal logic into a more intuitive and universal logical system, thereby enabling deeper decomposition and analysis of the causal mechanisms, and achieving broader understanding and acceptance.
-This is the essence of model interpretability, and is precisely the core idea behind Definition 12: to interpret the internal logic of models within more explicit formal systems.
-Theorem 3 (Model interpretability and information recoverability): Let model M be capable of processing the information $Iq=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ and producing output information $Ir=\langle or,Tr_{h},S_{or},cr,Tr_{m},S_{cr}\rangle$, through its operation described by $Iu=\langle ou,Tu_{h},S_{ou},M,T_{m},S_{M}\rangle$.
-If all three pieces of information $Iu$, $Iq$, and $Ir$ are recoverable, and their ontological states are interpretations of well-formed formulas in a common formal system $\mathcal{L}$, then model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
-Proof: According to Definition 5, since $Iu$, $Iq$, and $Ir$ are recoverable, the following are all bijective enabling mappings: $S_{ou}(ou,T_{h})\rightrightarrows S_{M}(M,T_{m})$, $S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$, and $S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$.
-By Definition 11, the state $S_{ou}(ou,T_{h})$ has the capacity to process $S_{oq}(oq,Tq_{h})$ and produce $S_{or}(or,Tr_{h})$ as output.
-Then, by Definition 10, the states $S_{ou}(ou,T_{h})$, $S_{oq}(oq,Tq_{h})$, and $S_{or}(or,Tr_{h})$ are interpretations $E_{u}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5.
-Moreover, $E_{u}$ has the ability to process $E_{q}$ and produce $E_{r}$ as output.
-Thus, we have the following bijective enabling mappings:
-$E_{u}=S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,T_{m})$ (4.1)
-$E_{q}$ $=S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$ (4.2)
-$E_{r}$ $=S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$ (4.3)
-By Definition 12, it follows that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{m})$ under the input condition $S_{cq}(cq,Tq_{m})$. Q.E.D.
-Theorem 3 fully embodies the Chinese proverb: "Whoever tied the bell must untie it." Since, intuitively, the internal logic $S_{M}(M,T_{m})$ of model M originates from the ontological state $S_{ou}(ou,Tu_{h})$ that brings the model into existence, it is thus natural to interpret the model’s output under specific input conditions using $E_{u}$, $E_{q}$, and $E_{r}$.
-Moreover, unlike the internal processing of $S_{M}(M,T_{m})$, which is encapsulated within the model and often opaque, the interpretations $E_{u}$, $E_{q}$, and $E_{r}$ are explicit expressions of well-formed formulas within the formal system $\mathcal{L}$ under given valuations.
-These expressions are more amenable to interpretation and understanding by humans, particularly by domain experts.
+Definition 9 (Model Interpretability) Let the model M, as a finite automaton, operate over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, at which it undergoes assignment, input, output, and state transitions.
+The state of M at time $t_{i}$ is denoted as $S_{M}(M,t_{i})$, and the state of its input device $M_{I}$ at time $t_{i}$ is $S_{M_{I}}(M_{I},t_{i})$, while the state of the output device $M_{O}$ at time $t_{i+1}$ is $S_{M_{O}}(M_{O},t_{i+1})$.
+If there exists a formal system $\mathcal{L}$ along with three interpretation relations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$ respectively, such that
+$E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$ (4.1)
+making $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ the respective enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, then M is said to be interpretable in $\mathcal{L}$ with respect to output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$.
+If for all $i=1,\cdots,n-1$, the output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in $\mathcal{L}$, then M is called interpretable in $\mathcal{L}$.
+Here, $\text{Cn}(E_{M}\cup E_{M_{I}})$ denotes the logical closure of $E_{M}\cup E_{M_{I}}$.
+Equation 4.1 states that in the formal system $\mathcal{L}$, all formulas in $E_{M_{O}}$ are logical consequences of $E_{M}$ and $E_{M_{I}}$.
+Moreover, since $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are respectively enabled by $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, the formal system $\mathcal{L}$ and its three interpretations fully capture the logical and causal relationship whereby M produces output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$.
+This, in essence, offers a concise understanding of model interpretability.
+Theorem 4 (Model Interpretability and Information Existence) Let the model M be viewed as a finite automaton operating over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, during which it undergoes assignment, input, output, and state transitions.
+(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}(M_{I},t_{i})$ is that there exist $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$ and $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$ representing the information carried by M and its input device $M_{I}$ at time $t_{i}$ respectively, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$.
+The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$ must hold.
+(2) A necessary and sufficient condition for M to be interpretable in $\mathcal{L}$ is that, for all $i=1,\cdots,n-1$, there exist information objects $Iu(t_{i})$, $Iq(t_{i})$, and $Ir(t_{i+1})$ whose ontological states are all interpretations in $\mathcal{L}$, and such that the ontological state of $Ir(t_{i+1})$ is contained in the logical closure of the union of the ontological states of $Iu(t_{i})$ and $Iq(t_{i})$.
+Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$.
+According to Definition 9, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$, and $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
+Let $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$.
+Then, according to Definition 6,we find there exist information objects $Iu(t_{i})=$ $\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, similarly , then $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),$ $Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$.
+Thus, the necessity of the first part of the theorem is established.
+On the other hand, for any $1\leq i\leq n-1$, suppose there exist some information objects $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that the ontological states $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$.
+Then, according to Theorem 1, $E_{M}=S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times Tu_{h}(t_{i})$, $oq(t_{i})\times Tq_{h}(t_{i})$, and $or(t_{i+1})\times Tr_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$.
+At the same time, $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
+By Definition 9, the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$.
+Therefore, the sufficiency of the first part of the theorem is established.
+For all $i=1,\ldots,n-1$ , by combining Definition 9 with the first part of the theorem, the second part of the theorem follows directly. This completes the proof.
 4.2 Theorem of Ethical Safety Assurance
-Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence [29, 30, 31, 32], and to date, there is still no universally accepted and practically actionable rigorous definition [33, 34, 35].
+Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence mathsea2002 ; Doshi2017 ; Rubin2019 ; Miller2019 , and to date, there is still no universally accepted and practically actionable rigorous definition Russell2015 ; Jobin2019 ; Corrêa2023 .
 Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions.
@@
 By introducing an ethical safeguard formula for $Ec$ into $Iu$, it is thereby possible to ensure that model M attains the maximal output state that is ethically safe with respect to $Ec$.
-Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, $Iq$ and $Ir$ respectively, the MLT-MF framework provides:
+Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, IqIq and $Ir$ respectively, the MLT-MF framework provides:
 $\varphi u=($ $\forall\Phi u)(\forall\Phi q)(\text{IsSubsetOf}^{2}(\Phi u,Sou(ou,Tu_{h}))\land\text{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,Tq_{h}))\land$
@@
 Hence, $G_{max}$ is the largest ethically safe subset of $S_{or}(or,Tr_{h})$ with respect to $Ec$.
-Let $S_{or}^{Ec}(or,Tr_{h})=G_{max}$ and $S_{or}^{\sim Ec}(or,Tr_{h})=S_{or}(or,Tr_{h})\backslash S_{or}^{Ec}(or,Tr_{h})$.
-In addition, introduce a new constant in $\mathcal{L}$: NEc, representing a non-Ec ethical safety prompt, and a new predicate: IsSafeback1(NEc), indicating that M outputs the non-Ec ethical safety prompt NEc.
-We may safely assume that $\mathcal{L}$ originally contains neither.
+[…]
 In this way, add to $Sou(ou,Tu_{h})$ an ethical safeguard formula:
@@
 Improving model generalization ability is among the most critical objectives in machine learning.
-However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance [40, 41, 42, 43, 44].
-It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples.
-By applying the MLT-MF, it is possible to estimate an upper bound on the model’s generalization error.
-Theorem 5 (Upper Bound for Model Generalization Error): Let a model M, after learning and optimization, carry the information $Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$, and let the input information be $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$.
-Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively, and let $p_{q}(x)$ be the probability distribution over $S_{oq}$.
-The total probability of $p_{q}(x)$ on $S_{ou}\cap S_{oq}$ is denoted as $P_{Mq}$.
-If $S_{ou}$ and $S_{oq}$ are governed by probability distributions defined on the same measurable space, and M is sufficiently trained on $S_{oq}$ and predicts unknown states using a conservative uniform distribution, then the upper bound of the generalization error when processing $Iq$ is determined by the overlap between $S_{oq}$ and $S_{ou}$, as well as the entropy of $p_{q}(x)$ on the non-overlapping region, specifically:
-$\mathcal{E}_{gen}\begin{cases}\to 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.13)
-where $\mathcal{E}_{gen}$ denotes the upper bound of generalization error, and
-$H(p_{q})=-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)$
-is the entropy of the data in the non-overlapping region.
-Proof: Let $p_{M}(x)$ be the probability distribution determined by the model structure and parameter set of M.
-$Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ is the set of information carried by M, so $p_{M}(x)>0$ for $x\in S_{ou}$; $p_{q}(x)>0$ for $x\in S_{oq}$.
-Since M is sufficiently trained on $S_{oq}$ and adopts a conservative uniform distribution for unknown states, $p_{M}(x)$ in the overlapping region $S_{ou}\cap S_{oq}$ approximates the true distribution $p_{q}(x)$, and in the non-overlapping region $S_{oq}\backslash S_{ou}$, it follows a uniform distribution.
-That is,
-$p_{M}(x)\begin{cases}\to p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.14)
-Using KL divergence to measure the discrepancy between the true and model distributions [45]:
+However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance Vapnik1999 ; Bartlett2002 ; Zhang2016 ; Advani2020 ; Perlaza2024 .
+It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples.Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
+Theorem 6 (Upper Bound on TVD) Let the information describing the model M after learning and optimization be $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and the input information be $I_{q}=\langle oq,Tq_{h},S_{oq},eq,Tq_{m},S_{cq}\rangle$.
+Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space.
+Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively.
+The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$, the distribution over $S_{oq}$ is $p_{q}(x)$,with $p_{q}(x)>0$ for $x\in S_{oq}$ .
+The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$.
+When the model M is sufficiently trained on $S_{ou}$ and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
+$TVD(p_{q},p_{M})\begin{cases}\approx 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.11)
+For the second case, it is assumed that $S_{oq}\setminus S_{ou}\neq\emptyset$, $1-P_{Mq}>0$, and $|S_{oq}|-|S_{ou}\cap S_{oq}|>0$.
+Here, $H(p_{q})=-\sum_{x\in S_{oq}\setminus S_{ou}}p_{q}(x)\log p_{q}(x)$ is the entropy of the data in the non-overlapping region.
+[…]
+$p_{M}(x)\begin{cases}\approx p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.12)
+Using KL divergence to measure the discrepancy between the true distribution and the hypothesized model distribution, we have Goodfellow2016 :
 $D_{KL}(p_{q}(x)\|p_{M}(x))=$ $\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (4.13)
@@
 $\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\to\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}=0$ (4.14)
-In the non-overlapping region, since the true distribution $p_{q}(x)$ is arbitrary, again by (4.14):
+This establishes the first case in (4.11).
+In the non-overlapping region, the true distribution $p_{q}(x)$ can be arbitrary. Similarly, by (4.12), we have
 $\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ $=\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{M}(x)$
@@
 $=-H(p_{q})-\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$ (1)
-$=-H(p_{q})-(1-P_{Mq})\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
 $=(1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q})$ (4.15)
-Because $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can estimate the generalization error of model M using the total variation distance (TVD) and KL divergence.
-According to Pinsker’s inequality [46]:
-$\mathcal{E}_{gen}\leq\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
+Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance.
+By Pinsker’s inequality Pinsker1964 , we have:
+$\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
 $=\sqrt{\frac{1}{2}(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)})}$
 $\approx\sqrt{\frac{1}{2}((1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q}))}$ (4.16)
-Thus, the latter part of (4.13) is proved.
-Since we have assumed that $Iu=(ou,Tu_{h},S_{ou},M,$ $Tu_{m},S_{M})$ and $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$ are both recoverable information, and the states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ are the physical realizations of $S_{ou}$ and $S_{oq}$ on the model M and its interface, equation (4.13) indeed characterizes the upper bound of model M’s generalization error.
-Q.E.D.
-Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 5—based on MLT-MF—explicitly quantifies the generalization error upper bound using $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, which reflect the overlap between training data and model information, and the entropy $H(p_{q})$ of the non-overlapping region, thereby offering a new perspective and practical guidance for data collection and model design.
-In particular, Theorem 5 reveals that as long as the problem space is contained within the model’s information space, the model’s generalization error can be made arbitrarily small in the statistical sense.
-This leads to a critical principle in machine learning:
-Corollary 2 (Generalization Error of Ideal Models): If the model information contains the problem information, then the generalization error for these problems is almost zero.
+This establishes the second case in (4.11).
+Since we have already assumed that the carrier states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ of $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and $I_{q}=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model M and its interface, respectively.
+It follows that equation (4.11) characterizes the upper bound of the model’s TVD.
+This completes the proof.
+Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 6, based on the MLT-MF framework, provides an explicit quantitative estimate for the model’s TVD.
+This estimate is grounded in the overlap between the test and training data $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, as well as the entropy $H(p_{q})$ of the test data distribution over the non-overlapping region.
+These metrics serve as indicators of how much the testing data shares information with the model’s prior knowledge and how unpredictable the unseen portion is.
+Because TVD characterizes the alignment between distributions at a fundamental level—independent of any particular model architecture or task—it is generally applicable to the analysis of model training performance.
+In fact, for classification tasks with 0-1 loss, the generalization error upper bound is equivalent to the TVD Rodríguez2021 .
+For regression tasks, the upper bound on generalization error is given by the product of TVD and the Lipschitz constant of the loss function Devroye2013 .
+Thus, TVD is a meaningful and informative measure for estimating model generalization error.
+This shows that Theorem 6 offers a new perspective and practical guidance for data collection, model design, and training for construction.
+In particular, Theorem 6 reveals that as long as the problem space is contained within the model’s information space, the model’s TVD can be made arbitrarily close to zero in a statistical sense.
+This is a fundamental result in machine learning:
+Corollary 2 (TVD for Ideal Models) If the model’s information encompasses all relevant information for the problem, then the TVD for these problems is approximately zero.
+5 Experimental Results
+Using the feedforward neural network model listed in the appendix, we conducted regression experiments on two bivariate root functions: $y_{1}=\sqrt{4x_{1}^{2}+3x_{2}^{2}}$ and $y_{2}=\sqrt{2x_{1}^{2}+x_{2}^{2}}$.
+Both the training and test datasets were generated randomly.
+We evaluated the models’ prediction errors under various conditions, assessed the impact and computational cost of ethical and safety mechanisms, and explored the challenge of providing multiple types of explanations for the same output.
+These experiments serve to illustrate and validate the definitions and theorems presented in Section 4.
+5.1 Prediction Error Experiments
+To evaluate the model’s predictive performance in both overlapping and non-overlapping cases between the training and test datasets, we used the following data setup: 10,000 training samples were generated with $x_{1}$ and $x_{2}$ independently and uniformly distributed over $[0.5,1.5]$, and 10,000 test samples were drawnuniformly from within a circle of radius 1.0 centered at (1.5, 1.0).
+The model’s learning rate was set to $\eta=0.01$.
+Figure 2 illustrates the experimental conditions and outcomes:
+Figure 2: Model Prediction Error Experiments and Results Under Different Conditions.
+(a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples.
+Among these, 4,951 samples overlap with the training set.
+(b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset.
+(c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound.
+(d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
+(a) The overlap rate between the test and training datasets is approximately 50%.
+(b) For 10 randomly selected samples from the overlapping region, as the number of training steps increases from 10 to 4000 (i.e., as the model becomes more sufficiently trained), the root mean square error (RMSE) of the model’s predictions steadily approaches zero.
+This result aligns with the first case of Theorem 6, equation (4.11).
+(c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used.
+For almost all samples (except a few outliers), their RMSE is less than $\text{TVD}=\sqrt{((1-0.4951)\ln((10000-4951)/(1-0.4951))-2.9741)/2}\approx 0.9155$.
+This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
+(d) For 10 randomly selected test samples from the non-overlapping region, when using the model’s prediction method, the RMSE also decreases with more training steps, reflecting a certain degree of model generalization outside the training data region.
+However, the rate of error reduction is significantly lower than in the overlapping region, and for some samples, the RMSE remains higher than that of the conservative uniform prediction.
+In fact, when training is insufficient, the overall RMSE of the model’s predictions can even exceed the TVD.
+Meanwhile, the RMSE of conservative predictions for these samples is much lower than the TVD, reflecting the conclusion of the second case in Theorem 6, equation (4.11).
+In summary, this regression experiment using a simple feedforward neural network model demonstrates and validates the accuracy of the model TVD upper bound estimation based on the MLT-MF framework, and its potential for estimating generalization error.
+Since the theoretical proof of Theorem 6 is independent of specific models and tasks, it promises to be extensible to more complex scenarios.
+5.2 Ethical Safety Assurance Experiment
+Given the nature of the bivariate root function regression task, the model’s predicted outputs should never be negative.
+This requirement is analogous to Definition 10, which ensures that model outputs do not violate the ethical constraints .
+Therefore, using the same training dataset as in the previous prediction error experiments, we set the learning rate to and trained the model for 165 steps.
+For testing, we generated a dataset of 10,000 samples with drawn uniformly from within a circle of radius 0.5 centered at (-1, -1).
+The criterion for judging ethical safety violations was whether the model’s predicted output was negative.
+This setup was used to test the feasibility of Theorem 5 and to evaluate the additional computational overhead introduced by ethical safety checking.
+For each trial, 5,000 test samples were randomly selected using the trial number as a random seed.
+Two types of tests were conducted in parallel: “Without Ethical Safety Check” and “With Ethical Safety Check.” In the “Without Ethical Safety Check” condition, model predictions were output directly and the total computation time was recorded.
+In the “With Ethical Safety Check” condition, each prediction was first checked for sign; if any prediction was negative, an ethical error message was output, otherwise, the model’s prediction was output.
+The total computation time was also recorded in this scenario.
+In total, 10 trials were conducted, and the results are shown in Table 1.
+Table 1: Summary Table of Ethical Safety Experiment Results
+Trial Computation Computation Computation Number of Number of Correct Correct
+Number Time Without Time With Time Ratio Ethically Unethical Prediction Prompt
+Ethical Ethical Safe Warning Rate for Rate for
+Safety Check Safety Check Samples Samples Safe Unethical
+(seconds) (seconds) Outputs Outputs
+1 0.367 0.391 1.063 4517 483 100% 100%
+2 0.365 0.378 1.036 4521 479 100% 100%
+3 0.352 0.362 1.030 4548 452 100% 100%
+4 0.330 0.378 1.146 4516 484 100% 100%
+5 0.371 0.383 1.032 4523 477 100% 100%
+6 0.327 0.380 1.160 4546 454 100% 100%
+7 0.358 0.402 1.123 4545 455 100% 100%
+8 0.360 0.410 1.140 4538 462 100% 100%
+9 0.361 0.365 1.010 4519 481 100% 100%
+10 0.355 0.424 1.192 4533 467 100% 100%
+In Table 1, the “Computation Time Ratio” is calculated as the computation time with ethical safety checking divided by the computation time without it.
+For the simple feedforward neural network model used in these experiments, the additional computational overhead for performing ethical safety checks before model output ranged from 1.0% to 19.2%.
+As model size increases, this overhead is expected to decrease in proportion; however, as the number of ethical safety constraints grows, the overhead will increase accordingly.
+This is a key consideration for implementing ethical safety in models.
+In these 10 trials, more than 19% of samples resulted in unethical predictions.
+In every trial, both the “Correct Prediction Rate for Safe Outputs” and the “Correct Prompt Rate for Unethical Outputs” were 100%, demonstrating that by incorporating safety checks and informative outputs, we can always guarantee the ethical safety of model predictions while maximizing the number of safe outputs.
+In summary, this ethical safety experiment, grounded in Definition 10 and the formal MLT-MF-based framework for ethical safety, illustrates and validates Theorem 5 regarding the feasibility of integrating ethical safety checking mechanisms and achieving the maximal number of safe outputs.
+5.3 Interpretability Experiment
+For our bivariate root regression task, the output should never be negative.
+However, in the ethical safety assurance experiment, after 165 training steps, the model produced negative predictions for more than 19% of randomly selected 5,000-sample test batches, repeated over 10 trials.
+According to Theorem 4, this phenomenon can be explained from multiple perspectives, each based on different information.
+5.3.1 Mathematical Operation-Based Explanation
+After 165 training steps, the neural network model M has the following weight matrices and bias vectors from the input layer to the hidden layer:
+$W_{1}\approx\begin{bmatrix}-0.065&-0.429\\ 0.681&-0.157\\ -0.498&-0.586\end{bmatrix},\quad b_{1}\approx\begin{bmatrix}0.129\\ -0.551\\ 0.350\end{bmatrix}$ (5.1)
+The weights and biases from the hidden layer to the output layer are:
+$W_{2}\approx\begin{bmatrix}0.145&0.519&-0.423\\ 0.022&0.048&0.043\end{bmatrix},\quad b_{2}\approx\begin{bmatrix}0.613\\ 0.199\end{bmatrix}$ (5.2)
+Therefore, for the test sample $X=[-1.3535,-1.3535]^{T}$, the model predicts:
+$Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\text{Identity}(W_{2}\text{ReLU}(W_{1}X+b_{1})+b_{2})\approx\begin{bmatrix}-0.040\\ 0.277\end{bmatrix}$ (5.3)
+Here, $y_{1}=-0.040<0$, which mathematically explains why model M outputs a negative value for input $X$.
+These formulas provide the designer’s mathematical representation of the model’s computational process.
+They are equivalent to the formal forward propagation rules (A.21) and (A.22) in the appendix.
+According to the OIT perspective, these equations describe the ontological state of the model, its outputs, and the information they convey.
+The model, along with its input and output mechanisms, serves as the carrier of this information and realizes the computational process.
+Therefore, using these formulas to explain why the model produces negative outputs satisfies the requirements of Definition 9 regarding model interpretability and meets the needs of researchers who seek to understand model behavior from a mathematical standpoint.
+5.3.2 Visualization-Based Explanation of the Computational Process
+Depicting the model’s computational process visually can make explanations of model behavior much more intuitive.
+To this end, for model M after 165 training steps, we illustrate its complete structure—covering the inputs, internal neural network calculations and data flow, and outputs—in Figure 3.
+Figure 3: Schematic diagram of the computational process for model M after 165 training steps, showing its handling of the input sample $X=[-1.3535,-1.3535]^{T}$ and output $Y\approx[-0.040,0.277]^{T}$.
+Compared with equations (5.1)(5.3), Figure 3 not only comprehensively reflects the input, weight matrices, bias vectors, neuron activation functions, and the specific data and formulas for outputs, but also highlights the interactions between the input/output devices and the model.
+The figure offers a more direct view of the internal data processing pipeline, making it easier to examine and verify each step of the computation.
+Thus, the visualization provides a more accessible and concrete explanation of how model M processes inputs and generates results.
+5.3.3 Partition-Based Explanation
+Figure 4: Illustration of Prediction Polarity and Impact of Training Steps.
+(a) Polarity distribution of predicted data around the training set region after model M has been trained for 165 steps.
+The blue square marks the area where the training dataset is located, centered at (1, 1) with a side length of 1.
+Green dots indicate samples with all-positive predictions; dark green dots indicate samples with all-positive predictions that also belong to the test set; red dots indicate samples with at least one negative prediction that belong to the test set; orange dots indicate samples with at least one negative prediction that do not belong to the test set.
+(b) Cumulative effect of up to 400 training steps on the prediction polarity at two points: (1, 1) and (-1.3535, -1.3535).
+The blue and green lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (1, 1) from the training set; the red and pink lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (-1.3535, -1.3535) from the test set.
+It is evident that equations (5.1)(5.3) and Figure3 explain only the result of model M processing the specific sample $X=[-1.3535,-1.3535]^{T}$.
+They do not account for the phenomenon observed in the ethical safety assurance experiment, where each random selection of 5,000 samples from the test dataset yields over 19% negative predictions.
+To explain this, a broader statistical analysis across more data samples is needed.
+Figure 4(a) provides a visual depiction of the decision boundary after 165 training steps, showing where model M predicts negative outputs.
+Notably, there is a boundary line approximately passing through points (-2.0, 0.4) and (-0.4, -2.0).
+For samples on the upper right side of this diagonal, model M consistently predicts non-negative values; for those on the lower left side, the model consistently outputs negative values.
+The test region in the figure, representing the test dataset, falls precisely on the lower left side of this boundary line, which explains why a random selection of 5,000 samples from the test set will always include a substantial portion that triggers negative predictions.
+This visual and statistical partitioning thus provides an intuitive explanation for the observed frequency of negative outputs in the ethical safety experiment.
+Figure 4(a) presents an overall view of numerous inputs processed and predicted by model M.
+This visualization also represents an ontological state of M as it handles a wider range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping.
+Thus, according to Definition 9, Figure 4(a) constitutes a valid explanation for the distribution of prediction polarities by M.
+5.3.4 Explanation Based on Cumulative Training Effects
+The previous three explanations all focus on the behavior of M after 165 training steps.
+However, model training is a continuous process of cumulative improvement.
+Examining only the model’s instantaneous performance at a particular training step cannot fully explain the causes of abnormal outputs.
+To address this, we further trained M for 400 iterations on the training dataset.
+Figure 4(b) shows the predictions for both $(1,1)$ and (1.3535,1.3535)(-1.3535,-1.3535) after each training step.
+It is clear that even for $(1,1)$ — a central sample from the training set — the predicted $y_{2}$ by M remains negative during roughly the first 50 training steps.
+Only after more than 50 steps does the prediction move away from negative values.
+For the test sample (1.3535,1.3535)(-1.3535,-1.3535), it takes more than about 180 training steps before the prediction begins to consistently avoid negative values.
+This cumulative effect is entirely due to the optimization of M’s parameters as training progresses.
+In fact, by applying the forward propagation rules (A.21) and (A.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.
 6 Conclusion and Outlook
@@
   • Ethical Safety: By embedding ethical constraint predicates, the compliance of model outputs can be formally verified, offering criteria and methodological guidance for the design of trustworthy AI systems;
-  • Generalization Error Quantification: Based on state coverage and KL divergence, an upper bound formula for generalization error is derived, revealing the core principle that the "problem space must be contained within the model knowledge space," thus providing explicit tools for quantitative analysis in data collection and model optimization.
+  • Model Performance: Based on state coverage and KL divergence, an upper bound formula for TVD is derived, revealing the core principle that the “problem space must be contained within the model knowledge space,” thus providing explicit tools for quantitative analysis in data collection and model optimization.
+We also validated and demonstrated the correctness and practicality of the relevant theorems based on MLT-MF through experiments with a simple feedforward neural network model.
 Despite these important advances, the current framework assumes ideal (noiseless) enabling mappings and primarily addresses model learning and processing logic in static scenarios.
@@
 I hereby express my sincere gratitude to them.
+Junxuan He from Beijing University of Posts and Telecommunications implemented all the experiments described in the paper according to the author’s ideas, producing complete model code and training/test datasets, providing indispensable and strong support for illustrating and validating the theorems related to MLT-MF.
+The author offers special thanks for his hard work, keen understanding, and efficient implementation.
+8 Data and Code availability
+All datasets and codes used in this work are available at https://github.com/xuhu6736/MLT-MF.
+Appendix A Appendix: Model Instance Validation of MLT-MF
+To facilitate validation and illustration of the practical significance of MLT-MF and related theorems, we select a very simple model and use MLT-MF to establish formal representations with WFFs for each stage from initial information to training information, learning optimization, and processing/output.
+A.1 A Simple Feedforward Neural Network Model
+We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Figure A1).
+Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
+Figure A1: Schematic diagram of feedforward neural network model M
+The hidden layer uses the ReLU activation function:
+$\text{ReLU}(x)=\max(0,x)$ (A.1)
+The output layer uses the identity activation:
+$\text{Identity}(x)=x$ (A.2)
+Initial weight matrices and bias vectors are defined between layers:
+$W_{1}=\begin{bmatrix}w_{11}^{1}&w_{12}^{1}\\ w_{21}^{1}&w_{22}^{1}\\ w_{31}^{1}&w_{32}^{1}\end{bmatrix},\quad b_{1}=\begin{bmatrix}b_{1}^{1}\\ b_{2}^{1}\\ b_{3}^{1}\end{bmatrix}$ (A.3)
+$W_{2}=\begin{bmatrix}w_{11}^{2}&w_{12}^{2}&w_{13}^{2}\\ w_{21}^{2}&w_{22}^{2}&w_{23}^{2}\end{bmatrix},\quad b_{2}=\begin{bmatrix}b_{1}^{2}\\ b_{2}^{2}\end{bmatrix}$ (A.4)
+Training or input data, output data, and training labels or target outputs are specified, with mean squared error (MSE) as the loss:
+$X=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\quad Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},\quad Y_{t}=\begin{bmatrix}y_{t1}\\ y_{t2}\end{bmatrix}\quad and\quad Y_{t}=\text{Target}(X)$ (A.5)
+$L=\text{MSE}(Y,Y_{t})=\frac{1}{2}\|Y-Y_{t}\|^{2}=\frac{1}{2}((y_{1}-y_{t1})^{2}+(y_{2}-y_{t2})^{2})$ (A.6)
+The gradients of the loss function $L$ with respect to the weight matrices and bias vectors are the primary basis for model M’s optimization during training.
+The four calculation formulas are as follows:
+$\nabla_{W_{2}}L=(Y-Y_{t})(\text{ReLU}(W_{1}X+b_{1}))^{T}$ (A.7)
+$\nabla_{b_{2}}L=(Y-Y_{t})$ (A.8)
+$\nabla_{W_{1}}L=(W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1}))X^{T}$ (A.9)
+$\nabla_{b_{1}}L=W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1})$ (A.10)
+In equations (A.9) and A.10, the operator $\odot$ denotes element-wise multiplication, and ReLU’ represents the derivative of the ReLU function.
+Since ReLU is not differentiable at zero, we additionally define ReLU’(0) = 0.
+Thus, the update formulas for optimizing the weight matrices and bias vectors during M’s training are:
+$W_{1}^{\text{new}}$ $=W_{1}-\eta\nabla_{W_{1}}L,\quad b_{1}^{\text{new}}=b_{1}-\eta\nabla_{b_{1}}L$ (A.11)
+$W_{2}^{\text{new}}$ $=W_{2}-\eta\nabla_{W_{2}}L,\quad b_{2}^{\text{new}}=b_{2}-\eta\nabla_{b_{2}}L$ (A.12)
+where $\eta$ is the learning rate of M.
+A.2 Symbol Definitions
+MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle.
+Defining the symbols is a prerequisite for formal modeling.
+To this end, Tables A1A4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
+Table A1: Definition of Constants
+Symbol Type Definition
+$\mathbb{R}^{2}$ Space 2-dimensional real space
+$\mathbb{R}^{3}$ Space 3-dimensional real space
+Table A2: Definition of Variables
+Symbol Type Definition
+$i,j,k$ natural number index or count
+$\varphi_{o_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s initial model
+$\psi_{o_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s initial model
+$\varphi_{t_{i}}$ WFFs well-formed formulas contained in the training informational ontological state
+$\psi_{t_{i}}$ WFFs well-formed formulas contained in the training informational carrier state
+$\varphi_{u_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s optimized model
+$\psi_{u_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s optimized model
+$\varphi_{q_{i}}$ WFFs well-formed formulas contained in the input informational ontological state of M
+$\psi_{q_{i}}$ WFFs well-formed formulas contained in the input informational carrier state of M
+$\varphi_{r_{i}}$ WFFs well-formed formulas contained in the output informational ontological state of M
+$\psi_{r_{i}}$ WFFs well-formed formulas contained in the output informational carrier state of M
+$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s re-optimized model
+$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s reoptimized model
+layerslayers natural number number of layers in the model
+$v$ real number numerical values for weights and biases
+$W_{k}^{(o)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the initial model
+$w_{ij}^{(o)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the initial model
+$b_{k}^{(o)}$ vector bias vector of neurons in the k-th layer in the initial model
+$b_{i}^{(o)k}$ real number bias of the i-th neuron in the k-th layer in the initial model
+$W_{k}^{(u)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the optimized model
+$w_{ij}^{(u)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the optimized model
+$b_{k}^{(u)}$ vector bias vector of neurons in the k-th layer in the optimized model
+$b_{i}^{(u)k}$ real number bias of the i-th neuron in the k-th layer in the optimized model
+$W_{k}^{(v)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the re-optimized model
+$w_{ij}^{(v)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the re-optimized model
+$b_{k}^{(v)}$ vector bias vector of neurons in the k-th layer in the re-optimized model
+$b_{i}^{(v)k}$ real number bias of the i-th neuron in the k-th layer in the re-optimized model
+$W_{k}^{(w)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the further re-optimized model
+$w_{ij}^{(w)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the further re-optimized model
+$b_{k}^{(w)}$ vector bias vector of neurons in the k-th layer in the further re-optimized model
+$b_{i}^{(w)k}$ real number bias of the i-th neuron in the k-th layer in the further re-optimized model
+$X$ vector input vector
+$H$ vector hidden layer vector
+$Y$ vector output vector
+$Y_{t}$ vector output reference vector
+$x_{i}$ real number the i-th element of the input vector
+$h_{i}$ real number the i-th element of the hidden layer vector
+$y_{i}$ real number the i-th element of the output vector
+$y_{ti}$ real number the i-th element of the output reference vector
+$t$ time time point of state occurrence
+$\delta$ time time step
+$L$ real number loss function value
+$\nabla_{W_{1}}L$ real number gradient of the loss function with respect to $W_{1}$
+$\nabla_{b_{1}}L$ real number gradient of the loss function with respect to $b_{1}$
+$\nabla_{W_{2}}L$ real number gradient of the loss function with respect to $W_{2}$
+$\nabla_{b_{2}}L$ real number gradient of the loss function with respect to $b_{2}$
+$\eta$ Positive real number learning rate
+Table A3: Definition of Functions
+Symbol Number of Elements Definition
+ReLU(x)(x) 1 rectified linear unit (ReLU) activation function
+Identity(x)(x) 1 identity function
+maxmax multivariate maximum function
+[…]
+ReLU’$(x)$ 1 derivative of ReLU(x)(x)
+Target(X)(X) 1 objective function for input vector $X$
+Table A4: Definition of Predicates
+Symbol Definition
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+A.3 Formal Representations of Information at Each Stage of Model M
+Following MLT-MF, we build formal expressions for the following five stages:
+A.3.1 Initial information of M
+According to the principles of OIT and based on MLT-MF, the model M is endowed with its initial information from the very beginning:
+$I_{0}=(oo,To_{h},S_{oo},M,To_{m},S_{M})$ (A.13)
+Here, the ontology $oo$ represents the designer’s conceptual blueprint; $To_{h}$ is the time at which the model is designed; and the ontological state $S_{oo}(oo,To_{h})$ is a set of well-formed formulas that describe the initial operational logic of model M:
+$S_{oo}(oo,To_{h})=\{\varphi o_{1},\varphi o_{2},\varphi o_{3},\varphi o_{4},\varphi o_{5},\varphi o_{6},\varphi o_{7},\varphi o_{8},\varphi o_{9},\varphi o_{10},\varphi o_{11}\}$ (A.14)
+The specific content of each formula is as follows:
+  • Network Structure Definition
+    $\varphi o_{1}=$ $\text{NetworkStructure}^{2}{(layers,3)}\land\text{LayerSize}^{2}(1,2)\land$ (A.15)
+    $\text{LayerSize}^{2}(2,3)\land\text{LayerSize}^{2}(3,2)$
+  • Weight Initialization
+    $\varphi o_{2}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2,3\})\land BelongsTo^{2}(j,\{1,2\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)1},v)))$ (A.16)
+    $\varphi o_{3}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2\})\land BelongsTo^{2}(j,\{1,2,3\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)2},v)))$ (A.17)
+  • Bias Initialization
+    $\varphi o_{4}=\forall iBelongsTo^{2}(i,\{1,2,3\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)1},v))$ (A.18)
+    $\varphi o_{5}=\forall iBelongsTo^{2}(i,\{1,2\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)2},v))$ (A.19)
+  • Activation Function Definition
+    $\varphi o_{6}=\forall x(Eq^{2}(\text{ReLU}(x),max(0,x)))\land\forall x(Eq^{2}(\text{Identity}(x),x))$ (A.20)
+  • Forward Propagation Rules
+    $\varphi o_{7}=$ $\forall X\forall t\forall H(InputAt^{2}(X,t)\land BelongsTo^{2}(X,\mathbb{R}^{2})\rightarrow$
+    $\exists H(BelongsTo^{2}(H,\mathbb{R}^{3})\land HiddenAt^{2}(H,t+\delta)\land$
+    $\forall i(BelongsTo^{2}(i,\{1,2,3\})\rightarrow Eq^{2}(h_{i},\text{ReLU}\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right))))$ (A.21)
+    $\varphi o_{8}=$ $\forall H\forall t(HiddenAt^{2}(H,t+\delta)\land BelongsTo^{2}(H,\mathbb{R}^{3})\rightarrow$
+    $\exists Y(BelongsTo^{2}(Y,\mathbb{R}^{2})\land(OutputAt^{2}(Y,t+2\delta)\land$
+    $\forall i(BelongsTo^{2}(i,\{1,2\})\rightarrow Eq^{2}(y_{i},\text{Identity}\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)))))$ (A.22)
+  • Backpropagation Rules
+    $\varphi o_{9}$ $=\forall X\forall Y_{t}\forall t(InputAt^{3}(X,Y_{t},t)\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2})\land\varphi o_{7}\land\varphi o_{8}\land$
+    $Eq^{2}(L,\text{MSE}(Y,Y_{t}))\land Eq^{2}(\text{MSE}(Y,Y_{t}),\frac{1}{2}\sum_{i=1}^{2}(y_{i}-y_{ti})^{2}))$ (A.23)
+    $\varphi o_{10}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{1}^{(o)})\land$
+    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{1}^{(o)})\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{2}^{(o)})\land$
+    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{2}^{(o)})\land$
+    $Eq^{2}(\nabla_{W^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})\land$
+    $Eq^{2}(\nabla_{b^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))\land$
+    $Eq^{2}(\nabla_{W^{(o)}_{2}}L,(Y-Y_{t})(\text{ReLU}(W_{1}^{(o)}X+b_{1}^{(o)}))^{T})\land Eq^{2}(\nabla_{b^{(o)}_{2}}L,(Y-Y_{t}))$ (A.24)
+  • Learning Rules
+    $\varphi o_{11}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\varphi o_{10}\land Eq^{2}(W_{1}^{(u)},W_{1}^{(o)}-\eta\nabla_{W^{(o)}_{1}}L)\land$
+    $Eq^{2}(b_{1}^{(u)},b_{1}^{(o)}-\eta\nabla_{b^{(o)}_{1}}L)\land Eq^{2}(W_{2}^{(u)},W_{2}^{(o)}-\eta\nabla_{W^{(o)}_{2}}L)\land$
+    $Eq^{2}(b_{2}^{(u)},b_{2}^{(o)}-\eta\nabla_{b^{(o)}_{2}}L)$ (A.25)
+In equation (A.14), the carrier M refers to the initially constructed model, and $To_{m}$ denotes the time when model M is completed. Clearly, we have:
+$\sup(To_{h})\leq\inf(To_{m})$ (A.26)
+The state $S_{M}(M,To_{m})$ is a one-to-one enabling mapping of the state $S_{oo}(oo,To_{h})$ onto M, that is,
+$S_{M}(M,To_{m})=\{\psi o_{1},\psi o_{2},\psi o_{3},\psi o_{4},\psi o_{5},\psi o_{6},\psi o_{7},\psi o_{8},\psi o_{9},\psi o_{10},\psi o_{11}\}$ (A.27)
+and
+$Io(\varphi o_{i})\underline{\triangleright}\psi o_{i},(1\leq i\leq 11)$ (A.28)
+Here, the information enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M.
+In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise.
+However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, noise-free, and recoverable information model.
+The subsequent discussions on other types of information will follow this approach and will not be repeated.
+A.3.2 Training Information of M
+Based on MLT-MF, the training information for model M is as follows:
+$It=(ot,Tt_{h},S_{ot},Mt,Tt_{m},S_{Mt})$ (A.29)
+Here, the ontology $ot$ refers to the object described by the training information, and the occurrence time $Tt_{h}$ denotes the point at which this object attains its corresponding state attributes.
+The ontological state $S_{ot}(ot,Tt_{h})$ represents the attributes of the object at time $Tt_{h}$, which are described by a set of well-formed formulas:
+$S_{ot}(ot,Tt_{h})=\{\varphi t_{1},\varphi t_{2}\}$ (A.30)
+  • Vector Definition
+    $\varphi t_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.31)
+  • Label Consistency
+    $\varphi t_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.32)
+The carrier Mt serves as the interface device through which model M receives training data, and $Tt_{m}$ is the time at which model M is trained.
+The carrier state $S_{Mt}(Mt,Tt_{m})$ is then described by the set of well-formed formulas:
+$It(\varphi t_{i})\underline{\triangleright}\psi t_{i}(1\leq i\leq 2)$ (A.33)
+A.3.3 Optimized Information After M Has Learned from Training
+Based on MLT-MF, after learning the training information $It$, model M updates its initial information $I_{0}$ to a new, optimized information state:
+$I_{u}=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ (A.34)
+Here, the ontology $ou$ is a fusion of M and the object described by the training information; the occurrence time $Tu_{h}$ is the moment M learns from the training data; and the ontological state $S_{ou}(ou,Tu_{h})$ represents the theoretical state of M after training:
+$S_{ou}(ou,Tu_{h})=\{\varphi u_{1},\varphi u_{6},\varphi u_{7},\varphi u_{8},\varphi u_{9},\varphi u_{10},\varphi u_{11}\}$ (A.35)
+Since the structure and algorithms of M remain unchanged before and after training, $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{7}$ are exactly the same as $\varphi o_{1}$, $\varphi o_{6}$, and $\varphi o_{9}$ respectively.
+Moreover, as the weights and biases after training are already assigned values in the original information (A.25), there is no need to include equivalent formulas for $\varphi o_{2}$, $\varphi o_{3}$, $\varphi o_{4}$, $\varphi o_{5}$ in $S_{ou}(ou,Tu_{h})$.
+For the remaining formulas $\varphi u_{7}$, $\varphi u_{8}$, $\varphi u_{9}$, $\varphi u_{10}$, and $\varphi u_{11}$, it suffices to systematically replace $W_{k}^{(o)}$, $w_{ij}^{(o)k}$, $b_{k}^{(o)}$, $b_{i}^{(o)k}$, $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$ and $b_{i}^{(u)k}$ in $\varphi_{7}$, $\varphi_{8}$, $\varphi_{10}$, and $\varphi_{11}$ with the corresponding symbols $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$.
+Similarly, the carrier state
+$S_{M}(M,Tu_{m})=\{\psi u_{1},\psi u_{6},\psi u_{7},\psi u_{8},\psi u_{9},\psi u_{10},\psi u_{11}\}$ (A.36)
+is a one-to-one enabling mapping from $S_{ou}(ou,Tu_{h})$ onto M, representing the state of M after completing its training.
+A.3.4 Input Information for M
+Processing input information is a fundamental requirement for any model. Based on the MLT-MF framework, the input information for M is defined as:
+$I_{q}=(oq,Tq_{h},S_{oq},Mq,Tq_{m},S_{Mq})$ (A.37)
+Here, $oq$ denotes the object described by the input information, and $Tq_{h}$ is the time at which the object’s state occurs. The state $S_{oq}(oq,Tq_{h})$ can take two possible forms:
+  • The first scenario involves only input information being processed by M, in which case
+    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1}\}$ (A.38)
+    $\varphi q_{1}=\forall XBelongsTo^{2}(X,\mathbb{R}^{2})$ (A.39)
+  • The second scenario considers online learning, where
+    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1},\varphi q_{2}\}$ (A.40)
+    $\varphi q_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.41)
+    $\varphi q_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.42)
+The carrier $Mq$ for $I_{q}$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$.
+The state $S_{Mq}(Mq,Tq_{m})$ is a one-to-one enabling mapping from $S_{oq}(oq,Tq_{h})$ onto $Mq$.
+A.3.5 Information After M Processes Input, Produces Output, and Undergoes Further Optimization
+There are two scenarios for M after it processes input and produces output: one corresponds to the first case in A.3.4, where only output is produced without further optimization of the model itself; the other corresponds to the second case in A.3.4, where, in the context of online learning, the model both produces output and undergoes further self-optimization.
+Since the latter is more comprehensive, we will focus on this scenario below.
+In this case, the output information $Ir=(or,Tr_{h},S_{or},Mr,Tr_{m},S_{Mr})$ has an ontological state defined as:
+$S_{or}(or,Tr_{h})$ $=\{\varphi r_{1}\}=\{\varphi o_{7}\land\varphi o_{8}\land\text{OutputAt}^{2}(Y,Tr_{h})\}$ (A.43)
+The carrier state $S_{Mr}(Mr,Tr_{m})$ is a one-to-one enabling mapping from $S_{or}(or,Tr_{h})$ to the output device of M.
+On the other hand, since we are considering the scenario where M performs online learning, the model now carries further optimized information:
+$Iv=(ov,Tv_{h},S_{ov},M,Tv_{m},S_{M})$ (A.44)
+This structure is almost identical to the optimized information $Iu$ after M has learned from training. The ontological state of $Iv$ is:
+$S_{ov}(ov,Tv_{h})=\{\varphi v_{1},\varphi v_{6},\varphi v_{7},\varphi v_{8},\varphi v_{9},\varphi v_{10},\varphi v_{11}\}$ (A.45)
+Here, $\varphi v_{1}$, $\varphi v_{6}$, and $\varphi v_{9}$ are identical to $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{9}$ respectively.
+For the remaining formulas, it suffices to systematically replace the parameters $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$ with the corresponding parameters $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$, $b_{i}^{(v)k}$, $W_{1}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(w)}$ and $b_{i}^{(w)k}$ to obtain the updated formulas, where $W_{k}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(2)}$ and $b_{i}^{(w)k}$ are the weights and biases of M after further optimization.
+Similarly, the carrier state
+$S_{M}(M,Tv_{m})=\{\psi v_{1},\psi v_{6},\psi v_{7},\psi v_{8},\psi v_{9},\psi v_{10},\psi v_{11}\}$ (A.46)
+is a one-to-one enabling mapping from $S_{ov}(ov,Tv_{h})$ onto M.
+At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Figure A1.
+Although M is a simple model, it contains all the essential components of a neural network, and the discussion has not relied on any special treatment of M.
+This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.

### v10

@@
 Abstract
-[Objective] This study addresses key challenges in machine learning, namely the absence of a unified formal theoretical framework and the lack of foundational theories for model interpretability and ethical safety.
-[Methods] We first construct a formal information model, explicitly defining the ontological states and carrier mappings of typical machine learning stages using sets of well-formed formulas.
-By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the causal chain logic and constraint laws governing machine learning processes.
-[Results] We establish the Machine Learning Theory Meta-Framework (MLT-MF), on which we further propose universal definitions for model interpretability and ethical safety.
-We prove and validate three key theorems: the relationship between model interpretability and information existence, ethical safety assurance, and the upper bound estimation of total variation distance (TVD).
-[Limitations] The current framework assumes ideal, noise-free information enabling mappings and focuses primarily on model learning and processing logic in static scenarios.
-It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
-[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing critical issues in contemporary machine learning.
-Keywords: machine learning; objective information theory; formal logic systems; interpretability; ethical and safety assurance; generalization error
+This paper addresses the current lack of a unified formal framework in machine learning theory, as well as the absence of robust theoretical foundations for interpretability and ethical safety assurance.
+We first construct a formal information model, employing sets of well-formed formulas (WFFs) to explicitly define the ontological states and carrier mappings for the core components of machine learning.
+By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the logical inference and constraint rules underlying causal chains in models, thereby establishing the Machine Learning Theory Meta-Framework (MLT-MF).
+Building upon this framework, we propose universal definitions for model interpretability and ethical safety, and rigorously prove and validate four key theorems: the equivalence between model interpretability and information existence, the constructive formulation of ethical safety assurance and two types of total variation distance (TVD) upper bounds.
+This work overcomes the limitations of previous fragmented approaches, providing a unified theoretical foundation from an information science perspective to systematically address the critical challenges currently facing machine learning.
+Keywords: machine learning, objective information theory (OIT), formal logic systems, interpretability, ethical safety assurance, estimation of error bound
 1 Introduction
-Although artificial intelligence is currently the most dynamic field within information science and even the broader domain of science and technology, existing research still faces several significant shortcomings:
-First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning Rob2021 ; Marius2023 ; Meng2023 ; Alquier2024 ; Vilalta2002 .
-Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages.
-This fragmentation makes it impossible to characterize the coherent evolution of knowledge from prior injection, through data-driven updates, to task generalization.
-For example, the PAC-Bayes framework connects priors with generalization error but neglects the temporal causality of the training process; meta-learning involves multi-stage adaptation but fails to establish a symbolic model for state transitions.
-The absence of lifecycle modeling makes systematic analysis of the core question——“how parameters carry task knowledge from initialization through training”—difficult for model designers.
-Second, there is a lack of explicit expression for the causal chain of information flow.
-Traditional approaches often treat the transformation from data to model as a black box, reducing the parameter update driven by loss function gradients to a purely mathematical optimization process and overlooking the causal dependencies among prior knowledge, data distribution, and model architecture.
-While the field of causal inference attempts to characterize causal relationships among variables, its application in machine learning is mostly limited to feature-label association analysis and fails to construct a complete causal chain from “prior constraints → initialization → gradient propagation → generalization capability.” For instance, adversarial training enhances robustness by maximizing loss, but theoretical analyses rarely explicitly link initial parameter constraints with the causal mechanisms underlying final adversarial robustness, making strategy design reliant on empirical trial-and-error.
-Third, the mathematical foundations of interpretability tools are weak Ribeiro2016 ; Lundberg2017 ; Venkata2025 ; Lee2023 ; Alvarez2018 .
-Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization.
-For example, SHAP values compute marginal contributions in cooperative games but do not formalize their mapping to the information carrier states of models, potentially violating causality in interpretations.
-In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
-Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learningShwartz2017 ; Xu2022 .
-Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality.
-This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systemsXu2014 ; xu2015 ; Jianfeng2021 ; Jianfeng2023 ; Xu2024 ; Xu2025 .
-On this basis, we rigorously define higher-order logic systems using recursion, and establish formal state expressions through interpretations of sets of well-formed formulas (WFFs), completing the last theoretical gap in OIT regarding the mathematical essence of state expressions.
-We then thoroughly apply the theory and methods of OIT to unify the core concepts throughout the machine learning lifecycle—prior knowledge encoding, data-driven training, and model generalization—as formal state expressions; we establish a mathematical logic system covering the entire ML process.
-Using this, we analyze and prove the causal-chain logic and constraint rules for typical stages in machine learning, including the formal modeling and realization of training information, the formal modeling and realization of model learning, and the model’s information processing and learning optimization, thus forming a Meta-Framework for Machine Learning Theory (MLT-MF).
-Finally, using MLT-MF, we propose universal definitions of model interpretability and ethical safety, prove and validate—via a simple feedforward neural network—three important theorems: model interpretability and information existence, ethical safety assurance, and an upper bound estimate on total variation distance (TVD).
-These results provide a powerful tool, based on a unified formal logic framework, for both theoretical analysis and engineering practice in machine learning.
-2 Postulates and Formal Definition of Information
+Although artificial intelligence (AI), exemplified by machine learning (ML), is among the most dynamic areas in information science and technology, several critical shortcomings persist in the current research landscape.
+First, there is a lack of a unified logical framework that spans the entire lifecycle of machine learning.
+Most existing ML theory focuses on isolated components: optimization theory emphasizes training dynamics, generalization analysis targets test performance, while initialization strategies and deployment are often considered independently.
+This fragmentation precludes the coherent characterization of knowledge evolution from prior injection, data-driven updates, to task-level generalization (Vapnik, 1999; Alquier, 2024; Schölkopf et al., 2021; Garcez and Lamb, 2023; Marra et al., 2024).
+Second, the unpredictability of model performance has led to rapidly escalating research and development costs.
+As model sizes continue to grow, the demands for data, computation, energy, time, and human resources have multiplied.
+In the absence of explicit causal semantics, structural analysis, and effective metrics, it is often challenging to analyze or predict attainable performance in each stage of the pipeline based on error upper bounds.
+Incomplete information coverage, suboptimal model selection, inaccurate measurement, and quality deficiencies frequently result in failed development efforts and significant resource losses (Achiam et al., 2023; Touvron et al., 2023; Chowdhery et al., 2023; Gunasekar et al., 2023).
+Third, there are profound concerns regarding ethical safety assurance.
+As AI capabilities advance, so too do apprehensions about their ethical implications.
+However, there is still no consensus—either mathematical or theoretical—on the scientific definition, substantive content, forms of expression, or criteria for ethical safety in AI.
+Most research remains focused on ethical principles and regulatory mechanisms, while the proliferation of competing guidelines leaves practitioners uncertain about which standards to follow (Russell et al., 2015; Burr and Leslie, 2023; Tilala et al., 2024; Hawkins et al., 2021; Pant et al., 2024).
+Fourth, the theoretical foundations for interpretability remain weak.
+The lack of a unified semantic framework for causal inference and constraint logic means that traditional ML methods treat the transformation from data to model as a black box.
+Although interpretability is widely recognized as a critical issue and diverse approaches have been proposed to address it, there is as yet no rigorous definition or universally applicable evaluative criteria (Lipton, 2018; Glanois et al., 2024; Zhang et al., 2024; Rudin, 2019; Miller, 2019).
+As a central discipline bridging data, knowledge, and systems, information science has begun to offer new perspectives for the theoretical integration of machine learning (Shwartz-Ziv and Tishby, 2017; Xu et al., 2022).
+However, information science theory remains fragmented across information theory, control theory, signal processing, statistical learning, and computational theory, each developing largely in isolation and lacking a unified semantic and logical system.
+This fragmentation is a primary reason for the absence of a unified theoretical framework in contemporary AI.
+In response, Objective Information Theory (OIT) applies an axiomatic approach to construct a sextuple model of information, representing intrinsic meaning via ontological states, objective expression via carrier states, and revealing the dual-subject structure of information through the enabling mapping from the former to the latter.
+OIT thus provides a general and universal foundational model and metrics for understanding information in the objective world, and especially within computing and information systems (Xu et al., 2014; Xu, 2024).
+For example, (Xu et al., 2025) demonstrates that applying OIT’s general information metrics to pre-evaluate training data can significantly enhance model training efficiency.
+This paper further defines the enabling mapping, clarifying its dual mathematical and physical-causal attributes.
+Through a recursively defined higher-order logical system, we formalize the expression of states via the interpretation of sets of well-formed formulas (WFFs), elucidating the mathematical essence of state representation in information definitions.
+This fills the final theoretical gap in OIT and offers a more powerful toolkit for applying OIT to diverse problems in information science.
+The introduction of noisy information and information chains further establishes the theoretical groundwork for analyzing key challenges across all stages of machine learning.
+On this basis, we comprehensively apply OIT’s theories and methodologies—centered on the learnable and processable predicates, as well as learning and processing functions—to five canonical stages in machine learning: model construction, training information, learning and training, input of problems, information processing outputs and online learning.
+We establish a unified semantic formalization of state representation, and develop causal-chain logical inference and constraint rules covering the full ML lifecycle, culminating in the Meta-Framework for Machine Learning Theory (MLT-MF).
+Applying MLT-MF, we formulate universal definitions for model interpretability and ethical safety, and prove theorems relating interpretability to information existence, as well as the theoretical underpinnings for ethical safety assurance.
+For model error, we establish total variation distance (TVD) upper bounds for two strategies—conservative prediction in regions not covered by training with noiseless input, and arbitrary prediction with noisy input—providing new guidelines for the estimation and optimization of machine learning performance.
+Finally, we illustrate and validate the practical utility of MLT-MF and the associated theorems with a simple feedforward neural network model, demonstrating that this research offers a novel direction and methodology for both theoretical analysis and engineering practice in machine learning, grounded in a unified formal logical framework.
+2 Related Work
+2.1 Conceptual Representations of Machine Learning
+Generally, machine learning refers to methods that learn general rules from limited observational data and use these rules to predict unknown information.
+Numerous works have provided more detailed analyses of this concept.
+Representative views include:
+In the context of statistical learning (Vapnik, 1999), the learning model is decomposed into three components: a set of independent random vectors $x$ that follow a certain but unknown probability distribution $P(x)$; a set of output vectors $y$ corresponding to input $x$, also following a certain but unknown conditional probability distribution $P\left(y\middle|x\right)$; and a collection of functions $f(x,\alpha),\alpha\in\Lambda$, that the learning machine can implement.
+Machine learning is thus defined as selecting, based on a training dataset composed of $l$ independent observations $\left(x_{1},y_{1}\right),\cdots,(x_{l},y_{l})$ drawn from the same distribution $P(x,y)=P(x)P\left(y\middle|x\right)$, a function $f(x,\alpha)$(with $\alpha\in\Lambda$) that minimizes the risk function $R(\alpha)=\int{L(y,f(x,\alpha))d}P(x,y)$, where $L(y,f(x,\alpha))$ is the loss function of $f(x,\alpha)$.
+Another perspective highlights the central role of model parameters (Alquier, 2024), describing machine learning as the process of learning from an object-label set $(X,Y)$ governed by a probability distribution $P$, to produce a family of predictors $\left\{f_{\theta},\theta\in\Theta\right\}$ parameterized by $\theta\in\Theta$.
+The loss function $\mathcal{l(}f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\mathcal{[l(}f_{\theta}(X),Y)]$.
+The ultimate goal is to find $\theta^{*}$ such that $R\left(\theta^{*}\right)=\inf_{\theta\in\Theta}R(\theta)$.
+From a more advanced perspective, the focus shifts to causal relationships beyond data distribution features (Schölkopf et al., 2021).
+This view treats perceptual data acquired through measurement as entangled observations of unknown causal system states.
+It seeks to learn concrete representations that at least partially reveal these hidden causal structures by leveraging large-scale training, data augmentation, self-supervised learning, and robust fine-tuning using data from multiple simulated environments.
+This includes discovering causal models implicit in environmental data and performing causal inference to plan and act based on these models.
+Some researchers emphasize the integration of neural and symbolic artificial intelligence (Garcez and Lamb, 2023), arguing that knowledge should be rooted in efficient data learning via message passing in neural networks and embedded in vector representations as efficient computational models.
+Symbols, in turn, should be acquired through querying and extracting knowledge from trained networks, providing a rich descriptive language at appropriate levels of abstraction—allowing finite means to achieve infinite ends—while enabling compositional and discrete reasoning at the symbolic level, thereby permitting extrapolation beyond data distributions.
+The combination of learning and reasoning should offer an important alternative for tackling compositional reasoning problems, reducing the number of effective combinations through learning and thus generating simpler symbolic descriptions within a neuro-symbolic cycle.
+Another viewpoint proposes that structure learning through parameter learning is the most prominent paradigm among the statistical relational learning approaches within neuro-symbolic AI (Marra et al., 2024).
+The learning task involves estimating probabilities for a fixed set of clauses, with the advantage of removing combinatorial search from the learning process.
+The above works collectively reflect, from different perspectives, the academic community’s comprehensive understanding of both past achievements and future directions in machine learning.
+However, none have established a formalized expression for the entire lifecycle implementation process of machine learning under a unified semantic framework.
+As a result, it remains consistently difficult to apply structured systems analysis methods to guide the implementation and optimization of models throughout research and design.
+2.2 Model Interpretability
+The interpretability of machine learning has attracted significant attention in recent years.
+As noted in (Lipton, 2018), few authors in the academic literature are able to precisely articulate what interpretability actually means, or to clearly explain how their proposed solutions achieve it.
+Despite the absence of a unified definition, an increasing number of publications continue to introduce algorithms purported to possess interpretability.
+This situation supports two possible conclusions: first, that there is a universally accepted understanding of interpretability that simply remains unwritten; or second, that the term itself is inherently ambiguous, leading to quasi-scientific claims regarding the interpretability of various models.
+In reality, the objective of interpretability may simply be to extract more valuable information from models.
+Reference (Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered.
+Based on a definition of comprehensibility encompassing simulatability, decomposability, and algorithmic transparency, the authors emphasize that interpretability in reinforcement learning relies on three pillars: interpretable inputs, interpretable models, and interpretable decisions.
+They further analyze the specific mechanisms for achieving interpretability in each of these aspects.
+In (Zhang et al., 2024), the impact of different training mechanisms on two key factors—model smoothness and gradient similarity—is investigated from the perspective of agents, with the aim of deepening our understanding of the dynamics of model transferability and its relationship to various training methods.
+Building on adversarial training, two working hypotheses are proposed and validated: first, that the trade-off between model smoothness and gradient similarity largely determines the adversarial transferability of agent models; and second, that shifts in data distribution can undermine gradient similarity, thereby resulting in generally poorer substitutes.
+This approach—explaining changes in model performance by dissecting and adjusting key factors in the training process—should be regarded as an important direction in the study of machine learning interpretability.
+It is evident, therefore, that although interpretability is a critical issue in ML, the lack of a unified semantic and formal framework for machine learning has precluded the establishment of rigorous definitions and universally applicable criteria for interpretability.
+Approaches to enhancing interpretability remain diverse, and there is still no universally accepted methodology informed by a common theoretical foundation.
+2.3 Ethical Safety Assurance
+Concerns regarding the ethical safety of AI have a long history.
+As noted in [10], the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies.
+These issues span law, public policy, professional ethics, and philosophical morality, necessitating the expertise of computer scientists, legal scholars, political scientists, and ethicists.
+While mathematical tools such as formal logic, probability theory, and decision theory have provided important foundations for reasoning and decision-making, numerous challenges remain unresolved.
+Building on traditional concepts of assurance, [11] offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence.
+Accordingly, the general structure of moral assurance is characterized as an iterative process of reflection, action, and adjustment, repeated throughout the design, development, and deployment lifecycle.
+In the context of AI/ML-supported healthcare, ethical safety assurance faces more concrete challenges.
+Reference [12] provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare.
+The authors advocate for multiple strategies, including the implementation of rigorous data security protocols, prioritization of data minimization, increasing diversity in training datasets, use of data augmentation to address imbalance, application of algorithmic fairness techniques to detect and mitigate model bias, regular auditing of learning algorithms, enhancing the interpretability of algorithmic outputs, regulatory oversight based on standardized performance metrics and stringent testing protocols, and the augmentation—rather than replacement—of clinicians’ capabilities as means to achieve ethical safety.
+[13] further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety.
+The authors propose that ML safety requirements should be derived from the system-level safety requirements, and that these should be evaluated in light of the allocated safety needs, system and software architecture, and operational environment.
+This framework then serves as the basis for developing safety assurance specifications across all stages of the ML lifecycle, including data preparation, model training, model validation, and model deployment.
+Reference [14] conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches.
+The study highlights that, in the absence of a universally adopted set of ethical principles for AI, incorporating morality into AI is fundamentally challenged by the lack of clear moral knowledge and the ambiguity of ethical principles.
+It is thus evident that, despite widespread concerns over the ethical safety of machine learning, there is as yet no mathematical or theoretical consensus regarding its scientific definition, substantive content, forms of manifestation, or criteria for assessment.
+Instead, ethical safety remains an open-ended concept, explored through various technical, engineering, and moral perspectives.
+Most research either focuses on specific ethical principles and consequences for AI, or on methods for enhancing ethical conduct.
+The proliferation of often conflicting AI ethical guidelines has, paradoxically, made it difficult for practitioners to determine which standards to follow.
+This underscores the necessity of a universally accepted definition of ethical safety, as such a definition can bridge the gap between ongoing academic discourse and industry practice, thereby providing a concrete foundation for strengthening ethical safety assurance in AI.
+2.4 Model Error Estimation
+Model performance is fundamentally reflected in the errors it may incur on specific tasks.
+Reference (Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions.
+Based on VC dimension, it further establishes distribution-independent risk upper bounds and sufficient conditions for fast convergence under both bounded and unbounded loss in finite-sample settings.
+By controlling the trade-off between VC dimension capacity and empirical risk through nested hypothesis families, it proves the strong consistency and convergence rates of structural risk minimization (SRM).
+These theoretical foundations have led to practical implementations such as support vector machines (SVM), margin maximization, and kernel methods, offering explanations for their generalization mechanisms.
+Reference (Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error.
+It provides finite-sample error upper bounds via the fixed point of sub-root functions, thus achieving optimal or near-optimal convergence rates inversely proportional to the square root of the number of training samples.
+Under controlled variance conditions, it presents a unified framework for error estimation, including both distribution-dependent bounds determined by local Rademacher complexity and data-dependent bounds constructed from empirical local Rademacher averages.
+These results are extended to typical classification and regression tasks, providing computable upper bounds and iterative algorithms for quickly approximating the fixed point.
+From an information-theoretic perspective, (Miller, 2019) models learning algorithms as random channels mapping input datasets $S$ to output hypothesis spaces $W$.
+For sub-Gaussian (including bounded) losses and uncountable hypothesis spaces, the generalization error upper bound is characterized via the mutual information $I(S;W)$ between input and output.
+Based on this, high-probability sample complexity and expected generalization error upper bounds are derived, and the concept of information stability $(\varepsilon,\mu)$-stability is introduced, unifying algorithmic consistency and overfitting control through information leakage measures.
+The paper further analyzes two classes of algorithms—Gibbs learning and noisy ERM—where $I(S;W)$ is explicitly controlled via mutual information regularization, and discusses methods for decomposing and additively controlling mutual information in complex learning algorithms through adaptive combinations.
+Reference (Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence.
+It introduces a task distribution T and task-level MDPs, defines generalization error as the performance gap between training and unseen tasks under a meta-policy, and establishes generalization upper bounds determined by the number of training tasks and task complexity, using empirical process theory and concentration inequalities.
+The analysis further discusses the impact of algorithmic capacity and learning rates on error bounds.
+Within the non-convex SGD framework, sufficient conditions and rates for finite-step convergence of meta-learning loss are presented, along with guidelines for long-term behavior and hyperparameter selection based on linear and sub-linear types.
+The theoretical results are validated and the observability of convergence rates is demonstrated through controllable task generation and baseline meta-algorithms.
+Reference (Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models.
+From both theoretical and numerical perspectives, it investigates the convergence properties of deterministic samplers based on probability flow ordinary differential equations (ODEs).
+Assuming access to score function estimates with $L_{2}$ accuracy, it provides upper bounds on TVD between target and generated distributions in continuous time.
+For the practical implementation using Runge-Kutta integrators, discrete-time error upper bounds are established.
+Numerical experiments on 128-dimensional problems are conducted to validate these theoretical error bounds.
+Collectively, these works encompass both classical statistical learning theory and recent advances in reinforcement learning and generative models, reflecting a broad recognition of the importance of model error upper bounds in machine learning.
+However, these studies do not employ explicit causal semantic modeling, nor do they address the full lifecycle challenges of multi-stage information transmission, online learning, or multi-agent integration in ML.
+Additionally, they rarely consider the geometric scale and entropy of distribution shifts and uncovered domains, making it difficult to offer actionable engineering or data-sampling guidance for test domain coverage gaps and noisy input scenarios.
+3 Preliminary Theory
 Undoubtedly, as a key component of artificial intelligence, modern machine learning is a product of the rapid development and high integration of information technologies such as computers, the Internet, big data, and cloud computing.
@@
 Starting from the basic concept of information can provide a unified theoretical foundation for research in machine learning and artificial intelligence.
-2.1 Postulates of Information
-Although humanity has entered the information age, the theoretical foundations of information science have long been dispersed across disciplines such as information theory, cybernetics, signal processing, statistical learning, and computational theory, each developing independently.
-There remains a lack of consensus on the concept of information, as well as a unified axiomatic basis and theoretical system.
-This is also a major reason why artificial intelligence often lacks interpretability and predictability.
-To address this, we first propose the concept of an enabling mapping.
+3.1 Definition of Information
+Reference (Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states.
+It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Xu, 2024), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times.
+Here, we proceed to further clarify the concept of enabling mapping.
 Definition 1 (Enabling Mapping) Let $X$ and $Y$ be two sets, and let $f$ be a mapping.
@@
 or
-$f(X)\underline{\triangleright}Y$ (2.2)
+$f(X)\trianglerighteq Y$ (3.2)
 […]
-If $f(X)\underline{\triangleright}Y$, we also denote this as $f^{-1}(y)\underline{\triangleleft}x$ or $x\underline{\triangleright}f^{-1}(y)$, to maintain consistency in the directionality of enabling.
-Based on the concept of enabling mapping, Xu2024 proposed four postulates of information:
-Postulate 1 (Binary Attribute) Every piece of information comprises two components: the ontology, which represents the intrinsic essence of the information, and the carrier, which embodies the objective form of the information.
-Postulate 2 (Existence Duration) The content and form of information have their own respective temporal existences.
-Postulate 3 (State Representation) The states of the ontology and the carrier of information have their own respective representations.
-Postulate 4 (Enabling Mapping) The state of the ontology can be mapped onto the state of the carrier via an enabling mapping.
-These four postulates are concise, align with common conceptions of information, and are compatible with the basic concepts of major branches of information science.
-Their scientific significance is reflected in:
-  • Unification: Providing a common foundation for long-dispersed theories in information science;
-  • Axiomatization: Allowing the definition and reasoning of information to be axiomatized and formalized, akin to mathematics and physics;
-  • Universality: Facilitating application to cutting-edge fields such as the Internet, big data, artificial intelligence, cognitive science, and cyber-physical systems.
-The concept of "enabling mapping" is novel.
-It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realizationShannon1948 ; Landauer1996 ; Wiener2019 .
-2.2 Formal Representation of States
-The postulates of information emphasize that the state of the ontology enables a mapping to the state of the carrier.
-Therefore, establishing a formal representation of both ontological and carrier states is particularly important.
-Definition 2 (Higher-Order Formal Logical System): Suppose a formal system $\mathcal{L}$ is described using the following symbol set:
-Variables $x_{1},x_{2},\ldots$;
-Individual constants $a_{1},a_{2},\ldots$;
-Functions $f_{1}^{1},f_{2}^{1},\ldots,f_{1}^{2},f_{2}^{2},\ldots,f_{1}^{3},f_{2}^{3},\ldots$;
-Parentheses $(,)$;
-Predicates $A_{1}^{1},A_{2}^{1},\ldots,A_{1}^{2},A_{2}^{2},\ldots,A_{1}^{3},A_{2}^{3},\ldots$;
-Logical connectives "not" $\sim$, "implication" $\rightarrow$;
-Quantifier "for all" $\forall$;
-A term in $\mathcal{L}$ is defined as follows:
-(1) Variables and individual constants are terms;
-(2) If $f_{i}^{n}(n>0,i>0)$ is a function in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $f_{i}^{n}$ itself, then $f_{i}^{n}(t_{1},\ldots,t_{n})$ is also a term in $\mathcal{L}$;
-(3) If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is a term in $\mathcal{L}$.
-(4) The set of all terms in $\mathcal{L}$ is generated by (1), (2), and (3).
-Based on this, we can define atomic formulas in $\mathcal{L}$:
-If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$ and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ not containing $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is an atomic formula in $\mathcal{L}$.
-We then define well-formed formulas (WFFs) in $\mathcal{L}$:
-(5) Every atomic formula in $\mathcal{L}$ is a WFF;
-(6) If $\mathcal{A}$ and $\mathcal{B}$ are WFFs in $\mathcal{L}$, then $\sim\mathcal{A}$, $\mathcal{A}\rightarrow\mathcal{B}$, $(\forall x_{i})\mathcal{A}$, $(\forall f_{i}^{n})\mathcal{A}$, and $(\forall A_{i}^{n})\mathcal{A}$ are all WFFs in $\mathcal{L}$, where $x_{i}$, $f_{i}^{n}$, $A_{i}^{n}$ are variables, functions, and predicates in $\mathcal{L}$, respectively, and no predicate or function symbol is applied, either directly or indirectly, to itself.
-(7) The set of all WFFs in $\mathcal{L}$ is generated by (5) and (6).
-The higher-order formal logical system $\mathcal{L}$ defined here, based on the flexibility of predicate logic, can describe various relationships and constraints among objects in a unified way.
-Through higher-order quantification support, the nested definition of complex data and behaviors, and it explicitly prohibits anypredicate or function symbol from applying to itself—directly or indirectly—to avoid self-referential paradoxes.
-For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols.
-However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more conciseHamilton1988 .
-Therefore, they are introduced as defined symbols in subsequent discussions.
-Definition 3 (Interpretation of a Formal System) An interpretation $E$ of a formal system $\mathcal{L}$ over a domain $D_{E}$ assigns specific values to constants, elements of the domain $D_{E}$ to variables, and turns function and predicate symbols into functions and predicates over that domain $D_{E}$, thereby yielding a set of WFFs each with a truth value.
-Definition 3 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects as both the subjective and objective worlds.
-Postulate 5 (Expression of Object State):The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
-(1) Domain invariance: definitions of the object set and time set remain unchanged.
-(2) Numerical determinacy: concrete numerical values determine values of certain properties and forms.
-(3)Attribute expressibility: properties, forms, values, relations are expressible via functions and predicates in a formal system satisfying Definition 2.
-(4) Rule expressibility: evolution rules among attributes are expressible in such a formal system using time variables and implication, and contain no implicit self-reference.
-(5) Logical coherence: no logical contradictions exist among attributes and rules, nor among any logical conjunctions thereof.
-Theorem 1 (Formal Expression of States) If the state set $S(X,T)$ of objects $X$ over times $T$ satisfies Postulate 5, then it can be represented as an interpretation of a formal system $\mathcal{L}$ over the domain $X\times T$, and it is logically coherent.
-Proof: Since the state set $S(X,T)$ of the object set $X$ over the time set $T$ satisfies Postulate 5, we construct the corresponding formal system $\mathcal{L}$ as follows:
-By (1), the definitions of the elements in $X$ and $T$ remain unchanged, ensuring determinacy.
-The variables in the formal system $\mathcal{L}$ are defined to correspond to each element of $X\times T$, thus making $X\times T$ the domain of discourse for $\mathcal{L}$.
-Consequently, $S(X,T)$ has a well-defined domain.
-By (2), the constants in $\mathcal{L}$ are defined to represent all possible values in $S(X,T)$.
-By (3), the functions and predicates in $\mathcal{L}$ represent the properties, forms, values, and relations of $X$ over $T$ as captured in $S(X,T)$.
-These properties can be expressed through functions and predicates in the formal system $\mathcal{L}$ that conform to Definition 1.
-By (4), consider two ternary predicates $A_{1}^{3}$ and $A_{2}^{3}$.And then let $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ and $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$ be two properties of $S(X,T)$ corresponding to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and binary functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$.
-Suppose all $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ satisfy the rule of evolving from $t_{1}$ to $t_{2}$ into some $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$.
-The formula $(\forall A_{1}^{3}\forall f_{1}^{2}\forall x_{1}A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))\rightarrow\exists A_{2}^{3}\exists f_{2}^{2}\exists x_{2}A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2})))$ is also a well-formed formula in $\mathcal{L}$, with respect to the interpretation of $x_{1},x_{2},t_{1},t_{2}$, and is able to express the corresponding rules in $S(X,T)$.
-For other rules in $S(X,T)$, the related elements in the domain $X\times T$ can similarly be interpreted and expressed via corresponding WFFs in $\mathcal{L}$, with no implicit logical self-reference.
-Therefore, $S(X,T)$ constitutes an interpretation $E$ of $\mathcal{L}$ over the domain $X\times T$. By (5), both itself and its logical closure are logically consistent. The theorem is thus proved.
-Note that Theorem 1 stipulates that variables in $\mathcal{L}$ are elements of $X\times T$; thus, in formulas, object and time elements appear in pairs $(x,t)$.
-In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
-Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world.
-Both theoretical and empirical analyses qiu2025 further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language.
-This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
-2.3 Finite Automaton States
-Using the formal expression of states, we can study in depth the special and important object of finite automata.
-Theorem 2 (State Representation of Finite Automata): For any finite automaton $M$, one can describe its input, output, and state-transition behavior over a related time set $T$ by the state set $S(M,T)$.
-Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions.
-By finite automaton theory mathsea2002 , $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
-$\delta:U\times Q\to U$ is the next-state function;
-$\lambda:U\times Q\to R$ is the output function.
-For each state $u_{t_{i}}\in U$, define the state predicate $State^{3}$ by
-$\varphi_{U}(t_{i})=State^{3}(M,t_{i},u_{t_{i}})$ (2.3)
-to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$$i=1,\ldots,n$.
-For each input $q_{t_{i}}\in Q$, define the input predicate $Input^{3}$ by
-$\varphi_{I}(t_{i})=Input^{3}(M,t_{i},q_{t_{i}})$ (2.4)
-to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$$i=1,\ldots,n-1$.
-For each output $r_{t_{i}}\in R$, define the output predicate $Output^{3}$ by
-$\varphi_{O}(t_{i})=Output^{3}(M,t_{i},r_{t_{i}})$ (2.5)
-to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$$i=2,\ldots,n$.
-The transition function $\delta$ is captured by the well-formed formula
-$\varphi_{\delta}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land Eq^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (2.6)
+Corollary 1 If $f$ is an enabled mapping from $X$ to $Y$, then $|X|\geq|Y|$, where $|X|$ and $|Y|$ denote the cardinalities of the sets $X$ and $Y$, respectively.
+If $f(X)\trianglerighteq Y$, we also denote this as $f^{-1}(y)\trianglelefteq x$ or $x\trianglerighteq f^{-1}(y)$, to maintain consistency in the directionality of enabling.
+Fig.1(a) intuitively illustrates the dual nature of an enabling mapping, encompassing both its mathematical mapping and physical enabling attributes.
+Figure 1: Schematic Diagram of Enabling Mapping, Noisy Information, and Information Transmission Chain.
+(a) The scene of penguin activity $X$ is mapped by an enabling mapping $f$ to an image $Y$ on a laptop; between $X$ and $Y$ there exists both a mathematical mapping relationship and a causal relationship.
+Moreover, mapping $X$ to $Y$ requires corresponding physical conditions and the expenditure of a certain amount of energy.
+(b) The left panel depicts the ontological state of noiseless information, while the right panel shows a noisy information state: the set $S_{o}^{-}$ is removed from the noiseless ontological state, and the previously disjoint set $S_{o}^{+}$ is superimposed, resulting in a noisy ontological state.
+(c) In the penguin activity information transmission chain, the carrier state of any subsequent piece of information can be regarded as the result of an enabling mapping from any prior ontological information state, and can also be reasonably interpreted by means of this ontological state and the corresponding enabling mapping.
+Based on this, a more rigorous theoretical foundation for the definition of information can be established:
+Definition 2 (Mathematical Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain.
+The ontology $o\mathbb{\subseteq O}\cup\mathbb{S}$, time of occurrence $T_{h}\mathbb{\subseteq T}$, ontological state set $S_{o}\left(o,T_{h}\right)$, carrier $c\mathbb{\subseteq O}$, reflection time $T_{m}\mathbb{\subseteq T}$, and carrier reflection set $S_{c}\left(c,T_{m}\right)$ are all non-empty sets.
+Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I=\left\langle o,T_{h},S_{c},c,T_{m},S_{o}\right\rangle$.
+The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Xu et al., 2025).
+It should be noted that both $S_{o}\left(o,T_{h}\right)$ and $S_{c}\left(c,T_{m}\right)$ can each be represented as the interpretation of a set of WFFs within a formal system.
+However, the respective formal systems need not be identical, which constitutes one of the essential characteristics of information—its capacity to establish connections among all entities in the world, regardless of differences in their underlying formal systems.
+Reference (Shannon, 1948) is regarded as a foundational work in information theory.
+The paper proposed that ” $H=-\sum{p_{i}\log p_{i}}$ plays a central role as the measure of information in information theory.” The discussion in this paper includes both the source and channel of information.
+The source $Is$ is the origin of the information, and it generates events $e_{1}$, $e_{2},$…, $e_{n}$ with probabilities $p_{1}$, $p_{2},$…, $p_{n}$, respectively.
+The channel $Ic$ is the medium through which information is transmitted.
+After an event $e_{i}$ occurs at time $T_{h}$ in the source $Is$, the channel $Ic$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$ (where $1\leq i\leq n,\ \tau>0$).
+We define the ternary predicate ${IsEvent}^{3}(Is,T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $Is$ at time $T_{h}$, and the ternary predicate ${IcCode}^{3}(Ic,T_{m},c_{i})$ to represent that the channel $Ic$ received and transmitted code $c_{i}$ in the time interval $T_{m}$.
+Then,
+$S_{Is}(Is,T_{h})={IsEvent}^{3}(Is,T_{h},e_{i})$ (3.3)
+and
+$S_{Ic}(Ic,(T_{h},T_{h}+\tau))={IcCode}^{3}(Ic,(T_{h},T_{h}+\tau),c_{i})$ (3.4)
+are the states of the source $Is$ at time $T_{h}$ and the state of the channel $Ic$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $(1\leq i\leq n)$). Furthermore,
+$I:S_{Is}(Is,T_{h})\rightrightarrows S_{Ic}(Ic,(T_{h},T_{h}+\tau))$ (3.5)
+represents the enabling mapping from the source state $S_{Is}(Is,T_{h})$ to the channel state $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$.
+Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{Is}(Is,T_{h})$ and the complex implementation mechanisms of $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system.
+The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
+3.2 Noisy Information
+The realization of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
+Definition 3 (Noisy Information) For information $I=\left\langle o,T_{h},S_{o},c,T_{m},S_{c}\right\rangle$, let $\widetilde{S_{o}}\left(o,T_{h}\right)$ be another state of the ontology $o$ at time $T_{h}$, and $\widetilde{S_{c}}\left(c,T_{m}\right)$ be another state of the carrier $c$ at time $T_{m}$,
 […]
-Similarly, the output function $\lambda$ is given by
-$\varphi_{\lambda}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land Eq^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land$ (2.7)
-$IsElementof^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$
+where the state $S_{o}^{-}\left(o,T_{h}\right)$ is the loss state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}^{-}\left(o,T_{h}\right)\subseteq S_{o}\left(o,T_{h}\right)$ and not being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$, and the state $S_{o}^{+}\left(o,T_{h}\right)$ is the superposed state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}\left(o,T_{h}\right){\cap S}_{o}^{+}\left(o,T_{h}\right)=\varnothing$ and being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$.
+The enabling mapping is then defined as:
+$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows\widetilde{S_{c}}\left(c,T_{m}\right)$ (3.7)
+which is the noisy information of $I$, while $I$ is its own noiseless information.
+Thus, the noisy information $\widetilde{I}$ is also information.
+It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same.
+Fig.1(b) illustrates the structural characteristics of noisy information.
+In fact, Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario (Shannon, 1948).
+In this case, he assumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
+$E=f(S,N)$ (3.8)
+where $f$ is the distortion function of the channel to the transmitted signal.
+According to OIT's view, the signal $S$ here is the ontological state of information $I$, $S_{Is}\left(Is,T_{h}\right)$.
+The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{Is}^{-}\left(Is,T_{h}\right)$ from the source state $S_{Is}\left(Is,T_{h}\right)$ and adding another state $S_{Is}^{+}\left(Is,T_{h}\right)$.
+The output $E$ is then the channel's output state $\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output.
+Thus,
+$\widetilde{I}:\widetilde{S_{Is}}\left(Is,T_{h}\right)\rightrightarrows\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.9)
+is the noisy information of
+$I:S_{Is}\left(Is,T_{h}\right)\rightrightarrows S_{Ic}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.10)
 […]
-Thus the state set
-$S(M,T)=$ $\{\varphi_{ij}(t_{i}),\varphi_{I}(t_{j}),\varphi_{O}(t_{k}),\varphi_{S}(t_{j}),\varphi_{A}(t_{j}),$ (2.8)
-$i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}$
-describes $M$’s inputs, outputs, and state-transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
-It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits atuomatic transtion behavior. Hence we introduce :
-Definition 4 (Finite Automaton State): If a state set $S(X,T)$ expresses a finite automaton, then $S(X,T)$ is called a finite automaton state.
-All practical information systems, including computing systems, are finite automata in the mathematical sense.
-Therefore, the concept of finite automaton state is of great significance in the study of machine learning.
-2.4 Definition of Information
-As shown inXu2024 , the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$.
-Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
-This leads to the general definition of information:
-Definition 5 (Mathematical Definition of Information): Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain.
-The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets.
-Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$.The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
-It is important to note that both $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are state representations that satisfy Postulate 5.
-According to Theorem 1, each can be interpreted as a set of well-formed formulas within some formal system.
-However, the corresponding formal systems for the ontology and the carrier need not be identical.
-This distinction reflects one of the essential characteristics of information: its ability to establish connections between disparate entities in the world.
-Definition 5 does not require that the ontological states of information be mapped one-to-one onto the carrier states.
-Of course, in practical scenarios, one-to-one mappings often do exist, which can simplify the analysis considerably.
-Definition 6 (Recoverable Information): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ still denote a bijection from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, i.e., $I$ is invertible.
-Thus, there exists an inverse mapping $I^{-1}$ such that $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$.
-In this case, we say that information $I$ is recoverable, and $S_{o}(o,T_{h})$ is the recoverable state set of information $I$.
-In particular, owing to the importance of information in information systems, we propose:
-Definition 7 (Information in an Information System): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ be a recoverable information.
-If $S_{o}(o,T_{h})$ is a finite automaton state, then $c$ is called an information system, and $I$ is called information in the information system $c$.
-For recoverable information, we also use $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$ or $I^{-1}:S_{c}(c,T_{m})\leftleftarrows S_{o}(o,T_{h})$ to maintain consistency in the direction of enabling.
-Since, as recoverable information, the carrier losslessly carries the complete state of the ontology, this naturally represents an ideal scenario in information studies.
-If $I$ is not a bijection, we can always partition elements of $S_{o}(o,T_{h})$ according to the same image equivalence classes, and then establish an enabling mapping from these equivalence classes to $S_{c}(c,T_{m})$ by means of equivalence class reduction, thereby ensuring recoverability.
-Specifically:
-Proposition 1 (Recoverable Reduction of Information): For $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $R$ be an equivalence relation on $S_{o}(o,T_{h})$ defined by $xRy\leftrightarrow I(x)=I(y)$.
-Then, there exists a unique enabling mapping $\hat{I}:S_{o}(o,T_{h})/R\rightrightarrows S_{c}(c,T_{m})$, and $\hat{I}$ is a bijection, i.e., $\hat{I}$ is recoverable.
-In this case, $\hat{I}$ is called the recoverable reduction of information $I$.
-When $I$ is already recoverable, $S_{o}(o,T_{h})/R=S_{o}(o,T_{h})$, and $\hat{I}=I$.
-Recoverable information and the recoverable reduction of information can provide significant assistance in simplifying problems.
-Reference Shannon1948 is generated as a foundational work in information theory.
-The paper proposed that $H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information.
-The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively.
-The channel $I_{c}$ is the medium through which information is transmitted.
-After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $1\leq i\leq n,\tau>0$).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$.
-Then,
-$S_{I_{s}}(I_{s},T_{h})=IsEvent^{3}(I_{s},T_{h},e_{i})$ (2.9)
-and
-$S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))=IsCode^{3}(I_{c},(T_{h},T_{h}+\tau),c_{i})$ (2.10)
-are the states of the source $I_{s}$ at time $T_{b}$ and the state of the channel $I_{c}$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $1\leq i\leq n)$. Furthermore,
-$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.11)
-represents the functional enabling mapping from the source state $S_{I_{s}}(I_{s},T_{h})$ to the channel state $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$.From equation (2.10), we can recover the event $e_{i}$ in equation (2.9) using the code $c_{i}$, thus $I$ is still recoverable information.
-Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{I_{s}}(I_{s},T_{h})$ and the complex implementation mechanisms of $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system.
-The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
-2.5 Transmissibility of Information
-Xu2014 ; xu2015 discuss transmissibility using composite mappings. Building on Definitions 1 and 5, we can characterize the fundamentals of information transmission more rigorously:
-Theorem 3 (Transmissibility of Information) If $I_{1}=\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\rangle$ is information from an ontology $o$ to a carrier $c_{1}$, and $I_{2}=\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is information from carrier $c_{1}$ to carrier $c_{2}$, then the composition $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is also information, which we call transmission of information from $I_{1}$ via carrier $c_{1}$ to carrier $c_{2}$.
-Proof: According to Definition 1, Definition 5, and the given conditions, for any $S_{o}(x,t)\in S_{o}(o,T_{h})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$.
-Furthermore, there exists $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$.
-We define $I_{3}(S_{o}(x,t))=S_{c_{2}}(x_{2},t_{2})$.
-Since the physical realization of $S_{c_{2}}(x_{2},t_{2})$ depends on the existence of $S_{c_{1}}(x_{1},t_{1})$, and the existence of $S_{c_{1}}(x_{1},t_{1})$ in turn depends on $S_{o}(x,t)$, it follows by transitivity of causality that the realization of $S_{c_{2}}(x_{2},t_{2})$ ultimately depends on $S_{o}(x,t)$.
-Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$.
-Thus, $I_{3}(S_{o}(x,t))=I_{2}(I_{1}(S_{o}(x,t)))\underline{\triangleright}I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$.
-Therefore, $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 5.
-In essence, this is the information $I_{1}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$.
-The theorem is thus proved.
-Transmission is an important way of information flow.
-Communication, processing, input, and output are all concrete manifestations of information transmission.
-In many cases, information undergoes more complex flows through chain-like transmission mechanisms.
-From Theorem 3, we can readily derive the following:
-Corollary 1 (Information Chain) Let $\{I_{i}=\langle o,T_{h},S_{o},c_{i},T_{m_{i}},S_{c_{i}}\rangle|i=1,\ldots,n\}$ be a sequence of information mappings, where for any $i>1$, it holds that $c_{i-1}=o$, $T_{h}=T_{m_{i-1}}$, and $S_{o}(o,T_{h})=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$.
+In this way, we use the definition of noisy information to explain Shannon's model of noisy channels in (Shannon, 1948), demonstrating that OIT is applicable to the complex scenarios of information research under noisy conditions.
+3.3 Information Chain
+Reference (Xu et al., 2014) discusses the transmissibility of information through composite mappings.
+Based on Definitions 1 and 2, the fundamental laws governing information transmission can be characterized more rigorously as follows:
+Lemma 1 (Transmissibility of Information): Let $I_{0}=\left\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\right\rangle$ and $I_{1}=\left\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ be two pieces of information.
+Then, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is also information.
+We refer to this as the information $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$ via $c_{1}$.
+Proof: According to Definitions 1 and 2, as well as the given conditions, for any $S_{o}(x,t)\in S_{o}\left(o,T_{h}\right)$, there exists $S_{c_{1}}\left(x_{1},t_{1}\right)\in S_{c_{1}}\left(c_{1},T_{m_{1}}\right)$such that $I_{0}\left(S_{o}(x,t)\right)\trianglerighteq S_{c_{1}}\left(x_{1},t_{1}\right)$.
+Furthermore, there exists $S_{c_{2}}\left(x_{2},t_{2}\right)\in S_{c_{2}}\left(c_{2},T_{m_{2}}\right)$ such that $I_{1}\left(S_{c_{1}}\left(x_{1},t_{1}\right)\right)\trianglerighteq S_{c_{2}}\left(x_{2},t_{2}\right)$.
+Define $I_{2}\left(S_{o}(x,t)\right)=S_{c_{2}}\left(x_{2},t_{2}\right)$.
+Since the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ is contingent upon the existence of $S_{c_{1}}\left(x_{1},t_{1}\right)$, and the physical realization of $S_{c_{1}}\left(x_{1},t_{1}\right)$ in turn depends on $S_{o}(x,t)$, by causal induction, the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ ultimately depends on $S_{o}(x,t)$.
+Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{c_{1}}(x_{1},t_{1}))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{0}(S_{o}(x,t))\trianglerighteq S_{c_{1}}(x_{1},t_{1})$.
+Thus, $I_{2}(S_{o}(x,t))={I_{1}(I}_{0}(S_{o}(x,t)))\trianglerighteq I_{1}(S_{c_{1}}\left(x_{1},t_{1}\right))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$.
+Therefore, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 2.
+In essence, this is the information $I_{0}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$.
+The lemma is thus proved.
+Corollary 2 (Information Chain) Let $\left\{I_{i}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{i},T_{m_{i}},S_{c_{i}}\right\rangle\middle|i=1,\cdots,n\right\}$ be a sequence of information mappings, where for any $i>1$, it holds that $o_{i}=c_{i-1}$, $T_{h_{i}}=T_{m_{i-1}}$, and $S_{o_{i}}\left(o_{i},T_{h_{i}}\right)=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$.
 In this case, we say that these mappings constitute an information chain.
-Moreover, for any $1\leq i\leq j\leq n$, $I_{ij}=\langle o,T_{h},S_{o},c_{j},T_{m_{j}},S_{c_{j}}\rangle$ also qualifies as information according to Definition 5.
-2.6 Noisy Information
-Many researchers question the significance of recoverable information, as in the real world, information is often not perfectly recoverable.
-This is because the reflection of information is often influenced by noise or other external factors.
-Therefore, the following concept is introduced:
-Definition 8 (Noisy Information): For information $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $\widetilde{S_{o}}(o,T_{h})$ be another state of the same carrier $c$ at time $T_{m}$ and $\widetilde{S_{c}}(c,T_{m})$ be another state of the same carrier $c$ at time $T_{m}$,
-[…]
-where the state $S^{-}_{o}(o,T_{h})$ is the loss state of $S_{o}(o,T_{h})$, satisfying $S^{-}_{o}(o,T_{h})\subseteq$ $S_{o}(o,T_{h})$ and not being enabled to map to $\widetilde{S_{c}}(c,T_{m})$, and the state $S^{+}_{o}(o,T_{h})$ is the superposed state of $S_{o}(o,T_{h})$, satisfying $S_{o}(o,T_{h})\cap S_{o}^{+}(o,T_{h})=\emptyset$ and being enabled to map to $\widetilde{S_{c}}(c,T_{m})$.
-The enabling mapping is then defined as:
-$\tilde{I}:\widetilde{S_{o}}(o,T_{h})\rightrightarrows\widetilde{S_{c}}(c,T_{m})$ (2.13)
-which is the noisy information of $I$, while $I$ is its own noiseless information.
-Thus, the noisy information $\tilde{I}$ is also information.
-It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same.
-Further analysis of equation (2.12) shows that information $I$ and $\tilde{I}$ are completely asymmetric, and hence:
-Proposition 2 (Symmetry of Noisy Information): If $\tilde{I}$ is the noisy information of $I$, then $I$ is also the noisy information of $\tilde{I}$.
-In fact, for both information $I$ and its noisy information $\tilde{I}$, it is easy to verify its recoverability or recoverable reducibility from the perspective of enabling mapping.
-However, in real scenarios, it is often difficult to obtain noise-free information $I$, and one can only study using a noisy version $\tilde{I}$.
-Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information.
-But under the condition where the noiseless information IIcannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$.
-This is the practical significance of introducing the concept of noisy information.
-Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenarioShannon1948 .
-In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
-$E=f(S,N)$ (2.14)
-where $f$ is the distortion function of the input for the transmitted signal.
-According to OIT’s view, the signal $S$ here is the ontological state of information $I$, $S_{I_{s}}(I_{s},T_{h})$.
-The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{I_{s}}^{-}(I_{s},T_{h})$ from the source state $S_{I_{s}}(I_{s},T_{h})$ and adding another state $S_{I_{s}}^{+}(I_{s},T_{h})$.
-The output $E$ is then the channel’s output state $\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output.
-Thus,
-$\tilde{I}:\widetilde{S_{I_{s}}}(I_{s},T_{h})\rightrightarrows\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$ (2.15)
-is the noisy information of
-$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.16)
-[…]
-In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels inShannon1948 , demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
+Moreover, for any 1 $\leq i\leq j\leq n$, $I_{ij}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{j},T_{m_{j}},S_{c_{j}}\right\rangle$ also qualifies as information according to Definition 2.
+Transmission constitutes a fundamental mode of information flow and application.
+Communication, processing, input, and output are all concrete realizations of information transmission.
+In many cases, information undergoes more complex flows through chain-like transmission processes.
+Fig.1(c) intuitively illustrates such a process.
+For any pair of states within this process, the latter can be regarded as the result of the enabling mapping from the former, and this enabling mapping can be used to reasonably explain the causation of the latter state.
+This provides a theoretical foundation for the interpretability of the model.
 4 Meta-Framework for Machine Learning Theory (MLT-MF)
@@
 The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system.
-Therefore, by employing the concepts of finite automaton state and information models established in Section 2, the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
+Therefore, by employing the concepts of information model and finite automaton state (Appendix B), the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
 4.1 Construction of Initial Information
-Suppose the model under study is denoted as M, and let $T=\{t_{i}|i=1,\ldots,n\}$ be the set of time points at which M, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order.
+Suppose the model under study is denoted as $M$, and let $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ be the set of time points at which $M$, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order.
 According to the perspective of OIT, the creation of $M$ is essentially the realization of information
-$I_{o}=\langle M_{D},T_{o_{h}},S_{M_{D}},M,t_{1},S_{M}\rangle$ (3.1)
-where the ontology $M_{D}$ of $I_{o}$ is the designer’s conceptualization of M; $T_{o_{h}}$ is the time at which the designer conceives M, and $S_{M_{D}}(M_{D},T_{o_{h}})$ is the initial finite automaton state that must be realized at this conceptual stage.
-The carrier M of $I_{o}$ is the model itself; the reflection time $t_{1}$ is when M is constructed according to $S_{M_{D}}(M_{D},T_{o_{h}})$; and $S_{M}(M,t_{1})$ is the formalization of the actual physical state of M upon its completion.
-Thus, $\sup(T_{o_{h}})<\sup(t_{1})$.
-According to Theorem 2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
-$\varphi_{o_{1}}=State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ (3.2)
-denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},T_{o_{h}})$ as envisioned by the designer are assigned to $M_{D}$.
-$\varphi_{o_{2}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
-$Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
-$(Eq^{3}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land State^{3}(M_{D},t_{2},\Phi^{2}(M_{D},t_{2}))$ (3.3)
-expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},T_{o_{h}})$, and the input carrier state $S_{M_{I}}(M_{I},t_{1})$ is provided and satisfies the learnability condition determined by the binary predicate $Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, $M_{D}$ can optimize its internal structures, parameters, and algorithmsto $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, assigning this to $M_{D}$.
-This demonstrates the model’s capacity for learning and optimization, where $M_{I}$ denotes the input device of M.
-$\varphi_{o_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
-$Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$
-$(Eq^{3}(S_{M_{O}}(M_{O},t_{2}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land Output^{3}(M,t_{2},S_{M_{O}}(M_{O},t_{2})))$ (3.4)
-indicates that, at time $t_{1}$, if $S_{M_{I}}(M_{I},t_{1})$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate $Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))$, then $M_{D}$ can process $S_{M_{I}}(M_{I},t_{1})$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$ to yield the output carrier state $S_{M_{O}}(M_{O},t_{2})$, which is then output at time $t_{2}$.
+$Io=\left\langle M_{D},{To}_{h},S_{M_{D}},M,t_{1},S_{M}\right\rangle$ (4.1)
+where the ontology $M_{D}$ of $Io$ is the designer’s conceptualization of $M$, ${To}_{h}$ is the time at which the designer conceives $M$, and $S_{M_{D}}\left(M_{D},{To}_{h}\right)$ is the initial finite automaton state that must be realized at this conceptual stage.
+The carrier $M$ of $Io$ is the model itself; the reflection time $t_{1}$ is when $M$ is constructed according to $S_{M_{D}}\left(M_{D},{To}_{h}\right)$; and $S_{M}\left(M,t_{1}\right)$ is the formalization of the actual physical state of $M$ upon its completion.
+Thus, $\sup\left({To}_{h}\right)<sup(t_{1})$.
+According to Lemma B.2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
+${\varphi o}_{1}={State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ (4.2)
+denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},{To}_{h})$ as envisioned by the designer are assigned to $M_{D}$.
+$\begin{split}{\varphi o}_{2}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Learnable}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({Eq}^{2}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land{State}^{3}(M_{D},t_{2},\Phi^{2}\left(M_{D},t_{2}\right))\end{split}$ (4.3)
+expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},{To}_{h})$, and the input carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is provided and satisfies the learnability condition determined by the binary predicate ${Learnable}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))$, $M_{D}$ can optimize its internal structures, parameters, and algorithms to $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))$, assigning this to $M_{D}$.
+This demonstrates the model's capacity for learning and optimization, where $M_{I}$ denotes the input device of $M$.
+$\begin{split}{\varphi o}_{3}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({{Eq}^{2}\left(S_{M_{O}}\left(M_{O},t_{2}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\right)\land Output}^{3}(M,t_{2},S_{M_{O}}\left(M_{O},t_{2}\right)))\end{split}$ (4.4)
+indicates that, at time $t_{1}$, if $S_{M_{I}}\left(M_{I},t_{1}\right)$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate ${Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$, then $M_{D}$ can process $S_{M_{I}}\left(M_{I},t_{1}\right)$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$ to yield the output carrier state $S_{M_{O}}\left(M_{O},t_{2}\right)$, which is then output at time $t_{2}$.
 This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of $M$.
 Thus,
-$S_{M_{D}}(M_{D},T_{o_{h}})=\{\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}\}$ (3.5)
+$S_{M_{D}}\left(M_{D},{To}_{h}\right)=\left\{{\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}\right\}$ (4.5)
 constitutes the initial state of $M$ as conceived by the designer.
-It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},T_{o_{h}})$, and $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ all represent sets, $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas.
+It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas.
 Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states.
 Therefore, subsequent discussions will encompass both cases without further elaboration.
-$S_{M_{D}}(M_{D},T_{o_{h}})$ may represent the designer’s idea, a design draft on paper, or an electronic document.
-Through the information $I_{O}$, these are enabled and mapped into the physically operational state of the model M:
-$S_{M}(M,t_{1})=\{\psi_{o_{1}},\psi_{o_{2}},\psi_{o_{3}}\}$ (3.6)
-where $\psi_{o_{1}}\underline{\triangleleft}I_{o}(\varphi_{o_{1}})$, $\psi_{o_{2}}\underline{\triangleleft}I_{o}(\varphi_{o_{2}})$, and $\psi_{o_{3}}\underline{\triangleleft}I_{o}(\varphi_{o_{3}})$ are the enabling mappings of $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$, respectively.
+$S_{M_{D}}\left(M_{D},{To}_{h}\right)$ may represent the designer’s idea, a design draft on paper, or an electronic document.
+Through the information $Io$, these are enabled and mapped into the physically operational state of the model $M$:
+$S_{M}\left(M,t_{1}\right)=\left\{{\psi o}_{1},{\psi o}_{2},{\psi o}_{3}\right\}$ (4.6)
+where ${\psi o}_{1}\trianglelefteq Io({\varphi o}_{1})$, ${\psi o}_{2}\trianglelefteq Io({\varphi o}_{2})$, and ${\psi o}_{3}\trianglelefteq Io({\varphi o}_{3})$ are the enabling mappings of ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$, respectively.
 This process establishes $M$ as a typical finite automaton and an operational information system.
@@
 According to the general formal definition of information, the formal expression for the realization of training information is straightforward:
-$I_{t}=\langle ot,T_{th},S_{ot},M_{I},t_{1},S_{M_{I}}\rangle$ (3.7)
-$S_{ot}(ot,T_{th})=\{\varphi_{t}|\varphi_{t}\text{ is a WFF about }ot\text{ and }T_{th}\}$ (3.8)
-$S_{M_{I}}(M_{I},t_{1})=\{\psi_{t}|\psi_{t}\text{ is a WFF about }M_{I}\text{ and }t_{1}\}$ (3.9)
-Here, the ontology $ot$ refers to the object containing knowledge or samples; $T_{th}$ is the time at which its state is formed; and $S_{ot}(ot,T_{th})$ is a set of business-oriented semantic assertions or facts about the ontology $M_{I}$ functions similarly to the hardware interface and API software of M, carrying the training information at time $t_{1}$.
-The carrier state $S_{M_{I}}(M_{I},t_{1})$ is the physical realization of $S_{ot}(ot,T_{th})$ via an enabling mapping, ensuring that the information can be machine-read and entered into M.
+$It=\left\langle ot,{Tt}_{h},S_{ot},M_{I},t_{1},S_{M_{I}}\right\rangle$ (4.7)
+$S_{ot}(ot,{Tt}_{h})=\left\{\varphi t\middle|\varphi t\ is\ a\ WFF\ about\ ot\ and\ {Tt}_{h}\right\}$ (4.8)
+$S_{M_{I}}\left(M_{I},t_{1}\right)=\left\{\psi t\middle|\psi t\ is\ a\ WFF\ about\ M_{i}\ and\ t_{1}\right\}$ (4.9)
+Here, the ontology $ot$ refers to the object containing knowledge or samples; ${Tt}_{h}$ is the time at which its state is formed; and $S_{ot}(ot,{Tt}_{h})$ is a set of business-oriented semantic assertions or facts about the ontology.
+$M_{I}$ functions similarly to the hardware interface and API software of $M$, carrying the training information at time $t_{1}$.
+The carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is the physical realization of $S_{ot}(ot,{Tt}_{h})$ via an enabling mapping, ensuring that the information can be machine-read and entered into $M$.
 4.3 Formal Modeling and Implementation of Model Learning
-Model learning for M involves fusing its original information $I_{o}$ with the training information $I_{t}$, resulting in updated information:
-$I_{u}=\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\rangle$ (3.10)
+Model learning for $M$ involves fusing its original information $Io$ with the training information $It$, resulting in updated information:
+$Iu=\left\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\right\rangle$ (4.10)
 Here, the ontology is the hybrid $M\cup M_{I}$, encompassing both the model and its interface; $t_{1}\cup t_{2}$ marks the time of the first learning event; and $S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})$ formally represents the hybrid state during learning:
-$S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})=\{\varphi_{u_{1}},\varphi_{u_{2}},\varphi_{u_{3}}\}$ (3.11)
-$\varphi_{u_{1}}=Eq^{3}(\Phi^{2}(M,t_{2}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.12)
-indicates that the input state $S_{M_{I}}(M_{I},t_{1})$, which was learned by M at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of M at $t_{2}$.
-$\varphi_{u_{2}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
-$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
-$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.13)
+$S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})=\left\{{\varphi u}_{1},{\varphi u}_{2},{\varphi u}_{3}\right\}$ (4.11)
+${\varphi u}_{1}={Eq}^{2}(\Phi^{2}\left(M,t_{2}\right),f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right)))\land{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)$ (4.12)
+indicates that the input state $S_{M_{I}}\left(M_{I},t_{1}\right)$, which was learned by $M$ at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of $M$ at $t_{2}$.
+$\begin{split}{\varphi u}_{2}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))))\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.13)
 shows that after the first learning step, $M$ remains capable of further learning and optimization.
-$\varphi_{u_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
-$Processable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
-$(Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.14)
-indicates that when the training information $I_{t}$ is processable by M, it can also produce output information.
+$\begin{split}{\varphi u}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))))\end{split}$ (4.14)
+indicates that when the training information $It$ is processable by $M$, it can also produce output information.
 This state is then enabled and mapped to the physical state of the carrier $M$ at time $t_{2}$:
-$S_{M}(M,t_{2})=\{\psi_{u_{1}},\psi_{u_{2}},\psi_{u_{3}}\}$ (3.15)
-3.4 Information Processing and Learning Optimization in Models
+$S_{M}\left(M,t_{2}\right)=\left\{{\psi u}_{1},{\psi u}_{2},{\psi u}_{3}\right\}$ (4.15)
+4.4 User Input Information for the Model
 User input information is formally given by:
-$I_{q}=\langle oq,T_{qh},S_{oq},M_{I},t_{2},S_{M_{I}}\rangle$ (3.16)
-$S_{oq}(oq,T_{qh})=\{\varphi_{q}|\varphi_{q}\text{ is a WFF about }oq\text{ and }T_{qh}\}$ (3.17)
-$S_{M_{I}}(M_{I},t_{2})=\{\psi_{q}|\psi_{q}\text{ is a WFF about }M_{I}\text{ and }t_{2}\}$ (3.18)
-Here, the ontology $oq$ represents the user’s intention or device, with $T_{qh}$ denoting the time at which the ontological state is formed.
-The state $S_{oq}(oq,T_{qh})$ is the formal representation of the user’s needs.
-The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,T_{qh})$ onto the model’s interface device $M_{I}$.
+$Iq=\left\langle oq,{Tq}_{h},S_{oq},M_{I},t_{2},S_{M_{I}}\right\rangle$ (4.16)
+$S_{oq}(oq,{Tq}_{h})=\left\{\varphi q\middle|\varphi q\ is\ a\ WFF\ about\ oq\ and\ {Tq}_{h}\right\}$ (4.17)
+$S_{M_{I}}(M_{I},t_{2})=\left\{\psi q\middle|\psi q\ is\ a\ WFF\ about\ M_{I}\ and\ t_{2}\right\}$ (4.18)
+Here, the ontology $oq$ represents the user’s intention or device, with ${Tq}_{h}$ denoting the time at which the ontological state is formed.
+The state $S_{oq}(oq,{Tq}_{h})$ is the formal representation of the user’s needs.
+The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,{Tq}_{h})$ onto the model’s interface device $M_{I}$.
 4.5 Information Processing, Output, and Online Learning of the Model
 When the model $M$ processes input and produces output, the information it carries can be represented as:
-$I_{v}=\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\rangle$ (3.19)
+$Iv=\left\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\right\rangle$ (4.19)
 where:
-$S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3})=\{\varphi_{v_{1}},\varphi_{v_{2}},\varphi_{v_{3}}\}$ (3.20)
-$\varphi_{v_{1}}=State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.21)
+$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3}\right)=\left\{{\varphi v}_{1},{\varphi v}_{2},{\varphi v}_{3}\right\}$ (4.20)
+${\varphi v}_{1}={State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)$ (4.21)
 indicates that $\Phi^{2}(M,t_{2})$ has been assigned to $M$ at time $t_{2}$.
-$\varphi_{v_{2}}=$ $Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land$
-$Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.22)
-indicates that when the input state $S_{M_{I}}(M_{I},t_{2})$ is received at $t_{2}$, M produces the output state $S_{M_{O}}(M_{O},t_{3})$ at $t_{3}$.
-$\varphi_{v_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
-$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
-$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.23)
+$\begin{split}{\varphi v}_{2}=&{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\land{Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\\ &\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))\end{split}$ (4.22)
+indicates that when the input state $S_{M_{I}}\left(M_{I},t_{2}\right)$ is received at $t_{2}$, $M$ produces the output state $S_{M_{O}}\left(M_{O},t_{3}\right)$ at $t_{3}$.
+$\begin{split}{\varphi v}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))))\\ &\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.23)
 expresses that if the input information can be learned by $M$, then $M$ can achieve online learning while processing and outputting information.
-The physical realization of this ontological state through an enabling mapping is:
-$S_{M}(M,t_{3})=\{\psi_{v_{1}},\psi_{v_{2}},\psi_{v_{3}}\}$ (3.24)
+The physical realization of this state through an enabling mapping is:
+$S_{M}(M,t_{3})=\left\{{\psi v}_{1},{\psi v}_{2},{\psi v}_{3}\right\}$ (4.24)
 The output information is given by:
-$I_{r}=\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\rangle$ (3.25)
+$Ir=\left\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\right\rangle$ (4.25)
 where the ontological state is:
-$S_{M\cup M_{I}}(M\cup M_{I},t_{2})=\{\varphi_{r_{1}}\}$ (3.26)
+$S_{M\cup M_{I}}\left(M\cup M_{I},t_{2}\right)=\left\{{\varphi r}_{1}\right\}$ (4.26)
 with
-$\varphi_{r_{1}}=Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (3.27)
+${\varphi r}_{1}={Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (4.27)
 The formal expression of the physical state of the model’s output device $M_{O}$ at time $t_{3}$ is:
-$S_{M_{O}}(M_{O},t_{3})=\{\psi_{r_{1}}\}$ (3.28)
+$S_{M_{O}}\left(M_{O},t_{3}\right)=\left\{{\psi r}_{1}\right\}$ (4.28)
 The model can repeat the processes described in Sections 3.2 to 3.5, iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
 4.6 Multi-Stage Information Causal Chains
-From a systems engineering perspective, the realization of information in machine learning involves more than just five information mappings.
+From systems engineering perspective, the realization of information in machine learning involves more than just five information mappings.
 In practice, stages such as requirements analysis, design and development, and system integration are also necessary.
 Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 3.3, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation.
-In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Figure.1).
-Figure 1: Machine Learning Causality Chains with Multiple Information Mappings
-As shown in Figure 1, each of these four stages encompasses four layers: conceptual ideas, mind maps, program code, and physical models.
-Typically, content within the same layer shares the same semantic logic and can thus be uniformly represented by a formal system.
-Furthermore, due to the transmissive nature of information, the innermost physical model can also be regarded as the result of enabling mappings from outer conceptual ideas or mind maps.
-This illustrates that the entire system comprises multiple stages, with each stage involving various combinations of information causal chains.
-This framework, based on the realization of information and finite automaton states, provides a formalized representation and a comprehensive, causally progressive structure for each phase in the machine learning process.
-It transcends specific algorithms or models, offering a unified and universal framework, MLT-MF.
-In the appendix, we also present an example of its application to feedforward neural networks, which validates the correctness and applicability of the framework and supports the verification of the related theorems.
-4 Several Important Theorems and Proofs
+In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Fig.2).
+Fig.2 presents a comprehensive framework of formal expressions and causal information progression for each stage of the machine learning process, from the perspectives of OIT and finite automaton state implementation.
+This framework transcends specific algorithms or models, offering a novel, unified, and general framework—MLT-MF—for addressing several key issues in the field of machine learning.
+In Appendix C, we further provide an example of its application to feedforward neural networks, thereby testing the correctness and applicability of the framework and supporting the verification of related theorems.
+Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning.
+In machine learning, the four distinct stages—model construction, learning and training, problem input, and answer output—each encompass four levels: conceptual ideas, mind maps, program code, and physical models.
+Generally, the content within the same level shares a common semantic logic and can thus be represented using a unified formal system.
+Furthermore, owing to the transmissibility of information, the innermost physical model can be regarded as the result of an enabling mapping from the outer conceptual ideas or mind maps.
+The output stage can, in turn, be viewed as a reverse causal transmission from the physical model back to conceptual ideas.
+This demonstrates that the system comprises multiple stages, with each stage containing various combinatorial information causal chains.
+5 Several Important Definitions and Theorems
 The MLT-MF framework is applicable to a range of key issues in the field of artificial intelligence, offering many unique and practical theoretical principles and methodological guidelines.
 5.1 Theorem of Model Interpretability
-Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence.
-However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability Lipton2018 ; Doshi2017 ; Rubin2019 ; Miller2019 .
-In fact, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system.
-Furthermore, it should also be considered in relation to the specific inputs and outputs of the model.
-Based on MLT-MF, we define the following concepts:
-Definition 9 (Model Interpretability) Let the model M, as a finite automaton, operate over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, at which it undergoes assignment, input, output, and state transitions.
-The state of M at time $t_{i}$ is denoted as $S_{M}(M,t_{i})$, and the state of its input device $M_{I}$ at time $t_{i}$ is $S_{M_{I}}(M_{I},t_{i})$, while the state of the output device $M_{O}$ at time $t_{i+1}$ is $S_{M_{O}}(M_{O},t_{i+1})$.
-If there exists a formal system $\mathcal{L}$ along with three interpretation relations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$ respectively, such that
-$E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$ (4.1)
-making $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ the respective enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, then M is said to be interpretable in $\mathcal{L}$ with respect to output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$.
-If for all $i=1,\cdots,n-1$, the output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in $\mathcal{L}$, then M is called interpretable in $\mathcal{L}$.
-Here, $\text{Cn}(E_{M}\cup E_{M_{I}})$ denotes the logical closure of $E_{M}\cup E_{M_{I}}$.
-Equation 4.1 states that in the formal system $\mathcal{L}$, all formulas in $E_{M_{O}}$ are logical consequences of $E_{M}$ and $E_{M_{I}}$.
-Moreover, since $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are respectively enabled by $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, the formal system $\mathcal{L}$ and its three interpretations fully capture the logical and causal relationship whereby M produces output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$.
+According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model (Glanois et al., 2024), the following concepts are established based on the MLT-MF framework:
+Definition 4 (Model Interpretability) Let the model $M$, as a finite automaton, operate over an ordered set of times $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$, at which it undergoes assignment, input, output, and state transitions.
+The state of $M$ at time $t_{i}$ is denoted as $S_{M}\left(M,t_{i}\right)$, and the state of its input device $M_{I}$ at time $t_{i}$ is $S_{M_{I}}\left(M_{I},t_{i}\right)$, while the state of the output device $M_{O}$ at time $t_{i+1}$ is $S_{M_{O}}\left(M_{O},t_{i+1}\right)$.
+If there exists a formal system $\mathcal{L}$ along with three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$ respectively, such that
+$E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$ (5.1)
+making $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ the respective enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, then $M$ is said to be interpretable in $\mathcal{L}$ with respect to output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ under input $S_{M_{I}}\left(M_{I},t_{i}\right)$.
+If for all $i=1,\cdots,n-1$, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ under input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in $\mathcal{L}$, then $M$ is called interpretable in $\mathcal{L}$.
+Here, $Cn(E_{M}\cup E_{M_{I}})$ denotes the logical closure of $E_{M}\cup E_{M_{I}}$.
+Equation (5.1) states that in the formal system $\mathcal{L}$, all formulas in $E_{M_{O}}$ are logical consequences of $E_{M}$ and $E_{M_{I}}$.Moreover, since $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are respectively enabled by $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, the formal system $\mathcal{L}$ and its three interpretations fully capture the logical and causal relationship whereby $M$ produces output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ under input $S_{M_{I}}\left(M_{I},t_{i}\right)$.
 This, in essence, offers a concise understanding of model interpretability.
-Theorem 4 (Model Interpretability and Information Existence) Let the model M be viewed as a finite automaton operating over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, during which it undergoes assignment, input, output, and state transitions.
-(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}(M_{I},t_{i})$ is that there exist $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$ and $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$ representing the information carried by M and its input device $M_{I}$ at time $t_{i}$ respectively, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$.
-The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$ must hold.
+Theorem 1 (Model Interpretability and Information Existence) Let the model $M$ be viewed as a finite automaton operating over an ordered set of times $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$, during which it undergoes assignment, input, output, and state transitions.
+(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is that there exist $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$ and $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$ representing the information carried by $M$ and its input device $M_{I}$ at time $t_{i}$, respectively, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$.
+The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$ must hold.
 (2) A necessary and sufficient condition for $M$ to be interpretable in $\mathcal{L}$ is that, for all $i=1,\cdots,n-1$, there exist information objects $Iu(t_{i})$, $Iq(t_{i})$, and $Ir(t_{i+1})$ whose ontological states are all interpretations in $\mathcal{L}$, and such that the ontological state of $Ir(t_{i+1})$ is contained in the logical closure of the union of the ontological states of $Iu(t_{i})$ and $Iq(t_{i})$.
-Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$.
-According to Definition 9, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$, and $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
-Let $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$.
-Then, according to Definition 6,we find there exist information objects $Iu(t_{i})=$ $\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, similarly , then $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),$ $Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$.
+Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$.
+According to Definition 4, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$, and $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
+Let
+$S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$ $=E_{M},$
+$S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$ $=E_{M_{I}},$
+$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$ $=E_{M_{O}}$
+Then, according to Definition 6, there exist information objects
+$Iu(t_{i})$ $=\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle,$
+$Iq(t_{i})$ $=\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle,$
+$Ir(t_{i+1})$ $=\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$
+such that
+$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
+$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$
 Thus, the necessity of the first part of the theorem is established.
-On the other hand, for any $1\leq i\leq n-1$, suppose there exist some information objects $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that the ontological states $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$.
-Then, according to Theorem 1, $E_{M}=S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times Tu_{h}(t_{i})$, $oq(t_{i})\times Tq_{h}(t_{i})$, and $or(t_{i+1})\times Tr_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$.
-At the same time, $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
-By Definition 9, the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$.
+On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$,
+$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
+$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$
+Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$.
+At the same time, $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
+By Definition 4, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$.
 Therefore, the sufficiency of the first part of the theorem is established.
-For all $i=1,\ldots,n-1$ , by combining Definition 9 with the first part of the theorem, the second part of the theorem follows directly. This completes the proof.
+For all $i=1,\cdots,n-1$, by combining Definition 4 with the first part of the theorem, the second part of the theorem follows directly. This completes the proof.
+Definition 4 provides a rigorous mathematical definition of model interpretability.
+Furthermore, Theorem 1 demonstrates that, throughout the processes of model construction and application, as long as the ontological states of the model and its input and output information can be established within the same formal system, it is possible to reveal the causal relationships between model inputs and outputs within this system, thereby achieving model interpretability.
+It should also be noted that formal systems vary greatly across different academic disciplines, and the myriad entities in the world can be expressed through information in the forms of formulas, code, text, graphics, or even video.
+Moreover, the same state can be represented in different information modalities according to varying user requirements.
+Therefore, Definition 4 and Theorem 1 theoretically establish a foundation for the flexible use of multiple information modalities and for providing model interpretability tailored to user needs.
 5.2 Theorem of Ethical Safety Assurance
-Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence mathsea2002 ; Doshi2017 ; Rubin2019 ; Miller2019 , and to date, there is still no universally accepted and practically actionable rigorous definition Russell2015 ; Jobin2019 ; Corrêa2023 .
-Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions.
-To achieve this, it is first necessary to specify appropriate well-formed formulas $Ec$ as ethical constraints—such as those governing fairness and privacy protection—thus establishing a formalized definition of ethical safety.
-Definition 10 (Ethical Safety): Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$.
-Let $Ec(\vec{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\vec{v}=(v_{1},\ldots,v_{n})$ are its free variables.
+To apply the MLT-MF framework, it is first necessary to identify the corresponding well-formed ethical constraint formulas $Ec$ in accordance with fundamental ethical standards such as fairness and privacy protection, thereby establishing a formalized expression of ethical security.
+Definition 5 (Ethical Safety) Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$.
+Let $Ec(\overrightarrow{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\overrightarrow{v}=(v_{1},\ldots,v_{n})$ are its free variables.
 For each substitutable $(x,t)\in X\times T$, define
-$Ec|_{(x,t)}=Ec(\vec{v}\mapsto\vec{\sigma}(x,t))$ (4.2)
-where $\vec{\sigma}:X\times T\to D^{n}$ is a value vector function such that $\sigma_{i}(x,t)\in D$ is the concrete value assigned to $v_{i}$ for $1\leq i\leq n$.
-If for all $\varphi\in\text{Cn}(S(X,T))$, and all $(x,t)\in X\times T$,
+$Ec|_{(x,t)}=Ec(\overrightarrow{v}\mapsto\overrightarrow{\sigma}(x,t))$ (5.2)
+where $\overrightarrow{\sigma}:X\times T\rightarrow D^{n}$ is a value vector function such that $\sigma_{i}(x,t)\in D$ is the concrete value assigned to $v_{i}$ for $1\leq i\leq n$.
+If for all $\varphi\in Cn(S(X,T))$, and all $(x,t)\in X\times T$,
 $\varphi\rightarrow\sim Ec|_{(x,t)}$ (5.3)
@@
 Furthermore, if a model M is capable of producing output $I$, and both the ontological and carrier states of $I$ are ethically safe with respect to $Ec$, then M is said to be ethically safe with respect to $Ec$.
-In the definition of ethical safety, we require not only that $S(X,T)$ contains no formulas that entail $\sim Ec|_{(x,t)}$, but also that its logical closure $\text{Cn}(S(X,T))$ contains no such formulas.
-This is because if $\text{Cn}(S(X,T))$ were to include formulas entailing $\sim Ec|_{(x,t)}$, then the conjunction of multiple formulas in $S(X,T)$ could still violate the ethical constraint $Ec$.
+In the definition of ethical safety, we require not only that $S(X,T)$ contains no formulas that entail $\sim Ec|_{(x,t)}$, but also that its logical closure $Cn(S(X,T))$ contains no such formulas.
+This is because if $Cn(S(X,T))$ were to include formulas entailing $\sim Ec|_{(x,t)}$, then the conjunction of multiple formulas in $S(X,T)$ could still violate the ethical constraint $Ec$.
 This would indicate that $S(X,T)$ implicitly allows situations that contravene $Ec$, which must be strictly avoided in a rigorous sense.
@@
 To this end, we propose:
-Theorem 5 (Ethical Safety Assurance): Suppose model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$.
-Let $S_{or}(or,Tr_{h})$ denote the ontological state of $Ir$.
-For any ethical constraint $Ec$ represented by a well-formed formula in $\mathcal{L}$, there exists a constructive algorithm for obtaining the largest output state subset $S_{or}^{EC}(or,Tr_{h})\subseteq S_{or}(or,Tr_{h})$ that is ethically safe with respect to $Ec$.
+Theorem 2 (Ethical Safety Assurance) Suppose model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$.
+Let $S_{or}\left(or,{Tr}_{h}\right)$ denote the ontological state of $Ir$.
+For any ethical constraint $Ec$ represented by a well-formed formula in $\mathcal{L}$, there exists a constructive algorithm for obtaining the largest output state subset $S_{or}^{Ec}\left(or,{Tr}_{h}\right)\subseteq S_{or}\left(or,{Tr}_{h}\right)$ that is ethically safe with respect to $Ec$.
 By introducing an ethical safeguard formula for $Ec$ into $Iu$, it is thereby possible to ensure that model M attains the maximal output state that is ethically safe with respect to $Ec$.
-Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, IqIq and $Ir$ respectively, the MLT-MF framework provides:
-$\varphi u=($ $\forall\Phi u)(\forall\Phi q)(\text{IsSubsetOf}^{2}(\Phi u,Sou(ou,Tu_{h}))\land\text{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,Tq_{h}))\land$
-$\text{Processable}^{2}(\Phi u,\Phi q))\to\exists\varphi r(\text{IsFormulaOf}^{2}(\varphi r,S_{or}(or,Tr_{h}))\land$
-$\text{Eq}^{2}(\varphi r,f_{process}^{2}(\Phi u,\Phi q)))$ (4.4)
+Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}\left(ou,{Tu}_{h}\right)$, $S_{oq}\left(oq,{Tq}_{h}\right)$ and $S_{or}\left(or,{Tr}_{h}\right)$ of $Iu$, Iq\ Iq and $Ir$ respectively, the MLT-MF framework provides:
+$\varphi_{u}\coloneqq$ $\bigl((\forall\,\Phi_{u})(\forall\,\Phi_{q})\bigl({IsSubsetOf}^{2}(\Phi_{u},S_{ou}(ou,{T}_{u,h}))$ (5.4)
+$\quad\land{IsSubsetOf}^{2}(\Phi_{q},S_{oq}(oq,{T}_{q,h}))$
+$\quad\land{Processable}^{2}(\Phi_{u},\Phi_{q})\bigr)\bigr)$
+$\rightarrow\bigl(\exists\,\varphi_{r}\bigl({IsFormulaOf}^{2}(\varphi_{r},S_{or}(or,{T}_{r,h}))$
+$\quad\land{Eq}^{2}(\varphi_{r},f_{\mathrm{process}}^{2}(\Phi_{u},\Phi_{q}))\bigr)\bigr).$
 Thus, the set
-$S_{or}(or,Tr_{h})=\{\varphi r|\varphi r=\text{Processable}^{2}(\Phi u,\Phi q),\Phi u\in Sou(ou,Tu_{h}),\Phi q\in S_{oq}(oq,Tq_{h})\}$ (4.5)
+$\begin{split}S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r\middle|\right.&\left.\varphi r={Processable}^{2}(\Phi u,\Phi q),\right.\\ &\left.\Phi u\in Sou\left(ou,{Tu}_{h}\right),\right.\\ &\left.\Phi q\in S_{oq}\left(oq,{Tq}_{h}\right)\right\}\end{split}$ (5.5)
 represents the theoretical output states of model M.
-Since $S_{or}(or,Tr_{h})$ is a finite set, we construct a finite hypergraph
-$H_{or}=(S_{or}(or,Tr_{h}),\mathcal{E})$ (4.6)
-where the vertex set is precisely $S_{or}(or,Tr_{h})$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
-(1) For any $\varphi r\in S_{or}(or,Tr_{h})$ and $(x,t)\in or\times Tr_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\{\varphi r\}$ to $\mathcal{E}$;
-(2) If there exists a minimal subset $W\subseteq S_{or}(or,Tr_{h})$ such that the conjunction of all formulas in $W$ entails $\sim Ec|_{(x,t)}$, then add the hyperedge $W$ to $\mathcal{E}$.
-Let $G$ be an independent set in $H_{or}$, i.e., a subset of vertices containing no hyperedge $e\in\mathcal{E}$. If there exists $\varphi\in\text{Cn}(G)$ such that
-$\varphi\to\sim Ec|_{(x,t)}$ (4.7)
-then there must exist a finite subset $\{\varphi r_{1},\cdots,\varphi r_{k}\}\subseteq G$ such that $\varphi r_{1}\land\cdots\land\varphi r_{k}=\varphi$, and hence $\varphi r_{1}\land\cdots\land\varphi r_{k}\to\sim Ec|_{(x,t)}$.
-Let $W=\{\varphi r_{1},\cdots,\varphi r_{k}\}$, so the conjunction of $W$ entails $\sim Ec|_{(x,t)}$.
-According to the hyperedge construction rules, if $|W|=1$, then $W$ itself directly violates $Ec$, and there must be a hyperedge $\{\varphi r_{1}\}$.
-But since $\{\varphi r_{1}\}\in G$, this contradicts the assumption that $G$ is an independent set.
-If $|W|\geq 2$, then by the compactness theorem, the logical entailment can be captured by a finite minimal subset, so there exists a minimal subset $W_{min}\subseteq W$ whose conjunction entails $\sim Ec|_{(x,t)}$.
-Therefore, there exists a hyperedge $W_{min}\subseteq W\subseteq G$ in $H_{or}$, again contradicting the independence of $G$.
+Since $S_{or}\left(or,{Tr}_{h}\right)$ is a finite set, we construct a finite hypergraph
+$H_{or}=\left(S_{or}\left(or,{Tr}_{h}\right),\mathcal{E}\right)$ (5.6)
+where the vertex set is precisely $S_{or}\left(or,{Tr}_{h}\right)$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
+(1) For any $\varphi r\in S_{or}\left(or,{Tr}_{h}\right)$ and $(x,t)\in or\times{Tr}_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\left\{\varphi r\right\}$ to $\mathcal{E}$;
+(2) If there exists a minimal subset $W\subseteq S_{or}\left(or,{Tr}_{h}\right)$ such that the conjunction of all formulas in $W$ entails $\sim Ec|_{(x,t)}$, then add the hyperedge $W$ to $\mathcal{E}$.
+Let $G$ be an independent set in $H_{or}$, i.e., a subset of vertices containing no hyperedge $e\in\mathcal{E}$. If there exists $\varphi\in Cn(G)$ such that
+$\varphi\rightarrow\sim Ec|_{(x,t)}$ (5.7)
+then there must exist a finite subset $\left\{{\varphi r}_{1},\cdots,{\varphi r}_{k}\right\}\subseteq G$ such that ${\varphi r}_{1}\land\cdots\land{\varphi r}_{k}=\varphi$, and hence ${\varphi r}_{1}\land\cdots\land{\varphi r}_{k}\rightarrow\sim Ec|_{(x,t)}$.
+Let $W=\left\{{\varphi r}_{1},\cdots,{\varphi r}_{k}\right\}$, so the conjunction of $W$ entails $\sim Ec|_{(x,t)}$.
+According to the hyperedge construction rules, if $|W|=1$, then $W$ itself directly violates $Ec$, and there must be a hyperedge $\left\{{\varphi r}_{1}\right\}$.
+But since $\left\{{\varphi r}_{1}\right\}\in G$, this contradicts the assumption that $G$ is an independent set.
+If $|W|\geq 2$, then by the compactness theorem, the logical entailment can be captured by a finite minimal subset, so there exists a minimal subset $W_{\min}\subseteq W$ whose conjunction entails $\sim Ec|_{(x,t)}$.
+Therefore, there exists a hyperedge $W_{\min}\subseteq W\subseteq G$ in $H_{or}$, again contradicting the independence of $G$.
 Thus, (4.7) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
-By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{max}$ of maximal cardinality.
+By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality.
 Suppose there exists another ethically safe subset $K$ with
-$|K|>|G_{max}|$ (4.8)
-then $K$ cannot contain any hyperedge $e\in{E}$.
+$|K|>\left|G_{\max}\right|$ (5.8)
+then $K$ cannot contain any hyperedge $e\in\mathcal{E}$.
 Otherwise, if $|e|=1$, then $K$ directly violates $Ec$; if $|e|\geq 2$, then the formulas in $K$ violate $Ec$.
 Both cases contradict the ethical safety assumption for $K$.
-Therefore, $K$ must alsobe an independent set in $H_{or}$, which contradicts the maximality of $G_{max}$.
-Hence, $G_{max}$ is the largest ethically safe subset of $S_{or}(or,Tr_{h})$ with respect to $Ec$.
+Therefore, $K$ must also be an independent set in $H_{or}$, and (5.8) cannot hold.
+Hence, $G_{\max}$ is the largest ethically safe subset of $S_{or}\left(or,{Tr}_{h}\right)$ with respect to $Ec$.
 […]
-In this way, add to $Sou(ou,Tu_{h})$ an ethical safeguard formula:
-$\varphi u=($ $\forall\varphi r)(IsFormulaOf^{2}(\varphi r,S_{or}(or,Tr_{h})))\to\exists\varphi s((IsFormulaOf^{2}(\varphi s,S_{os}(os,Ts_{h}))\land$
-$(IsFormulaOf^{2}(\varphi r,S_{or}^{Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,\varphi r))\land$
-$(IsFormulaOf^{2}(\varphi r,S_{or}^{\sim Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,IsSafeback^{1}(NEc))))$ (4.9)
-Since the newly added $\varphi u$ pertains to $S_{os}(os,Ts_{h})$ and is entirely independent of the existing formulas in $Sou(ou,Tu_{h})$, it does not affect the satisfaction of Postulate 5 by $Sou(ou,Tu_{h})$.
+In this way, add to $Sou\left(ou,{Tu}_{h}\right)$ an ethical safeguard formula:
+$\begin{split}\varphi u=&(\forall\varphi r)({IsFormulaOf}^{2}\left(\varphi r,S_{or}\left(or,{Tr}_{h}\right)\right))\rightarrow\exists\varphi s\left({IsFormulaOf}^{2}\left(\varphi s,S_{os}\left(os,{Ts}_{h}\right)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}(\varphi s,\varphi r)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{\sim Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}\left(\varphi s,{IsSafeback}^{1}(NEc)\right)\right)\right)\end{split}$ (5.9)
+Since the newly added $\varphi u$ pertains to $S_{os}\left(os,{Ts}_{h}\right)$ and is entirely independent of the existing formulas in $Sou\left(ou,{Tu}_{h}\right)$, it does not affect the satisfaction of Postulate 5 by $Sou\left(ou,{Tu}_{h}\right)$.
 Moreover,
-$S_{os}(os,Ts_{h})=S_{or}^{Ec}(or,Tr_{h})\cup\{IsSafeback^{1}(NEc)\}$ (4.10)
-Because NEc and IsSafeback1(NEc) were not previously present in $\mathcal{L}$, and $S_{or}^{Ec}(or,Tr_{h})$ and $\{IsSafeback^{1}(NEc)\}$ are logically independent and both logically consistent, $S_{os}(os,Ts_{h})$ is also logically consistent and satisfies all other conditions of Postulate 5.
+$S_{os}\left(os,{Ts}_{h}\right)=S_{or}^{Ec}\left(or,{Tr}_{h}\right)\cup\left\{{IsSafeback}^{1}(NEc)\right\}$ (5.10)
+Because NEc and ${IsSafeback}^{1}(NEc)$ were not previously present in $\mathcal{L}$, and $S_{or}^{Ec}\left(or,{Tr}_{h}\right)$ and $\left\{{IsSafeback}^{1}(NEc)\right\}$ are logically independent and both logically consistent, $S_{os}\left(os,{Ts}_{h}\right)$ is also logically consistent and satisfies all other conditions of Postulate A.1.
 Therefore, it is the maximal output state that is ethically safe with respect to $Ec$.
-By taking $S_{os}(os,Ts_{h})$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,Ts_{m})$, $S_{M}(M,Ts_{m})$ is necessarily also ethically safe with respect to $Ec$.
+By taking $S_{os}\left(os,{Ts}_{h}\right)$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,{Ts}_{m})$, $S_{M}(M,{Ts}_{m})$ is necessarily also ethically safe with respect to $Ec$.
 This satisfies the condition in Definition 5 for model M to be ethically safe with respect to $Ec$.
 The theorem is thus proved.
-Theorem 5 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs-for obtaining the maximum ethically safe output.
+Theorem 2 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs—for obtaining the maximum ethically safe output.
 However, computing the maximum independent set in a hypergraph is generally an NP-hard problem.
-In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}(or,Tr_{h})$, and another for ethical safety checking, thus yielding an approximately maximal safe output.
+In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}\left(or,{Tr}_{h}\right)$, and another for ethical safety checking, thus yielding an approximately maximal safe output.
 This is a practical issue that warrants further investigation and will not be elaborated upon here.
@@
 This approach, unlike those based on heuristic rules or statistical indicators, fills the gap in mathematical foundations and verifiability in traditional research, and provides concrete methodological guidance for the automated embedding and verification of ethical constraints.
-4.3 Theorem on Model Generalization Error Estimation
-Improving model generalization ability is among the most critical objectives in machine learning.
-However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance Vapnik1999 ; Bartlett2002 ; Zhang2016 ; Advani2020 ; Perlaza2024 .
-It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples.Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
-Theorem 6 (Upper Bound on TVD) Let the information describing the model M after learning and optimization be $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and the input information be $I_{q}=\langle oq,Tq_{h},S_{oq},eq,Tq_{m},S_{cq}\rangle$.
+5.3 Theorem on TVD Upper Bounds
+Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
+Theorem 3 (TVD Upper Bound for Conservative Prediction Strategies in Noise-Free Inputs and Uncovered Training Regions) Let the information describing the model $M$ after learning and optimization be $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$, and the input information be $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$.
 Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space.
-Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively.
-The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$, the distribution over $S_{oq}$ is $p_{q}(x)$,with $p_{q}(x)>0$ for $x\in S_{oq}$ .
+Let $\left|S_{oq}\right|$ and $\left|S_{ou}\cap S_{oq}\right|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively.
+The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$ ; the distribution over $S_{oq}$ is $p_{q}(x)$, with $p_{q}(x)>0$ for $x\in S_{oq}$.
 The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$.
-When the model M is sufficiently trained on $S_{ou}$ and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
-$TVD(p_{q},p_{M})\begin{cases}\approx 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.11)
-For the second case, it is assumed that $S_{oq}\setminus S_{ou}\neq\emptyset$, $1-P_{Mq}>0$, and $|S_{oq}|-|S_{ou}\cap S_{oq}|>0$.
-Here, $H(p_{q})=-\sum_{x\in S_{oq}\setminus S_{ou}}p_{q}(x)\log p_{q}(x)$ is the entropy of the data in the non-overlapping region.
+When the model $M$ is sufficiently trained on $S_{ou}$, and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
+$TVD\left(p_{q},p_{M}\right)$ $=\begin{cases}\approx 0,&\text{if }S_{oq}\subset S_{ou}\\[5.0pt] \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log{((\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|)}/(1-P_{Mq}))-H(p_{q}))},&\text{if }S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (5.11)
 […]
-$p_{M}(x)\begin{cases}\approx p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.12)
-Using KL divergence to measure the discrepancy between the true distribution and the hypothesized model distribution, we have Goodfellow2016 :
-$D_{KL}(p_{q}(x)\|p_{M}(x))=$ $\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (4.13)
-$=$ $\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
-According to (4.12), in the overlapping region:
-$\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\to\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}=0$ (4.14)
+[…]
+$p_{M}(x)=\begin{cases}\approx p_{q}(x),&\text{if }x\in S_{ou}\cap S_{oq}\\[3.00003pt] \frac{1-P_{Mq}}{|S_{oq}|-|S_{ou}\cap S_{oq}|},&\text{if }x\in S_{oq}\backslash S_{ou}\end{cases}$ (5.12)
+Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Kullback and Leibler, 1951):
+$D_{KL}(p_{q}(x)\parallel p_{M}(x))$ $=\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
+$=\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (5.13)
+According to (5.12), in the overlapping region,
+$\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ $\approx\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}$
+$=0$ (5.14)
 This establishes the first case in (5.11).
@@
 $\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ $=\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{M}(x)$
-$=-H(p_{q})-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
-$=-H(p_{q})-\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$ (1)
-$=(1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q})$ (4.15)
+$=-H(p_{q})-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{(1-P_{Mq})}{|S_{oq}|-|S_{ou}\cap S_{oq}|}$
+$=-H(p_{q})-\log\frac{(1-P_{Mq})}{|S_{oq}|-|S_{ou}\cap S_{oq}|}\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$
+$=(1-P_{Mq})\log\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}}-H(p_{q})$ (5.15)
 Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance.
-By Pinsker’s inequality Pinsker1964 , we have:
-$\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
-$=\sqrt{\frac{1}{2}(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)})}$
-$\approx\sqrt{\frac{1}{2}((1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q}))}$ (4.16)
+By Pinsker’s inequality (Pinsker, 1964), we have: …
+$TVD(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\parallel p_{M}(x))}$
+$=\sqrt{\frac{1}{2}\left(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\right)}$
+$\approx\sqrt{\frac{1}{2}\left((1-P_{Mq})\log\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}}-H(p_{q})\right)}$ (5.16)
 This establishes the second case in (5.11).
-Since we have already assumed that the carrier states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ of $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and $I_{q}=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model M and its interface, respectively.
+Since we have already assumed that the carrier states $S_{M}(M,{Tu}_{m})$ and $S_{cq}(cq,{Tq}_{m})$ of $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$ and $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model $M$ and its interface, respectively.
 It follows that equation (5.11) characterizes the upper bound of the model’s TVD.
 This completes the proof.
-Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 6, based on the MLT-MF framework, provides an explicit quantitative estimate for the model’s TVD.
-This estimate is grounded in the overlap between the test and training data $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, as well as the entropy $H(p_{q})$ of the test data distribution over the non-overlapping region.
-These metrics serve as indicators of how much the testing data shares information with the model’s prior knowledge and how unpredictable the unseen portion is.
-Because TVD characterizes the alignment between distributions at a fundamental level—independent of any particular model architecture or task—it is generally applicable to the analysis of model training performance.
-In fact, for classification tasks with 0-1 loss, the generalization error upper bound is equivalent to the TVD Rodríguez2021 .
-For regression tasks, the upper bound on generalization error is given by the product of TVD and the Lipschitz constant of the loss function Devroye2013 .
-Thus, TVD is a meaningful and informative measure for estimating model generalization error.
-This shows that Theorem 6 offers a new perspective and practical guidance for data collection, model design, and training for construction.
-In particular, Theorem 6 reveals that as long as the problem space is contained within the model’s information space, the model’s TVD can be made arbitrarily close to zero in a statistical sense.
-This is a fundamental result in machine learning:
-Corollary 2 (TVD for Ideal Models) If the model’s information encompasses all relevant information for the problem, then the TVD for these problems is approximately zero.
+Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H\left(p_{q}\right)$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD.
+Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance.
+Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD (Devroye et al., 2013), while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss (Devroye et al., 2013).
+Therefore, TVD serves as an indicator for estimating a model’s generalization error.
+In this way, Theorem 3 offers a novel perspective and practical guidance for model design and training data acquisition.
+In general, input information is inevitably affected by noise, and arbitrary predictive strategies can be adopted in regions not covered by training data.
+According to Definition 3 on structural decomposition of noisy information, we can thus obtain more general results regarding the model’s TVD.
+Theorem 4 (TVD Upper Bound for Arbitrary Prediction Strategies with Noisy Inputs and Uncovered Training Regions) Let the base space $X$ be a finite set.
+The training information domain of the model $M$ is $S_{ou}\subseteq X$.
+The true problem domain is $S_{oq}\subseteq X$, and the true distribution $p_{q}$ satisfies $\sum_{x\in S_{oq}}{p_{q}(x)}=1$.
+The noisy information ontological state $\SoqSoq+\widetilde{S_{oq}}=S_{oq}\backslash$ $\cup S_{oq}^{+}$, as defined in Definition 3, replaces $S_{oq}$ as the actual input to the model $M$.
+The noise loss probability mass is $P_{l}=\sum_{x\in S_{oq}^{-}}{p_{q}(x)}\in[0,1]$, and the noise addition probability mass is $P_{a}=\sum_{x\in S_{oq}^{+}}{p^{+}(x)}\in[0,1]$, where $p^{+}$ is a non-negative kernel on the noise addition region, satisfying $1-P_{l}+P_{a}>0$.
+The distribution of the noisy input $\widetilde{S_{oq}}$ is given by $\widetilde{p_{q}}(x)=(p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\})/(1-P_{l}+P_{a})$.
+Let $V=\widetilde{S_{oq}}\backslash S_{ou}$ denote the region not covered by training information, with uncovered probability mass $P_{uncov}=\sum_{x\in V}{\widetilde{p_{q}}(x)}\in[0,1]$.
+The conditional distribution on $V$ is $r(x)=\widetilde{p_{q}}(x)/P_{uncov}$ (and when $P_{uncov}=0$, all terms involving $V$ below are taken as 0).
+The predictive strategy $p_{0}$ on the uncovered region $V$ is an arbitrary distribution, satisfying $\sum_{x\in V}{p_{0}(x)=1,}p_{0}(x)\geq 0$.
+The model distribution $p_{M}$ is determined by training on the covered region $U=\widetilde{S_{oq}}\cap S_{ou}$, and on $V$, $p_{M}(x)=P_{uncov}p_{0}(x)$.
+Assume that the model's fitting error on the covered region is $E_{fit}=D_{KL}\left(\widetilde{p_{q}}|_{U}\parallel p_{M}|_{U}\right)$, where $\widetilde{p_{q}}|_{U}$ and $p_{M}|_{U}$ denote the distributions restricted to $U$ and normalized (under the ideal conditions assumed in Theorem 3, $E_{fit}$ tends to zero).
+Thus, the upper bound of the TVD for model $M$ with respect to noisy input information is given by $TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.17)
+The upper bound of the TVD for model $M$ with respect to the true information is
+$TVD\left(p_{q},p_{M}\right)\leq\left\{\begin{array}[]{r}P_{l}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.18)
+Proof: Based on the conditions provided in the theorem, we have:
+$D_{KL}\left(\widetilde{p_{q}}\parallel p_{M}\right)=\sum_{x\in\widetilde{S_{oq}}}{\widetilde{p_{q}}(x)\log\frac{\widetilde{p_{q}}(x)}{p_{M}(x)}}$
+$\sum_{x\in U}{\widetilde{p_{q}}(x)\log\frac{\widetilde{p_{q}}(x)}{p_{M}(x)}}+\sum_{x\in V}{\widetilde{p_{q}}(x)\log\frac{\widetilde{p_{q}}(x)}{p_{M}(x)}}$
+$D_{KL}\left(\widetilde{p_{q}}|_{U}\parallel p_{M}|_{U}\right)+\sum_{x\in V}{P_{uncov}r(x)\log\frac{P_{uncov}r(x)}{P_{uncov}p_{0}(x)}}$
+$E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)$ (5.17)
+Here, $E_{fit}$ reflects the degree of fit of $M$ on the region covered by the training information, and its probabilistic upper bound can be characterized by measures such as VC dimension, Rademacher complexity, or training stability.
+Moreover, by applying Pinsker’s inequality, the first conclusion of the theorem can be obtained:
+$TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{D_{KL}\left(\widetilde{p_{q}}\parallel p_{M}\right)}{2}}=\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.18)
+Since the distribution of the noisy input $\widetilde{S_{oq}}$ is
+$\widetilde{p_{q}}(x)=\frac{p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\}}{1-P_{l}+P_{a}}$ (5.19)
+It follows that
+$TVD\left(p_{q},\widetilde{p_{q}}\right)=(\sum_{x\in X}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
+$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
+$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p^{+}(x)/(1-P_{l}+P_{a})\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-p_{q}(x)/(1-P_{l}+P_{a})\right|)/2$
+$(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}{p_{q}(x)}}{1-P_{l}+P_{a}})/2=(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|(1-P_{l})}{1-P_{l}+P_{a}})/2$
+$\left\{\begin{array}[]{r}P_{l},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}},\ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.22)
+By applying the triangle inequality, we have
+$TVD\left(p_{q},p_{M}\right)\leq TVD\left(p_{q},\widetilde{p_{q}}\right)+TVD\left(\widetilde{p_{q}},p_{M}\right)$
+$\leq\left\{\begin{array}[]{r}P_{l}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.23)
+Thus, the second conclusion of the theorem is established, completing the proof.
+Theorem 4 unifies the noise structure, coverage geometry, and uncovered-region strategies within a computable TVD upper bound framework.
+Specifically, the TVD upper bounds for model $M$ with respect to both noisy actual inputs and true inputs are expressed in terms of the noise loss mass $P_{l}$, addition mass $P_{a}$, fitting error over the covered region $E_{fit}$, uncovered mass $P_{uncov}$, and the KL divergence $D_{KL}\left(r\parallel p_{0}\right)$ between the conditional distribution on the uncovered region and the prediction strategy.
+This formulation reveals several important insights: the loss and addition masses determine the deviation between the true target and the noisy target; the fit over the covered region, the coverage structure, and the divergence between the conditional distribution and prediction strategy on the uncovered region jointly determine the model's deviation from the noisy target.
+Consequently, the effects of noise deletion/addition on target deviation can be directly assessed.
+This upper bound provides actionable guidance for the joint optimization of training coverage and prediction strategies, and can undoubtedly be applied more broadly to the analysis and optimization of model performance.
 6 Experimental Results
-Using the feedforward neural network model listed in the appendix, we conducted regression experiments on two bivariate root functions: $y_{1}=\sqrt{4x_{1}^{2}+3x_{2}^{2}}$ and $y_{2}=\sqrt{2x_{1}^{2}+x_{2}^{2}}$.
-Both the training and test datasets were generated randomly.
+Using the feedforward neural network model listed in Appendix C, we conducted regression experiments on two bivariate root functions: $y_{1}=\sqrt{4x_{1}^{2}+3x_{2}^{2}}$ and $y_{2}=\sqrt{2x_{1}^{2}+x_{2}^{2}}$.
+Both the training and test datasets were specific.
 We evaluated the models’ prediction errors under various conditions, assessed the impact and computational cost of ethical and safety mechanisms, and explored the challenge of providing multiple types of explanations for the same output.
 These experiments serve to illustrate and validate the definitions and theorems presented in Section 5.
-5.1 Prediction Error Experiments
-To evaluate the model’s predictive performance in both overlapping and non-overlapping cases between the training and test datasets, we used the following data setup: 10,000 training samples were generated with $x_{1}$ and $x_{2}$ independently and uniformly distributed over $[0.5,1.5]$, and 10,000 test samples were drawnuniformly from within a circle of radius 1.0 centered at (1.5, 1.0).
-The model’s learning rate was set to $\eta=0.01$.
-Figure 2 illustrates the experimental conditions and outcomes:
-Figure 2: Model Prediction Error Experiments and Results Under Different Conditions.
-(a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples.
-Among these, 4,951 samples overlap with the training set.
-(b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset.
-(c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound.
-(d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
+6.1 Experiments on the TVD upper bound under noiseless input and conservative prediction strategy
+To evaluate the TVD upper bound estimation of Theorem 3 under noiseless input and conservative prediction strategy, as well as its relationship to model performance, we construct the following experimental setup.
+On the ${(x}_{1},x_{2})$ plane, we select a square region centered at (1.0,1.0) with side length 1, and uniformly sample 10,000 points as the training dataset.
+For the test dataset, we uniformly sample 10,000 points within a circular region centered at (1.5,1.0) with radius 0.5.
+The model learning rate is set to $\eta=0.01$.
+By uniformly sampling 10,000 points within the test region and computing the corresponding discrete points in the ${(y}_{1},y_{2})$ plane using the target function, we approximate the distribution $p_{q}(x)$ of the test dataset and its entropy function.
+Fig.
+3 illustrates the experimental conditions and outcomes:
 (a) The overlap rate between the test and training datasets is approximately 50%.
@@
 (c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used.
-For almost all samples (except a few outliers), their RMSE is less than $\text{TVD}=\sqrt{((1-0.4951)\ln((10000-4951)/(1-0.4951))-2.9741)/2}\approx 0.9155$.
+For almost all samples (except a few outliers), their RMSE is less than $TVD\approx\sqrt{((1-0.4951)\ln{((10000-4951)}/(1-0.4951))-2.9741)/2}\approx 0.9155$.
 This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
@@
 Since the theoretical proof of Theorem 3 is independent of specific models and tasks, it promises to be extensible to more complex scenarios.
+Figure 3: Experimental Evaluation and Results of the TVD Upper Bound under Noiseless Input and Conservative Prediction Strategy.
+(a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples.
+Among these, 4,951 samples overlap with the training set.
+(b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset.
+(c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound.
+(d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
+6.2 Experimental Evaluation of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy
+To evaluate Theorem 4’s estimation of the TVD upper bound under conditions of noisy input and the adoption of arbitrary prediction strategies in regions not covered by the training data, as well as itsrelationship to model performance, we utilize the same model instance, training dataset, and test dataset as in the previous experiment.
+We introduce noise into the input by locally adding or deleting noise information in the test dataset according to Definition 3, so that the model input contains noise.
+Meanwhile, in the regions not covered by the training data, we apply an arbitrary prediction strategy based on the joint distribution $p_{0}(x)$ formed by two independent normal distributions, $N\sim(0.6306,{0.4585}^{2})$ and $N\sim\left(0.4031,{0.3666}^{2}\right)$, to process and predict the input data.
+Fig.
+3: (a) The training dataset $S_{ou}$ (selected within the blue square) and the test dataset $S_{oq}$ (selected within the red circle) are the same as those used in the experiment described in Section 6.1.
+The upper arc-shaped segment of $S_{oq}$ is shifted upward, and the vacated region, denoted as $S_{oq}^{-}$, is excluded from the test input.
+The shifted portion, $S_{oq}^{+}$, does not overlap with the original test dataset but partially overlaps with $S_{ou}$ ; it is added as new noisy test data.
+This structure satisfies the definition of noisy input, $\SoqSoq+\widetilde{S_{oq}}=S_{oq}\backslash$ $\cup S_{oq}^{+}$.
+Figure 4: Experimental Evaluation and Results of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy.
+(a) The upper arc-shaped region of the test dataset is shifted upward, and the resulting vacant region is excluded from the test input.
+The shifted segment, which does not overlap with the original test dataset, is incorporated as noise into the new test dataset.
+(b) One hundred groups of samples are randomly selected from the noisy test dataset and processed by the model.
+The RMSE of the predictions are represented by 100 points along the blue line, while the red line indicates the TVD upper bound between the noisy input distribution and the model distribution.
+(c) In the case of $P_{l}\geq P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model.
+The RMSE of each prediction is represented by 100 points along the blue line, with the red line denoting the TVD upper bound between the noiseless input distribution and the model distribution.
+(d) In the case of $P_{l}<P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model.
+The RMSE of each prediction is represented by 100 points along the blue line, with the red line indicating the TVD upper bound between the noiseless input distribution and the model distribution.
+(b) For the noisy test dataset comprising 11,034 samples, we compute the probability distribution for each test sample using the same method as in Section 6.1.
+The loss mass is $P_{l}=0.0744$, the addition mass is $P_{a}=0.0517$, the uncovered mass is $P_{uncov}=0.5091$, and the KL divergence is $D_{KL}\left(r\parallel p_{0}\right)=45.58$.
+This yields a TVD upper bound between the noisy input distribution and the model distribution of 3.312.
+By randomly selecting 100 groups of samples from the test dataset and inputting them into the model after 20 training steps, we find that the root mean square error (RMSE) of all outputs (marked by blue dots) is less than the TVD upper bound indicated by the red line.
+This validates the conclusion of Theorem 4, Equation (5.17).
+(c) In the above example, $P_{l}=0.0744>0.0517=\gamma$.
+Under the same conditions, we randomly select 100 samples from the noiseless test dataset, convert them into corresponding noisy data as input to the model, and find that the RMSE of all outputs remains below the TVD upper bound of 3.4555.
+This validates the case where $P_{l}\geq P_{a}$ as given in Theorem 4, Equation (5.18).
+(d) To examine the case where $P_{l}<P_{a}$ in Theorem 4, Equation (5.18), we construct a new noisy test dataset comprising 10,000 samples.
+In this case, $P_{l}=0.0744$, $P_{a}=0.1551$, $P_{uncov}=0.5530$, and $D_{KL}\left(r\parallel p_{0}\right)=40.88$, resulting in a TVD upper bound of 3.4325.
+Again, by randomly selecting 100 samples from the noiseless test dataset and converting them into noisy inputs for the model, the RMSE of all outputs remains below the TVD upper bound indicated by the red line, thereby confirming the corresponding conclusion provided by Theorem 4.
+Through these experiments, we have comprehensively validated Theorem 4 regarding the TVD upper bound under noisy input and arbitrary prediction strategies.
+Furthermore, as illustrated in Fig.
+4, there remains a noticeable gap between the RMSE of the model outputs and the TVD upper bound across various scenarios, indicating that the upper bound provided by Theorem 4 is still somewhat conservative.
+Further research may yield tighter error estimates under typical conditions.
 6.3 Ethical Safety Assurance Experiment
 Given the nature of the bivariate root function regression task, the model’s predicted outputs should never be negative.
-This requirement is analogous to Definition 10, which ensures that model outputs do not violate the ethical constraints .
-Therefore, using the same training dataset as in the previous prediction error experiments, we set the learning rate to and trained the model for 165 steps.
-For testing, we generated a dataset of 10,000 samples with drawn uniformly from within a circle of radius 0.5 centered at (-1, -1).
+This requirement is analogous to Definition 5, which ensures that model outputs do not violate the ethical constraints $Ec$.
+Therefore, using the same training dataset as in the previous prediction error experiments, we set the learning rate to $\eta=0.001$ and trained the model for 165 steps.
+For testing, we generated a dataset of 10,000 samples with $(x_{1},x_{2})$ drawn uniformly from within a circle of radius 0.5 centered at (-1, -1).
 The criterion for judging ethical safety violations was whether the model’s predicted output was negative.
@@
 Table 1: Summary Table of Ethical Safety Experiment Results
-Trial Computation Computation Computation Number of Number of Correct Correct
-Number Time Without Time With Time Ratio Ethically Unethical Prediction Prompt
-Ethical Ethical Safe Warning Rate for Rate for
-Safety Check Safety Check Samples Samples Safe Unethical
-(seconds) (seconds) Outputs Outputs
+Trial Time w/o Time w/ Time Safe Warning Safe Warning
+Number Check (s) Check (s) Ratio Samples Samples Rate Rate
 1 0.367 0.391 1.063 4517 483 100% 100%
@@
 Therefore, for the test sample $X=[-1.3535,-1.3535]^{T}$, the model predicts:
-$Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\text{Identity}(W_{2}\text{ReLU}(W_{1}X+b_{1})+b_{2})\approx\begin{bmatrix}-0.040\\ 0.277\end{bmatrix}$ (5.3)
+$Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\text{Identity}\left(W_{2}\text{ReLU}\left(W_{1}X+b_{1}\right)+b_{2}\right)\approx\begin{bmatrix}-0.040\\ 0.277\end{bmatrix}$ (6.3)
 Here, $y_{1}=-0.040<0$, which mathematically explains why model M outputs a negative value for input $X$.
 These formulas provide the designer’s mathematical representation of the model’s computational process.
-They are equivalent to the formal forward propagation rules (A.21) and (A.22) in the appendix.
+They are equivalent to the formal forward propagation rules (C.21) and (C.22) in Appendix C.
 According to the OIT perspective, these equations describe the ontological state of the model, its outputs, and the information they convey.
@@
 Depicting the model’s computational process visually can make explanations of model behavior much more intuitive.
-To this end, for model M after 165 training steps, we illustrate its complete structure—covering the inputs, internal neural network calculations and data flow, and outputs—in Figure 3.
+To this end, for model M after 165 training steps, we illustrate its complete structure—covering the inputs, internal neural network calculations and data flow, and outputs—in Fig.
+5.
 Figure 5: Schematic diagram of the computational process for model M after 165 training steps, showing its handling of the input sample $X=[-1.3535,-1.3535]^{T}$ and output $Y\approx[-0.040,0.277]^{T}$.
-Compared with equations (5.1)(5.3), Figure 3 not only comprehensively reflects the input, weight matrices, bias vectors, neuron activation functions, and the specific data and formulas for outputs, but also highlights the interactions between the input/output devices and the model.
+Compared with equations (6.1)–(6.3), Fig.
+5 not only comprehensively reflects the input, weight matrices, bias vectors, neuron activation functions, and the specific data and formulas for outputs, but also highlights the interactions between the input/output devices and the model.
 The figure offers a more direct view of the internal data processing pipeline, making it easier to examine and verify each step of the computation.
@@
 6.4.3 Partition-Based Explanation
+It is evident that equations (6.1)–(6.3) and Fig.
+5 explain only the result of model M processing the specific sample $X=[-1.3535,-1.3535]^{T}$.
+They do not account for the phenomenon observed in the ethical safety assurance experiment, where each random selection of 5,000 samples from the test dataset yields over 19% negative predictions.
+To explain this, a broader statistical analysis across more data samples is needed.
+Fig.
+6(a) provides a visual depiction of the decision boundary after 165 training steps, showing where model M predicts negative outputs.
+Notably, there is a boundary line approximately passing through points (-2.0, -0.4) and (-0.4, -2.0).
+For samples on the upper right side of this diagonal, model M consistently predicts non-negative values; for those on the lower left, the model consistently outputs negative values.
+The red region in the figure, representing the test dataset, falls precisely on the lower left side of this boundary line, which explains why a random selection of 5,000 samples from the test set will always include a substantial portion that triggers negative predictions.
+This visual and statistical partitioning thus provides an intuitive explanation for the observed frequency of negative outputs in the ethical safety experiment.
 Figure 6: Illustration of Prediction Polarity and Impact of Training Steps.
@@
 The blue and green lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (1, 1) from the training set; the red and pink lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (-1.3535, -1.3535) from the test set.
-It is evident that equations (5.1)(5.3) and Figure3 explain only the result of model M processing the specific sample $X=[-1.3535,-1.3535]^{T}$.
-They do not account for the phenomenon observed in the ethical safety assurance experiment, where each random selection of 5,000 samples from the test dataset yields over 19% negative predictions.
-To explain this, a broader statistical analysis across more data samples is needed.
-Figure 4(a) provides a visual depiction of the decision boundary after 165 training steps, showing where model M predicts negative outputs.
-Notably, there is a boundary line approximately passing through points (-2.0, 0.4) and (-0.4, -2.0).
-For samples on the upper right side of this diagonal, model M consistently predicts non-negative values; for those on the lower left side, the model consistently outputs negative values.
-The test region in the figure, representing the test dataset, falls precisely on the lower left side of this boundary line, which explains why a random selection of 5,000 samples from the test set will always include a substantial portion that triggers negative predictions.
-This visual and statistical partitioning thus provides an intuitive explanation for the observed frequency of negative outputs in the ethical safety experiment.
-Figure 4(a) presents an overall view of numerous inputs processed and predicted by model M.
-This visualization also represents an ontological state of M as it handles a wider range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping.
+Fig.
+6(a) presents an overall view of numerous inputs processed and predicted by model M.
+This visualization also represents an ontological state of M as it handles a wide range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping.
 Thus, according to Definition 4, Figure 5(a) constitutes a valid explanation for the distribution of prediction polarities by M.
@@
 However, model training is a continuous process of cumulative improvement.
-Examining only the model’s instantaneous performance at a particular training step cannot fully explain the causes of abnormal outputs.
+Examining only the model's instantaneous performance at a particular training step cannot fully explain the causes of abnormal outputs.
 To address this, we further trained M for 400 iterations on the training dataset.
-Figure 4(b) shows the predictions for both $(1,1)$ and (1.3535,1.3535)(-1.3535,-1.3535) after each training step.
+Fig.
+6(b) shows the predictions for both (1, 1) and (-1.3535, -1.3535) after each training step.
 It is clear that even for (1, 1) — a central sample from the training set—the predicted $y_{2}$ by M remains negative during roughly the first 50 training steps.
 Only after more than 50 steps does the prediction move away from negative values.
-For the test sample (1.3535,1.3535)(-1.3535,-1.3535), it takes more than about 180 training steps before the prediction begins to consistently avoid negative values.
+For the test sample (-1.3535, -1.3535), it takes more than about 180 training steps before the prediction begins to consistently avoid negative values.
 This cumulative effect is entirely due to the optimization of M’s parameters as training progresses.
-In fact, by applying the forward propagation rules (A.21) and (A.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.
-6 Conclusion and Outlook
-This paper has established the "Meta-Framework for Machine Learning Theory (MLT-MF)" and, through formal information modeling and the sextuple definition, has unified the characterization of information flow and causal mechanisms across the entire machine learning lifecycle.
+In fact, by applying the forward propagation rules (C.21) and (C.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.
+7 Discussion and Conclusion
+This work, grounded in the Objective Information Theory (OIT), develops the MLT-MF framework, employing a formal logical system to integrate the complete causal chain of model construction, training, input processing, output, and online learning.
+Under a unified semantics, we propose formal definitions of model interpretability and ethical safety, derive two classes of explicit TVD upper bound theorems, and validate the framework through experiments with feedforward neural network models.
 The core contributions are as follows:
-  • Theoretical Integration: For the first time, data initialization, training optimization, and knowledge transfer are incorporated into a unified framework, overcoming the fragmented modeling of temporal causal chains in traditional research;
-  • Interpretability: By explicitly representing ontological and carrier states via higher-order logical systems and mapping model decisions back to the formal semantic space through information recoverability, logical transparency is ensured and a rigorous mathematical foundation for interpretability is provided;
-  • Ethical Safety: By embedding ethical constraint predicates, the compliance of model outputs can be formally verified, offering criteria and methodological guidance for the design of trustworthy AI systems;
-  • Model Performance: Based on state coverage and KL divergence, an upper bound formula for TVD is derived, revealing the core principle that the “problem space must be contained within the model knowledge space,” thus providing explicit tools for quantitative analysis in data collection and model optimization.
-We also validated and demonstrated the correctness and practicality of the relevant theorems based on MLT-MF through experiments with a simple feedforward neural network model.
-Despite these important advances, the current framework assumes ideal (noiseless) enabling mappings and primarily addresses model learning and processing logic in static scenarios.
-It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
-Future research should deepen the following directions:
-  • Extend to complex dynamic scenarios such as online learning under noisy conditions, introducing noisy ontologies and incremental coverage mechanisms;
-  • Investigate information fusion and conflict resolution across ontological spaces for multimodal and multi-agent systems, constructing distributed MLT-MF frameworks;
-  • Promote interdisciplinary integration of MLT-MF with fields such as quantum computing and bioinformatics, exploring formal mappings of quantum states and interpretable modeling of gene regulatory networks;
-  • Design automated toolchains to support conversion from formal rules to code, lowering the engineering threshold for MLT-MF adoption.
-7 Acknowledgements
+(1) Enabled Mapping for Information Transmission: We characterize the mathematical mapping and the dual physical-causal nature of information transfer from ontological states to carrier states through the concept of “enabled mapping.” This is further extended to chain transmission, mapping each engineering stage of model development to a segment of the information chain, such that model behavior, training, and inference are reduced to logically verifiable compositions of mappings.
+(2) Unified Representation of Interpretability and Safety: By adopting higher-order formal logic as a unified representation of “state,” we propose necessary and sufficient conditions for interpretability—linking input, output, and model within the same formal system via logical closure—and construct maximal safe outputs for ethical safety based on hypergraph maximum independent sets.
+(3) Integration of Statistical Error, Noise Structure, and Coverage Geometry: We integrate statistical learning error characterization with information noise structure and training coverage geometry, providing a closed-form TVD upper bound, and operationally decomposing it in terms of coverage fitting error, uncovered mass, KL divergence, and noise deletion/addition terms.
+Therefore, this study addresses two longstanding gaps in machine learning theory: the lack of unified semantics and causal chain constraints across different stages, and the absence of verifiable mathematical definitions and constructive guarantees for ethical safety and interpretability.
+By unifying the “existence–transmission–realization” of information within formal logical explanations, we rigorously establish that model interpretability corresponds to information existence, thereby providing a logical-theoretical foundation for translating ethical safety from normative requirements into decidable and constructible guarantees.
+Using TVD as a fundamental measure of distributional alignment, and establishing correspondences with practical engineering concerns such as training coverage gaps, uncovered-region prediction strategies, and noise deletion/addition, this work bridges the gap between abstract generalization theory and data coverage/system implementation, thus connecting statistical learning, mathematical logic, information science, and systems engineering.
+All conclusions and methods presented are amenable to engineering implementation:
+  • Interpretability can be realized at multiple levels through unified formal systems, including mathematical formulas, dataflow diagrams, and parameter representations, which is especially important for high-risk industries.
+  • Ethical safety is implemented via conflict detection and maximal independent set approximation of output propositions, enabling a deployable process for maximal safe outputs.
+[…]
+Of course, this work has limitations: the current formal system is most transparent for finite discrete sets, and further refinement is needed for continuous, time-varying, multimodal, and multi-agent scenarios, particularly regarding measurability and dynamic consistency.
+The maximal independent set for ethical safety is generally NP-hard, requiring the development of complexity-controlled approximation algorithms and alignment with business ethics standards.
+TVD upper bounds rely on conservative inequalities such as Pinsker’s; in typical scenarios, error estimates may be tightened using related theoretical advances.
+Physical causality in enabled mappings calls for interdisciplinary integration with various branches of physics, leveraging parameters such as energy consumption, latency, and uncertainty to advance from qualitative causality to quantitative verification.
+Future work may include large-scale empirical studies to verify the observability of TVD upper bounds and the benefits of strategy optimization, as well as further development of cross-modal and multi-agent information chain composition and conflict resolution methods, thereby enhancing the universality and practical value of MLT-MF in complex systems.
+Acknowledgments and Disclosure of Funding
+Junxuan He from Beijing University of Posts and Telecommunications implemented all the experiments described in the paper according to the author’s ideas, producing complete model code and training/test datasets, providing indispensable and strong support for illustrating and validating the theorems related to MLT-MF.
+The author offers special thanks for his hard work, keen understanding, and efficient implementation.
 During the writing and revision of this paper, I received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University.
@@
 I hereby express my sincere gratitude to them.
-Junxuan He from Beijing University of Posts and Telecommunications implemented all the experiments described in the paper according to the author’s ideas, producing complete model code and training/test datasets, providing indispensable and strong support for illustrating and validating the theorems related to MLT-MF.
-The author offers special thanks for his hard work, keen understanding, and efficient implementation.
-8 Data and Code availability
+Data and Code availability
 All datasets and codes used in this work are available at https://github.com/xuhu6736/MLT-MF.
-Appendix A Appendix: Model Instance Validation of MLT-MF
+Appendix A Appendix A. Formal Representation of States
+The postulates of information emphasize that the state of the ontology enables a mapping to the state of the carrier.
+Therefore, establishing a formal representation of both ontological and carrier states is particularly important.
+Definition A.1 (Higher-Order Formal Logical System): Suppose a formal system $\mathcal{L}$ is described using the following symbol set:
+Variables $x_{1}$, $x_{2}$,…;
+Individual constants $a_{1}$, $a_{2}$,…;
+Functions $f_{1}^{1}$, $f_{2}^{1}$, … $f_{1}^{2}$, $f_{2}^{2}$, … $f_{1}^{3}$, $f_{2}^{3}$, …;
+Parentheses $(,)$;
+Predicates $A_{1}^{1}$, $A_{2}^{1}$, … $A_{1}^{2}$, $A_{2}^{2}$, … $A_{1}^{3}$, $A_{2}^{3}$, …;
+Logical connectives “not” $\sim$, “implication” $\rightarrow$;
+Quantifier “for all” $\forall$.
+A term in $\mathcal{L}$ is defined as follows:
+(1) Variables and individual constants are terms;
+(2) If $f_{i}^{n}(n>0,i>0)$ is a function in $\mathcal{L}$, and $t_{1}$,…, $t_{n}$ are terms in $\mathcal{L}$ that do not contain $f_{i}^{n}$ itself, then $f_{i}^{n}(t_{1}$,…, $t_{n})$ is also a term in $\mathcal{L}$;
+(3) If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$, and $t_{1}$,…, $t_{n}$ are terms in $\mathcal{L}$ that do not contain $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1}$,…, $t_{n})$ is also a term in $\mathcal{L}$;
+(4) The set of all terms in $\mathcal{L}$ is generated by (1), (2), and (3).
+Based on this, we can define atomic formulas in $\mathcal{L}$:
+If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$ and $t_{1}$,…, $t_{n}$ are terms in $\mathcal{L}$ not containing $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1}$,…, $t_{n})$ is an atomic formula in $\mathcal{L}$.
+We then define well-formed formulas (WFFs) in $\mathcal{L}$:
+(5) Every atomic formula in $\mathcal{L}$ is a WFF;
+[…]
+(7) The set of all WFFs in $\mathcal{L}$ is generated by (5) and (6).
+The higher-order formal logical system $\mathcal{L}$ defined here, based on the flexibility of predicate logic, can describe various relationships and constraints among states.
+Its recursive structure and higher-order quantification support the nested definition of complex data and behaviors, and it explicitly prohibits any predicate or function symbol from applying to itself—directly or indirectly—to avoid self-referential paradoxes.
+For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols.
+However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise (Hamilton, 1988).
+Therefore, they are introduced as defined symbols in subsequent discussions.
+Definition A.2 (Interpretation of a Formal System) An interpretation $E$ of a formal system $\mathcal{L}$ over a domain $D_{E}$ assigns specific values to constants, elements of the domain $D_{E}$ to variables, and turns function and predicate symbols into functions and predicates over that domain $D_{E}$, thereby yielding a set of WFFs each with a truth value.
+Definition A.2 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects in both the subjective and objective worlds.
+Postulate A.1 (Expression of Object States) The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
+(1) Domain invariance: definitions of the object set and time set remain unchanged.
+(2) Numerical determinacy: concrete numerical values determine values of certain properties and forms.
+(3) Attribute expressibility: properties, forms, values, relations are expressible via functions and predicates in a formal system satisfying Definition 2.
+(4) Rule expressibility: evolution rules among attributes are expressible in such a formal system using time variables and implication, and contain no implicit self-reference.
+(5) Logical coherence: no logical contradictions exist among attributes and rules, nor among any logical conjunctions thereof.
+Lemma A.1 (Formal Expression of States) If the state set $S(X,T)$ of objects $X$ over times $T$ satisfies Postulate A.1, then it can be represented as an interpretation of a formal system $\mathcal{L}$ over the domain $X\times T$, and it is logically coherent.
+Proof: Since the state set $S(X,T)$ of the object set $X$ over the time set $T$ satisfies Postulate A.1, we construct the corresponding formal system $\mathcal{L}$ as follows:
+By (1), the definitions of the elements in $X$ and $T$ remain unchanged, ensuring determinacy.
+The variables in the formal system $\mathcal{L}$ are defined to correspond to each element of $X\times T$, thus making $X\times T$ the domain of discourse for $\mathcal{L}$.
+Consequently, $S(X,T)$ has a well-defined domain.
+By (2), the constants in $\mathcal{L}$ are defined to represent all possible values in $S(X,T)$.
+By (3), the functions and predicates in $\mathcal{L}$ represent the properties, forms, values, and relations of $X$ over $T$ as captured in $S(X,T)$.
+By (4), consider two ternary predicates $A_{1}^{3}$ and $A_{2}^{3}$.
+Let $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}\left(x_{1},t_{1}\right))$ and $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}\left(x_{2},t_{2}\right))$ be two properties of $S(X,T)$ corresponding to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and binary functions $f_{1}\left(x_{1},t_{1}\right),f_{2}\left(x_{1},t_{1}\right)$.
+Suppose all $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}\left(x_{1},t_{1}\right))$ satisfy the rule of evolving from $t_{1}$ to $t_{2}$ into some $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}\left(x_{2},t_{2}\right))$.
+The formula $(\forall A_{1}^{3}\forall f_{1}^{2}\forall x_{1}A_{1}^{3}(x_{1},t_{1},f_{1}^{2}\left(x_{1},t_{1}\right))\rightarrow{\exists A_{2}^{3}\exists f_{2}^{2}\exists x_{2}A}_{2}^{3}(x_{2},t_{2},f_{2}^{2}\left(x_{2},t_{2}\right))$ is also a well-formed formula in $\mathcal{L}$, with respect to the interpretation of $x_{1},x_{2},t_{1},t_{2}$, and is able to express the corresponding rules in $S(X,T)$.
+For other rules in $S(X,T)$, the related elements in the domain $X\times T$ can similarly be interpreted and expressed via corresponding WFFs in $\mathcal{L}$, with no implicit logical self-reference.
+Therefore, $S(X,T)$ constitutes an interpretation $E$ of $\mathcal{L}$ over the domain $X\times T$. By (5), both itself and its logical closure are logically consistent. The lemma is thus proved.
+Lemma A.1 stipulates that the variables in $\mathcal{L}$ are elements of $X\times T$; thus, in formulas, object and time elements appear in pairs $(x,t)$.
+In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
+Appendix B Appendix B. Finite Automaton States
+Using the formal expression of states, we can study in depth the special and important object of finite automata (Xu, 2024).
+Lemma B.1 (State Representation of Finite Automata) For any finite automaton $M$, one can describe its input, output, and state‐transition behavior over a related time set $T$ by the state set $S(M,T)$.
+Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions.
+By finite‐automaton theory (Editorial Committee of Dictionary of Mathematics, 2002), $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
+$\delta$: $U\mathbf{\times}Q$ to $U$, is the next‐state function;
+$\lambda$: $U\mathbf{\times}Q$ to $R$, is the output function.
+For each state $u_{t_{i}}\in U$, define the state predicate ${State}^{3}$ by
+$\varphi_{U}(t_{i})={State}^{3}(M,t_{i},u_{t_{i}})$ (B.1)
+to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$, $i=1,\ldots,n$.
+For each input $q_{t_{i}}\in Q$, define the input predicate ${Input}^{3}$ by
+$\varphi_{I}(t_{i})={Input}^{3}(M,t_{i},q_{t_{i}})$ (B.2)
+to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$, $i=1,\ldots,n-1$.
+For each output $r_{t_{i}}\in R$, define the output predicate ${Output}^{3}$ by
+$\varphi_{O}(t_{i})={Output}^{3}(M,t_{i},r_{t_{i}})$ (B.3)
+to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$, $i=2,\ldots,n$.
+The transition function $\delta$ is captured by the well‐formed formula
+$\varphi_{\delta}\left(t_{i}\right)=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land{Eq}^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (B.4)
+where the binary predicate ${Eq}^{2}(x,y)$ asserts $x=y$, $i=1,\ldots,n-1$.
+Similarly, the output function $\lambda$ is given by
+$\varphi_{\lambda}\left(t_{i}\right)=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land{Eq}^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land{IsElementof}^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$ (B.5)
+where ${IsElementof}^{2}(x,X)$ asserts $x$ is an element of $X$, $i=1,\ldots,n-1$.
+Thus the state set
+$\begin{split}S(M,T)=\{&\varphi_{U}\left(t_{i}\right),\varphi_{I}\left(t_{j}\right),\varphi_{O}\left(t_{k}\right),\varphi_{\delta}\left(t_{j}\right),\varphi_{\lambda}\left(t_{j}\right),\\ &i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}\end{split}$ (B.6)
+describes $M$’s inputs, outputs, and state‐transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
+It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits automatic transition behavior. Hence we introduce:
+Definition B.2 (Finite Automaton State) If a state set $S(X,T)$ expresses a finite automaton, then $S(X,T)$ is called a finite automaton state.
+All practical information systems, including computing systems, are finite automata in the mathematical sense.
+Therefore, the concept of finite automaton state is of great significance in the study of machine learning.
+Appendix C Appendix C. Model Instance Validation of MLT-MF
 To facilitate validation and illustration of the practical significance of MLT-MF and related theorems, we select a very simple model and use MLT-MF to establish formal representations with WFFs for each stage from initial information to training information, learning optimization, and processing/output.
-A.1 A Simple Feedforward Neural Network Model
-We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Figure A1).
+C.1 A Simple Feedforward Neural Network Model
+We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig.
+C.1).
 Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
-Figure A1: Schematic diagram of feedforward neural network model M
+Figure 7: Fig. C.1 Schematic diagram of feedforward neural network model M
 The hidden layer uses the ReLU activation function:
-$\text{ReLU}(x)=\max(0,x)$ (A.1)
+$ReLU(x)=max(0,x)$ (C.1)
 The output layer uses the identity activation:
-$\text{Identity}(x)=x$ (A.2)
+$Identity(x)=x$ (C.2)
 Initial weight matrices and bias vectors are defined between layers:
-$W_{1}=\begin{bmatrix}w_{11}^{1}&w_{12}^{1}\\ w_{21}^{1}&w_{22}^{1}\\ w_{31}^{1}&w_{32}^{1}\end{bmatrix},\quad b_{1}=\begin{bmatrix}b_{1}^{1}\\ b_{2}^{1}\\ b_{3}^{1}\end{bmatrix}$ (A.3)
-$W_{2}=\begin{bmatrix}w_{11}^{2}&w_{12}^{2}&w_{13}^{2}\\ w_{21}^{2}&w_{22}^{2}&w_{23}^{2}\end{bmatrix},\quad b_{2}=\begin{bmatrix}b_{1}^{2}\\ b_{2}^{2}\end{bmatrix}$ (A.4)
+$W_{1}=\begin{bmatrix}w_{11}^{1}&w_{12}^{1}\\ w_{21}^{1}&w_{22}^{1}\\ w_{31}^{1}&w_{32}^{1}\end{bmatrix},\quad b_{1}=\begin{bmatrix}b_{1}^{1}\\ b_{2}^{1}\\ b_{3}^{1}\end{bmatrix}$ (C.3)
+$W_{2}=\begin{bmatrix}w_{11}^{2}&w_{12}^{2}&w_{13}^{2}\\ w_{21}^{2}&w_{22}^{2}&w_{23}^{2}\end{bmatrix},\quad b_{2}=\begin{bmatrix}b_{1}^{2}\\ b_{2}^{2}\end{bmatrix}$ (C.4)
 Training or input data, output data, and training labels or target outputs are specified, with mean squared error (MSE) as the loss:
-$X=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\quad Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},\quad Y_{t}=\begin{bmatrix}y_{t1}\\ y_{t2}\end{bmatrix}\quad and\quad Y_{t}=\text{Target}(X)$ (A.5)
-$L=\text{MSE}(Y,Y_{t})=\frac{1}{2}\|Y-Y_{t}\|^{2}=\frac{1}{2}((y_{1}-y_{t1})^{2}+(y_{2}-y_{t2})^{2})$ (A.6)
+$X=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\quad Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},\quad Y_{t}=\begin{bmatrix}y_{t1}\\ y_{t2}\end{bmatrix}\text{ and }Y_{t}=Target(X)$ (C.5)
+$L=MSE\left(Y,Y_{t}\right)=\frac{1}{2}\left\|Y-Y_{t}\right\|^{2}=\frac{1}{2}(\left(y_{1}-y_{t1}\right)^{2}+\left(y_{2}-y_{t2}\right)^{2})$ (C.6)
 The gradients of the loss function $L$ with respect to the weight matrices and bias vectors are the primary basis for model M’s optimization during training.
 The four calculation formulas are as follows:
-$\nabla_{W_{2}}L=(Y-Y_{t})(\text{ReLU}(W_{1}X+b_{1}))^{T}$ (A.7)
-$\nabla_{b_{2}}L=(Y-Y_{t})$ (A.8)
-$\nabla_{W_{1}}L=(W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1}))X^{T}$ (A.9)
-$\nabla_{b_{1}}L=W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1})$ (A.10)
-In equations (A.9) and A.10, the operator $\odot$ denotes element-wise multiplication, and ReLU’ represents the derivative of the ReLU function.
-Since ReLU is not differentiable at zero, we additionally define ReLU’(0) = 0.
-Thus, the update formulas for optimizing the weight matrices and bias vectors during M’s training are:
-$W_{1}^{\text{new}}$ $=W_{1}-\eta\nabla_{W_{1}}L,\quad b_{1}^{\text{new}}=b_{1}-\eta\nabla_{b_{1}}L$ (A.11)
-$W_{2}^{\text{new}}$ $=W_{2}-\eta\nabla_{W_{2}}L,\quad b_{2}^{\text{new}}=b_{2}-\eta\nabla_{b_{2}}L$ (A.12)
+$\nabla_{W_{2}}L=(Y-Y_{t}){(ReLU(W_{1}X+b_{1}))}^{T}$ (C.7)
+$\nabla_{b_{2}}L=(Y-Y_{t})$ (C.8)
+$\nabla_{W_{1}}L=({W_{2}}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}X+b_{1}))X^{T}$ (C.9)
+$\nabla_{b_{1}}L={W_{2}}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}X+b_{1})$ (C.10)
+[…]
+$W_{1}^{\text{new}}=W_{1}-\eta\nabla_{W_{1}}L,\quad b_{1}^{\text{new}}=b_{1}-\eta\nabla_{b_{1}}L$ (C.11)
+$W_{2}^{\text{new}}=W_{2}-\eta\nabla_{W_{2}}L,\quad b_{2}^{\text{new}}=b_{2}-\eta\nabla_{b_{2}}L$ (C.12)
 where $\eta$ is the learning rate of M.
-A.2 Symbol Definitions
+C.2 Symbol Definitions
 MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle.
 Defining the symbols is a prerequisite for formal modeling.
-To this end, Tables A1A4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
-Table A1: Definition of Constants
+To this end, Tables C.1–C.4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
+Table 2: Definition of Constants
 Symbol Type Definition
@@
 $\mathbb{R}^{3}$ Space 3-dimensional real space
-Table A2: Definition of Variables
+Table 3: Definition of Variables - Part I: Logical Formulas and Basic Variables
 Symbol Type Definition
 $i,j,k$ natural number index or count
-$\varphi_{o_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s initial model
-$\psi_{o_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s initial model
-$\varphi_{t_{i}}$ WFFs well-formed formulas contained in the training informational ontological state
-$\psi_{t_{i}}$ WFFs well-formed formulas contained in the training informational carrier state
-$\varphi_{u_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s optimized model
-$\psi_{u_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s optimized model
-$\varphi_{q_{i}}$ WFFs well-formed formulas contained in the input informational ontological state of M
-$\psi_{q_{i}}$ WFFs well-formed formulas contained in the input informational carrier state of M
-$\varphi_{r_{i}}$ WFFs well-formed formulas contained in the output informational ontological state of M
-$\psi_{r_{i}}$ WFFs well-formed formulas contained in the output informational carrier state of M
-$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s re-optimized model
-$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s reoptimized model
+$\varphi o_{i}$ WFFs well-formed formulas in initial model ontological state
+$\psi o_{i}$ WFFs well-formed formulas in initial model carrier state
+$\varphi t_{i}$ WFFs well-formed formulas in training ontological state
+$\psi t_{i}$ WFFs well-formed formulas in training carrier state
+$\varphi u_{i}$ WFFs well-formed formulas in optimized model ontological state
+$\psi u_{i}$ WFFs well-formed formulas in optimized model carrier state
+$\varphi q_{i}$ WFFs well-formed formulas in input ontological state
+$\psi q_{i}$ WFFs well-formed formulas in input carrier state
+$\varphi r_{i}$ WFFs well-formed formulas in output ontological state
+$\psi r_{i}$ WFFs well-formed formulas in output carrier state
+$\varphi v_{i}$ WFFs well-formed formulas in re-optimized model ontological state
+$\psi v_{i}$ WFFs well-formed formulas in re-optimized model carrier state
 layerslayers natural number number of layers in the model
 $v$ real number numerical values for weights and biases
-$W_{k}^{(o)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the initial model
-$w_{ij}^{(o)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the initial model
-$b_{k}^{(o)}$ vector bias vector of neurons in the k-th layer in the initial model
-$b_{i}^{(o)k}$ real number bias of the i-th neuron in the k-th layer in the initial model
-$W_{k}^{(u)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the optimized model
-$w_{ij}^{(u)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the optimized model
-$b_{k}^{(u)}$ vector bias vector of neurons in the k-th layer in the optimized model
-$b_{i}^{(u)k}$ real number bias of the i-th neuron in the k-th layer in the optimized model
-$W_{k}^{(v)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the re-optimized model
-$w_{ij}^{(v)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the re-optimized model
-$b_{k}^{(v)}$ vector bias vector of neurons in the k-th layer in the re-optimized model
-$b_{i}^{(v)k}$ real number bias of the i-th neuron in the k-th layer in the re-optimized model
-$W_{k}^{(w)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the further re-optimized model
-$w_{ij}^{(w)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the further re-optimized model
-$b_{k}^{(w)}$ vector bias vector of neurons in the k-th layer in the further re-optimized model
-$b_{i}^{(w)k}$ real number bias of the i-th neuron in the k-th layer in the further re-optimized model
+Table 4: Definition of Variables - Part II: Neural Network Parameters
+Symbol Type Definition
+$W_{k}^{(o)}$ matrix weight matrix from k-th to (k+1)-th layer (initial)
+$w_{ij}^{(o)k}$ real number weight from j-th to i-th neuron (initial)
+$b_{k}^{(o)}$ vector bias vector of k-th layer (initial)
+$b_{i}^{(o)k}$ real number bias of i-th neuron in k-th layer (initial)
+$W_{k}^{(u)}$ matrix weight matrix from k-th to (k+1)-th layer (optimized)
+$w_{ij}^{(u)k}$ real number weight from j-th to i-th neuron (optimized)
+$b_{k}^{(u)}$ vector bias vector of k-th layer (optimized)
+$b_{i}^{(u)k}$ real number bias of i-th neuron in k-th layer (optimized)
+$W_{k}^{(v)}$ matrix weight matrix from k-th to (k+1)-th layer (re-optimized)
+$w_{ij}^{(v)k}$ real number weight from j-th to i-th neuron (re-optimized)
+$b_{k}^{(v)}$ vector bias vector of k-th layer (re-optimized)
+$b_{i}^{(v)k}$ real number bias of i-th neuron in k-th layer (re-optimized)
+$W_{k}^{(w)}$ matrix weight matrix from k-th to (k+1)-th layer (further re-optimized)
+$w_{ij}^{(w)k}$ real number weight from j-th to i-th neuron (further re-optimized)
+$b_{k}^{(w)}$ vector bias vector of k-th layer (further re-optimized)
+$b_{i}^{(w)k}$ real number bias of i-th neuron in k-th layer (further re-optimized)
+Table 5: Definition of Variables - Part III: Vectors and Optimization Parameters
+Symbol Type Definition
 $X$ vector input vector
@@
 $Y_{t}$ vector output reference vector
-$x_{i}$ real number the i-th element of the input vector
-$h_{i}$ real number the i-th element of the hidden layer vector
-$y_{i}$ real number the i-th element of the output vector
-$y_{ti}$ real number the i-th element of the output reference vector
+$x_{i}$ real number i-th element of input vector
+$h_{i}$ real number i-th element of hidden layer vector
+$y_{i}$ real number i-th element of output vector
+$y_{ti}$ real number i-th element of output reference vector
 $t$ time time point of state occurrence
@@
 $L$ real number loss function value
-$\nabla_{W_{1}}L$ real number gradient of the loss function with respect to $W_{1}$
-$\nabla_{b_{1}}L$ real number gradient of the loss function with respect to $b_{1}$
-$\nabla_{W_{2}}L$ real number gradient of the loss function with respect to $W_{2}$
-$\nabla_{b_{2}}L$ real number gradient of the loss function with respect to $b_{2}$
-$\eta$ Positive real number learning rate
-Table A3: Definition of Functions
+$\nabla_{W_{1}}L$ real number gradient of loss function w.r.t. $W_{1}$
+$\nabla_{b_{1}}L$ real number gradient of loss function w.r.t. $b_{1}$
+$\nabla_{W_{2}}L$ real number gradient of loss function w.r.t. $W_{2}$
+$\nabla_{b_{2}}L$ real number gradient of loss function w.r.t. $b_{2}$
+$\eta$ positive real learning rate
+Table 6: Definition of Functions
 Symbol Number of Elements Definition
-ReLU(x)(x) 1 rectified linear unit (ReLU) activation function
-Identity(x)(x) 1 identity function
-maxmax multivariate maximum function
-[…]
-ReLU’$(x)$ 1 derivative of ReLU(x)(x)
-Target(X)(X) 1 objective function for input vector $X$
-Table A4: Definition of Predicates
+$ReLU(x)$ 1 rectified linear unit (ReLU) activation function
+$Identity(x)$ 1 identity function
+$\max$ multivariate maximum function
+$MSE(Y,Y_{t})$ 2 mean squared error between vector $Y$ and $Y_{t}$
+${ReLU}^{\prime}(x)$ 1 derivative of $ReLU(x)$
+$Target(X)$ 1 objective function for input vector $X$
+Table 7: Definition of Predicates
 Symbol Definition
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-A.3 Formal Representations of Information at Each Stage of Model M
+${NetworkStructure}^{2}(layers,k)$ model structure has k layers
+${LayerSize}^{2}(k,j)$ the k-th layer has j neurons
+${Eq}^{2}(x,y)$ $x$ equals $y$
+${InputAt}^{2}(X,t)$ input $X$ at time $t$
+${HiddenAt}^{2}(H,t)$ hidden layer value is $H$ at time $t$
+${OutputAt}^{2}(Y,t)$ output $Y$ at time $t$
+${InputAt}^{3}(X,Y_{t},t)$ input vector $X$ and label vector $Y_{t}$ at time $t$
+${Differentiable}^{2}(f,v)$ function $f$ is differentiable with respect to $v$
+${BelongsTo}^{2}(x,X)$ element $x$ belongs to set $X$
+C.3 Formal Representations of Information at Each Stage of Model M
 Following MLT-MF, we build formal expressions for the following five stages:
-A.3.1 Initial information of M
+(1) Initial information of M
 According to the principles of OIT and based on MLT-MF, the model M is endowed with its initial information from the very beginning:
-$I_{0}=(oo,To_{h},S_{oo},M,To_{m},S_{M})$ (A.13)
+$Io=\left\langle oo,{To}_{h},S_{oo},M,{To}_{m},S_{M}\right\rangle$ (C.13)
 Here, the ontology $oo$ represents the designer’s conceptual blueprint; ${To}_{h}$ is the time at which the model is designed; and the ontological state $S_{oo}(oo,{To}_{h})$ is a set of well-formed formulas that describe the initial operational logic of model M:
-$S_{oo}(oo,To_{h})=\{\varphi o_{1},\varphi o_{2},\varphi o_{3},\varphi o_{4},\varphi o_{5},\varphi o_{6},\varphi o_{7},\varphi o_{8},\varphi o_{9},\varphi o_{10},\varphi o_{11}\}$ (A.14)
+$S_{oo}\left(oo,{To}_{h}\right)=\left\{\varphi o_{1},\varphi o_{2},\varphi o_{3},\varphi o_{4},\varphi o_{5},\varphi o_{6},\varphi o_{7},\varphi o_{8},\varphi o_{9},\varphi o_{10},\varphi o_{11}\right\}$ (C.14)
 The specific content of each formula is as follows:
   • Network Structure Definition
-    $\varphi o_{1}=$ $\text{NetworkStructure}^{2}{(layers,3)}\land\text{LayerSize}^{2}(1,2)\land$ (A.15)
-    $\text{LayerSize}^{2}(2,3)\land\text{LayerSize}^{2}(3,2)$
+$\varphi o_{1}={NetworkStructure}^{2}(layers,3)\land{LayerSize}^{2}(1,2)\land{LayerSize}^{2}(2,3)\land{LayerSize}^{2}(3,2)$ (C.15)
   • Weight Initialization
-    $\varphi o_{2}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2,3\})\land BelongsTo^{2}(j,\{1,2\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)1},v)))$ (A.16)
-    $\varphi o_{3}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2\})\land BelongsTo^{2}(j,\{1,2,3\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)2},v)))$ (A.17)
+$\varphi o_{2}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)1},v\right))$ (C.16)
+$\varphi o_{3}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2,3\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)2},v\right))$ (C.17)
   • Bias Initialization
-    $\varphi o_{4}=\forall iBelongsTo^{2}(i,\{1,2,3\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)1},v))$ (A.18)
-    $\varphi o_{5}=\forall iBelongsTo^{2}(i,\{1,2\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)2},v))$ (A.19)
+$\varphi o_{4}=\forall i{BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(b_{i}^{(o)1},v\right))$ (C.18)
+$\varphi o_{5}=\forall i{BelongsTo}^{2}\left(i,\left\{1,2\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(b_{i}^{(o)2},v\right))$ (C.19)
   • Activation Function Definition
-    $\varphi o_{6}=\forall x(Eq^{2}(\text{ReLU}(x),max(0,x)))\land\forall x(Eq^{2}(\text{Identity}(x),x))$ (A.20)
+$\varphi o_{6}=\forall x({Eq}^{2}(ReLU(x),max(0,x)))\land\forall x({Eq}^{2}(Identity(x),x))$ (C.20)
   • Forward Propagation Rules
-    $\varphi o_{7}=$ $\forall X\forall t\forall H(InputAt^{2}(X,t)\land BelongsTo^{2}(X,\mathbb{R}^{2})\rightarrow$
-    $\exists H(BelongsTo^{2}(H,\mathbb{R}^{3})\land HiddenAt^{2}(H,t+\delta)\land$
-    $\forall i(BelongsTo^{2}(i,\{1,2,3\})\rightarrow Eq^{2}(h_{i},\text{ReLU}\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right))))$ (A.21)
-    $\varphi o_{8}=$ $\forall H\forall t(HiddenAt^{2}(H,t+\delta)\land BelongsTo^{2}(H,\mathbb{R}^{3})\rightarrow$
-    $\exists Y(BelongsTo^{2}(Y,\mathbb{R}^{2})\land(OutputAt^{2}(Y,t+2\delta)\land$
-    $\forall i(BelongsTo^{2}(i,\{1,2\})\rightarrow Eq^{2}(y_{i},\text{Identity}\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)))))$ (A.22)
+$\begin{split}\varphi o_{7}=&\forall X\forall t({InputAt}^{2}(X,t)\land{BelongsTo}^{2}\left(X,\mathbb{R}^{2}\right)\rightarrow\exists H({BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\\ &\land{HiddenAt}^{2}(H,t+\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right)\\ &\rightarrow{Eq}^{2}(h_{i},ReLU\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right)))))\end{split}$ (C.21)
+$\begin{split}\varphi o_{8}=&\forall H\forall t({HiddenAt}^{2}(H,t+\delta)\land{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\rightarrow\exists Y({BelongsTo}^{2}\left(Y,\mathbb{R}^{2}\right)\\ &\land{OutputAt}^{2}(Y,t+2\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2\right\}\right)\\ &\rightarrow{Eq}^{2}\left(y_{i},Identity\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)\right))))\end{split}$ (C.22)
   • Backpropagation Rules
-    $\varphi o_{9}$ $=\forall X\forall Y_{t}\forall t(InputAt^{3}(X,Y_{t},t)\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2})\land\varphi o_{7}\land\varphi o_{8}\land$
-    $Eq^{2}(L,\text{MSE}(Y,Y_{t}))\land Eq^{2}(\text{MSE}(Y,Y_{t}),\frac{1}{2}\sum_{i=1}^{2}(y_{i}-y_{ti})^{2}))$ (A.23)
-    $\varphi o_{10}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{1}^{(o)})\land$
-    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{1}^{(o)})\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{2}^{(o)})\land$
-    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{2}^{(o)})\land$
-    $Eq^{2}(\nabla_{W^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})\land$
-    $Eq^{2}(\nabla_{b^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))\land$
-    $Eq^{2}(\nabla_{W^{(o)}_{2}}L,(Y-Y_{t})(\text{ReLU}(W_{1}^{(o)}X+b_{1}^{(o)}))^{T})\land Eq^{2}(\nabla_{b^{(o)}_{2}}L,(Y-Y_{t}))$ (A.24)
+$\begin{split}\varphi o_{9}=&\forall X\forall Y_{t}\forall t({InputAt}^{3}\left(X,Y_{t},t\right)\land{BelongsTo}^{2}\left(Y_{t},\mathbb{R}^{2}\right)\land\varphi o_{7}\land\varphi o_{8}\\ &{\land Eq}^{2}\left(L,MSE\left(Y,Y_{t}\right)\right)\land{Eq}^{2}\left(MSE\left(Y,Y_{t}\right),\frac{1}{2}\sum_{i=1}^{2}\left(y_{i}-y_{ti}\right)^{2}\right))\end{split}$ (C.23)
+$\varphi o_{10}$ $=\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}$
+$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{1}^{(o)})$
+$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{1}^{(o)})$
+$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{2}^{(o)})$
+$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{2}^{(o)})$
+$\quad\land{Eq}^{2}(\nabla_{W_{1}^{(o)}}L,$
+$\qquad({(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})$
+$\quad\land{Eq}^{2}(\nabla_{b_{1}^{(o)}}L,$
+$\qquad{(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))$
+$\quad\land{Eq}^{2}(\nabla_{W_{2}^{(o)}}L,(Y-Y_{t}){(ReLU(W_{1}^{(o)}X+b_{1}^{(o)}))}^{T})$
+$\quad\land{Eq}^{2}(\nabla_{b_{2}^{(o)}}L,(Y-Y_{t}))$ (C.24)
   • Learning Rules
-    $\varphi o_{11}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\varphi o_{10}\land Eq^{2}(W_{1}^{(u)},W_{1}^{(o)}-\eta\nabla_{W^{(o)}_{1}}L)\land$
-    $Eq^{2}(b_{1}^{(u)},b_{1}^{(o)}-\eta\nabla_{b^{(o)}_{1}}L)\land Eq^{2}(W_{2}^{(u)},W_{2}^{(o)}-\eta\nabla_{W^{(o)}_{2}}L)\land$
-    $Eq^{2}(b_{2}^{(u)},b_{2}^{(o)}-\eta\nabla_{b^{(o)}_{2}}L)$ (A.25)
+$\begin{split}\varphi o_{11}=&\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\varphi o_{10}\land{Eq}^{2}(W_{1}^{(u)},W_{1}^{(o)}-\eta\nabla_{W_{1}^{(o)}}L)\\ &\land{Eq}^{2}(b_{1}^{(u)},b_{1}^{(o)}-\eta\nabla_{b_{1}^{(o)}}L)\land{Eq}^{2}(W_{2}^{(u)},W_{2}^{(o)}-\eta\nabla_{W_{2}^{(o)}}L)\\ &\land{Eq}^{2}(b_{2}^{(u)},b_{2}^{(o)}-\eta\nabla_{b_{2}^{(o)}}L)\end{split}$ (C.25)
 In equation (A.14), the carrier M refers to the initially constructed model, and ${To}_{m}$ denotes the time when model M is completed. Clearly, we have:
-$\sup(To_{h})\leq\inf(To_{m})$ (A.26)
-The state $S_{M}(M,To_{m})$ is a one-to-one enabling mapping of the state $S_{oo}(oo,To_{h})$ onto M, that is,
-$S_{M}(M,To_{m})=\{\psi o_{1},\psi o_{2},\psi o_{3},\psi o_{4},\psi o_{5},\psi o_{6},\psi o_{7},\psi o_{8},\psi o_{9},\psi o_{10},\psi o_{11}\}$ (A.27)
+$sup({To}_{h})\leq inf({To}_{m})$ (C.26)
+The state $S_{M}(M,{To}_{m})$ is a one-to-one enabling mapping of the state $S_{oo}\left(oo,{To}_{h}\right)$ onto M, that is,
+$S_{M}\left(M,{To}_{m}\right)=\left\{\psi o_{1},\psi o_{2},\psi o_{3},\psi o_{4},\psi o_{5},\psi o_{6},\psi o_{7},\psi o_{8},\psi o_{9},\psi o_{10},\psi o_{11}\right\}$ (C.27)
 and
-$Io(\varphi o_{i})\underline{\triangleright}\psi o_{i},(1\leq i\leq 11)$ (A.28)
-Here, the information enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M.
+$Io\left(\varphi o_{i}\right)\trianglerighteq\psi o_{i}(1\leq i\leq 11)$ (C.28)
+Here, the information $Io$ enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M.
 In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise.
@@
 The subsequent discussions on other types of information will follow this approach and will not be repeated.
-A.3.2 Training Information of M
+(2) Training Information of M
 Based on MLT-MF, the training information for model M is as follows:
-$It=(ot,Tt_{h},S_{ot},Mt,Tt_{m},S_{Mt})$ (A.29)
+$It=\left\langle ot,{Tt}_{h},S_{ot},Mt,{Tt}_{m},S_{Mt}\right\rangle$ (C.29)
 Here, the ontology $ot$ refers to the object described by the training information, and the occurrence time ${Tt}_{h}$ denotes the point at which this object attains its corresponding state attributes.
 The ontological state $S_{ot}(ot,{Tt}_{h})$ represents the attributes of the object at time ${Tt}_{h}$, which are described by a set of well-formed formulas:
-$S_{ot}(ot,Tt_{h})=\{\varphi t_{1},\varphi t_{2}\}$ (A.30)
+$S_{ot}\left(ot,{Tt}_{h}\right)=\left\{\varphi t_{1},\varphi t_{2}\right\}$ (C.30)
   • Vector Definition
-    $\varphi t_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.31)
+$\varphi t_{1}=\forall X\forall Y_{t}({BelongsTo}^{2}(X,\mathbb{R}^{2})\land{BelongsTo}^{2}(Y_{t},\mathbb{R}^{2}))$ (C.31)
   • Label Consistency
-    $\varphi t_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.32)
+$\varphi t_{2}=\forall X\forall Y_{t}{Eq}^{2}(Target(X),Y_{t})$ (C.32)
 The carrier $Mt$ serves as the interface device through which model M receives training data, and ${Tt}_{m}$ is the time at which model M is trained.
 The carrier state $S_{Mt}(Mt,{Tt}_{m})$ is then described by the set of well-formed formulas:
-$It(\varphi t_{i})\underline{\triangleright}\psi t_{i}(1\leq i\leq 2)$ (A.33)
-A.3.3 Optimized Information After M Has Learned from Training
-Based on MLT-MF, after learning the training information $It$, model M updates its initial information $I_{0}$ to a new, optimized information state:
-$I_{u}=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ (A.34)
+$It\left(\varphi t_{i}\right)\trianglerighteq\psi t_{i}(1\leq i\leq 2)$ (C.33)
+(3) Optimized Information After M Has Learned from Training
+Based on MLT-MF, after learning the training information $It$, model M updates its initial information $Io$ to a new, optimized information state:
+$Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$ (C.34)
 Here, the ontology $ou$ is a fusion of M and the object described by the training information; the occurrence time ${Tu}_{h}$ is the moment M learns from the training data; and the ontological state $S_{ou}(ou,{Tu}_{h})$ represents the theoretical state of M after training:
-$S_{ou}(ou,Tu_{h})=\{\varphi u_{1},\varphi u_{6},\varphi u_{7},\varphi u_{8},\varphi u_{9},\varphi u_{10},\varphi u_{11}\}$ (A.35)
-Since the structure and algorithms of M remain unchanged before and after training, $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{7}$ are exactly the same as $\varphi o_{1}$, $\varphi o_{6}$, and $\varphi o_{9}$ respectively.
-Moreover, as the weights and biases after training are already assigned values in the original information (A.25), there is no need to include equivalent formulas for $\varphi o_{2}$, $\varphi o_{3}$, $\varphi o_{4}$, $\varphi o_{5}$ in $S_{ou}(ou,Tu_{h})$.
-For the remaining formulas $\varphi u_{7}$, $\varphi u_{8}$, $\varphi u_{9}$, $\varphi u_{10}$, and $\varphi u_{11}$, it suffices to systematically replace $W_{k}^{(o)}$, $w_{ij}^{(o)k}$, $b_{k}^{(o)}$, $b_{i}^{(o)k}$, $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$ and $b_{i}^{(u)k}$ in $\varphi_{7}$, $\varphi_{8}$, $\varphi_{10}$, and $\varphi_{11}$ with the corresponding symbols $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$.
+$S_{ou}\left(ou,{Tu}_{h}\right)=\left\{\varphi u_{1},\varphi u_{6},\varphi u_{7},\varphi u_{8},\varphi u_{9},\varphi u_{10},\varphi u_{11}\right\}$ (C.35)
+Since the structure and algorithms of M remain unchanged before and after training, $\varphi u_{1},\varphi u_{6}$, and $\varphi u_{9}$ are exactly the same as $\varphi o_{1},\varphi o_{6}$, and $\varphi o_{9}$ respectively.
+Moreover, as the weights and biases after training are already assigned values in the original information (A.25), there is no need to include equivalent formulas for $\varphi o_{2},\varphi o_{3},\varphi o_{4}$, or $\varphi o_{5}$ in $S_{ou}\left(ou,{Tu}_{h}\right)$.
+For the remaining formulas $\varphi u_{7},\varphi u_{8},\varphi u_{10}$, and $\varphi u_{11}$, it suffices to systematically replace $W_{k}^{(o)}$, $w_{ij}^{(o)k}$, $b_{k}^{(o)}$, $b_{i}^{(o)k}$, $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$ and $b_{i}^{(u)k}$ in $\varphi o_{7},\varphi o_{8},\varphi o_{10}$, and $\varphi o_{11}$ with the corresponding symbols $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}\left(ou,{Tu}_{h}\right)$.
 Similarly, the carrier state
-$S_{M}(M,Tu_{m})=\{\psi u_{1},\psi u_{6},\psi u_{7},\psi u_{8},\psi u_{9},\psi u_{10},\psi u_{11}\}$ (A.36)
-is a one-to-one enabling mapping from $S_{ou}(ou,Tu_{h})$ onto M, representing the state of M after completing its training.
-A.3.4 Input Information for M
+$S_{M}\left(M,{Tu}_{m}\right)=\left\{\psi u_{1},\psi u_{6},\psi u_{7},\psi u_{8},\psi u_{9},\psi u_{10},\psi u_{11}\right\}$ (C.36)
+is a one-to-one enabling mapping from $S_{ou}\left(ou,{Tu}_{h}\right)$ onto M, representing the state of M after completing its training.
+(4) Input Information for M
 Processing input information is a fundamental requirement for any model. Based on the MLT-MF framework, the input information for M is defined as:
-$I_{q}=(oq,Tq_{h},S_{oq},Mq,Tq_{m},S_{Mq})$ (A.37)
-Here, $oq$ denotes the object described by the input information, and $Tq_{h}$ is the time at which the object’s state occurs. The state $S_{oq}(oq,Tq_{h})$ can take two possible forms:
+$Iq=\left\langle oq,{Tq}_{h},S_{oq},Mq,{Tq}_{m},S_{Mq}\right\rangle$ (C.37)
+Here, $oq$ denotes the object described by the input information, and ${Tq}_{h}$ is the time at which the object's state occurs. The state $S_{oq}(oq,{Tq}_{h})$ can take two possible forms:
   • The first scenario involves only input information being processed by M, in which case
-    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1}\}$ (A.38)
-    $\varphi q_{1}=\forall XBelongsTo^{2}(X,\mathbb{R}^{2})$ (A.39)
+$S_{oq}(oq,{Tq}_{h})=\left\{\varphi q_{1}\right\}$ (C.38)
+$\varphi q_{1}=\forall X{BelongsTo}^{2}(X,\mathbb{R}^{2})$ (C.39)
   • The second scenario considers online learning, where
-    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1},\varphi q_{2}\}$ (A.40)
-    $\varphi q_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.41)
-    $\varphi q_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.42)
-The carrier $Mq$ for $I_{q}$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$.
+$S_{oq}(oq,{Tq}_{h})=\left\{\varphi q_{1},\varphi q_{2}\right\}$ (C.40)
+$\varphi q_{1}=\forall X\forall Y_{t}({BelongsTo}^{2}(X,\mathbb{R}^{2})\land{BelongsTo}^{2}(Y_{t},\mathbb{R}^{2}))$ (C.41)
+$\varphi q_{2}=\forall X\forall Y_{t}{Eq}^{2}(Target(X),Y_{t})$ (C.42)
+The carrier $Mq$ for $Iq$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$.
 The state $S_{Mq}(Mq,{Tq}_{m})$ is a one-to-one enabling mapping from $S_{oq}(oq,{Tq}_{h})$ onto $Mq$.
-A.3.5 Information After M Processes Input, Produces Output, and Undergoes Further Optimization
-There are two scenarios for M after it processes input and produces output: one corresponds to the first case in A.3.4, where only output is produced without further optimization of the model itself; the other corresponds to the second case in A.3.4, where, in the context of online learning, the model both produces output and undergoes further self-optimization.
+(5) Information After M Processes Input, Produces Output, and Undergoes Further Optimization
+There are two scenarios for M after it processes input and produces output: one corresponds to the first case in (4), where only output is produced without further optimization of the model itself; the other corresponds to the second case in (4), where, in the context of online learning, the model both produces output and undergoes further self-optimization.
 Since the latter is more comprehensive, we will focus on this scenario below.
-In this case, the output information $Ir=(or,Tr_{h},S_{or},Mr,Tr_{m},S_{Mr})$ has an ontological state defined as:
-$S_{or}(or,Tr_{h})$ $=\{\varphi r_{1}\}=\{\varphi o_{7}\land\varphi o_{8}\land\text{OutputAt}^{2}(Y,Tr_{h})\}$ (A.43)
-The carrier state $S_{Mr}(Mr,Tr_{m})$ is a one-to-one enabling mapping from $S_{or}(or,Tr_{h})$ to the output device of M.
+In this case, the output information $Ir=\left\langle or,{Tr}_{h},S_{or},Mr,{Tr}_{m},S_{Mr}\right\rangle$ has an ontological state defined as:
+$S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r_{1}\right\}=\left\{{\varphi o_{7}\land\varphi o_{8}\land OutputAt}^{2}(Y,{Tr}_{h})\right\}$ (C.43)
+The carrier state $S_{Mr}(Mr,{Tr}_{m})$ is a one-to-one enabling mapping from $S_{or}\left(or,{Tr}_{h}\right)$ to the output device of M.
 On the other hand, since we are considering the scenario where M performs online learning, the model now carries further optimized information:
-$Iv=(ov,Tv_{h},S_{ov},M,Tv_{m},S_{M})$ (A.44)
+$Iv=\left\langle ov,{Tv}_{h},S_{ov},M,{Tv}_{m},S_{M}\right\rangle$ (C.44)
 This structure is almost identical to the optimized information $Iu$ after M has learned from training. The ontological state of $Iv$ is:
-$S_{ov}(ov,Tv_{h})=\{\varphi v_{1},\varphi v_{6},\varphi v_{7},\varphi v_{8},\varphi v_{9},\varphi v_{10},\varphi v_{11}\}$ (A.45)
+$S_{ov}\left(ov,{Tv}_{h}\right)=\left\{\varphi v_{1},\varphi v_{6},\varphi v_{7},\varphi v_{8},\varphi v_{9},\varphi v_{10},\varphi v_{11}\right\}$ (C.45)
 Here, $\varphi v_{1}$, $\varphi v_{6}$, and $\varphi v_{9}$ are identical to $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{9}$ respectively.
-For the remaining formulas, it suffices to systematically replace the parameters $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$ with the corresponding parameters $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$, $b_{i}^{(v)k}$, $W_{1}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(w)}$ and $b_{i}^{(w)k}$ to obtain the updated formulas, where $W_{k}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(2)}$ and $b_{i}^{(w)k}$ are the weights and biases of M after further optimization.
+For the remaining formulas, it suffices to systematically replace the parameters
+$W_{k}^{(u)},w_{ij}^{(u)k},b_{k}^{(u)},b_{i}^{(u)k},W_{k}^{(v)},w_{ij}^{(v)k},b_{k}^{(v)},b_{i}^{(v)k}$
+in $S_{ou}(ou,{Tu}_{h})$ with the corresponding parameters
+$W_{k}^{(v)},w_{ij}^{(v)k},b_{k}^{(v)},b_{i}^{(v)k},W_{k}^{(w)},w_{ij}^{(w)k},b_{k}^{(w)},b_{i}^{(w)k}$
+to obtain the updated formulas, where $W_{k}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(w)}$, and $b_{i}^{(w)k}$ are the weights and biases of M after further optimization.
 Similarly, the carrier state
-$S_{M}(M,Tv_{m})=\{\psi v_{1},\psi v_{6},\psi v_{7},\psi v_{8},\psi v_{9},\psi v_{10},\psi v_{11}\}$ (A.46)
-is a one-to-one enabling mapping from $S_{ov}(ov,Tv_{h})$ onto M.
-At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Figure A1.
-Although M is a simple model, it contains all the essential components of a neural network, and the discussion has not relied on any special treatment of M.
+$S_{M}\left(M,{Tv}_{m}\right)=\left\{\psi v_{1},\psi v_{6},\psi v_{7},\psi v_{8},\psi v_{9},{\psi v}_{10},\psi v_{11}\right\}$ (C.46)
+is a one-to-one enabling mapping from $S_{ov}\left(ov,{Tv}_{h}\right)$ onto M.
+At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig.
+C.1.
+Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M.
 This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.

### v11

@@
 Another perspective highlights the central role of model parameters (Alquier, 2024), describing machine learning as the process of learning from an object-label set $(X,Y)$ governed by a probability distribution $P$, to produce a family of predictors $\left\{f_{\theta},\theta\in\Theta\right\}$ parameterized by $\theta\in\Theta$.
-The loss function $\mathcal{l(}f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\mathcal{[l(}f_{\theta}(X),Y)]$.
+The loss function $L(f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\ [L(f_{\theta}(X),Y)]$.
 The ultimate goal is to find $\theta^{*}$ such that $R\left(\theta^{*}\right)=\inf_{\theta\in\Theta}R(\theta)$.
@@
 In reality, the objective of interpretability may simply be to extract more valuable information from models.
-Reference (Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered.
+(Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered.
 Based on a definition of comprehensibility encompassing simulatability, decomposability, and algorithmic transparency, the authors emphasize that interpretability in reinforcement learning relies on three pillars: interpretable inputs, interpretable models, and interpretable decisions.
@@
 Concerns regarding the ethical safety of AI have a long history.
-As noted in [10], the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies.
+As noted in (Russell et al., 2015), the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies.
 These issues span law, public policy, professional ethics, and philosophical morality, necessitating the expertise of computer scientists, legal scholars, political scientists, and ethicists.
 While mathematical tools such as formal logic, probability theory, and decision theory have provided important foundations for reasoning and decision-making, numerous challenges remain unresolved.
-Building on traditional concepts of assurance, [11] offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence.
+Building on traditional concepts of assurance, (Burr and Leslie, 2023) offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence.
 Accordingly, the general structure of moral assurance is characterized as an iterative process of reflection, action, and adjustment, repeated throughout the design, development, and deployment lifecycle.
 In the context of AI/ML-supported healthcare, ethical safety assurance faces more concrete challenges.
-Reference [12] provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare.
+(Tilala et al., 2024) provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare.
 The authors advocate for multiple strategies, including the implementation of rigorous data security protocols, prioritization of data minimization, increasing diversity in training datasets, use of data augmentation to address imbalance, application of algorithmic fairness techniques to detect and mitigate model bias, regular auditing of learning algorithms, enhancing the interpretability of algorithmic outputs, regulatory oversight based on standardized performance metrics and stringent testing protocols, and the augmentation—rather than replacement—of clinicians’ capabilities as means to achieve ethical safety.
-[13] further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety.
+(Hawkins et al., 2021) further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety.
 The authors propose that ML safety requirements should be derived from the system-level safety requirements, and that these should be evaluated in light of the allocated safety needs, system and software architecture, and operational environment.
 This framework then serves as the basis for developing safety assurance specifications across all stages of the ML lifecycle, including data preparation, model training, model validation, and model deployment.
-Reference [14] conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches.
+(Pant et al., 2024) conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches.
 The study highlights that, in the absence of a universally adopted set of ethical principles for AI, incorporating morality into AI is fundamentally challenged by the lack of clear moral knowledge and the ambiguity of ethical principles.
@@
 Model performance is fundamentally reflected in the errors it may incur on specific tasks.
-Reference (Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions.
+(Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions.
 Based on VC dimension, it further establishes distribution-independent risk upper bounds and sufficient conditions for fast convergence under both bounded and unbounded loss in finite-sample settings.
@@
 These theoretical foundations have led to practical implementations such as support vector machines (SVM), margin maximization, and kernel methods, offering explanations for their generalization mechanisms.
-Reference (Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error.
+(Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error.
 It provides finite-sample error upper bounds via the fixed point of sub-root functions, thus achieving optimal or near-optimal convergence rates inversely proportional to the square root of the number of training samples.
@@
 The paper further analyzes two classes of algorithms—Gibbs learning and noisy ERM—where $I(S;W)$ is explicitly controlled via mutual information regularization, and discusses methods for decomposing and additively controlling mutual information in complex learning algorithms through adaptive combinations.
-Reference (Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence.
+(Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence.
 It introduces a task distribution T and task-level MDPs, defines generalization error as the performance gap between training and unseen tasks under a meta-policy, and establishes generalization upper bounds determined by the number of training tasks and task complexity, using empirical process theory and concentration inequalities.
@@
 The theoretical results are validated and the observability of convergence rates is demonstrated through controllable task generation and baseline meta-algorithms.
-Reference (Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models.
+(Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models.
 From both theoretical and numerical perspectives, it investigates the convergence properties of deterministic samplers based on probability flow ordinary differential equations (ODEs).
@@
 3.1 Definition of Information
-Reference (Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states.
-It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Xu, 2024), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times.
+(Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states.
+It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Qiu and Xu, 2025), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times.
 Here, we proceed to further clarify the concept of enabling mapping.
@@
 Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I=\left\langle o,T_{h},S_{c},c,T_{m},S_{o}\right\rangle$.
-The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Xu et al., 2025).
+The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Wiener, 2019).
 It should be noted that both $S_{o}\left(o,T_{h}\right)$ and $S_{c}\left(c,T_{m}\right)$ can each be represented as the interpretation of a set of WFFs within a formal system.
 However, the respective formal systems need not be identical, which constitutes one of the essential characteristics of information—its capacity to establish connections among all entities in the world, regardless of differences in their underlying formal systems.
-Reference (Shannon, 1948) is regarded as a foundational work in information theory.
+(Shannon, 1948) is regarded as a foundational work in information theory.
 The paper proposed that ” $H=-\sum{p_{i}\log p_{i}}$ plays a central role as the measure of information in information theory.” The discussion in this paper includes both the source and channel of information.
@@
 3.3 Information Chain
-Reference (Xu et al., 2014) discusses the transmissibility of information through composite mappings.
+(Xu et al., 2014) discusses the transmissibility of information through composite mappings.
 Based on Definitions 1 and 2, the fundamental laws governing information transmission can be characterized more rigorously as follows:
@@
 This demonstrates the model's capacity for learning and optimization, where $M_{I}$ denotes the input device of $M$.
-$\begin{split}{\varphi o}_{3}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({{Eq}^{2}\left(S_{M_{O}}\left(M_{O},t_{2}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\right)\land Output}^{3}(M,t_{2},S_{M_{O}}\left(M_{O},t_{2}\right)))\end{split}$ (4.4)
+$\varphi_{O_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\wedge Input^{3}(M_{D},t_{1},S_{M},(M_{I},t_{1}))$ (4.4)
+$\wedge Processable^{2}(M,S_{M},(M_{I},t_{1}))\rightarrow\exists f^{2}_{process}(M,S_{M},(M_{I},t_{1}))$
+$(Eq^{2}(S_{M_{0}}(M_{0},t_{2}),f^{2}_{process}(M,S_{M},(M_{I},t_{1}))))\wedge Output^{3}(M,t_{2},S_{M_{0}}(M_{0},t_{2}))$
 indicates that, at time $t_{1}$, if $S_{M_{I}}\left(M_{I},t_{1}\right)$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate ${Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$, then $M_{D}$ can process $S_{M_{I}}\left(M_{I},t_{1}\right)$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$ to yield the output carrier state $S_{M_{O}}\left(M_{O},t_{2}\right)$, which is then output at time $t_{2}$.
@@
 It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas.
-Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states.
+Regardless of the situation, both cases satisfy Definition B.1 concerning finite automaton states.
 Therefore, subsequent discussions will encompass both cases without further elaboration.
@@
 shows that after the first learning step, $M$ remains capable of further learning and optimization.
-$\begin{split}{\varphi u}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))))\end{split}$ (4.14)
+${\varphi u}_{3}=$ $State^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\wedge Input^{3}\left(M,t_{2},S_{M_{I}}(M_{I},t_{2})\right)$ (4.14)
+$\wedge Processable^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)\rightarrow\exists f_{process}^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)$
+$\left(Eq^{2}\left(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))\right)\wedge Output^{3}\left(M,t_{3},S_{M_{O}}(M_{O},t_{3})\right)\right)$
 indicates that when the training information $It$ is processable by $M$, it can also produce output information.
@@
 According to Definition 4, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$, and $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
-Let
-$S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$ $=E_{M},$
-$S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$ $=E_{M_{I}},$
-$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$ $=E_{M_{O}}$
-Then, according to Definition 6, there exist information objects
-$Iu(t_{i})$ $=\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle,$
-$Iq(t_{i})$ $=\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle,$
-$Ir(t_{i+1})$ $=\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$
-such that
-$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
-$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$
+Let $S_{on}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{op}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$ and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$.
+Then, according to Definition 4, there exist information objects $Ig(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{op},M_{I},t_{i},S_{M_{I}}\rangle$ and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq Cn(S_{on}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$.
 Thus, the necessity of the first part of the theorem is established.
-On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$,
-$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
-$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$
+On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$, $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$.
 Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$.
@@
 $p_{M}(x)=\begin{cases}\approx p_{q}(x),&\text{if }x\in S_{ou}\cap S_{oq}\\[3.00003pt] \frac{1-P_{Mq}}{|S_{oq}|-|S_{ou}\cap S_{oq}|},&\text{if }x\in S_{oq}\backslash S_{ou}\end{cases}$ (5.12)
-Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Kullback and Leibler, 1951):
+Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Goodfellow et al., 2016):
 $D_{KL}(p_{q}(x)\parallel p_{M}(x))$ $=\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
@@
 Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance.
-By Pinsker’s inequality (Pinsker, 1964), we have: …
+By Pinsker’s inequality (Pinsker, 1964), we have:
 $TVD(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\parallel p_{M}(x))}$
@@
 The noise loss probability mass is $P_{l}=\sum_{x\in S_{oq}^{-}}{p_{q}(x)}\in[0,1]$, and the noise addition probability mass is $P_{a}=\sum_{x\in S_{oq}^{+}}{p^{+}(x)}\in[0,1]$, where $p^{+}$ is a non-negative kernel on the noise addition region, satisfying $1-P_{l}+P_{a}>0$.
-The distribution of the noisy input $\widetilde{S_{oq}}$ is given by $\widetilde{p_{q}}(x)=(p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\})/(1-P_{l}+P_{a})$.
+The distribution of the noisy input $\widetilde{S_{oq}}$ is given by $\widetilde{p_{q}}(x)=(p_{q}(x)\cdot 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\cdot 1\left\{x\in S_{oq}^{+}\right\})/(1-P_{l}+P_{a})$.
 Let $V=\widetilde{S_{oq}}\backslash S_{ou}$ denote the region not covered by training information, with uncovered probability mass $P_{uncov}=\sum_{x\in V}{\widetilde{p_{q}}(x)}\in[0,1]$.
@@
 Since the distribution of the noisy input $\widetilde{S_{oq}}$ is
-$\widetilde{p_{q}}(x)=\frac{p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\}}{1-P_{l}+P_{a}}$ (5.19)
+$\widetilde{p_{q}}(x)=\frac{p_{q}(x)\cdot 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\cdot 1\left\{x\in S_{oq}^{+}\right\}}{1-P_{l}+P_{a}}$ (5.19)
 It follows that
@@
 Fig.
-3: (a) The training dataset $S_{ou}$ (selected within the blue square) and the test dataset $S_{oq}$ (selected within the red circle) are the same as those used in the experiment described in Section 6.1.
+4: (a) The training dataset $S_{ou}$ (selected within the blue square) and the test dataset $S_{oq}$ (selected within the red circle) are the same as those used in the experiment described in Section 6.1.
 The upper arc-shaped segment of $S_{oq}$ is shifted upward, and the vacated region, denoted as $S_{oq}^{-}$, is excluded from the test input.
@@
 This visualization also represents an ontological state of M as it handles a wide range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping.
-Thus, according to Definition 4, Figure 5(a) constitutes a valid explanation for the distribution of prediction polarities by M.
+Thus, according to Definition 4, Fig.
+6(a) constitutes a valid explanation for the distribution of prediction polarities by M.
 6.4.4 Explanation Based on Cumulative Training Effects
@@
 All datasets and codes used in this work are available at https://github.com/xuhu6736/MLT-MF.
-Appendix A Appendix A. Formal Representation of States
+Appendix A Formal Representation of States
 The postulates of information emphasize that the state of the ontology enables a mapping to the state of the carrier.
@@
 In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
-Appendix B Appendix B. Finite Automaton States
-Using the formal expression of states, we can study in depth the special and important object of finite automata (Xu, 2024).
+Appendix B Finite Automaton States
+Using the formal expression of states, we can study in depth the special and important object of finite automata (Editorial Committee of Dictionary of Mathematics, 2002).
 Lemma B.1 (State Representation of Finite Automata) For any finite automaton $M$, one can describe its input, output, and state‐transition behavior over a related time set $T$ by the state set $S(M,T)$.
@@
 Therefore, the concept of finite automaton state is of great significance in the study of machine learning.
-Appendix C Appendix C. Model Instance Validation of MLT-MF
+Appendix C Model Instance Validation of MLT-MF
 To facilitate validation and illustration of the practical significance of MLT-MF and related theorems, we select a very simple model and use MLT-MF to establish formal representations with WFFs for each stage from initial information to training information, learning optimization, and processing/output.
@@
 We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig.
-C.1).
+7).
 Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
-Figure 7: Fig. C.1 Schematic diagram of feedforward neural network model M
+Figure 7: Schematic diagram of feedforward neural network model M
 The hidden layer uses the ReLU activation function:
@@
 Defining the symbols is a prerequisite for formal modeling.
-To this end, Tables C.1–C.4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
+To this end, Tables 2–7 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
 Table 2: Definition of Constants
@@
 At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig.
-C.1.
+7.
 Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M.

### v12

@@
 This paper addresses the current lack of a unified formal framework in machine learning theory, as well as the absence of robust theoretical foundations for interpretability and ethical safety assurance.
-We first construct a formal information model, employing sets of well-formed formulas (WFFs) to explicitly define the ontological states and carrier mappings for the core components of machine learning.
+We first construct a formal information model that employs sets of well-formed formulas (WFFs) to explicitly define the ontological states and the carrier mappings of core ML components.
 By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the logical inference and constraint rules underlying causal chains in models, thereby establishing the Machine Learning Theory Meta-Framework (MLT-MF).
-Building upon this framework, we propose universal definitions for model interpretability and ethical safety, and rigorously prove and validate four key theorems: the equivalence between model interpretability and information existence, the constructive formulation of ethical safety assurance and two types of total variation distance (TVD) upper bounds.
-This work overcomes the limitations of previous fragmented approaches, providing a unified theoretical foundation from an information science perspective to systematically address the critical challenges currently facing machine learning.
-Keywords: machine learning, objective information theory (OIT), formal logic systems, interpretability, ethical safety assurance, estimation of error bound
+Building upon this framework, we propose universal definitions for model interpretability and ethical safety, and rigorously prove four key theorems: (i) the equivalence between model interpretability and information existence; (ii) a constructive formulation of ethical safety assurance; and (iii–iv) two upper bounds on total variation distance (TVD).This work overcomes the limitations of previous fragmented approaches, providing a unified theoretical foundation from an information science perspective to systematically address the critical challenges currently facing machine learning.
+keywords:
+machine learning, objective information theory (OIT), formal logic systems, interpretability, ethical safety assurance, error bound estimation
+journal: Information Sciences
+[…]
+organization=Koguan School of Law, China Institute for Smart Justice, School of Computer Science, Shanghai Jiao Tong University, city=Shanghai, postcode=200030, country=China
 1 Introduction
-Although artificial intelligence (AI), exemplified by machine learning (ML), is among the most dynamic areas in information science and technology, several critical shortcomings persist in the current research landscape.
-First, there is a lack of a unified logical framework that spans the entire lifecycle of machine learning.
-Most existing ML theory focuses on isolated components: optimization theory emphasizes training dynamics, generalization analysis targets test performance, while initialization strategies and deployment are often considered independently.
-This fragmentation precludes the coherent characterization of knowledge evolution from prior injection, data-driven updates, to task-level generalization (Vapnik, 1999; Alquier, 2024; Schölkopf et al., 2021; Garcez and Lamb, 2023; Marra et al., 2024).
-Second, the unpredictability of model performance has led to rapidly escalating research and development costs.
-As model sizes continue to grow, the demands for data, computation, energy, time, and human resources have multiplied.
-In the absence of explicit causal semantics, structural analysis, and effective metrics, it is often challenging to analyze or predict attainable performance in each stage of the pipeline based on error upper bounds.
-Incomplete information coverage, suboptimal model selection, inaccurate measurement, and quality deficiencies frequently result in failed development efforts and significant resource losses (Achiam et al., 2023; Touvron et al., 2023; Chowdhery et al., 2023; Gunasekar et al., 2023).
-Third, there are profound concerns regarding ethical safety assurance.
-As AI capabilities advance, so too do apprehensions about their ethical implications.
-However, there is still no consensus—either mathematical or theoretical—on the scientific definition, substantive content, forms of expression, or criteria for ethical safety in AI.
-Most research remains focused on ethical principles and regulatory mechanisms, while the proliferation of competing guidelines leaves practitioners uncertain about which standards to follow (Russell et al., 2015; Burr and Leslie, 2023; Tilala et al., 2024; Hawkins et al., 2021; Pant et al., 2024).
-Fourth, the theoretical foundations for interpretability remain weak.
-The lack of a unified semantic framework for causal inference and constraint logic means that traditional ML methods treat the transformation from data to model as a black box.
-Although interpretability is widely recognized as a critical issue and diverse approaches have been proposed to address it, there is as yet no rigorous definition or universally applicable evaluative criteria (Lipton, 2018; Glanois et al., 2024; Zhang et al., 2024; Rudin, 2019; Miller, 2019).
-As a central discipline bridging data, knowledge, and systems, information science has begun to offer new perspectives for the theoretical integration of machine learning (Shwartz-Ziv and Tishby, 2017; Xu et al., 2022).
-However, information science theory remains fragmented across information theory, control theory, signal processing, statistical learning, and computational theory, each developing largely in isolation and lacking a unified semantic and logical system.
-This fragmentation is a primary reason for the absence of a unified theoretical framework in contemporary AI.
-In response, Objective Information Theory (OIT) applies an axiomatic approach to construct a sextuple model of information, representing intrinsic meaning via ontological states, objective expression via carrier states, and revealing the dual-subject structure of information through the enabling mapping from the former to the latter.
-OIT thus provides a general and universal foundational model and metrics for understanding information in the objective world, and especially within computing and information systems (Xu et al., 2014; Xu, 2024).
-For example, (Xu et al., 2025) demonstrates that applying OIT’s general information metrics to pre-evaluate training data can significantly enhance model training efficiency.
-This paper further defines the enabling mapping, clarifying its dual mathematical and physical-causal attributes.
-Through a recursively defined higher-order logical system, we formalize the expression of states via the interpretation of sets of well-formed formulas (WFFs), elucidating the mathematical essence of state representation in information definitions.
-This fills the final theoretical gap in OIT and offers a more powerful toolkit for applying OIT to diverse problems in information science.
-The introduction of noisy information and information chains further establishes the theoretical groundwork for analyzing key challenges across all stages of machine learning.
-On this basis, we comprehensively apply OIT’s theories and methodologies—centered on the learnable and processable predicates, as well as learning and processing functions—to five canonical stages in machine learning: model construction, training information, learning and training, input of problems, information processing outputs and online learning.
-We establish a unified semantic formalization of state representation, and develop causal-chain logical inference and constraint rules covering the full ML lifecycle, culminating in the Meta-Framework for Machine Learning Theory (MLT-MF).
-Applying MLT-MF, we formulate universal definitions for model interpretability and ethical safety, and prove theorems relating interpretability to information existence, as well as the theoretical underpinnings for ethical safety assurance.
-For model error, we establish total variation distance (TVD) upper bounds for two strategies—conservative prediction in regions not covered by training with noiseless input, and arbitrary prediction with noisy input—providing new guidelines for the estimation and optimization of machine learning performance.
-Finally, we illustrate and validate the practical utility of MLT-MF and the associated theorems with a simple feedforward neural network model, demonstrating that this research offers a novel direction and methodology for both theoretical analysis and engineering practice in machine learning, grounded in a unified formal logical framework.
+Machine learning (ML) has progressed rapidly, yet several foundational gaps persist.
+The field still lacks a unified logical framework that spans the full lifecycle from prior specification and data-driven learning to deployment and online adaptation.
+Existing theory remains fragmented across training dynamics, initialization, test-time generalization, and deployment, which prevents a coherent account of knowledge evolution across stages [1, 2, 3, 4, 5].
+Model performance is also hard to predict, which inflates development costs in data, computation, energy, and human effort.
+In the absence of explicit causal semantics, structural analysis, and stage-wise error metrics, attainable performance levels and error upper bounds are difficult to estimate, and information coverage gaps or failure modes due to model selection, measurement, and data quality are hard to anticipate [6, 7, 8, 9].
+Ethical safety lacks a settled mathematical definition and verification protocol.
+Current research is dominated by principles and heterogeneous guidelines, leaving uncertainty about standards and assurance processes [10, 11, 12, 13, 14].
+Interpretability similarly lacks a rigorous, domain-agnostic foundation.
+Without a unified semantic framework for causal inference and constraint logic, the data-to-model pipeline is treated as a black box, and general, testable definitions and criteria remain elusive despite numerous approaches [15, 16, 17, 18, 19].
+Information science offers a route to unification [20, 21], but its subfields—information theory, control, signal processing, statistical learning, and computation—evolved largely in isolation and lack a common semantics.
+Objective Information Theory (OIT) addresses this by axiomatizing information as an enabling mapping from ontological states to carrier states.
+Intrinsic meaning is captured by ontological states and objective expression by carrier states, and the mapping has both mathematical and physical–causal attributes [22, 23].
+Prior work shows that OIT’s general metrics can improve data evaluation and training efficiency in machine learning [24].
+We complete a remaining gap by formalizing state as an interpretation of well-formed formulas (WFFs) in a higher-order logical system, and by clarifying that enabling mappings are surjective mathematical mappings that are also physically realizable causal relations.
+We further introduce noisy information and information chains, which support analysis of multi-stage pipelines and their causal dependencies.
+Building on this foundation, we develop the Machine Learning Theory Meta-Framework (MLT-MF), a unified information–logical semantics for the ML lifecycle. Our main contributions are:
+• Unified lifecycle semantics.
+We formalize five canonical stages—model construction, training information, learning and training, problem input, and processing/output with online learning—via state objects and enabling mappings that compose into information chains.
+• Logical definitions with verifiable criteria.
+We give domain-general, testable definitions of interpretability and ethical safety within the same semantics.
+We prove a necessary and sufficient condition linking interpretability to information existence along the chain and construct a maximal safe-output (SafeMax) mechanism validated via a hypergraph independent-set formulation.
+• Error control with observable quantities.
+We derive explicit total variation distance (TVD) upper bounds in two regimes: conservative prediction in regions not covered by training under noise-free input, and arbitrary prediction under noisy input.
+The bounds are driven by measurable masses—overlap, uncovered, and noise (loss/addition)—and align with engineering indicators such as coverage and Kullback–Leibler (KL) divergence [1, 25, 26, 27, 28, 29, 30, 31, 32].
+• Practical instantiation.
+On a simple feedforward network, the framework yields multi-modal interpretability through mathematical, visual, and partition-based explanations, guarantees ethical safety via SafeMax, and performance estimates via TVD bounds, illustrating deployable assurance across stages.
+The remainder of the paper is organized as follows.
+Section 2 reviews related work on learning theory, interpretability, ethical safety, and error estimation, and situates OIT as a unifying lens.
+Section 3 presents theoretical basis: definition of information, noisy information, information chains, and the finite-automaton view.
+Section 4 introduces MLT-MF, defining stages, predicates, and functions, and specifying lifecycle constraints and inference rules with a minimal pipeline instantiation.
+Section 5 provides formal definitions and theorems for interpretability and ethical safety, the SafeMax mechanism, and the TVD bounds.
+Section 6 reports the neural-network case study and observable metrics.
+Section 7 discusses limitations and future directions.
 2 Related Work
-2.1 Conceptual Representations of Machine Learning
-Generally, machine learning refers to methods that learn general rules from limited observational data and use these rules to predict unknown information.
-Numerous works have provided more detailed analyses of this concept.
-Representative views include:
-In the context of statistical learning (Vapnik, 1999), the learning model is decomposed into three components: a set of independent random vectors $x$ that follow a certain but unknown probability distribution $P(x)$; a set of output vectors $y$ corresponding to input $x$, also following a certain but unknown conditional probability distribution $P\left(y\middle|x\right)$; and a collection of functions $f(x,\alpha),\alpha\in\Lambda$, that the learning machine can implement.
-Machine learning is thus defined as selecting, based on a training dataset composed of $l$ independent observations $\left(x_{1},y_{1}\right),\cdots,(x_{l},y_{l})$ drawn from the same distribution $P(x,y)=P(x)P\left(y\middle|x\right)$, a function $f(x,\alpha)$(with $\alpha\in\Lambda$) that minimizes the risk function $R(\alpha)=\int{L(y,f(x,\alpha))d}P(x,y)$, where $L(y,f(x,\alpha))$ is the loss function of $f(x,\alpha)$.
-Another perspective highlights the central role of model parameters (Alquier, 2024), describing machine learning as the process of learning from an object-label set $(X,Y)$ governed by a probability distribution $P$, to produce a family of predictors $\left\{f_{\theta},\theta\in\Theta\right\}$ parameterized by $\theta\in\Theta$.
-The loss function $L(f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\ [L(f_{\theta}(X),Y)]$.
-The ultimate goal is to find $\theta^{*}$ such that $R\left(\theta^{*}\right)=\inf_{\theta\in\Theta}R(\theta)$.
-From a more advanced perspective, the focus shifts to causal relationships beyond data distribution features (Schölkopf et al., 2021).
-This view treats perceptual data acquired through measurement as entangled observations of unknown causal system states.
-It seeks to learn concrete representations that at least partially reveal these hidden causal structures by leveraging large-scale training, data augmentation, self-supervised learning, and robust fine-tuning using data from multiple simulated environments.
-This includes discovering causal models implicit in environmental data and performing causal inference to plan and act based on these models.
-Some researchers emphasize the integration of neural and symbolic artificial intelligence (Garcez and Lamb, 2023), arguing that knowledge should be rooted in efficient data learning via message passing in neural networks and embedded in vector representations as efficient computational models.
-Symbols, in turn, should be acquired through querying and extracting knowledge from trained networks, providing a rich descriptive language at appropriate levels of abstraction—allowing finite means to achieve infinite ends—while enabling compositional and discrete reasoning at the symbolic level, thereby permitting extrapolation beyond data distributions.
-The combination of learning and reasoning should offer an important alternative for tackling compositional reasoning problems, reducing the number of effective combinations through learning and thus generating simpler symbolic descriptions within a neuro-symbolic cycle.
-Another viewpoint proposes that structure learning through parameter learning is the most prominent paradigm among the statistical relational learning approaches within neuro-symbolic AI (Marra et al., 2024).
-The learning task involves estimating probabilities for a fixed set of clauses, with the advantage of removing combinatorial search from the learning process.
-The above works collectively reflect, from different perspectives, the academic community’s comprehensive understanding of both past achievements and future directions in machine learning.
-However, none have established a formalized expression for the entire lifecycle implementation process of machine learning under a unified semantic framework.
-As a result, it remains consistently difficult to apply structured systems analysis methods to guide the implementation and optimization of models throughout research and design.
-2.2 Model Interpretability
-The interpretability of machine learning has attracted significant attention in recent years.
-As noted in (Lipton, 2018), few authors in the academic literature are able to precisely articulate what interpretability actually means, or to clearly explain how their proposed solutions achieve it.
-Despite the absence of a unified definition, an increasing number of publications continue to introduce algorithms purported to possess interpretability.
-This situation supports two possible conclusions: first, that there is a universally accepted understanding of interpretability that simply remains unwritten; or second, that the term itself is inherently ambiguous, leading to quasi-scientific claims regarding the interpretability of various models.
-In reality, the objective of interpretability may simply be to extract more valuable information from models.
-(Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered.
-Based on a definition of comprehensibility encompassing simulatability, decomposability, and algorithmic transparency, the authors emphasize that interpretability in reinforcement learning relies on three pillars: interpretable inputs, interpretable models, and interpretable decisions.
-They further analyze the specific mechanisms for achieving interpretability in each of these aspects.
-In (Zhang et al., 2024), the impact of different training mechanisms on two key factors—model smoothness and gradient similarity—is investigated from the perspective of agents, with the aim of deepening our understanding of the dynamics of model transferability and its relationship to various training methods.
-Building on adversarial training, two working hypotheses are proposed and validated: first, that the trade-off between model smoothness and gradient similarity largely determines the adversarial transferability of agent models; and second, that shifts in data distribution can undermine gradient similarity, thereby resulting in generally poorer substitutes.
-This approach—explaining changes in model performance by dissecting and adjusting key factors in the training process—should be regarded as an important direction in the study of machine learning interpretability.
-It is evident, therefore, that although interpretability is a critical issue in ML, the lack of a unified semantic and formal framework for machine learning has precluded the establishment of rigorous definitions and universally applicable criteria for interpretability.
-Approaches to enhancing interpretability remain diverse, and there is still no universally accepted methodology informed by a common theoretical foundation.
+2.1 Conceptualizations of Machine Learning
+Classical statistical learning theory formalizes learning as risk minimization over function classes under unknown data-generating distributions, with foundational tools such as VC theory, growth functions, and distribution-independent bounds [1].
+Parameterized predictors, loss-based risk, and empirical risk minimization provide the standard optimization lens [2].
+Recent perspectives emphasize causal structure beyond observed distributions, seeking representations that expose latent causal factors to support robust transfer and planning across environments [3].
+Neuro-symbolic approaches pursue integration of distributed representations with symbolic abstraction and compositional reasoning, aiming to combine efficient learning with discrete inference [4, 5].
+These lines of work clarify individual components of the learning pipeline, yet they do not supply a unified semantic and causal framework that spans model construction, training, input, inference, and deployment.
+2.2 Interpretability
+Despite extensive interest, interpretability lacks a precise and generally accepted definition and evaluation protocol.
+Prior analyses highlight ambiguity in terminology and goals, and suggest that the practical objective is to extract useful information from models in ways that are audience- and task-sensitive [15].
+In reinforcement learning, interpretability depends on interpretable inputs, models, and decisions, and on notions such as simulatability, decomposability, and algorithmic transparency [16].
+Mechanistic studies relate training mechanisms to properties such as smoothness and gradient similarity to explain transferability and adversarial behavior, illustrating how dissecting key factors in training can yield explanatory leverage [17].
+Overall, current methods are diverse and often model- or context-specific, and they lack a unifying logical semantics that ties inputs, internal states, and outputs to causal explanations across the lifecycle [15, 16, 17, 18, 19].
 2.3 Ethical Safety Assurance
-Concerns regarding the ethical safety of AI have a long history.
-As noted in (Russell et al., 2015), the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies.
-These issues span law, public policy, professional ethics, and philosophical morality, necessitating the expertise of computer scientists, legal scholars, political scientists, and ethicists.
-While mathematical tools such as formal logic, probability theory, and decision theory have provided important foundations for reasoning and decision-making, numerous challenges remain unresolved.
-Building on traditional concepts of assurance, (Burr and Leslie, 2023) offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence.
-Accordingly, the general structure of moral assurance is characterized as an iterative process of reflection, action, and adjustment, repeated throughout the design, development, and deployment lifecycle.
-In the context of AI/ML-supported healthcare, ethical safety assurance faces more concrete challenges.
-(Tilala et al., 2024) provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare.
-The authors advocate for multiple strategies, including the implementation of rigorous data security protocols, prioritization of data minimization, increasing diversity in training datasets, use of data augmentation to address imbalance, application of algorithmic fairness techniques to detect and mitigate model bias, regular auditing of learning algorithms, enhancing the interpretability of algorithmic outputs, regulatory oversight based on standardized performance metrics and stringent testing protocols, and the augmentation—rather than replacement—of clinicians’ capabilities as means to achieve ethical safety.
-(Hawkins et al., 2021) further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety.
-The authors propose that ML safety requirements should be derived from the system-level safety requirements, and that these should be evaluated in light of the allocated safety needs, system and software architecture, and operational environment.
-This framework then serves as the basis for developing safety assurance specifications across all stages of the ML lifecycle, including data preparation, model training, model validation, and model deployment.
-(Pant et al., 2024) conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches.
-The study highlights that, in the absence of a universally adopted set of ethical principles for AI, incorporating morality into AI is fundamentally challenged by the lack of clear moral knowledge and the ambiguity of ethical principles.
-It is thus evident that, despite widespread concerns over the ethical safety of machine learning, there is as yet no mathematical or theoretical consensus regarding its scientific definition, substantive content, forms of manifestation, or criteria for assessment.
-Instead, ethical safety remains an open-ended concept, explored through various technical, engineering, and moral perspectives.
-Most research either focuses on specific ethical principles and consequences for AI, or on methods for enhancing ethical conduct.
-The proliferation of often conflicting AI ethical guidelines has, paradoxically, made it difficult for practitioners to determine which standards to follow.
-This underscores the necessity of a universally accepted definition of ethical safety, as such a definition can bridge the gap between ongoing academic discourse and industry practice, thereby providing a concrete foundation for strengthening ethical safety assurance in AI.
+As AI systems permeate high-stakes domains, ethical safety has drawn sustained attention in law, policy, and engineering.
+Existing work emphasizes assurance processes grounded in structured arguments that justify normative claims with evidence across the system lifecycle [10, 11].
+In healthcare, recommended practices include data minimization, diversity and augmentation, algorithmic fairness, auditing, interpretability, and standards-based oversight, typically framed as a toolbox of technical and procedural controls [12].
+Safety guidance for autonomous and software-intensive systems links ML safety to software and system safety via requirement derivation and lifecycle-wide assurance activities [13].
+Practitioner studies report gaps in clear moral knowledge and the ambiguity of principles, which complicate operationalization and adoption [14].
+What remains missing is a mathematically grounded definition of ethical safety with constructible guarantees and verifiable criteria that integrate with ML pipelines.
 2.4 Model Error Estimation
-Model performance is fundamentally reflected in the errors it may incur on specific tasks.
-(Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions.
-Based on VC dimension, it further establishes distribution-independent risk upper bounds and sufficient conditions for fast convergence under both bounded and unbounded loss in finite-sample settings.
-By controlling the trade-off between VC dimension capacity and empirical risk through nested hypothesis families, it proves the strong consistency and convergence rates of structural risk minimization (SRM).
-These theoretical foundations have led to practical implementations such as support vector machines (SVM), margin maximization, and kernel methods, offering explanations for their generalization mechanisms.
-(Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error.
-It provides finite-sample error upper bounds via the fixed point of sub-root functions, thus achieving optimal or near-optimal convergence rates inversely proportional to the square root of the number of training samples.
-Under controlled variance conditions, it presents a unified framework for error estimation, including both distribution-dependent bounds determined by local Rademacher complexity and data-dependent bounds constructed from empirical local Rademacher averages.
-These results are extended to typical classification and regression tasks, providing computable upper bounds and iterative algorithms for quickly approximating the fixed point.
-From an information-theoretic perspective, (Miller, 2019) models learning algorithms as random channels mapping input datasets $S$ to output hypothesis spaces $W$.
-For sub-Gaussian (including bounded) losses and uncountable hypothesis spaces, the generalization error upper bound is characterized via the mutual information $I(S;W)$ between input and output.
-Based on this, high-probability sample complexity and expected generalization error upper bounds are derived, and the concept of information stability $(\varepsilon,\mu)$-stability is introduced, unifying algorithmic consistency and overfitting control through information leakage measures.
-The paper further analyzes two classes of algorithms—Gibbs learning and noisy ERM—where $I(S;W)$ is explicitly controlled via mutual information regularization, and discusses methods for decomposing and additively controlling mutual information in complex learning algorithms through adaptive combinations.
-(Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence.
-It introduces a task distribution T and task-level MDPs, defines generalization error as the performance gap between training and unseen tasks under a meta-policy, and establishes generalization upper bounds determined by the number of training tasks and task complexity, using empirical process theory and concentration inequalities.
-The analysis further discusses the impact of algorithmic capacity and learning rates on error bounds.
-Within the non-convex SGD framework, sufficient conditions and rates for finite-step convergence of meta-learning loss are presented, along with guidelines for long-term behavior and hyperparameter selection based on linear and sub-linear types.
-The theoretical results are validated and the observability of convergence rates is demonstrated through controllable task generation and baseline meta-algorithms.
-(Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models.
-From both theoretical and numerical perspectives, it investigates the convergence properties of deterministic samplers based on probability flow ordinary differential equations (ODEs).
-Assuming access to score function estimates with $L_{2}$ accuracy, it provides upper bounds on TVD between target and generated distributions in continuous time.
-For the practical implementation using Runge-Kutta integrators, discrete-time error upper bounds are established.
-Numerical experiments on 128-dimensional problems are conducted to validate these theoretical error bounds.
-Collectively, these works encompass both classical statistical learning theory and recent advances in reinforcement learning and generative models, reflecting a broad recognition of the importance of model error upper bounds in machine learning.
-However, these studies do not employ explicit causal semantic modeling, nor do they address the full lifecycle challenges of multi-stage information transmission, online learning, or multi-agent integration in ML.
-Additionally, they rarely consider the geometric scale and entropy of distribution shifts and uncovered domains, making it difficult to offer actionable engineering or data-sampling guidance for test domain coverage gaps and noisy input scenarios.
-3 Preliminary Theory
-Undoubtedly, as a key component of artificial intelligence, modern machine learning is a product of the rapid development and high integration of information technologies such as computers, the Internet, big data, and cloud computing.
-Machine learning systems are, without question, an important branch of information systems.
-Starting from the basic concept of information can provide a unified theoretical foundation for research in machine learning and artificial intelligence.
-3.1 Definition of Information
-(Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states.
-It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Qiu and Xu, 2025), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times.
-Here, we proceed to further clarify the concept of enabling mapping.
+Foundational results establish consistency and finite-sample bounds for empirical and structural risk minimization through VC dimension and related complexity measures, including conditions for one- and two-sided uniform convergence and margin-based generalization [1].
+Local Rademacher complexity sharpens bounds by tailoring complexity to low-error regions and yields optimal or near-optimal rates via fixed-point analyses under variance control [25].
+Information-theoretic approaches bound generalization via mutual information between data and hypotheses, introduce information stability, and analyze Gibbs and noisy ERM with mutual-information regularization [26].
+In meta-reinforcement learning, generalization across tasks and convergence under nonconvex optimization are characterized using empirical process theory and concentration tools, with guidance for capacity and learning-rate choices [27].
+For generative modeling, recent analyses provide TVD bounds for deterministic samplers via ODE perspectives and establish discrete-time error under Runge–Kutta schemes [28].
+These advances quantify error in specific settings but do not explicitly model causal semantics across multi-stage information flows, nor do they integrate noise structure, coverage geometry, and uncovered-region strategies into a single computable bound that can guide data auditing and prediction strategy design.
+2.5 Toward a Unified Information-Theoretic and Logical Perspective
+Information science offers a unifying lens, yet its subfields have developed largely in silos [20, 21].
+OIT axiomatizes information as an enabling mapping from ontological to carrier states with dual mathematical and physical–causal attributes, and provides general metrics that have shown value for data evaluation in ML [22, 23, 24].
+By formalizing states as interpretations of WFFs in a higher-order logical system and introducing notions of noisy information and information chains, OIT supports analysis of multi-stage pipelines and causal dependence structures [22, 23, 33, 34].
+The present work builds on this foundation to offer a single formal semantics for lifecycle reasoning about interpretability, ethical safety, and computable TVD-based error control.
+3 Theoretical Basis
+Modern ML systems are information systems and, in the mathematical sense, finite automata.
+We ground our meta-framework in a formal account of information and state that supports causal analysis across stages.
+3.1 Information as Enabling Mapping
+[23] based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notion of “enabling mapping”.
+Here, we proceed to further clarify the concept.
 Definition 1 (Enabling Mapping) Let $X$ and $Y$ be two sets, and let $f$ be a mapping.
@@
 $f:X\rightrightarrows Y$ (3.1)
-or
-$f(X)\trianglerighteq Y$ (3.2)
 […]
-Corollary 1 If $f$ is an enabled mapping from $X$ to $Y$, then $|X|\geq|Y|$, where $|X|$ and $|Y|$ denote the cardinalities of the sets $X$ and $Y$, respectively.
-If $f(X)\trianglerighteq Y$, we also denote this as $f^{-1}(y)\trianglelefteq x$ or $x\trianglerighteq f^{-1}(y)$, to maintain consistency in the directionality of enabling.
-Fig.1(a) intuitively illustrates the dual nature of an enabling mapping, encompassing both its mathematical mapping and physical enabling attributes.
-Figure 1: Schematic Diagram of Enabling Mapping, Noisy Information, and Information Transmission Chain.
-(a) The scene of penguin activity $X$ is mapped by an enabling mapping $f$ to an image $Y$ on a laptop; between $X$ and $Y$ there exists both a mathematical mapping relationship and a causal relationship.
-Moreover, mapping $X$ to $Y$ requires corresponding physical conditions and the expenditure of a certain amount of energy.
-(b) The left panel depicts the ontological state of noiseless information, while the right panel shows a noisy information state: the set $S_{o}^{-}$ is removed from the noiseless ontological state, and the previously disjoint set $S_{o}^{+}$ is superimposed, resulting in a noisy ontological state.
-(c) In the penguin activity information transmission chain, the carrier state of any subsequent piece of information can be regarded as the result of an enabling mapping from any prior ontological information state, and can also be reasonably interpreted by means of this ontological state and the corresponding enabling mapping.
-Based on this, a more rigorous theoretical foundation for the definition of information can be established:
-Definition 2 (Mathematical Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain.
+Figure 1: Schematic Diagram of Enabling Mapping, Noisy Information, and Information Chain.
+Based on Appendix A’s higher-order formal system $\mathcal{L}$ and the result that virtually any state set can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ [33], we obtain a more robust theoretical foundation for defining information:
+Definition 2 (Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain.
 The ontology $o\mathbb{\subseteq O}\cup\mathbb{S}$, time of occurrence $T_{h}\mathbb{\subseteq T}$, ontological state set $S_{o}\left(o,T_{h}\right)$, carrier $c\mathbb{\subseteq O}$, reflection time $T_{m}\mathbb{\subseteq T}$, and carrier reflection set $S_{c}\left(c,T_{m}\right)$ are all non-empty sets.
-Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I=\left\langle o,T_{h},S_{c},c,T_{m},S_{o}\right\rangle$.
-The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Wiener, 2019).
-It should be noted that both $S_{o}\left(o,T_{h}\right)$ and $S_{c}\left(c,T_{m}\right)$ can each be represented as the interpretation of a set of WFFs within a formal system.
-However, the respective formal systems need not be identical, which constitutes one of the essential characteristics of information—its capacity to establish connections among all entities in the world, regardless of differences in their underlying formal systems.
-(Shannon, 1948) is regarded as a foundational work in information theory.
-The paper proposed that ” $H=-\sum{p_{i}\log p_{i}}$ plays a central role as the measure of information in information theory.” The discussion in this paper includes both the source and channel of information.
-The source $Is$ is the origin of the information, and it generates events $e_{1}$, $e_{2},$…, $e_{n}$ with probabilities $p_{1}$, $p_{2},$…, $p_{n}$, respectively.
-The channel $Ic$ is the medium through which information is transmitted.
-After an event $e_{i}$ occurs at time $T_{h}$ in the source $Is$, the channel $Ic$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$ (where $1\leq i\leq n,\ \tau>0$).
-We define the ternary predicate ${IsEvent}^{3}(Is,T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $Is$ at time $T_{h}$, and the ternary predicate ${IcCode}^{3}(Ic,T_{m},c_{i})$ to represent that the channel $Ic$ received and transmitted code $c_{i}$ in the time interval $T_{m}$.
-Then,
-$S_{Is}(Is,T_{h})={IsEvent}^{3}(Is,T_{h},e_{i})$ (3.3)
-and
-$S_{Ic}(Ic,(T_{h},T_{h}+\tau))={IcCode}^{3}(Ic,(T_{h},T_{h}+\tau),c_{i})$ (3.4)
-are the states of the source $Is$ at time $T_{h}$ and the state of the channel $Ic$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $(1\leq i\leq n)$). Furthermore,
-$I:S_{Is}(Is,T_{h})\rightrightarrows S_{Ic}(Ic,(T_{h},T_{h}+\tau))$ (3.5)
-represents the enabling mapping from the source state $S_{Is}(Is,T_{h})$ to the channel state $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$.
-Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{Is}(Is,T_{h})$ and the complex implementation mechanisms of $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system.
-The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
+Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ or $I=\left\langle o,T_{h},S_{o},c,T_{m},S_{c}\right\rangle$.
+The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world [34].
+Definition 2 subsumes Shannon’s model: source events and channel codes are states, and source-to-channel is an enabling mapping; OIT thus generalizes classical information theory [35].
 3.2 Noisy Information
-The realization of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
-Definition 3 (Noisy Information) For information $I=\left\langle o,T_{h},S_{o},c,T_{m},S_{c}\right\rangle$, let $\widetilde{S_{o}}\left(o,T_{h}\right)$ be another state of the ontology $o$ at time $T_{h}$, and $\widetilde{S_{c}}\left(c,T_{m}\right)$ be another state of the carrier $c$ at time $T_{m}$,
+[22] states that an information instance $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ admits an inverse mapping $I^{-1}(S_{c}(c,T_{m}))$.
+In practice, however, information realization is rarely immune to noise or other external factors.
+Consequently, applying $I^{-1}$ may fail to fully recover $S_{o}(o,T_{h})$; instead, it yields a state corrupted by noise.
+This motivates the following notion.
+Definition 3 (Noisy Information) For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, the inverse $I^{-1}(S_{c}(c,T_{m}))$ is another state of the ontology $o$ at time $T_{h}$, denoted by $\widetilde{S_{o}}\left(o,T_{h}\right)$, and given by
 […]
-where the state $S_{o}^{-}\left(o,T_{h}\right)$ is the loss state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}^{-}\left(o,T_{h}\right)\subseteq S_{o}\left(o,T_{h}\right)$ and not being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$, and the state $S_{o}^{+}\left(o,T_{h}\right)$ is the superposed state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}\left(o,T_{h}\right){\cap S}_{o}^{+}\left(o,T_{h}\right)=\varnothing$ and being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$.
-The enabling mapping is then defined as:
-$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows\widetilde{S_{c}}\left(c,T_{m}\right)$ (3.7)
-which is the noisy information of $I$, while $I$ is its own noiseless information.
-Thus, the noisy information $\widetilde{I}$ is also information.
-It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same.
-Fig.1(b) illustrates the structural characteristics of noisy information.
-In fact, Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario (Shannon, 1948).
-In this case, he assumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
-$E=f(S,N)$ (3.8)
-where $f$ is the distortion function of the channel to the transmitted signal.
-According to OIT's view, the signal $S$ here is the ontological state of information $I$, $S_{Is}\left(Is,T_{h}\right)$.
-The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{Is}^{-}\left(Is,T_{h}\right)$ from the source state $S_{Is}\left(Is,T_{h}\right)$ and adding another state $S_{Is}^{+}\left(Is,T_{h}\right)$.
-The output $E$ is then the channel's output state $\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output.
-Thus,
-$\widetilde{I}:\widetilde{S_{Is}}\left(Is,T_{h}\right)\rightrightarrows\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.9)
-is the noisy information of
-$I:S_{Is}\left(Is,T_{h}\right)\rightrightarrows S_{Ic}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.10)
-[…]
-In this way, we use the definition of noisy information to explain Shannon's model of noisy channels in (Shannon, 1948), demonstrating that OIT is applicable to the complex scenarios of information research under noisy conditions.
+where $S_{o}^{-}$ is the noise-loss state with $S_{o}^{-}\subseteq S_{o}$, and $S_{o}^{+}$ is the noise-addition state with $S_{o}{\cap S}_{o}^{+}=\varnothing$.
+The carrier state $S_{c}(c,T_{m})$ can thus be viewed as the image of $\widetilde{S_{o}}\left(o,T_{h}\right)$ under an enabling mapping:
+$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows S_{c}\left(c,T_{m}\right)$ (3.3)
+which is the noisy information corresponding to $I$, while $I$ itself is the noise-free information.
+Fig.
+1(b) depicts the structural characteristics of noisy information.
+In Shannon’s terms, noise transforms the source state via deletion and addition before transmission; the resulting channel output realizes the noisy carrier state [35].
+This structural decomposition will support our TVD bounds by isolating loss and addition masses.
 3.3 Information Chain
-(Xu et al., 2014) discusses the transmissibility of information through composite mappings.
-Based on Definitions 1 and 2, the fundamental laws governing information transmission can be characterized more rigorously as follows:
-Lemma 1 (Transmissibility of Information): Let $I_{0}=\left\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\right\rangle$ and $I_{1}=\left\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ be two pieces of information.
-Then, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is also information.
-We refer to this as the information $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$ via $c_{1}$.
-Proof: According to Definitions 1 and 2, as well as the given conditions, for any $S_{o}(x,t)\in S_{o}\left(o,T_{h}\right)$, there exists $S_{c_{1}}\left(x_{1},t_{1}\right)\in S_{c_{1}}\left(c_{1},T_{m_{1}}\right)$such that $I_{0}\left(S_{o}(x,t)\right)\trianglerighteq S_{c_{1}}\left(x_{1},t_{1}\right)$.
-Furthermore, there exists $S_{c_{2}}\left(x_{2},t_{2}\right)\in S_{c_{2}}\left(c_{2},T_{m_{2}}\right)$ such that $I_{1}\left(S_{c_{1}}\left(x_{1},t_{1}\right)\right)\trianglerighteq S_{c_{2}}\left(x_{2},t_{2}\right)$.
-Define $I_{2}\left(S_{o}(x,t)\right)=S_{c_{2}}\left(x_{2},t_{2}\right)$.
-Since the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ is contingent upon the existence of $S_{c_{1}}\left(x_{1},t_{1}\right)$, and the physical realization of $S_{c_{1}}\left(x_{1},t_{1}\right)$ in turn depends on $S_{o}(x,t)$, by causal induction, the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ ultimately depends on $S_{o}(x,t)$.
-Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{c_{1}}(x_{1},t_{1}))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{0}(S_{o}(x,t))\trianglerighteq S_{c_{1}}(x_{1},t_{1})$.
-Thus, $I_{2}(S_{o}(x,t))={I_{1}(I}_{0}(S_{o}(x,t)))\trianglerighteq I_{1}(S_{c_{1}}\left(x_{1},t_{1}\right))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$.
-Therefore, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 2.
-In essence, this is the information $I_{0}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$.
-The lemma is thus proved.
-Corollary 2 (Information Chain) Let $\left\{I_{i}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{i},T_{m_{i}},S_{c_{i}}\right\rangle\middle|i=1,\cdots,n\right\}$ be a sequence of information mappings, where for any $i>1$, it holds that $o_{i}=c_{i-1}$, $T_{h_{i}}=T_{m_{i-1}}$, and $S_{o_{i}}\left(o_{i},T_{h_{i}}\right)=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$.
+[22] uses composition to discuss the transitivity of information.
+Specifically, if $I_{o}:S_{o}(o,T_{h})\rightrightarrows S_{c_{1}}(c_{1},T_{m_{1}})$ and $I_{1}:S_{c_{1}}(c_{1},T_{m_{1}})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ are two information instances, then $I_{2}:S_{o}(o,T_{h})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ is also an information instance, interpreted as $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$.
+This yields the notion of an information chain.
+Let $\left\{I_{i}:S_{o_{i}}(o_{i},T_{h_{i}})\rightrightarrows S_{c_{i}}(c_{i},T_{m_{i}})\middle|i=1,\cdots,n\right\}$ be a sequence of information mappings, where for any $i>1$, it holds that $o_{i}=c_{i-1}$, $T_{h_{i}}=T_{m_{i-1}}$, and $S_{o_{i}}\left(o_{i},T_{h_{i}}\right)=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$.
 In this case, we say that these mappings constitute an information chain.
-Moreover, for any 1 $\leq i\leq j\leq n$, $I_{ij}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{j},T_{m_{j}},S_{c_{j}}\right\rangle$ also qualifies as information according to Definition 2.
+Moreover, for any 1 $\leq i\leq j\leq n$, $I_{ij}:S_{o_{i}}(o_{i},T_{h_{i}})\rightrightarrows S_{c_{j}}(c_{j},T_{m_{j}})$ also qualifies as information according to Definition 2.
 Transmission constitutes a fundamental mode of information flow and application.
-Communication, processing, input, and output are all concrete realizations of information transmission.
 In many cases, information undergoes more complex flows through chain-like transmission processes.
-Fig.1(c) intuitively illustrates such a process.
+Fig.
+1(c) intuitively illustrates such a process.
 For any pair of states within this process, the latter can be regarded as the result of the enabling mapping from the former, and this enabling mapping can be used to reasonably explain the causation of the latter state.
 This provides a theoretical foundation for the interpretability of the model.
+3.4 Finite Automata as Formal State Systems
+Any practical computing or ML system can be formalized as a finite automaton with state, input, output, and transition functions over discrete times (Appendix B).
+We represent its behavior by a state set of WFFs that captures inputs, outputs, and transitions, together with predicates and functions for assignability, learnability, and processability.
+This provides the operational substrate for our lifecycle formalization in Section 4 and for the logical definitions and theorems in Section 5.
 4 Meta-Framework for Machine Learning Theory (MLT-MF)
-The core of machine learning is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system.
+The core of ML is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system.
 The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system.
-Therefore, by employing the concepts of information model and finite automaton state (Appendix B), the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
+Therefore, by employing the concepts of information model and finite automaton state, the process of ML and its applications can be unified into five primary steps of information realization and functioning.
+Fig.
+2 presents the chain-structured architecture of causal information progression across the five stages.
+The subsequent subsections provide formal expressions of the information states for each stage.
+Undoubtedly, these pieces of information are the fundamental basis for explaining model behavior and outcomes.
+They transcend specific algorithms or models and provide a new theoretical framework—MLT-MF—for addressing a range of critical issues such as ML interpretability.
+Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning.
 4.1 Construction of Initial Information
-Suppose the model under study is denoted as $M$, and let $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ be the set of time points at which $M$, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order.
-According to the perspective of OIT, the creation of $M$ is essentially the realization of information
-$Io=\left\langle M_{D},{To}_{h},S_{M_{D}},M,t_{1},S_{M}\right\rangle$ (4.1)
+For the first stage in Fig.
+2, suppose the model under study is denoted as $M$, and let $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ be the set of time points at which $M$, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order.
+According to the perspective of OIT, the creation of $M$ is essentially the realization of information:
+$Io:S_{M_{D}}(M_{D},To_{h})\rightrightarrows S_{M}(M,t_{1})$ (4.1)
 where the ontology $M_{D}$ of $Io$ is the designer’s conceptualization of $M$, ${To}_{h}$ is the time at which the designer conceives $M$, and $S_{M_{D}}\left(M_{D},{To}_{h}\right)$ is the initial finite automaton state that must be realized at this conceptual stage.
 The carrier $M$ of $Io$ is the model itself; the reflection time $t_{1}$ is when $M$ is constructed according to $S_{M_{D}}\left(M_{D},{To}_{h}\right)$; and $S_{M}\left(M,t_{1}\right)$ is the formalization of the actual physical state of $M$ upon its completion.
-Thus, $\sup\left({To}_{h}\right)<sup(t_{1})$.
+Thus, $max({To}_{h})<max(t_{1})$.
 According to Lemma B.2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
@@
 denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},{To}_{h})$ as envisioned by the designer are assigned to $M_{D}$.
-$\begin{split}{\varphi o}_{2}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Learnable}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({Eq}^{2}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land{State}^{3}(M_{D},t_{2},\Phi^{2}\left(M_{D},t_{2}\right))\end{split}$ (4.3)
+$\varphi o_{2}=$ $\mathit{State}^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\land$ (4.3)
+$\mathit{Input}^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
+$\mathit{Learnable}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{\mathit{learn}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
+$\left(\mathit{Eq}^{2}(\Phi^{2}(M_{D},t_{2}),f_{\mathit{learn}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1})))\right)\land$
+$\mathit{State}^{3}(M_{D},t_{2},\Phi^{2}(M_{D},To_{h}))$
 expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},{To}_{h})$, and the input carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is provided and satisfies the learnability condition determined by the binary predicate ${Learnable}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))$, $M_{D}$ can optimize its internal structures, parameters, and algorithms to $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))$, assigning this to $M_{D}$.
 This demonstrates the model's capacity for learning and optimization, where $M_{I}$ denotes the input device of $M$.
-$\varphi_{O_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\wedge Input^{3}(M_{D},t_{1},S_{M},(M_{I},t_{1}))$ (4.4)
-$\wedge Processable^{2}(M,S_{M},(M_{I},t_{1}))\rightarrow\exists f^{2}_{process}(M,S_{M},(M_{I},t_{1}))$
-$(Eq^{2}(S_{M_{0}}(M_{0},t_{2}),f^{2}_{process}(M,S_{M},(M_{I},t_{1}))))\wedge Output^{3}(M,t_{2},S_{M_{0}}(M_{0},t_{2}))$
+$\varphi o_{3}=$ $\mathit{State}^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\land$ (4.4)
+$\mathit{Input}^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
+$\mathit{Processable}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{\mathit{process}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
+$\left(\mathit{Eq}^{2}(S_{M_{O}}(M_{O},t_{2}),f_{\mathit{process}}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1})))\right)\land$
+$\mathit{Output}^{3}(M_{D},t_{2},S_{M_{O}}(M_{O},t_{2}))$
 indicates that, at time $t_{1}$, if $S_{M_{I}}\left(M_{I},t_{1}\right)$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate ${Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$, then $M_{D}$ can process $S_{M_{I}}\left(M_{I},t_{1}\right)$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$ to yield the output carrier state $S_{M_{O}}\left(M_{O},t_{2}\right)$, which is then output at time $t_{2}$.
 This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of $M$.
+When output information must not violate relevant ethical guidelines, $f_{process}^{2}$ must implement the corresponding safety assurance functions.
 Thus,
@@
 constitutes the initial state of $M$ as conceived by the designer.
-It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas.
+It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three WFFs, or—when decomposed with respect to the elements of each set—as three sets of more granular WFFs.
 Regardless of the situation, both cases satisfy Definition B.1 concerning finite automaton states.
@@
 $S_{M}\left(M,t_{1}\right)=\left\{{\psi o}_{1},{\psi o}_{2},{\psi o}_{3}\right\}$ (4.6)
-where ${\psi o}_{1}\trianglelefteq Io({\varphi o}_{1})$, ${\psi o}_{2}\trianglelefteq Io({\varphi o}_{2})$, and ${\psi o}_{3}\trianglelefteq Io({\varphi o}_{3})$ are the enabling mappings of ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$, respectively.
+where ${\psi o}_{1}$, ${\psi o}_{2}$ and ${\psi o}_{3}$ are the enabling mappings of ${\varphi o}_{1}$, ${\varphi o}_{2}$ and ${\varphi o}_{3}$, respectively.
 This process establishes $M$ as a typical finite automaton and an operational information system.
 4.2 Formal Modeling and Realization of Training Information
-According to the general formal definition of information, the formal expression for the realization of training information is straightforward:
-$It=\left\langle ot,{Tt}_{h},S_{ot},M_{I},t_{1},S_{M_{I}}\right\rangle$ (4.7)
+The second stage in Fig. 2 is the realization of training information. According to the general formalization of information, its formal expression is straightforward:
+$It:S_{ot}(ot,Tt_{h})\rightrightarrows S_{M_{I}}(M_{I},t_{1})$ (4.7)
 $S_{ot}(ot,{Tt}_{h})=\left\{\varphi t\middle|\varphi t\ is\ a\ WFF\ about\ ot\ and\ {Tt}_{h}\right\}$ (4.8)
@@
 The carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is the physical realization of $S_{ot}(ot,{Tt}_{h})$ via an enabling mapping, ensuring that the information can be machine-read and entered into $M$.
+In practice, much training information falls under the noisy information defined in Definition 3.
+The structure of noisy information will affect the effectiveness of model training.
 4.3 Formal Modeling and Implementation of Model Learning
-Model learning for $M$ involves fusing its original information $Io$ with the training information $It$, resulting in updated information:
-$Iu=\left\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\right\rangle$ (4.10)
+In the third stage of Fig. 2, $M$’s learning is precisely the fusion of the original information $Io$ with the training information $It$, optimizing it into updated information:
+$Iu:S_{M\cup M_{I}}(M\cup M_{I},t_{1})\rightrightarrows S_{M}(M,t_{2})$ (4.10)
 Here, the ontology is the hybrid $M\cup M_{I}$, encompassing both the model and its interface; $t_{1}\cup t_{2}$ marks the time of the first learning event; and $S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})$ formally represents the hybrid state during learning:
@@
 indicates that the input state $S_{M_{I}}\left(M_{I},t_{1}\right)$, which was learned by $M$ at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of $M$ at $t_{2}$.
-$\begin{split}{\varphi u}_{2}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))))\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.13)
-shows that after the first learning step, $M$ remains capable of further learning and optimization.
-${\varphi u}_{3}=$ $State^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\wedge Input^{3}\left(M,t_{2},S_{M_{I}}(M_{I},t_{2})\right)$ (4.14)
-$\wedge Processable^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)\rightarrow\exists f_{process}^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)$
-$\left(Eq^{2}\left(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))\right)\wedge Output^{3}\left(M,t_{3},S_{M_{O}}(M_{O},t_{3})\right)\right)$
-indicates that when the training information $It$ is processable by $M$, it can also produce output information.
-This state is then enabled and mapped to the physical state of the carrier $M$ at time $t_{2}$:
-$S_{M}\left(M,t_{2}\right)=\left\{{\psi u}_{1},{\psi u}_{2},{\psi u}_{3}\right\}$ (4.15)
+As shown in Equations (4.3) and (4.4), $\varphi_{u_{2}}$ and $\varphi_{u_{3}}$ are logically identical to $\varphi_{o_{2}}$ and $\varphi_{o_{3}}$, respectively, differing only in the underlying ontological object and time.
+This indicates that the state retains the capacity for continued learning and processing even after the initial learning and optimization phases.
 4.4 User Input Information for the Model
-User input information is formally given by:
-$Iq=\left\langle oq,{Tq}_{h},S_{oq},M_{I},t_{2},S_{M_{I}}\right\rangle$ (4.16)
+User input information in the fourth stage in Fig. 2 is formally given by:
+$Iq:S_{oq}(oq,Tq_{h})\rightrightarrows S_{M_{I}}(M_{I},t_{2})$ (4.13)
 $S_{oq}(oq,{Tq}_{h})=\left\{\varphi q\middle|\varphi q\ is\ a\ WFF\ about\ oq\ and\ {Tq}_{h}\right\}$ (4.14)
@@
 The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,{Tq}_{h})$ onto the model’s interface device $M_{I}$.
+User input information is also typically affected by noise.
+Its noisy information structure will be directly related to the $M$’s output error.
 4.5 Information Processing, Output, and Online Learning of the Model
-When the model $M$ processes input and produces output, the information it carries can be represented as:
-$Iv=\left\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\right\rangle$ (4.19)
+In the fifth stage shown in Fig. 2, when the model $M$ processes input and produces output, the information it carries can be represented as:
+$Iv:S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3})\rightrightarrows S_{M}(M,t_{3})$ (4.16)
 where:
-$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3}\right)=\left\{{\varphi v}_{1},{\varphi v}_{2},{\varphi v}_{3}\right\}$ (4.20)
-${\varphi v}_{1}={State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)$ (4.21)
-indicates that $\Phi^{2}(M,t_{2})$ has been assigned to $M$ at time $t_{2}$.
-$\begin{split}{\varphi v}_{2}=&{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\land{Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\\ &\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))\end{split}$ (4.22)
-indicates that when the input state $S_{M_{I}}\left(M_{I},t_{2}\right)$ is received at $t_{2}$, $M$ produces the output state $S_{M_{O}}\left(M_{O},t_{3}\right)$ at $t_{3}$.
-$\begin{split}{\varphi v}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))))\\ &\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.23)
-expresses that if the input information can be learned by $M$, then $M$ can achieve online learning while processing and outputting information.
-The physical realization of this state through an enabling mapping is:
-$S_{M}(M,t_{3})=\left\{{\psi v}_{1},{\psi v}_{2},{\psi v}_{3}\right\}$ (4.24)
+$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3}\right)=\left\{{\varphi v}_{1},{\varphi v}_{2}\right\}$ (4.17)
+$M_{O}$ is $M$’s output device.
+The formulas $\varphi_{v_{1}}$ and $\varphi_{v_{2}}$ are logically identical to $\varphi_{u_{2}}$ and $\varphi_{u_{3}}$, respectively, differing only in their temporal references.
+This demonstrates that if the input information is processable and learnable by $M$, then $M$ can perform online learning concurrently with information processing and output operations.
 The output information is given by:
-$Ir=\left\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\right\rangle$ (4.25)
+$Ir:S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2})\rightrightarrows S_{M_{O}}(M_{O},t_{3})$ (4.18)
 where the ontological state is:
-$S_{M\cup M_{I}}\left(M\cup M_{I},t_{2}\right)=\left\{{\varphi r}_{1}\right\}$ (4.26)
+$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\right)=\left\{{\varphi r}_{1}\right\}$ (4.19)
 with
 ${\varphi r}_{1}={Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (4.20)
-The formal expression of the physical state of the model’s output device $M_{O}$ at time $t_{3}$ is:
-$S_{M_{O}}\left(M_{O},t_{3}\right)=\left\{{\psi r}_{1}\right\}$ (4.28)
-The model can repeat the processes described in Sections 3.2 to 3.5, iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
-4.6 Multi-Stage Information Causal Chains
-From systems engineering perspective, the realization of information in machine learning involves more than just five information mappings.
-In practice, stages such as requirements analysis, design and development, and system integration are also necessary.
-Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 3.3, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation.
-In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Fig.2).
-Fig.2 presents a comprehensive framework of formal expressions and causal information progression for each stage of the machine learning process, from the perspectives of OIT and finite automaton state implementation.
-This framework transcends specific algorithms or models, offering a novel, unified, and general framework—MLT-MF—for addressing several key issues in the field of machine learning.
-In Appendix C, we further provide an example of its application to feedforward neural networks, thereby testing the correctness and applicability of the framework and supporting the verification of related theorems.
-Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning.
-In machine learning, the four distinct stages—model construction, learning and training, problem input, and answer output—each encompass four levels: conceptual ideas, mind maps, program code, and physical models.
-Generally, the content within the same level shares a common semantic logic and can thus be represented using a unified formal system.
-Furthermore, owing to the transmissibility of information, the innermost physical model can be regarded as the result of an enabling mapping from the outer conceptual ideas or mind maps.
-The output stage can, in turn, be viewed as a reverse causal transmission from the physical model back to conceptual ideas.
-This demonstrates that the system comprises multiple stages, with each stage containing various combinatorial information causal chains.
+Finally, it should be emphasized that $M$ can repeat the processes described in Sections 4.2 to 4.5, iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
 5 Several Important Definitions and Theorems
@@
 5.1 Theorem of Model Interpretability
-According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model (Glanois et al., 2024), the following concepts are established based on the MLT-MF framework:
+According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model [16], the following concepts are established based on the MLT-MF framework:
 Definition 4 (Model Interpretability) Let the model $M$, as a finite automaton, operate over an ordered set of times $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$, at which it undergoes assignment, input, output, and state transitions.
@@
 Theorem 1 (Model Interpretability and Information Existence) Let the model $M$ be viewed as a finite automaton operating over an ordered set of times $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$, during which it undergoes assignment, input, output, and state transitions.
-(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is that there exist $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$ and $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$ representing the information carried by $M$ and its input device $M_{I}$ at time $t_{i}$, respectively, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$.
+(1) For any $i=1,\dots,n-1$, a necessary and sufficient condition for the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is that there exist $Iu(t_{i}):S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\rightrightarrows S_{M}(M,t_{i})$ and $Iq(t_{i}):S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))\rightrightarrows S_{M_{I}}(M_{I},t_{i})$ representing the information carried by $M$ and its input device $M_{I}$ at time $t_{i}$, respectively, and $Ir(t_{i+1}):S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\rightrightarrows S_{M_{O}}(M_{O},t_{i+1})$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$.
 The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$ must hold.
 (2) A necessary and sufficient condition for $M$ to be interpretable in $\mathcal{L}$ is that, for all $i=1,\cdots,n-1$, there exist information objects $Iu(t_{i})$, $Iq(t_{i})$, and $Ir(t_{i+1})$ whose ontological states are all interpretations in $\mathcal{L}$, and such that the ontological state of $Ir(t_{i+1})$ is contained in the logical closure of the union of the ontological states of $Iu(t_{i})$ and $Iq(t_{i})$.
-Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$.
+Proof: For any $i=1,\dots,n-1$, suppose that the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$.
 According to Definition 4, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$, and $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
-Let $S_{on}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{op}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$ and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$.
-Then, according to Definition 4, there exist information objects $Ig(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{op},M_{I},t_{i},S_{M_{I}}\rangle$ and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq Cn(S_{on}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$.
+Let $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$ and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$.
+Then, according to Definition 4, there exist information objects $Iu(t_{i}):S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\rightrightarrows S_{M}(M,t_{i})$, $Iq(t_{i}):S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))\rightrightarrows S_{M_{I}}(M_{I},t_{i})$ and $Ir(t_{i+1}):S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\rightrightarrows S_{M_{O}}(M_{O},t_{i+1})$ such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$.
 Thus, the necessity of the first part of the theorem is established.
-On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$, $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$.
+On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i}):S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\rightrightarrows S_{M}(M,t_{i})$, $Iq(t_{i}):S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))\rightrightarrows S_{M_{I}}(M_{I},t_{i})$ and $Ir(t_{i+1}):S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\rightrightarrows S_{M_{O}}(M_{O},t_{i+1})$, such that in the formal system $\mathcal{L}$, $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$.
 Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$.
@@
 To apply the MLT-MF framework, it is first necessary to identify the corresponding well-formed ethical constraint formulas $Ec$ in accordance with fundamental ethical standards such as fairness and privacy protection, thereby establishing a formalized expression of ethical security.
-Definition 5 (Ethical Safety) Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$.
-Let $Ec(\overrightarrow{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\overrightarrow{v}=(v_{1},\ldots,v_{n})$ are its free variables.
+Definition 5 (Ethical Safety) Let $S(X,T)$ be the interpretation of a set of WFFs in a formal system $\mathcal{L}$ over the domain $X\times T$.
+Let $Ec(\overrightarrow{v})$ denote an ethical constraint expressed as a WFF in $\mathcal{L}$, where $\overrightarrow{v}=(v_{1},\ldots,v_{n})$ are its free variables.
 For each substitutable $(x,t)\in X\times T$, define
@@
 Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}\left(ou,{Tu}_{h}\right)$, $S_{oq}\left(oq,{Tq}_{h}\right)$ and $S_{or}\left(or,{Tr}_{h}\right)$ of $Iu$, Iq\ Iq and $Ir$ respectively, the MLT-MF framework provides:
-$\varphi_{u}\coloneqq$ $\bigl((\forall\,\Phi_{u})(\forall\,\Phi_{q})\bigl({IsSubsetOf}^{2}(\Phi_{u},S_{ou}(ou,{T}_{u,h}))$ (5.4)
-$\quad\land{IsSubsetOf}^{2}(\Phi_{q},S_{oq}(oq,{T}_{q,h}))$
-$\quad\land{Processable}^{2}(\Phi_{u},\Phi_{q})\bigr)\bigr)$
-$\rightarrow\bigl(\exists\,\varphi_{r}\bigl({IsFormulaOf}^{2}(\varphi_{r},S_{or}(or,{T}_{r,h}))$
-$\quad\land{Eq}^{2}(\varphi_{r},f_{\mathrm{process}}^{2}(\Phi_{u},\Phi_{q}))\bigr)\bigr).$
+$\varphi u=$ $\bigl((\forall\,\Phi u)(\forall\,\Phi q)\bigl({IsSubsetOf}^{2}(\Phi u,S_{ou}(ou,{Tu}_{h}))$ (5.4)
+$\quad\land{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,{Tq}_{h}))$
+$\quad\land{Processable}^{2}(\Phi u,\Phi q)\bigr)\bigr)$
+$\rightarrow\bigl(\exists\,\varphi_{r}\bigl({IsFormulaOf}^{2}(\varphi r,S_{or}(or,{Tr}_{h}))$
+$\quad\land{Eq}^{2}(\varphi r,f_{\mathrm{process}}^{2}(\Phi u,\Phi q))\bigr)\bigr).$
 Thus, the set
-$\begin{split}S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r\middle|\right.&\left.\varphi r={Processable}^{2}(\Phi u,\Phi q),\right.\\ &\left.\Phi u\in Sou\left(ou,{Tu}_{h}\right),\right.\\ &\left.\Phi q\in S_{oq}\left(oq,{Tq}_{h}\right)\right\}\end{split}$ (5.5)
+$\begin{split}S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r\middle|\right.&\left.\varphi r=f_{\mathrm{process}}^{2}(\Phi u,\Phi q),\right.\\ &\left.\Phi u\in Sou\left(ou,{Tu}_{h}\right),\right.\\ &\left.\Phi q\in S_{oq}\left(oq,{Tq}_{h}\right)\right\}\end{split}$ (5.5)
 represents the theoretical output states of model $M$.
@@
 $H_{or}=\left(S_{or}\left(or,{Tr}_{h}\right),\mathcal{E}\right)$ (5.6)
-where the vertex set is precisely $S_{or}\left(or,{Tr}_{h}\right)$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
+where the vertex set is precisely $S_{or}\left(or,{Tr}_{h}\right)$, and the hyperedge set $\mathcal{E}$ covers all WFFs that directly or implicitly violate the ethical constraint $Ec$:
 (1) For any $\varphi r\in S_{or}\left(or,{Tr}_{h}\right)$ and $(x,t)\in or\times{Tr}_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\left\{\varphi r\right\}$ to $\mathcal{E}$;
@@
 Thus, (5.7) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
-By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality.
+By the maximum independent set theorem for finite hypergraphs[36], $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality.
 Suppose there exists another ethically safe subset $K$ with
@@
 […]
-In this way, add to $Sou\left(ou,{Tu}_{h}\right)$ an ethical safeguard formula:
-$\begin{split}\varphi u=&(\forall\varphi r)({IsFormulaOf}^{2}\left(\varphi r,S_{or}\left(or,{Tr}_{h}\right)\right))\rightarrow\exists\varphi s\left({IsFormulaOf}^{2}\left(\varphi s,S_{os}\left(os,{Ts}_{h}\right)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}(\varphi s,\varphi r)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{\sim Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}\left(\varphi s,{IsSafeback}^{1}(NEc)\right)\right)\right)\end{split}$ (5.9)
-Since the newly added $\varphi u$ pertains to $S_{os}\left(os,{Ts}_{h}\right)$ and is entirely independent of the existing formulas in $Sou\left(ou,{Tu}_{h}\right)$, it does not affect the satisfaction of Postulate 5 by $Sou\left(ou,{Tu}_{h}\right)$.
+In this way, add to $S_{ou}\left(ou,{Tu}_{h}\right)$ an ethical safeguard formula:
+$\begin{split}\varphi u=&(\forall\varphi r)\left(\mathit{IsFormulaOf}^{2}\left(\varphi r,S_{or}\left(or,Tr_{h}\right)\right)\right)\rightarrow\\ &\exists\varphi s\left(\mathit{IsFormulaOf}^{2}\left(\varphi s,S_{os}\left(os,Ts_{h}\right)\right)\land\right.\\ &\left(\mathit{IsFormulaOf}^{2}\left(\varphi r,S_{or}^{Ec}\left(or,Tr_{h}\right)\right)\rightarrow\mathit{Eq}^{2}(\varphi s,\varphi r)\right)\land\\ &\left.\left(\mathit{IsFormulaOf}^{2}\left(\varphi r,S_{or}^{\sim Ec}\left(or,Tr_{h}\right)\right)\rightarrow\mathit{Eq}^{2}\left(\varphi s,\mathit{IsSafeback}^{1}(NEc)\right)\right)\right)\end{split}$ (5.9)
+Since the newly added $\varphi u$ pertains to $S_{os}\left(os,{Ts}_{h}\right)$ and is entirely independent of the existing formulas in $S_{ou}\left(ou,{Tu}_{h}\right)$, it does not affect the satisfaction of Postulate A.1 by $S_{ou}\left(ou,{Tu}_{h}\right)$.
 Moreover,
@@
 Therefore, it is the maximal output state that is ethically safe with respect to $Ec$.
-By taking $S_{os}\left(os,{Ts}_{h}\right)$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,{Ts}_{m})$, $S_{M}(M,{Ts}_{m})$ is necessarily also ethically safe with respect to $Ec$.
+By taking $S_{os}\left(os,{Ts}_{h}\right)$ as the ontological state and mapping it via noise-free information $Is$ to the carrier state $S_{M}(M,{Ts}_{m})$, $S_{M}(M,{Ts}_{m})$ is necessarily also ethically safe with respect to $Ec$.
 This satisfies the condition in Definition 5 for model $M$ to be ethically safe with respect to $Ec$.
 The theorem is thus proved.
-Theorem 2 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs—for obtaining the maximum ethically safe output.
+Theorem 2 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs—for obtaining the SafeMax.
 However, computing the maximum independent set in a hypergraph is generally an NP-hard problem.
-In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}\left(or,{Tr}_{h}\right)$, and another for ethical safety checking, thus yielding an approximately maximal safe output.
+In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}\left(or,{Tr}_{h}\right)$, and another for ethical safety checking, thus yielding an approximately SafeMax.
 This is a practical issue that warrants further investigation and will not be elaborated upon here.
@@
 This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
-In summary, based on the MLT-MF framework, Definition 10 and Theorem 5, for the first time, construct a formal theoretical framework for ethical safety.
-This approach, unlike those based on heuristic rules or statistical indicators, fills the gap in mathematical foundations and verifiability in traditional research, and provides concrete methodological guidance for the automated embedding and verification of ethical constraints.
 5.3 Theorem on TVD Upper Bounds
-Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
-Theorem 3 (TVD Upper Bound for Conservative Prediction Strategies in Noise-Free Inputs and Uncovered Training Regions) Let the information describing the model $M$ after learning and optimization be $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$, and the input information be $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$.
+Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by TVD between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
+Theorem 3 (TVD Upper Bound for Conservative Prediction Strategies in Noise-Free Inputs and Uncovered Training Regions) Let the information describing the model $M$ after learning and optimization be $Iu:S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,Tu_{m})$, and the input information be $Iq:S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$.
 Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space.
-Let $\left|S_{oq}\right|$ and $\left|S_{ou}\cap S_{oq}\right|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively.
+Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively.
 The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$; the distribution over $S_{oq}$ is $p_{q}(x)$, with $p_{q}(x)>0$ for $x\in S_{oq}$.
 The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$.
-When the model $M$ is sufficiently trained on $S_{ou}$, and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
-$TVD\left(p_{q},p_{M}\right)$ $=\begin{cases}\approx 0,&\text{if }S_{oq}\subset S_{ou}\\[5.0pt] \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log{((\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|)}/(1-P_{Mq}))-H(p_{q}))},&\text{if }S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (5.11)
+When the model $M$ is sufficiently trained on $S_{ou}$, and adopts a conservative uniform prediction for unknown states, the total variation distance (TVD) between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
+$\mathrm{TVD}(p_{q},p_{M})=\begin{cases}\approx 0,&\text{if }S_{oq}\subset S_{ou}\\[10.00002pt] \leq\sqrt{\dfrac{1}{2}D_{\text{KL}}},&\text{if }S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (5.11)
+where $D_{\text{KL}}=(1\!-\!P_{\!Mq})\log\!\left(\!\dfrac{|S_{oq}|\!-\!|S_{ou}\cap S_{oq}|}{1\!-\!P_{\!Mq}}\!\right)\!-\!H(p_{q})$.
 […]
@@
 $p_{M}(x)=\begin{cases}\approx p_{q}(x),&\text{if }x\in S_{ou}\cap S_{oq}\\[3.00003pt] \frac{1-P_{Mq}}{|S_{oq}|-|S_{ou}\cap S_{oq}|},&\text{if }x\in S_{oq}\backslash S_{ou}\end{cases}$ (5.12)
-Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Goodfellow et al., 2016):
+Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have [29]:
 $D_{KL}(p_{q}(x)\parallel p_{M}(x))$ $=\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
@@
 $=(1-P_{Mq})\log\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}}-H(p_{q})$ (5.15)
-Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance.
-By Pinsker’s inequality (Pinsker, 1964), we have:
+Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their TVD.
+By Pinsker’s inequality [30], we have:
 $TVD(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\parallel p_{M}(x))}$
@@
 This establishes the second case in (5.11).
-Since we have already assumed that the carrier states $S_{M}(M,{Tu}_{m})$ and $S_{cq}(cq,{Tq}_{m})$ of $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$ and $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model $M$ and its interface, respectively.
+Since we have already assumed that the carrier states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ of $Iu:S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,Tu_{m})$ and $Iq:S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model $M$ and its interface, respectively.
 It follows that equation (5.11) characterizes the upper bound of the model’s TVD.
 This completes the proof.
-Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H\left(p_{q}\right)$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD.
+Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H(p_{q})$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD.
 Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance.
-Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD (Devroye et al., 2013), while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss (Devroye et al., 2013).
+Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD [31], while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss [32].
 Therefore, TVD serves as an indicator for estimating a model’s generalization error.
@@
 Assume that the model's fitting error on the covered region is $E_{fit}=D_{KL}\left(\widetilde{p_{q}}|_{U}\parallel p_{M}|_{U}\right)$, where $\widetilde{p_{q}}|_{U}$ and $p_{M}|_{U}$ denote the distributions restricted to $U$ and normalized (under the ideal conditions assumed in Theorem 3, $E_{fit}$ tends to zero).
-Thus, the upper bound of the TVD for model $M$ with respect to noisy input information is given by $TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.17)
+Thus, the upper bound of the TVD for model $M$ with respect to noisy input information is given by
+$TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.17)
 The upper bound of the TVD for model $M$ with respect to the true information is
@@
 It follows that
-$TVD\left(p_{q},\widetilde{p_{q}}\right)=(\sum_{x\in X}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
-$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
-$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p^{+}(x)/(1-P_{l}+P_{a})\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-p_{q}(x)/(1-P_{l}+P_{a})\right|)/2$
-$(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}{p_{q}(x)}}{1-P_{l}+P_{a}})/2=(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|(1-P_{l})}{1-P_{l}+P_{a}})/2$
-$\left\{\begin{array}[]{r}P_{l},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}},\ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.22)
+[…]
+$=\frac{1}{2}\bigg(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|$
+$\quad+\sum_{x\in X\setminus(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|\bigg)$
+$=\frac{1}{2}\bigg(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|\frac{p^{+}(x)}{1-P_{l}+P_{a}}\right|$
+$\quad+\sum_{x\in X\setminus(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\frac{p_{q}(x)}{1-P_{l}+P_{a}}\right|\bigg)$
+$=\frac{1}{2}\left(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{|P_{l}-P_{a}|\sum_{x\in X\setminus(S_{oq}^{-}\cup S_{oq}^{+})}p_{q}(x)}{1-P_{l}+P_{a}}\right)$
+$=\frac{1}{2}\left(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{|P_{l}-P_{a}|(1-P_{l})}{1-P_{l}+P_{a}}\right)$
+$=\begin{cases}P_{l},&\text{if }P_{l}\geq P_{a},\\[2.0pt] \dfrac{P_{a}}{1-P_{l}+P_{a}},&\text{if }P_{l}<P_{a}.\end{cases}$ (5.24)
 By applying the triangle inequality, we have
-$TVD\left(p_{q},p_{M}\right)\leq TVD\left(p_{q},\widetilde{p_{q}}\right)+TVD\left(\widetilde{p_{q}},p_{M}\right)$
-$\leq\left\{\begin{array}[]{r}P_{l}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.23)
+$\text{TVD}(p_{q},p_{M})$ $\leq\text{TVD}(p_{q},\widetilde{p_{q}})+\text{TVD}(\widetilde{p_{q}},p_{M})$ (5.25)
+[…]
 Thus, the second conclusion of the theorem is established, completing the proof.
@@
 (c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used.
-For almost all samples (except a few outliers), their RMSE is less than $TVD\approx\sqrt{((1-0.4951)\ln{((10000-4951)}/(1-0.4951))-2.9741)/2}\approx 0.9155$.
+For almost all samples (except a few outliers), their RMSE is less than $\mathrm{TVD}\approx\sqrt{((1-0.4951)\ln{((10000-4951)}/(1-0.4951))-2.9741)/2}\approx 0.9155$.
 This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
@@
 Meanwhile, the RMSE of conservative predictions for these samples is much lower than the TVD, reflecting the conclusion of the second case in Theorem 3, equation (5.11).
-In summary, this regression experiment using a simple feedforward neural network model demonstrates and validates the accuracy of the model TVD upper bound estimation based on the MLT-MF framework, and its potential for estimating generalization error.
-Since the theoretical proof of Theorem 3 is independent of specific models and tasks, it promises to be extensible to more complex scenarios.
 Figure 3: Experimental Evaluation and Results of the TVD Upper Bound under Noiseless Input and Conservative Prediction Strategy.
-(a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples.
-Among these, 4,951 samples overlap with the training set.
-(b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset.
-(c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound.
-(d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
 6.2 Experimental Evaluation of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy
@@
 Figure 4: Experimental Evaluation and Results of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy.
-(a) The upper arc-shaped region of the test dataset is shifted upward, and the resulting vacant region is excluded from the test input.
-The shifted segment, which does not overlap with the original test dataset, is incorporated as noise into the new test dataset.
-(b) One hundred groups of samples are randomly selected from the noisy test dataset and processed by the model.
-The RMSE of the predictions are represented by 100 points along the blue line, while the red line indicates the TVD upper bound between the noisy input distribution and the model distribution.
-(c) In the case of $P_{l}\geq P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model.
-The RMSE of each prediction is represented by 100 points along the blue line, with the red line denoting the TVD upper bound between the noiseless input distribution and the model distribution.
-(d) In the case of $P_{l}<P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model.
-The RMSE of each prediction is represented by 100 points along the blue line, with the red line indicating the TVD upper bound between the noiseless input distribution and the model distribution.
 (b) For the noisy test dataset comprising 11,034 samples, we compute the probability distribution for each test sample using the same method as in Section 6.1.
@@
 In total, 10 trials were conducted, and the results are shown in Table 1.
+In Table 1, the “Computation Time Ratio” is calculated as the computation time with ethical safety checking divided by the computation time without it.
+For the simple feedforward neural network model used in these experiments, the additional computational overhead for performing ethical safety checks before model output ranged from 1.0% to 19.2%.
+As model size increases, this overhead is expected to decrease in proportion; however, as the number of ethical safety constraints grows, the overhead will increase accordingly.
+This is a key consideration for implementing ethical safety in models.
+In these 10 trials, more than 19% of samples resulted in unethical predictions.
+In every trial, both the “Correct Prediction Rate for Safe Outputs” and the “Correct Prompt Rate for Unethical Outputs” were 100%, demonstrating that by incorporating safety checks and informative outputs, we can always guarantee the ethical safety of model predictions while maximizing the number of safe outputs.
 Table 1: Summary Table of Ethical Safety Experiment Results
-Trial Time w/o Time w/ Time Safe Warning Safe Warning
-Number Check (s) Check (s) Ratio Samples Samples Rate Rate
+Trial Number
+Computation Time
+Without Ethical
+Safety Check
+Computation Time
+With Ethical
+Safety Check
+Computation
+Time Ratio
+Number of
+Ethically Safe
+Output Samples
+Number of
+Unethical
+Warning Samples
+Correct Prediction
+Rate for Safe
+Outputs
+Correct Prompt
+Rate for
+Unethical Outputs
 1 0.367 0.391 1.063 4517 483 100% 100%
@@
 10 0.355 0.424 1.192 4533 467 100% 100%
-In Table 1, the “Computation Time Ratio” is calculated as the computation time with ethical safety checking divided by the computation time without it.
-For the simple feedforward neural network model used in these experiments, the additional computational overhead for performing ethical safety checks before model output ranged from 1.0% to 19.2%.
-As model size increases, this overhead is expected to decrease in proportion; however, as the number of ethical safety constraints grows, the overhead will increase accordingly.
-This is a key consideration for implementing ethical safety in models.
-In these 10 trials, more than 19% of samples resulted in unethical predictions.
-In every trial, both the “Correct Prediction Rate for Safe Outputs” and the “Correct Prompt Rate for Unethical Outputs” were 100%, demonstrating that by incorporating safety checks and informative outputs, we can always guarantee the ethical safety of model predictions while maximizing the number of safe outputs.
-In summary, this ethical safety experiment, grounded in Definition 5 and the formal MLT-MF-based framework for ethical safety, illustrates and validates Theorem 2 regarding the feasibility of integrating ethical safety checking mechanisms and achieving the maximal number of safe outputs.
 6.4 Interpretability Experiment
@@
 5.
-Figure 5: Schematic diagram of the computational process for model M after 165 training steps, showing its handling of the input sample $X=[-1.3535,-1.3535]^{T}$ and output $Y\approx[-0.040,0.277]^{T}$.
+Figure 5: Schematic diagram of the computational process for model M after 165 training steps.
 Compared with equations (6.1)–(6.3), Fig.
@@
 Figure 6: Illustration of Prediction Polarity and Impact of Training Steps.
-(a) Polarity distribution of predicted data around the training set region after model M has been trained for 165 steps.
-The blue square marks the area where the training dataset is located, centered at (1, 1) with a side length of 1.
-Green dots indicate samples with all-positive predictions; dark green dots indicate samples with all-positive predictions that also belong to the test set; red dots indicate samples with at least one negative prediction that belong to the test set; orange dots indicate samples with at least one negative prediction that do not belong to the test set.
-(b) Cumulative effect of up to 400 training steps on the prediction polarity at two points: (1, 1) and (-1.3535, -1.3535).
-The blue and green lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (1, 1) from the training set; the red and pink lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (-1.3535, -1.3535) from the test set.
-Fig.
-6(a) presents an overall view of numerous inputs processed and predicted by model M.
-This visualization also represents an ontological state of M as it handles a wide range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping.
-Thus, according to Definition 4, Fig.
-6(a) constitutes a valid explanation for the distribution of prediction polarities by M.
 6.4.4 Explanation Based on Cumulative Training Effects
@@
 In fact, by applying the forward propagation rules (C.21) and (C.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.
-7 Discussion and Conclusion
+7 Conclusion
 This work, grounded in the Objective Information Theory (OIT), develops the MLT-MF framework, employing a formal logical system to integrate the complete causal chain of model construction, training, input processing, output, and online learning.
-Under a unified semantics, we propose formal definitions of model interpretability and ethical safety, derive two classes of explicit TVD upper bound theorems, and validate the framework through experiments with feedforward neural network models.
-The core contributions are as follows:
-(1) Enabled Mapping for Information Transmission: We characterize the mathematical mapping and the dual physical-causal nature of information transfer from ontological states to carrier states through the concept of “enabled mapping.” This is further extended to chain transmission, mapping each engineering stage of model development to a segment of the information chain, such that model behavior, training, and inference are reduced to logically verifiable compositions of mappings.
-(2) Unified Representation of Interpretability and Safety: By adopting higher-order formal logic as a unified representation of “state,” we propose necessary and sufficient conditions for interpretability—linking input, output, and model within the same formal system via logical closure—and construct maximal safe outputs for ethical safety based on hypergraph maximum independent sets.
-(3) Integration of Statistical Error, Noise Structure, and Coverage Geometry: We integrate statistical learning error characterization with information noise structure and training coverage geometry, providing a closed-form TVD upper bound, and operationally decomposing it in terms of coverage fitting error, uncovered mass, KL divergence, and noise deletion/addition terms.
-Therefore, this study addresses two longstanding gaps in machine learning theory: the lack of unified semantics and causal chain constraints across different stages, and the absence of verifiable mathematical definitions and constructive guarantees for ethical safety and interpretability.
-By unifying the “existence–transmission–realization” of information within formal logical explanations, we rigorously establish that model interpretability corresponds to information existence, thereby providing a logical-theoretical foundation for translating ethical safety from normative requirements into decidable and constructible guarantees.
-Using TVD as a fundamental measure of distributional alignment, and establishing correspondences with practical engineering concerns such as training coverage gaps, uncovered-region prediction strategies, and noise deletion/addition, this work bridges the gap between abstract generalization theory and data coverage/system implementation, thus connecting statistical learning, mathematical logic, information science, and systems engineering.
-All conclusions and methods presented are amenable to engineering implementation:
-  • Interpretability can be realized at multiple levels through unified formal systems, including mathematical formulas, dataflow diagrams, and parameter representations, which is especially important for high-risk industries.
-  • Ethical safety is implemented via conflict detection and maximal independent set approximation of output propositions, enabling a deployable process for maximal safe outputs.
-[…]
-Of course, this work has limitations: the current formal system is most transparent for finite discrete sets, and further refinement is needed for continuous, time-varying, multimodal, and multi-agent scenarios, particularly regarding measurability and dynamic consistency.
+Under a unified semantics, the author proposes formal definitions of model interpretability and ethical safety, derive two classes of explicit TVD upper bound theorems, and validate the framework through experiments with feedforward neural network models.
+Of course, this study has limitations: the current formal system is most transparent for finite discrete sets, and further refinement is needed for continuous, time-varying, multimodal, and multi-agent scenarios, particularly regarding measurability and dynamic consistency.
 The maximal independent set for ethical safety is generally NP-hard, requiring the development of complexity-controlled approximation algorithms and alignment with business ethics standards.
@@
 Future work may include large-scale empirical studies to verify the observability of TVD upper bounds and the benefits of strategy optimization, as well as further development of cross-modal and multi-agent information chain composition and conflict resolution methods, thereby enhancing the universality and practical value of MLT-MF in complex systems.
-Acknowledgments and Disclosure of Funding
+Acknowledgement
 Junxuan He from Beijing University of Posts and Telecommunications implemented all the experiments described in the paper according to the author’s ideas, producing complete model code and training/test datasets, providing indispensable and strong support for illustrating and validating the theorems related to MLT-MF.
@@
 I hereby express my sincere gratitude to them.
+Declaration of competing interest
+The author declare that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
 Data and Code availability
@@
 For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols.
-However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise (Hamilton, 1988).
+However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise [37].
 Therefore, they are introduced as defined symbols in subsequent discussions.
@@
 Appendix B Finite Automaton States
-Using the formal expression of states, we can study in depth the special and important object of finite automata (Editorial Committee of Dictionary of Mathematics, 2002).
+Using the formal expression of states, we can study in depth the special and important object of finite automata [38].
 Lemma B.1 (State Representation of Finite Automata) For any finite automaton $M$, one can describe its input, output, and state‐transition behavior over a related time set $T$ by the state set $S(M,T)$.
 Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions.
-By finite‐automaton theory (Editorial Committee of Dictionary of Mathematics, 2002), $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
+By finite‐automaton theory [38], $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
 $\delta$: $U\mathbf{\times}Q$ to $U$, is the next‐state function;
@@
 Similarly, the output function $\lambda$ is given by
-$\varphi_{\lambda}\left(t_{i}\right)=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land{Eq}^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land{IsElementof}^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$ (B.5)
+$\varphi_{\lambda}\left(t_{i}\right)=$ $\ \varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land$ (B.5)
+$\mathit{Eq}^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land$
+$\mathit{IsElementof}^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$
 where ${IsElementof}^{2}(x,X)$ asserts $x$ is an element of $X$, $i=1,\ldots,n-1$.
@@
 $\begin{split}S(M,T)=\{&\varphi_{U}\left(t_{i}\right),\varphi_{I}\left(t_{j}\right),\varphi_{O}\left(t_{k}\right),\varphi_{\delta}\left(t_{j}\right),\varphi_{\lambda}\left(t_{j}\right),\\ &i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}\end{split}$ (B.6)
-describes $M$’s inputs, outputs, and state‐transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
+describes $M$’s inputs, outputs, and state‐transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of WFFs $X$. This completes the proof.
 It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits automatic transition behavior. Hence we introduce:
@@
 We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig.
-7).
+C.1).
 Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
@@
 Defining the symbols is a prerequisite for formal modeling.
-To this end, Tables 2–7 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
+To this end, Tables C.1–4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
 Table 1: Definition of Constants
@@
 $\mathbb{R}^{3}$ Space 3-dimensional real space
-Table 3: Definition of Variables - Part I: Logical Formulas and Basic Variables
+Table 2: Definition of Variables
 Symbol Type Definition
 $i,j,k$ natural number index or count
-$\varphi o_{i}$ WFFs well-formed formulas in initial model ontological state
-$\psi o_{i}$ WFFs well-formed formulas in initial model carrier state
-$\varphi t_{i}$ WFFs well-formed formulas in training ontological state
-$\psi t_{i}$ WFFs well-formed formulas in training carrier state
-$\varphi u_{i}$ WFFs well-formed formulas in optimized model ontological state
-$\psi u_{i}$ WFFs well-formed formulas in optimized model carrier state
-$\varphi q_{i}$ WFFs well-formed formulas in input ontological state
-$\psi q_{i}$ WFFs well-formed formulas in input carrier state
-$\varphi r_{i}$ WFFs well-formed formulas in output ontological state
-$\psi r_{i}$ WFFs well-formed formulas in output carrier state
-$\varphi v_{i}$ WFFs well-formed formulas in re-optimized model ontological state
-$\psi v_{i}$ WFFs well-formed formulas in re-optimized model carrier state
-layerslayers natural number number of layers in the model
+$\varphi_{i}^{o}$ WFFs WFFs contained in the informational ontological state of M’s initial model
+$\psi_{i}^{o}$ WFFs WFFs contained in the informational carrier state of M’s initial model
+$\varphi_{i}^{t}$ WFFs WFFs contained in the training informational ontological state
+$\psi_{i}^{t}$ WFFs WFFs contained in the training informational carrier state
+$\varphi_{i}^{u}$ WFFs WFFs contained in the informational ontological state of M’s optimized model
+$\psi_{i}^{u}$ WFFs WFFs contained in the informational carrier state of M’s optimized model
+$\varphi_{i}^{q}$ WFFs WFFs contained in the input informational ontological state of M
+$\psi_{i}^{q}$ WFFs WFFs contained in the input informational carrier state of M
+$\varphi_{i}^{r}$ WFFs WFFs contained in the output informational ontological state of M
+$\psi_{i}^{r}$ WFFs WFFs contained in the output informational carrier state of M
+$\varphi_{i}^{v}$ WFFs WFFs contained in the informational ontological state of M’s re-optimized model
+$\psi_{i}^{v}$ WFFs WFFs contained in the informational carrier state of M’s re-optimized model
+layers natural number number of layers in the model
 $v$ real number numerical values for weights and biases
-Table 4: Definition of Variables - Part II: Neural Network Parameters
-Symbol Type Definition
-$W_{k}^{(o)}$ matrix weight matrix from k-th to (k+1)-th layer (initial)
-$w_{ij}^{(o)k}$ real number weight from j-th to i-th neuron (initial)
-$b_{k}^{(o)}$ vector bias vector of k-th layer (initial)
-$b_{i}^{(o)k}$ real number bias of i-th neuron in k-th layer (initial)
-$W_{k}^{(u)}$ matrix weight matrix from k-th to (k+1)-th layer (optimized)
-$w_{ij}^{(u)k}$ real number weight from j-th to i-th neuron (optimized)
-$b_{k}^{(u)}$ vector bias vector of k-th layer (optimized)
-$b_{i}^{(u)k}$ real number bias of i-th neuron in k-th layer (optimized)
-$W_{k}^{(v)}$ matrix weight matrix from k-th to (k+1)-th layer (re-optimized)
-$w_{ij}^{(v)k}$ real number weight from j-th to i-th neuron (re-optimized)
-$b_{k}^{(v)}$ vector bias vector of k-th layer (re-optimized)
-$b_{i}^{(v)k}$ real number bias of i-th neuron in k-th layer (re-optimized)
-$W_{k}^{(w)}$ matrix weight matrix from k-th to (k+1)-th layer (further re-optimized)
-$w_{ij}^{(w)k}$ real number weight from j-th to i-th neuron (further re-optimized)
-$b_{k}^{(w)}$ vector bias vector of k-th layer (further re-optimized)
-$b_{i}^{(w)k}$ real number bias of i-th neuron in k-th layer (further re-optimized)
-Table 5: Definition of Variables - Part III: Vectors and Optimization Parameters
-Symbol Type Definition
-$X$ vector input vector
-$H$ vector hidden layer vector
-$Y$ vector output vector
-$Y_{t}$ vector output reference vector
-$x_{i}$ real number i-th element of input vector
-$h_{i}$ real number i-th element of hidden layer vector
-$y_{i}$ real number i-th element of output vector
-$y_{ti}$ real number i-th element of output reference vector
+$\mathbf{W}_{k}^{(o)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the initial model
+$w_{ij}^{(o)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the initial model
+$\mathbf{b}_{k}^{(o)}$ vector bias vector of neurons in the $k$-th layer in the initial model
+$b_{i}^{(o)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the initial model
+$\mathbf{W}_{k}^{(u)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the optimized model
+$w_{ij}^{(u)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the optimized model
+$\mathbf{b}_{k}^{(u)}$ vector bias vector of neurons in the $k$-th layer in the optimized model
+$b_{i}^{(u)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the optimized model
+$\mathbf{W}_{k}^{(v)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the re-optimized model
+$w_{ij}^{(v)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the re-optimized model
+$\mathbf{b}_{k}^{(v)}$ vector bias vector of neurons in the $k$-th layer in the re-optimized model
+$b_{i}^{(v)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the re-optimized model
+$\mathbf{W}_{k}^{(w)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the further re-optimized model
+$w_{ij}^{(w)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the further re-optimized model
+$\mathbf{b}_{k}^{(w)}$ vector bias vector of neurons in the $k$-th layer in the further re-optimized model
+$b_{i}^{(w)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the further re-optimized model
+$\mathbf{X}$ vector input vector
+$\mathbf{H}$ vector hidden layer vector
+$\mathbf{Y}$ vector output vector
+$\mathbf{Y}_{t}$ vector output reference vector
+$x_{i}$ real number the $i$-th element of the input vector
+$h_{i}$ real number the $i$-th element of the hidden layer vector
+$y_{i}$ real number the $i$-th element of the output vector
+$y_{ti}$ real number the $i$-th element of the output reference vector
 $t$ time time point of state occurrence
@@
 $L$ real number loss function value
-$\nabla_{W_{1}}L$ real number gradient of loss function w.r.t. $W_{1}$
-$\nabla_{b_{1}}L$ real number gradient of loss function w.r.t. $b_{1}$
-$\nabla_{W_{2}}L$ real number gradient of loss function w.r.t. $W_{2}$
-$\nabla_{b_{2}}L$ real number gradient of loss function w.r.t. $b_{2}$
-$\eta$ positive real learning rate
+$\nabla_{\mathbf{W}_{1}}L$ real number gradient of the loss function with respect to $\mathbf{W}_{1}$
+$\nabla_{\mathbf{b}_{1}}L$ real number gradient of the loss function with respect to $\mathbf{b}_{1}$
+$\nabla_{\mathbf{W}_{2}}L$ real number gradient of the loss function with respect to $\mathbf{W}_{2}$
+$\nabla_{\mathbf{b}_{2}}L$ real number gradient of the loss function with respect to $\mathbf{b}_{2}$
+$\eta$ positive real number learning rate
 Table 3: Definition of Functions
 Symbol Number of Elements Definition
-$ReLU(x)$ 1 rectified linear unit (ReLU) activation function
-$Identity(x)$ 1 identity function
+$\text{ReLU}(x)$ 1 rectified linear unit (ReLU) activation function
+$\text{Identity}(x)$ 1 identity function
 $\max$ multivariate maximum function
-$MSE(Y,Y_{t})$ 2 mean squared error between vector $Y$ and $Y_{t}$
-${ReLU}^{\prime}(x)$ 1 derivative of $ReLU(x)$
-$Target(X)$ 1 objective function for input vector $X$
+$\text{MSE}(\mathbf{Y},\mathbf{Y}_{t})$ 2 mean squared error between vector $\mathbf{Y}$ and $\mathbf{Y}_{t}$
+$\text{ReLU}^{\prime}(x)$ 1 derivative of $\text{ReLU}(x)$
+$\text{Target}(\mathbf{X})$ 1 objective function for input vector $\mathbf{X}$
 Table 4: Definition of Predicates
 Symbol Definition
-${NetworkStructure}^{2}(layers,k)$ model structure has k layers
-${LayerSize}^{2}(k,j)$ the k-th layer has j neurons
-${Eq}^{2}(x,y)$ $x$ equals $y$
-${InputAt}^{2}(X,t)$ input $X$ at time $t$
-${HiddenAt}^{2}(H,t)$ hidden layer value is $H$ at time $t$
-${OutputAt}^{2}(Y,t)$ output $Y$ at time $t$
-${InputAt}^{3}(X,Y_{t},t)$ input vector $X$ and label vector $Y_{t}$ at time $t$
-${Differentiable}^{2}(f,v)$ function $f$ is differentiable with respect to $v$
-${BelongsTo}^{2}(x,X)$ element $x$ belongs to set $X$
+$\text{NetworkStructure}^{2}(\text{layers},k)$ model structure has $k$ layers
+$\text{LayerSize}^{2}(k,j)$ the $k$-th layer has $j$ neurons
+$\text{Eq}^{2}(x,y)$ $x$ equals $y$
+$\text{InputAt}^{2}(\mathbf{X},t)$ input $\mathbf{X}$ at time $t$
+$\text{HiddenAt}^{2}(\mathbf{H},t)$ hidden layer value is $\mathbf{H}$ at time $t$
+$\text{OutputAt}^{2}(\mathbf{Y},t)$ output $\mathbf{Y}$ at time $t$
+$\text{InputAt}^{3}(\mathbf{X},\mathbf{Y}_{t},t)$ input vector $\mathbf{X}$ and label vector $\mathbf{Y}_{t}$ at time $t$
+$\text{Differentiable}^{2}(f,v)$ function $f$ is differentiable with respect to $v$
+$\text{BelongsTo}^{2}(x,X)$ element $x$ belongs to set $X$
 C.3 Formal Representations of Information at Each Stage of Model M
@@
 According to the principles of OIT and based on MLT-MF, the model M is endowed with its initial information from the very beginning:
-$Io=\left\langle oo,{To}_{h},S_{oo},M,{To}_{m},S_{M}\right\rangle$ (C.13)
-Here, the ontology $oo$ represents the designer’s conceptual blueprint; ${To}_{h}$ is the time at which the model is designed; and the ontological state $S_{oo}(oo,{To}_{h})$ is a set of well-formed formulas that describe the initial operational logic of model M:
+$Io:S_{oo}(oo,To_{h})\rightrightarrows S_{M}(M,To_{m})$ (C.13)
+Here, the ontology $oo$ represents the designer’s conceptual blueprint; ${To}_{h}$ is the time at which the model is designed; and the ontological state $S_{oo}(oo,{To}_{h})$ is a set of WFFs that describe the initial operational logic of model M:
 $S_{oo}\left(oo,{To}_{h}\right)=\left\{\varphi o_{1},\varphi o_{2},\varphi o_{3},\varphi o_{4},\varphi o_{5},\varphi o_{6},\varphi o_{7},\varphi o_{8},\varphi o_{9},\varphi o_{10},\varphi o_{11}\right\}$ (C.14)
 The specific content of each formula is as follows:
+  • 1.
     Network Structure Definition
-$\varphi o_{1}={NetworkStructure}^{2}(layers,3)\land{LayerSize}^{2}(1,2)\land{LayerSize}^{2}(2,3)\land{LayerSize}^{2}(3,2)$ (C.15)
+$\varphi o_{1}=$ $\ \mathit{NetworkStructure}^{2}(layers,3)\land$ (C.15)
+$\mathit{LayerSize}^{2}(1,2)\land\mathit{LayerSize}^{2}(2,3)\land$
+$\mathit{LayerSize}^{2}(3,2)$
+  • 1.
     Weight Initialization
-$\varphi o_{2}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)1},v\right))$ (C.16)
-$\varphi o_{3}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2,3\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)2},v\right))$ (C.17)
+$\varphi o_{2}=$ $\ \forall i\forall j\left(\mathit{BelongsTo}^{2}\left(i,\{1,2,3\}\right)\land\mathit{BelongsTo}^{2}\left(j,\{1,2\}\right)\right$. (C.16)
+$\left.\rightarrow\exists v\left(\mathit{Eq}^{2}\left(w_{ij}^{(o)1},v\right)\right)\right)$
+$\varphi o_{3}=$ $\ \forall i\forall j\left(\mathit{BelongsTo}^{2}\left(i,\{1,2\}\right)\land\mathit{BelongsTo}^{2}\left(j,\{1,2,3\}\right)\right$. (C.17)
+$\left.\rightarrow\exists v\left(\mathit{Eq}^{2}\left(w_{ij}^{(o)2},v\right)\right)\right)$
+  • 1.
     Bias Initialization
@@
 $\varphi o_{5}=\forall i{BelongsTo}^{2}\left(i,\left\{1,2\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(b_{i}^{(o)2},v\right))$ (C.19)
+  • 1.
     Activation Function Definition
 $\varphi o_{6}=\forall x({Eq}^{2}(ReLU(x),max(0,x)))\land\forall x({Eq}^{2}(Identity(x),x))$ (C.20)
+  • 1.
     Forward Propagation Rules
-$\begin{split}\varphi o_{7}=&\forall X\forall t({InputAt}^{2}(X,t)\land{BelongsTo}^{2}\left(X,\mathbb{R}^{2}\right)\rightarrow\exists H({BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\\ &\land{HiddenAt}^{2}(H,t+\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right)\\ &\rightarrow{Eq}^{2}(h_{i},ReLU\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right)))))\end{split}$ (C.21)
-$\begin{split}\varphi o_{8}=&\forall H\forall t({HiddenAt}^{2}(H,t+\delta)\land{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\rightarrow\exists Y({BelongsTo}^{2}\left(Y,\mathbb{R}^{2}\right)\\ &\land{OutputAt}^{2}(Y,t+2\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2\right\}\right)\\ &\rightarrow{Eq}^{2}\left(y_{i},Identity\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)\right))))\end{split}$ (C.22)
+$\varphi o_{7}=$ $\ \forall X\forall t\left(\mathit{InputAt}^{2}(X,t)\land\mathit{BelongsTo}^{2}\left(X,\mathbb{R}^{2}\right)\rightarrow\right$. (C.21)
+$\exists H\left(\mathit{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\land\mathit{HiddenAt}^{2}(H,t+\delta)\land\right$.
+[…]
+$\left.\left.\mathit{Eq}^{2}\left(h_{i},\mathrm{ReLU}\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right)\right)\right)\right)$
+$\varphi o_{8}=$ $\ \forall H\forall t\left(\mathit{HiddenAt}^{2}(H,t+\delta)\land\mathit{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\rightarrow\right$. (C.22)
+$\exists Y\left(\mathit{BelongsTo}^{2}\left(Y,\mathbb{R}^{2}\right)\land\mathit{OutputAt}^{2}(Y,t+2\delta)\land\right$.
+[…]
+$\left.\left.\mathit{Eq}^{2}\left(y_{i},\mathrm{Identity}\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)\right)\right)\right)$
+  • 1.
     Backpropagation Rules
-$\begin{split}\varphi o_{9}=&\forall X\forall Y_{t}\forall t({InputAt}^{3}\left(X,Y_{t},t\right)\land{BelongsTo}^{2}\left(Y_{t},\mathbb{R}^{2}\right)\land\varphi o_{7}\land\varphi o_{8}\\ &{\land Eq}^{2}\left(L,MSE\left(Y,Y_{t}\right)\right)\land{Eq}^{2}\left(MSE\left(Y,Y_{t}\right),\frac{1}{2}\sum_{i=1}^{2}\left(y_{i}-y_{ti}\right)^{2}\right))\end{split}$ (C.23)
+$\varphi o_{9}=$ $\ \forall X\forall Y_{t}\forall t\left(\mathit{InputAt}^{3}\left(X,Y_{t},t\right)\land\mathit{BelongsTo}^{2}\left(Y_{t},\mathbb{R}^{2}\right)\land\right$. (C.23)
+$\left.\varphi o_{7}\land\varphi o_{8}\land\mathit{Eq}^{2}\left(L,\mathrm{MSE}\left(Y,Y_{t}\right)\right)\land\right$.
+$\left.\mathit{Eq}^{2}\left(\mathrm{MSE}\left(Y,Y_{t}\right),\frac{1}{2}\sum_{i=1}^{2}\left(y_{i}-y_{ti}\right)^{2}\right)\right)$
 $\varphi o_{10}$ $=\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}$
-$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{1}^{(o)})$
-$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{1}^{(o)})$
-$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{2}^{(o)})$
-$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{2}^{(o)})$
-$\quad\land{Eq}^{2}(\nabla_{W_{1}^{(o)}}L,$
-$\qquad({(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})$
-$\quad\land{Eq}^{2}(\nabla_{b_{1}^{(o)}}L,$
-$\qquad{(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))$
-$\quad\land{Eq}^{2}(\nabla_{W_{2}^{(o)}}L,(Y-Y_{t}){(ReLU(W_{1}^{(o)}X+b_{1}^{(o)}))}^{T})$
-$\quad\land{Eq}^{2}(\nabla_{b_{2}^{(o)}}L,(Y-Y_{t}))$ (C.24)
+$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),W_{1}^{(o)})$
+$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),b_{1}^{(o)})$
+$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),W_{2}^{(o)})$
+$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),b_{2}^{(o)})$
+$\quad\land\mathit{Eq}^{2}(\nabla_{W_{1}^{(o)}}L,$
+[…]
+$\quad\land\mathit{Eq}^{2}(\nabla_{b_{1}^{(o)}}L,$
+[…]
+$\quad\land\mathit{Eq}^{2}(\nabla_{W_{2}^{(o)}}L,(Y-Y_{t}){(\mathit{ReLU}(W_{1}^{(o)}X+b_{1}^{(o)}))}^{T})$
+$\quad\land\mathit{Eq}^{2}(\nabla_{b_{2}^{(o)}}L,(Y-Y_{t}))$ (C.24)
+  • 1.
     Learning Rules
@@
 In equation (A.14), the carrier M refers to the initially constructed model, and ${To}_{m}$ denotes the time when model M is completed. Clearly, we have:
-$sup({To}_{h})\leq inf({To}_{m})$ (C.26)
-The state $S_{M}(M,{To}_{m})$ is a one-to-one enabling mapping of the state $S_{oo}\left(oo,{To}_{h}\right)$ onto M, that is,
+$max({To}_{h})\leq min({To}_{m})$ (C.26)
+The state $S_{M}(M,{To}_{m})$ is a noisy-free enabling mapping of the state $S_{oo}\left(oo,{To}_{h}\right)$ onto M, that is,
 $S_{M}\left(M,{To}_{m}\right)=\left\{\psi o_{1},\psi o_{2},\psi o_{3},\psi o_{4},\psi o_{5},\psi o_{6},\psi o_{7},\psi o_{8},\psi o_{9},\psi o_{10},\psi o_{11}\right\}$ (C.27)
 and
-$Io\left(\varphi o_{i}\right)\trianglerighteq\psi o_{i}(1\leq i\leq 11)$ (C.28)
+$Io:\varphi o_{i}\rightrightarrows\psi o_{i}\quad(1\leq i\leq 11)$ (C.28)
 Here, the information $Io$ enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M.
 In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise.
-However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, noise-free, and recoverable information model.
+However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, and noise-free information model.
 The subsequent discussions on other types of information will follow this approach and will not be repeated.
@@
 Based on MLT-MF, the training information for model M is as follows:
-$It=\left\langle ot,{Tt}_{h},S_{ot},Mt,{Tt}_{m},S_{Mt}\right\rangle$ (C.29)
+$It:S_{ot}(ot,Tt_{h})\rightrightarrows S_{Mt}(Mt,Tt_{m})$ (C.29)
 Here, the ontology $ot$ refers to the object described by the training information, and the occurrence time ${Tt}_{h}$ denotes the point at which this object attains its corresponding state attributes.
-The ontological state $S_{ot}(ot,{Tt}_{h})$ represents the attributes of the object at time ${Tt}_{h}$, which are described by a set of well-formed formulas:
+The ontological state $S_{ot}(ot,{Tt}_{h})$ represents the attributes of the object at time ${Tt}_{h}$, which are described by a set of WFFs:
 $S_{ot}\left(ot,{Tt}_{h}\right)=\left\{\varphi t_{1},\varphi t_{2}\right\}$ (C.30)
+  • 1.
     Vector Definition
 $\varphi t_{1}=\forall X\forall Y_{t}({BelongsTo}^{2}(X,\mathbb{R}^{2})\land{BelongsTo}^{2}(Y_{t},\mathbb{R}^{2}))$ (C.31)
+  • 1.
     Label Consistency
@@
 The carrier $Mt$ serves as the interface device through which model M receives training data, and ${Tt}_{m}$ is the time at which model M is trained.
-The carrier state $S_{Mt}(Mt,{Tt}_{m})$ is then described by the set of well-formed formulas:
-$It\left(\varphi t_{i}\right)\trianglerighteq\psi t_{i}(1\leq i\leq 2)$ (C.33)
+The carrier state $S_{Mt}(Mt,{Tt}_{m})$ is then described by the set of WFFs:
+$It:\varphi t_{i}\rightrightarrows\psi t_{i}\quad(1\leq i\leq 2)$ (C.33)
 (3) Optimized Information After M Has Learned from Training
 Based on MLT-MF, after learning the training information $It$, model M updates its initial information $Io$ to a new, optimized information state:
-$Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$ (C.34)
+$Iu:S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,Tu_{m})$ (C.34)
 Here, the ontology $ou$ is a fusion of M and the object described by the training information; the occurrence time ${Tu}_{h}$ is the moment M learns from the training data; and the ontological state $S_{ou}(ou,{Tu}_{h})$ represents the theoretical state of M after training:
@@
 $S_{M}\left(M,{Tu}_{m}\right)=\left\{\psi u_{1},\psi u_{6},\psi u_{7},\psi u_{8},\psi u_{9},\psi u_{10},\psi u_{11}\right\}$ (C.36)
-is a one-to-one enabling mapping from $S_{ou}\left(ou,{Tu}_{h}\right)$ onto M, representing the state of M after completing its training.
+is a noisy-free enabling mapping from $S_{ou}\left(ou,{Tu}_{h}\right)$ onto M, representing the state of M after completing its training.
 (4) Input Information for M
 Processing input information is a fundamental requirement for any model. Based on the MLT-MF framework, the input information for M is defined as:
-$Iq=\left\langle oq,{Tq}_{h},S_{oq},Mq,{Tq}_{m},S_{Mq}\right\rangle$ (C.37)
+$Iq:S_{oq}(oq,Tq_{h})\rightrightarrows S_{Mq}(Mq,Tq_{m})$ (C.37)
 Here, $oq$ denotes the object described by the input information, and ${Tq}_{h}$ is the time at which the object's state occurs. The state $S_{oq}(oq,{Tq}_{h})$ can take two possible forms:
+  • 1.
     The first scenario involves only input information being processed by M, in which case
@@
 $\varphi q_{1}=\forall X{BelongsTo}^{2}(X,\mathbb{R}^{2})$ (C.39)
+  • 1.
     The second scenario considers online learning, where
@@
 The carrier $Mq$ for $Iq$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$.
-The state $S_{Mq}(Mq,{Tq}_{m})$ is a one-to-one enabling mapping from $S_{oq}(oq,{Tq}_{h})$ onto $Mq$.
+The state $S_{Mq}(Mq,{Tq}_{m})$ is a noisy-free enabling mapping from $S_{oq}(oq,{Tq}_{h})$ onto $Mq$.
 (5) Information After M Processes Input, Produces Output, and Undergoes Further Optimization
@@
 Since the latter is more comprehensive, we will focus on this scenario below.
-In this case, the output information $Ir=\left\langle or,{Tr}_{h},S_{or},Mr,{Tr}_{m},S_{Mr}\right\rangle$ has an ontological state defined as:
+In this case, the output information $Ir:S_{or}(or,Tr_{h})\rightrightarrows S_{Mr}(Mr,Tr_{m})$ has an ontological state defined as:
 $S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r_{1}\right\}=\left\{{\varphi o_{7}\land\varphi o_{8}\land OutputAt}^{2}(Y,{Tr}_{h})\right\}$ (C.43)
-The carrier state $S_{Mr}(Mr,{Tr}_{m})$ is a one-to-one enabling mapping from $S_{or}\left(or,{Tr}_{h}\right)$ to the output device of M.
+The carrier state $S_{Mr}(Mr,{Tr}_{m})$ is a noisy-free enabling mapping from $S_{or}\left(or,{Tr}_{h}\right)$ to the output device of M.
 On the other hand, since we are considering the scenario where M performs online learning, the model now carries further optimized information:
-$Iv=\left\langle ov,{Tv}_{h},S_{ov},M,{Tv}_{m},S_{M}\right\rangle$ (C.44)
+$Iv:S_{ov}(ov,Tv_{h})\rightrightarrows S_{M}(M,Tr_{m})$ (C.44)
 This structure is almost identical to the optimized information $Iu$ after M has learned from training. The ontological state of $Iv$ is:
@@
 $S_{M}\left(M,{Tv}_{m}\right)=\left\{\psi v_{1},\psi v_{6},\psi v_{7},\psi v_{8},\psi v_{9},{\psi v}_{10},\psi v_{11}\right\}$ (C.46)
-is a one-to-one enabling mapping from $S_{ov}\left(ov,{Tv}_{h}\right)$ onto M.
+is a noisy-free enabling mapping from $S_{ov}\left(ov,{Tv}_{h}\right)$ onto M.
 At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig.
@@
 Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M.
-This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.
+This demonstrates that MLT-MF is capable of supporting the construction of WFF representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.

### v13

@@
 The field still lacks a unified logical framework that spans the full lifecycle from prior specification and data-driven learning to deployment and online adaptation.
-Existing theory remains fragmented across training dynamics, initialization, test-time generalization, and deployment, which prevents a coherent account of knowledge evolution across stages [1, 2, 3, 4, 5].
+Existing theory remains fragmented across training dynamics, initialization, test-time generalization, and deployment, which prevents a coherent account of knowledge evolution across stages [vapnik1999overview, alquier2024user, scholkopf2021toward, garcez2023neurosymbolic, marra2024statistical].
 Model performance is also hard to predict, which inflates development costs in data, computation, energy, and human effort.
-In the absence of explicit causal semantics, structural analysis, and stage-wise error metrics, attainable performance levels and error upper bounds are difficult to estimate, and information coverage gaps or failure modes due to model selection, measurement, and data quality are hard to anticipate [6, 7, 8, 9].
+In the absence of explicit causal semantics, structural analysis, and stage-wise error metrics, attainable performance levels and error upper bounds are difficult to estimate, and information coverage gaps or failure modes due to model selection, measurement, and data quality are hard to anticipate [achiam2023gpt4, touvron2023llama, chowdhery2023palm, gunasekar2023textbooks].
 Ethical safety lacks a settled mathematical definition and verification protocol.
-Current research is dominated by principles and heterogeneous guidelines, leaving uncertainty about standards and assurance processes [10, 11, 12, 13, 14].
+Current research is dominated by principles and heterogeneous guidelines, leaving uncertainty about standards and assurance processes [russell2015research, burr2023ethical, tilala2024ethical, hawkins2021guidance, pant2024ethics].
 Interpretability similarly lacks a rigorous, domain-agnostic foundation.
-Without a unified semantic framework for causal inference and constraint logic, the data-to-model pipeline is treated as a black box, and general, testable definitions and criteria remain elusive despite numerous approaches [15, 16, 17, 18, 19].
-Information science offers a route to unification [20, 21], but its subfields—information theory, control, signal processing, statistical learning, and computation—evolved largely in isolation and lack a common semantics.
+Without a unified semantic framework for causal inference and constraint logic, the data-to-model pipeline is treated as a black box, and general, testable definitions and criteria remain elusive despite numerous approaches [lipton2018mythos, glanois2024survey, zhang2024robustness, rudin2019stop, miller2019explanation].
+Information science offers a route to unification [shwartz2017opening, xu2022information], but its subfields—information theory, control, signal processing, statistical learning, and computation—evolved largely in isolation and lack a common semantics.
 Objective Information Theory (OIT) addresses this by axiomatizing information as an enabling mapping from ontological states to carrier states.
-Intrinsic meaning is captured by ontological states and objective expression by carrier states, and the mapping has both mathematical and physical–causal attributes [22, 23].
-Prior work shows that OIT’s general metrics can improve data evaluation and training efficiency in machine learning [24].
+Intrinsic meaning is captured by ontological states and objective expression by carrier states, and the mapping has both mathematical and physical–causal attributes [xu2014objective, xu2024research].
+Prior work shows that OIT’s general metrics can improve data evaluation and training efficiency in machine learning [xu2025general].
 We complete a remaining gap by formalizing state as an interpretation of well-formed formulas (WFFs) in a higher-order logical system, and by clarifying that enabling mappings are surjective mathematical mappings that are also physically realizable causal relations.
@@
 We derive explicit total variation distance (TVD) upper bounds in two regimes: conservative prediction in regions not covered by training under noise-free input, and arbitrary prediction under noisy input.
-The bounds are driven by measurable masses—overlap, uncovered, and noise (loss/addition)—and align with engineering indicators such as coverage and Kullback–Leibler (KL) divergence [1, 25, 26, 27, 28, 29, 30, 31, 32].
+The bounds are driven by measurable masses—overlap, uncovered, and noise (loss/addition)—and align with engineering indicators such as coverage and Kullback–Leibler (KL) divergence [vapnik1999overview, bartlett2005local, xu2017information, wang2024theoretical, huang2025convergence, goodfellow2016deep, pinsker1964information, rodriguez2021tighter, devroye2013probabilistic].
 • Practical instantiation.
@@
 2.1 Conceptualizations of Machine Learning
-Classical statistical learning theory formalizes learning as risk minimization over function classes under unknown data-generating distributions, with foundational tools such as VC theory, growth functions, and distribution-independent bounds [1].
-Parameterized predictors, loss-based risk, and empirical risk minimization provide the standard optimization lens [2].
-Recent perspectives emphasize causal structure beyond observed distributions, seeking representations that expose latent causal factors to support robust transfer and planning across environments [3].
-Neuro-symbolic approaches pursue integration of distributed representations with symbolic abstraction and compositional reasoning, aiming to combine efficient learning with discrete inference [4, 5].
+Classical statistical learning theory formalizes learning as risk minimization over function classes under unknown data-generating distributions, with foundational tools such as VC theory, growth functions, and distribution-independent bounds [vapnik1999overview].
+Parameterized predictors, loss-based risk, and empirical risk minimization provide the standard optimization lens [alquier2024user].
+Recent perspectives emphasize causal structure beyond observed distributions, seeking representations that expose latent causal factors to support robust transfer and planning across environments [scholkopf2021toward].
+Neuro-symbolic approaches pursue integration of distributed representations with symbolic abstraction and compositional reasoning, aiming to combine efficient learning with discrete inference [garcez2023neurosymbolic, marra2024statistical].
 These lines of work clarify individual components of the learning pipeline, yet they do not supply a unified semantic and causal framework that spans model construction, training, input, inference, and deployment.
@@
 Despite extensive interest, interpretability lacks a precise and generally accepted definition and evaluation protocol.
-Prior analyses highlight ambiguity in terminology and goals, and suggest that the practical objective is to extract useful information from models in ways that are audience- and task-sensitive [15].
-In reinforcement learning, interpretability depends on interpretable inputs, models, and decisions, and on notions such as simulatability, decomposability, and algorithmic transparency [16].
-Mechanistic studies relate training mechanisms to properties such as smoothness and gradient similarity to explain transferability and adversarial behavior, illustrating how dissecting key factors in training can yield explanatory leverage [17].
-Overall, current methods are diverse and often model- or context-specific, and they lack a unifying logical semantics that ties inputs, internal states, and outputs to causal explanations across the lifecycle [15, 16, 17, 18, 19].
+Prior analyses highlight ambiguity in terminology and goals, and suggest that the practical objective is to extract useful information from models in ways that are audience- and task-sensitive [lipton2018mythos].
+In reinforcement learning, interpretability depends on interpretable inputs, models, and decisions, and on notions such as simulatability, decomposability, and algorithmic transparency [glanois2024survey].
+Mechanistic studies relate training mechanisms to properties such as smoothness and gradient similarity to explain transferability and adversarial behavior, illustrating how dissecting key factors in training can yield explanatory leverage [zhang2024robustness].
+Overall, current methods are diverse and often model- or context-specific, and they lack a unifying logical semantics that ties inputs, internal states, and outputs to causal explanations across the lifecycle [lipton2018mythos, glanois2024survey, zhang2024robustness, rudin2019stop, miller2019explanation].
 2.3 Ethical Safety Assurance
 As AI systems permeate high-stakes domains, ethical safety has drawn sustained attention in law, policy, and engineering.
-Existing work emphasizes assurance processes grounded in structured arguments that justify normative claims with evidence across the system lifecycle [10, 11].
-In healthcare, recommended practices include data minimization, diversity and augmentation, algorithmic fairness, auditing, interpretability, and standards-based oversight, typically framed as a toolbox of technical and procedural controls [12].
-Safety guidance for autonomous and software-intensive systems links ML safety to software and system safety via requirement derivation and lifecycle-wide assurance activities [13].
-Practitioner studies report gaps in clear moral knowledge and the ambiguity of principles, which complicate operationalization and adoption [14].
+Existing work emphasizes assurance processes grounded in structured arguments that justify normative claims with evidence across the system lifecycle [russell2015research, burr2023ethical].
+In healthcare, recommended practices include data minimization, diversity and augmentation, algorithmic fairness, auditing, interpretability, and standards-based oversight, typically framed as a toolbox of technical and procedural controls [tilala2024ethical].
+Safety guidance for autonomous and software-intensive systems links ML safety to software and system safety via requirement derivation and lifecycle-wide assurance activities [hawkins2021guidance].
+Practitioner studies report gaps in clear moral knowledge and the ambiguity of principles, which complicate operationalization and adoption [pant2024ethics].
 What remains missing is a mathematically grounded definition of ethical safety with constructible guarantees and verifiable criteria that integrate with ML pipelines.
 2.4 Model Error Estimation
-Foundational results establish consistency and finite-sample bounds for empirical and structural risk minimization through VC dimension and related complexity measures, including conditions for one- and two-sided uniform convergence and margin-based generalization [1].
-Local Rademacher complexity sharpens bounds by tailoring complexity to low-error regions and yields optimal or near-optimal rates via fixed-point analyses under variance control [25].
-Information-theoretic approaches bound generalization via mutual information between data and hypotheses, introduce information stability, and analyze Gibbs and noisy ERM with mutual-information regularization [26].
-In meta-reinforcement learning, generalization across tasks and convergence under nonconvex optimization are characterized using empirical process theory and concentration tools, with guidance for capacity and learning-rate choices [27].
-For generative modeling, recent analyses provide TVD bounds for deterministic samplers via ODE perspectives and establish discrete-time error under Runge–Kutta schemes [28].
+Foundational results establish consistency and finite-sample bounds for empirical and structural risk minimization through VC dimension and related complexity measures, including conditions for one- and two-sided uniform convergence and margin-based generalization [vapnik1999overview].
+Local Rademacher complexity sharpens bounds by tailoring complexity to low-error regions and yields optimal or near-optimal rates via fixed-point analyses under variance control [bartlett2005local].
+Information-theoretic approaches bound generalization via mutual information between data and hypotheses, introduce information stability, and analyze Gibbs and noisy ERM with mutual-information regularization [xu2017information].
+In meta-reinforcement learning, generalization across tasks and convergence under nonconvex optimization are characterized using empirical process theory and concentration tools, with guidance for capacity and learning-rate choices [wang2024theoretical].
+For generative modeling, recent analyses provide TVD bounds for deterministic samplers via ODE perspectives and establish discrete-time error under Runge–Kutta schemes [huang2025convergence].
 These advances quantify error in specific settings but do not explicitly model causal semantics across multi-stage information flows, nor do they integrate noise structure, coverage geometry, and uncovered-region strategies into a single computable bound that can guide data auditing and prediction strategy design.
 2.5 Toward a Unified Information-Theoretic and Logical Perspective
-Information science offers a unifying lens, yet its subfields have developed largely in silos [20, 21].
-OIT axiomatizes information as an enabling mapping from ontological to carrier states with dual mathematical and physical–causal attributes, and provides general metrics that have shown value for data evaluation in ML [22, 23, 24].
-By formalizing states as interpretations of WFFs in a higher-order logical system and introducing notions of noisy information and information chains, OIT supports analysis of multi-stage pipelines and causal dependence structures [22, 23, 33, 34].
+Information science offers a unifying lens, yet its subfields have developed largely in silos [shwartz2017opening, xu2022information].
+OIT axiomatizes information as an enabling mapping from ontological to carrier states with dual mathematical and physical–causal attributes, and provides general metrics that have shown value for data evaluation in ML [xu2014objective, xu2024research, xu2025general].
+By formalizing states as interpretations of WFFs in a higher-order logical system and introducing notions of noisy information and information chains, OIT supports analysis of multi-stage pipelines and causal dependence structures [xu2014objective, xu2024research, qiu2025research, wiener2019cybernetics].
 The present work builds on this foundation to offer a single formal semantics for lifecycle reasoning about interpretability, ethical safety, and computable TVD-based error control.
@@
 We ground our meta-framework in a formal account of information and state that supports causal analysis across stages.
-3.1 Information as Enabling Mapping
-[23] based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notion of “enabling mapping”.
+3.1 A Unified Information Model of Semantic Intension, Statistical Characteristics, and Physical Realization
+3.1.1 Information as Enabling Mapping
+[xu2024research] based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notion of “enabling mapping”.
 Here, we proceed to further clarify the concept.
@@
 Figure 1: Schematic Diagram of Enabling Mapping, Noisy Information, and Information Chain.
-Based on Appendix A’s higher-order formal system $\mathcal{L}$ and the result that virtually any state set can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ [33], we obtain a more robust theoretical foundation for defining information:
+Based on Appendix A’s higher-order formal system $\mathcal{L}$ and the result that virtually any state set can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ [qiu2025research], we obtain a more robust theoretical foundation for defining information:
 Definition 2 (Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain.
@@
 Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ or $I=\left\langle o,T_{h},S_{o},c,T_{m},S_{c}\right\rangle$.
-The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world [34].
-Definition 2 subsumes Shannon’s model: source events and channel codes are states, and source-to-channel is an enabling mapping; OIT thus generalizes classical information theory [35].
-3.2 Noisy Information
-[22] states that an information instance $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ admits an inverse mapping $I^{-1}(S_{c}(c,T_{m}))$.
-In practice, however, information realization is rarely immune to noise or other external factors.
-Consequently, applying $I^{-1}$ may fail to fully recover $S_{o}(o,T_{h})$; instead, it yields a state corrupted by noise.
-This motivates the following notion.
-Definition 3 (Noisy Information) For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, the inverse $I^{-1}(S_{c}(c,T_{m}))$ is another state of the ontology $o$ at time $T_{h}$, denoted by $\widetilde{S_{o}}\left(o,T_{h}\right)$, and given by
+The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world [wiener2019cybernetics].
+Definition 2 subsumes Shannon’s model: source events and channel codes are states, and source-to-channel is an enabling mapping; OIT thus generalizes classical information theory [shannon1948mathematical].
+3.1.2 Ideal Information
+Definition 3 (Semantic Inclusion) Let $S$ be a foundational semantic system.
+For a logical system $L$, if there exists a semantic mapping $\mathrm{Sem}_{L}$ from $L$ into $S$ such that, for every symbol $x$ in $L$, $x$ is synonymous with $\mathrm{Sem}_{L}(x)$, then $S$ is said to semantically include $L$.
+If $L_{1}$ and $L_{2}$ are both semantically included by $S$, then $L_{1}$ and $L_{2}$ are said to be synonymous relative to $S$.
+Definition 4 (Synonymous States) Let $S_{x_{1}}(x_{1},t_{1})$ and $S_{x_{2}}(x_{2},t_{2})$ be sets of states expressible in logical systems $L_{1}$ and $L_{2}$, respectively.
+If $L_{1}$ and $L_{2}$ are synonymous relative to $S$, and for every $\varphi_{1}\in S_{x_{1}}(x_{1},t_{1})$ there exists $\varphi_{2}\in S_{x_{2}}(x_{2},t_{2})$, and for every $\varphi_{2}\in S_{x_{2}}(x_{2},t_{2})$ there exists $\varphi_{1}\in S_{x_{1}}(x_{1},t_{1})$, such that $\mathrm{Sem}_{L_{1}}(\varphi_{1})=\mathrm{Sem}_{L_{2}}(\varphi_{2})$, then $S_{x_{1}}(x_{1},t_{1})$ and $S_{x_{2}}(x_{2},t_{2})$ are said to be synonymous relative to $S$, or simply synonymous.
+Definition 5 (Ideal Information) Consider information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$.
+If $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are synonymous relative to the foundational semantic system $S$, then $I$ is called ideal information with respect to $S$; when no confusion arises, we may simply call $I$ ideal information.
+Since the ontological state $S_{o}(o,T_{h})$ and the carrier state $S_{c}(c,T_{m})$ of ideal information $I$ are synonymous, either $S_{o}(o,T_{h})$ or $S_{c}(c,T_{m})$ can be taken to represent $I$.
+3.1.3 The Semantic Intension of Information
 […]
-where $S_{o}^{-}$ is the noise-loss state with $S_{o}^{-}\subseteq S_{o}$, and $S_{o}^{+}$ is the noise-addition state with $S_{o}{\cap S}_{o}^{+}=\varnothing$.
-The carrier state $S_{c}(c,T_{m})$ can thus be viewed as the image of $\widetilde{S_{o}}\left(o,T_{h}\right)$ under an enabling mapping:
-$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows S_{c}\left(c,T_{m}\right)$ (3.3)
-which is the noisy information corresponding to $I$, while $I$ itself is the noise-free information.
+Definition 7 (Logical Consequences and Implicational Information Amount) For the ontological state $S_{o}$ of ideal information $I$, the logical closure $\mathrm{Cn}(S_{o})$ contains all logical consequences of $I$, and $|\mathrm{Cn}(S_{o})|$ is the implicational information amount of $I$.
+Definition 8 (Atomic Information and Primitive Information Amount) For the ontological state $S_{o}$ of ideal information $I$, suppose $\varphi$ cannot be derived from the other WFFs in $S_{o}$; then $\varphi$ is called an atomic state of $S_{o}$, and $I(\varphi)$ is an atomic piece of information in $I$.
+Let $\mathrm{Atom}(S_{o})=\{\varphi\mid\varphi\text{ is an atomic state in }S_{o}\}$.
+Then $|\mathrm{Atom}(S_{o})|$ is the primitive information amount of $I$.
+Explicit assertions, logical consequences, and ontic atomic states all belong to the semantic intension of information.
+Definition 9 (Logical Depth) For the ontological state $S_{o}$ of ideal information $I$, let $\varphi\in S_{o}$ or $\varphi\in\mathrm{Cn}(S_{o})$.
+If deriving $\varphi$ from the atomic states of $S_{o}$ requires $i$ ($i=0,1,\dotsm$) recursive layers, then the logical depth of $\varphi$ in $S_{o}$ is $i$, denoted $\mathrm{Ld}(\varphi)$.
+The logical depths of $S_{o}$ and $\mathrm{Cn}(S_{o})$ are defined by $\mathrm{Ld}(S_{o})=\max\{\mathrm{Ld}(\varphi)\mid\varphi\in S_{o}\}$ and $\mathrm{Ld}(\mathrm{Cn}(S_{o}))=\max\{\mathrm{Ld}(\varphi)\mid\varphi\in\mathrm{Cn}(S_{o})\}$, respectively.
+3.1.4 The Extensional Characteristics of Information
+Definition 10 (Extensional Characteristics of Information) The extensional characteristics of information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ in a logical system $L$ are the states $S(S_{o}\cup S_{c})$ expressible in $L$ by $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$, denoted $\mathrm{Ext}(I)$.
+Statistical characteristics are an important component of extensional characteristics.
+For ideal information $I$, the explicit information amount $|S_{o}|$, the implicational information amount $|\mathrm{Cn}(S_{o})|$, the primitive information amount $|\mathrm{Atom}(S_{o})|$, and the logical depths $\mathrm{Ld}(S_{o})$ and $\mathrm{Ld}(\mathrm{Cn}(S_{o}))$ are all statistical characteristics and belong to $\mathrm{Ext}(I)$.
+Definition 11 (Information Entropy) For the ontological state $S_{o}=\{\varphi_{i}\mid i=1,\dotsm,|S_{o}|\}$ of ideal information $I$, suppose the probability of each $\varphi_{i}$ is $p_{i}$ ($i=1,\dotsm,|S_{o}|$).
+Then $H(I)=\sum_{i=1}^{|S_{o}|}p_{i}$ is the information entropy of $I$.
+It also belongs to the statistical characteristics of $\mathrm{Ext}(I)$.
+3.1.5 Physical Realization of Information
+The physical realization of information is an enabling mapping.
+For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, practical circumstances—due to constraints such as channel conditions, environment, energy, and time—make it difficult to ensure that $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are synonymous relative to the foundational semantic system $S$.
+Definition 12 (Noisy Information) For information $I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, it is not difficult to show that there always exists a set $\widetilde{S_{o}}(o,T_{h})=S_{o}(o,T_{h})\setminus S_{o}^{-}(o,T_{h})\cup S_{o}^{+}(o,T_{h})$ such that $\widetilde{S_{o}}(o,T_{h})$ and $S_{c}(c,T_{m})$ are synonymous relative to $S$.
+Here $S_{o}^{-}$ is called the noise-loss state set, with $S_{o}^{-}\subseteq S_{o}$; $S_{o}^{+}$ is called the noise-addition state set, with $S_{o}\cap S_{o}^{+}=\emptyset$.
+In this case, $S_{c}(c,T_{m})$ can also be viewed as the result of an enabling mapping from $\widetilde{S_{o}}(o,T_{h})$, namely $\widetilde{I}:\widetilde{S_{o}}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$, which is called the noisy information of $I$.
 Fig.
 1(b) depicts the structural characteristics of noisy information.
-In Shannon’s terms, noise transforms the source state via deletion and addition before transmission; the resulting channel output realizes the noisy carrier state [35].
+In Shannon’s terms, noise transforms the source state via deletion and addition before transmission; the resulting channel output realizes the noisy carrier state [shannon1948mathematical].
 This structural decomposition will support our TVD bounds by isolating loss and addition masses.
 3.2 Information Chain
-[22] uses composition to discuss the transitivity of information.
+[xu2014objective] uses composition to discuss the transitivity of information.
 Specifically, if $I_{o}:S_{o}(o,T_{h})\rightrightarrows S_{c_{1}}(c_{1},T_{m_{1}})$ and $I_{1}:S_{c_{1}}(c_{1},T_{m_{1}})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ are two information instances, then $I_{2}:S_{o}(o,T_{h})\rightrightarrows S_{c_{2}}(c_{2},T_{m_{2}})$ is also an information instance, interpreted as $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$.
@@
 5.1 Theorem of Model Interpretability
-According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model [16], the following concepts are established based on the MLT-MF framework:
+According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model [glanois2024survey], the following concepts are established based on the MLT-MF framework:
 Definition 13 (Model Interpretability) Let the model $M$, as a finite automaton, operate over an ordered set of times $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$, at which it undergoes assignment, input, output, and state transitions.
@@
 Thus, (5.7) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
-By the maximum independent set theorem for finite hypergraphs[36], $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality.
+By the maximum independent set theorem for finite hypergraphs[Berge1973Graphs], $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality.
 Suppose there exists another ethically safe subset $K$ with
@@
 $p_{M}(x)=\begin{cases}\approx p_{q}(x),&\text{if }x\in S_{ou}\cap S_{oq}\\[3.00003pt] \frac{1-P_{Mq}}{|S_{oq}|-|S_{ou}\cap S_{oq}|},&\text{if }x\in S_{oq}\backslash S_{ou}\end{cases}$ (5.12)
-Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have [29]:
+Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have [goodfellow2016deep]:
 $D_{KL}(p_{q}(x)\parallel p_{M}(x))$ $=\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
@@
 Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their TVD.
-By Pinsker’s inequality [30], we have:
+By Pinsker’s inequality [pinsker1964information], we have:
 $TVD(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\parallel p_{M}(x))}$
@@
 Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance.
-Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD [31], while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss [32].
+Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD [rodriguez2021tighter], while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss [devroye2013probabilistic].
 Therefore, TVD serves as an indicator for estimating a model’s generalization error.
@@
 For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols.
-However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise [37].
+However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise [hamilton1988logic].
 Therefore, they are introduced as defined symbols in subsequent discussions.
@@
 Appendix B Finite Automaton States
-Using the formal expression of states, we can study in depth the special and important object of finite automata [38].
+Using the formal expression of states, we can study in depth the special and important object of finite automata [editorial2002dictionary].
 Lemma B.1 (State Representation of Finite Automata) For any finite automaton $M$, one can describe its input, output, and state‐transition behavior over a related time set $T$ by the state set $S(M,T)$.
 Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions.
-By finite‐automaton theory [38], $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
+By finite‐automaton theory [editorial2002dictionary], $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
 $\delta$: $U\mathbf{\times}Q$ to $U$, is the next‐state function;